### v1

Abstract
The rapid evolution of deep learning (DL) models and the ever-increasing size of available datasets have raised the interest of the research community in the always important field of vision-based hand gesture recognition (VHGR), and delivered a wide range of applications, such as sign language understanding and human-computer interaction using cameras. Despite the large volume of research works in the field, a structured and complete survey on VHGR is still missing, leaving researchers to navigate through hundreds of papers in order to find the right combination of data, model, and approach for each task. The current survey aims to fill this gap by presenting a comprehensive overview of this aspect of computer vision. With a systematic research methodology that identifies the state-of-the-art works and a structured presentation of the various methods, datasets, and evaluation metrics, this review aims to constitute a useful guideline for researchers, helping them to choose the right strategy for delving into a certain VHGR task. Starting with the methodology used for study selection, literature retrieval, and the analytical framing, the survey identifies and organizes key VHGR approaches using a taxonomy-based format in various dimensions such as input modality and application domain. The core of the survey provides an in-depth analysis of state-of-the-art techniques across three primary VHGR tasks: static gesture recognition, isolated dynamic gestures and continuous gesture recognition. For each task, the architectural trends and learning strategies are listed. Additionally, the study reviews commonly used datasets —emphasizing on annotation schemes— and evaluates standard performance metrics. It concludes by identifying major challenges in VHGR, including both general computer vision issues and domain-specific obstacles, and outlines promising directions for future research.
keywords:
Vision-Based Hand Gesture Recognition; Human-Robot Interaction; Human-Computer Interaction; Sign Language; Deep Learning; Gesture Recognition Taxonomy
1 Introduction
Understanding human behaviors—such as hand gestures [2], actions [156], and facial expressions [150]—is a key area of artificial intelligence research. There is growing interest in developing Intelligent User Interfaces (IUIs) that offer more natural and intuitive ways for humans to interact with machines, replacing traditional tools like the keyboard and mouse. This has led to increased focus on Hand Gesture Recognition (HGR) across various applications. For example, HGR is used in human-robot interaction in industrial settings (e.g., robotic arm control [134, 96]), healthcare (e.g., robot-assisted surgery [139]), and autonomous mobile robot navigation (e.g., UAVs [194] and UGVs [17, 105]). It also plays a key role in human-computer interaction, especially in AR/VR environments [32, 199, 5]. In addition, the large population of deaf and hard-of-hearing individuals has driven research into automatic sign language recognition systems, aiming to improve communication and inclusion for these communities [196, 26].
Given the importance of the applications mentioned above, Hand Gesture Recognition (HGR) has grown into a broad research field, generally divided into two main approaches: vision-based and sensor-based. Vision-based HGR relies on visual data such as RGB and depth images, as well as features derived from them—like pose heatmaps or optical flow. Sensor-based HGR, on the other hand, uses data from devices such as wearable gloves or surface EMG (sEMG) sensors, and even non-wearable sources like acoustic signals. Researchers have also explored multi-modal methods that combine inputs from different sources. In this study, we focus on vision-based HGR, as sensor-based methods often involve high cost and complex setup, making them less practical for real-world use. Vision-based approaches better align with the goal of HGR: to enable more natural, unobtrusive, and accessible human-machine interaction.
Many of the published surveys consider both sensor-based and vision-based methods together [140, 162, 64, 6]. As a result, the analysis of the proposed methods in these surveys tends to be superficial, because there are so many methods to cover. Other review papers also present the DAQ devices used by researchers [43], further broadening the range of related publications. The undeniable importance of sign language recognition has led some researchers to conduct surveys dedicated exclusively to this topic [2, 155, 170, 145, 174]. In addition, many authors consider both deep learning (DL) and traditional machine learning (ML) methods, or in some cases only traditional ML approaches [174, 140, 43, 132, 64, 6], while others include outdated methods [162, 123, 132, 7, 3, 84]. Although this broader approach helps lay the foundational understanding of VHGR and facilitates a historical perspective, it often limits the focus on current state-of-the-art methods, which are arguably more critical for driving the field forward.
Some surveys aim to provide a quantitative overview of the field [132, 64, 3, 84], offering bibliometric analyses on various aspects of VHGR. While such surveys present a statistical and objective perspective and help identify major subfields, they fall short in advancing the field due to their limited examination of the existing methodologies.
In contrast, to the aforementioned surveys, this survey emphasizes very recent publications and provides a detailed analysis of their contributions, aiming to identify current challenges and innovations that could serve as a valuable starting point for future research. A comparison with existing reviews is presented in Table 1.
Table 1: Comparison of this survey with existing ones. ML refers to traditional Machine Learning/Computer Vision techniques.
Publication Year Range # Papers Input Data Models Scope
[2] 2021 - - Vision-based DL CSLR
[123] 2021 2014-2020 98 Vision-based DL & ML General
[7] 2021 2014-2021 138 Sensor- & Vision-based DL SLR
[3] 2021 2001-2021 649 Vision-based DL & ML SLR
[43] 2022 - - Sensor- & Vision-based DL & ML SLR
[132] 2022 2000-2020 571 Sensor- & Vision-based DL & ML General
[84] 2022 2010-2021 80 Vision-based DL & ML SLR
[155] 2023 2015-2023 - Vision-based DL ISLR
[162] 2024 2014-2024 - Sensor- & Vision-based DL & ML General
[64] 2024 2018-2023 256 Sensor- & Vision-based ML General
[6] 2024 2015-2024 58 Sensor- & Vision-based DL & ML SLR
[174] 2024 - - Sensor- & Vision-based DL & ML SLR
[140] 2024 - - Sensor- & Vision-based DL & ML General
This survey 2025 2021-2024 290 Vision-based DL General
More specifically, the survey focuses on vision-based methods for three key Hand Gesture Recognition (HGR) tasks, which are discussed in detail in the following sections: static HGR [208, 131], isolated dynamic HGR (IDHGR) [82, 68, 214, 60, 206], and continuous dynamic HGR (CHGR) [122, 27, 74]. Static HGR involves recognizing gestures from single frames or spatial data only. Isolated dynamic HGR extends this by analyzing short, pre-segmented sequences to capture both spatial and temporal patterns. Continuous dynamic HGR goes a step further by detecting multiple gestures within long, unsegmented video sequences. Each task presents its own unique challenges. For example, CHGR often relies on weak supervision, making optimization more difficult. At the same time, all tasks share common issues such as interference from irrelevant background movements or variations in lighting, viewpoint, and hand appearance.
Vision-based Hand Gesture Recognition (HGR) has a long history, beginning with traditional computer vision techniques based on hand-crafted features and evolving into complex deep learning pipelines [124]. Early approaches typically followed a three-stage process: pre-processing, feature extraction, and classification. Pre-processing often involved segmenting the hand region using skin color or other visual cues. Feature extraction relied on descriptors like Histogram of Oriented Gradients (HOG), while classification was usually performed using models such as Support Vector Machines (SVMs) or Hidden Markov Models (HMMs), depending on the task. With the rise of deep learning, these hand-crafted features were gradually replaced by learned representations, which offer better generalization. Modern deep learning architectures have become increasingly deep and sophisticated, especially for dynamic HGR tasks. These models first capture short-term motion patterns and then aggregate them to understand long-term context [122, 73, 62, 75, 74]. Due to their superior performance, this study focuses primarily on deep learning-based methods.
Given the broad scope of vision-based Hand Gesture Recognition (HGR) within computer vision, it is crucial to identify the current state-of-the-art methods for the three tasks mentioned above, as well as the key challenges that have shaped the field’s development. This survey aims to 1) provide an overview of the field, enabling readers to quickly find the most relevant publications for their needs, 2) offer a detailed analysis of each method, highlighting their key contributions, and 3) present the state-of-the-art solutions, based on the previous analysis, and explain the main challenges that continue to drive research in HGR.
The main contributions of this study can be summarized as follows:
  1. 1.
    A structured and comprehensive methodology for reviewing the literature;
  2. 2.
    A broad overview of vision-based HGR, organized according to multiple taxonomy criteria;
  3. 3.
    A detailed analysis of recent methods, highlighting their key characteristics and innovations;
  4. 4.
    A synthesis of current state-of-the-art approaches along with a discussion of the key challenges facing the field.
The structure of this paper is organized to guide the reader through a comprehensive review of vision-based hand gesture recognition (VHGR). Section 2 presents the literature review methodology, outlining the systematic approach used to collect and evaluate relevant works. Section 3 introduces a taxonomy of VHGR methods based on diverse criteria, offering a structured classification of existing approaches. Sections 4, 5 and 6 delve into the three main VHGR tasks –—static, isolated dynamic, and continuous gesture recognition—– discussing representative methods and their underlying principles. Special attention is given to sign language recognition and semaphoric gestures in dedicated subsections. Section 7 presents the main datasets used for the evaluation of the presented methods and Section 8 summarizes the evaluation metrics commonly used in the field. Finally, Section 9 outlines key challenges in Human Gesture Recognition and highlights current solutions, while Section 10 concludes the survey by reflecting on recent trends and highlighting open challenges that shape the future of VHGR research.
2 Literature Review Methodology
The review begins with a clear definition of the research methodology, establishing a structured foundation for the comprehensive exploration of the VHGR landscape. This systematic approach is intended to serve as a valuable reference for researchers and practitioners seeking to navigate and contribute to the field. The initial step involved defining the scope of the study, which informed the formulation of inclusion and exclusion criteria used to guide the selection of relevant publications. Subsequently, targeted queries were conducted across major scientific databases, including Scopus and IEEE Xplore, to retrieve works aligned with the established criteria. The final stage entailed a thorough examination of the selected literature, with the objective of distilling the key innovations, contributions, and methodological insights of each publication. These elements are further elaborated in the following sections.
2.1 Survey Scope
Positioned at the intersection of computer vision and human-computer interaction, this work sets out to map the evolving landscape of vision-based hand gesture recognition. By distilling the most impactful recent contributions, it seeks to clarify the field’s current trajectory and highlight its most pressing research needs. The review is structured around five foundational pillars:
  • Advanced deep learning methods introduced in recent literature;
  • Benchmark datasets employed across various VHGR tasks;
  • Established evaluation metrics used to measure system effectiveness;
  • Key challenges that continue to shape the research agenda;
  • Innovative solutions and future directions aimed at overcoming existing limitations.
2.2 Selection Criteria
To ensure the quality, relevance, and completeness of this review, a well-defined set of inclusion and exclusion criteria was established prior to initiating the literature search. The selection process was guided by three key axes:
  • Relevance: Only works that directly address VHGR were considered. This includes methods focused on hand gesture recognition and sign language recognition based solely on visual modalities such as RGB images, infrared (IR), optical flow, motion history, depth maps, and pose representations (e.g., skeleton landmarks and heatmaps derived either from RGB-based pose estimators or visual data acquisition devices). Studies centered on Human Action Recognition (HAR), sensor-based techniques (e.g., using electromyography or radar), or multimodal approaches combining visual and non-visual data were excluded to maintain a focused scope.
  • Up-to-date: To reflect the current state of the field, priority was given to publications from 2021 onward, particularly those introducing new methods. Exceptions were made for seminal works—defined here as milestone contributions post-2015 with high citation counts—that have significantly shaped VHGR research. In the case of datasets, no strict publication date was imposed, but only those actively used by the community or released within the past four years were included. All survey papers considered were published after 2021.
  • Impact: To further focus on seminal and influential works, publications were sourced from highly cited venues (i.e. having an h5-index of 90 or above on Google Scholar), with a single exception for the IEEE International Conference on Automatic Face & Gesture Recognition, which is closely aligned with VHGR. In addition to the above, and aiming to provide the theoretical context of the study, a small number of milestone papers has also been referenced, though not analyzed in the main methods section.
2.3 Literature Search
To identify publications that satisfy the previously defined criteria, a comprehensive literature search was conducted using specialized queries across several well-established scientific databases, including Scopus, IEEE Xplore, ACM Digital Library, and arXiv for pre-prints. Google Scholar was also used to double-check for publications from major venues that might have been overlooked. The search queries were designed based on the previously outlined inclusion/exclusion criteria, utilizing specific commands (e.g., EXACTSRCTITLE, LIMIT-TO) and boolean operators (e.g., AND, OR, NOT) to refine the results. A list of the keywords that were applied in the title, abstract and index terms of each one publication, can be found in Table 2. This process retrieved over 900 publications in the Scopus database. Out of them, only 210 papers were related to vision-based HGR. A second query has been performed on IEEE Xplore, excluding the non-IEEE venues, and resulted in 633 articles, most of them already listed in Scopus. A few simpler queries were performed on Google Scholar and arXiv, and duplicates were ignored, resulting in 290 papers in total, including review papers, methods, and datasets.
Table 2: Keywords used in literature search
Keywords
gesture recognition, gesture classification, sign language recognition, sign language understanding, pointing recognition, human finger pointing, hand gestures, human gestures, gesture, hand gesture, body gesture
2.4 Categorizing the Identified Articles
After applying the inclusion and exclusion criteria, the remaining publications were carefully reviewed for relevance. Those that did not meet the criteria were excluded from further analysis. The selected papers were then categorized into three distinct groups: methods, datasets, and review papers. Each category was analyzed separately, and the corresponding information was organized into three dedicated tables. Publications relevant to more than one category were included in each applicable table as separate entries. Key information for each paper was systematically recorded in the appropriate table columns, as detailed below:
  • Methods: The name of the proposed method, datasets used for evaluation, a brief description of the architecture, the input modality, the learning strategy, and the evaluation metrics;
  • Datasets: For each dataset, we recorded the following details: name, type of gesture based on temporal characteristics (e.g., static, isolated dynamic, or continuous), data modality, total number of samples, number of subjects, and—when applicable to continuous HGR—the number of gesture instances. For sign language understanding (SLU) datasets, we also noted the language, vocabulary sizes (both sign and text), and total duration (for video-based datasets). Additional attributes include the source of the dataset (e.g., laboratory, television, or web), frame rate, resolution, and the environmental conditions under which the data was recorded.
  • Surveys: For survey papers, a concise description of the content was collected.
(a) The distribution of the selected papers on the top-10 venues, in terms of number of publications.
(b) The number of selected publications grouped by year.
Figure 1: Bibliometric statistics for the selected papers.
Additionally, for each publication, metadata such as title, authors’ names, venue, publication year, citation count, keywords, task, application domain, authors’ motivation, challenges, and future directions were also recorded. This meticulous process resulted in a curated collection of 290 papers. Some bibliometrics are provided in Fig. 1. Notably, the number of review papers is limited, highlighting the research gap that this survey aims to address.
3 Taxonomy of Hand Gesture Recognition Methods
Figure 2: Taxonomy of hand gesture recognition methods.
This section performs an overview of the landscape of vision-based hand gesture recognition methods, including various aspects, such as input, representation, applications, etc.. More specifically, a set of complementary and diverse criteria, each focusing on key aspects of a VHGR system, is considered, resulting into the classification of related works into a corresponding set of main categories (while each class can be further divided into sub-categories, as will be detailed below). The different criteria that are used to organize the research works in the field are graphically illustrated in Fig. 2 and are detailed as follows:
  • Application Domain: This criterion refers to the various domains in which VHGR (Visual Hand Gesture Recognition) solutions have been applied. The following categories are considered:
    • Sign Language Understanding (SLU): Involves the automatic interpretation of gestures in either a single sign language (monolingual SLU [102, 12]) or across multiple sign languages (multilingual SLU [126]).
    • Human-Computer Interaction (HCI): Encompasses any form of interaction between users and computers. Numerous studies focus on HCI [131, 193], with particular emphasis on AR/VR applications [199, 32, 5].
    • Human-Robot Interaction (HRI): Includes gesture-based control of mobile robots such as UGVs [17, 152] and UAVs [211, 194], industrial robotic arms [96], assistive robots [29], and surgical robots [139].
    • Human-Machine Interaction (HMI): Focuses on interfaces between humans and machines, especially in automotive contexts. VHGR enables recognition of traffic gestures (e.g., from police officers) [45, 59, 185] and supports the development of more intuitive in-vehicle user interfaces [124, 138, 53, 115, 52].
  • Task: Complementing the application domains, this criterion focuses on the nature of the gestures themselves, specifically their temporal characteristics. Gestures are broadly classified as either static or dynamic, depending on whether they involve a fixed hand posture or a sequence of movements over time. Dynamic gestures typically consist of three temporal phases: preparation, core (or nucleus), and retraction [124], which reflect the gesture’s evolution from initiation to completion. Illustrative examples are provided in Fig. 3. The recognition of dynamic gestures can be further divided into isolated and continuous, considering only one gesture per data sequence and identifying gestures in a continuous manner, respectively.
    Figure 3: Gesture recognition based on task: a) static [33] and b) dynamic [27] VHGR.
  • Input Modality: In continuation of the previous criteria, which focused on application contexts and gesture types, this criterion examines the nature and variety of visual input data used for VHGR. The choice of input modality directly influences the representation architecture and the deep learning models employed. While RGB frames remain the most commonly used modality, other forms of visual data have also been explored—such as depth maps, infrared (IR) images, skeleton joints (and derived features like bones and motion [38]), pose heatmaps [27, 214], optical flow, and motion history [119]. Representative examples are shown in Fig. 4.
    Based on the type of input, methods are broadly categorized as:
    • Pose-based (or skeleton-based): These approaches exclusively process skeleton data.
    • Appearance-based: These rely on visual cues from RGB, depth, IR, or similar modalities.
    (a) From left to right: RGB, Optical Flow, Depth and IR [124]
    (b) Skeleton joints [102]
    (c) Pose heatmap [27]
    (d) Motion history [119]
    Figure 4: Modalities that are considered for VHGR.
    In terms of the number of input modalities, methods can be classified as:
    • Uni-modal: Utilize a single type of input data.
    • Multi-modal: Integrate two or more modalities. These methods can be further categorized based on their multi-modality fusion strategy [177, 60] (see Fig. 5):
        Early (data-level) fusion: Combines modalities at the input level before processing [49, 180].
        Late (score-level) fusion: Fuses final prediction outputs from independent uni-modal models [119, 167, 51].
        Intermediate (feature-level) fusion: Merges features extracted from separate modalities at a single processing level  [172].
        Multi-level fusion: Enables interaction across modalities at multiple stages of processing [214, 60].
    Figure 5: Multi-modality fusion techniques: a) Early (data-level) fusion, b) Late (score-level) fusion, c) Intermediate (feature-level) fusion, and d) Multi-level fusion. Illustration from [60]
  • Type of Information Conveyed: Building upon the previous criterion, which focused on the nature of the input, this dimension considers the meaning or intent behind the gestures. Specifically, it examines the kind of information gestures are meant to transmit [16], resulting in the following categories (illustrated in Fig. 6):
    • Semaphoric gestures convey specific commands or instructions and are commonly used in gesture-based interfaces [124, 131, 17].
    • Language gestures function similarly to semaphoric ones but are exclusively used in the context of sign languages [12, 158].
    • Deictic (pointing) gestures are used to indicate objects, directions, or spatial targets [127].
    • Manipulative gestures aim to interact with or mimic the manipulation of physical objects [50].
    (a) Semaphoric [86]
    (b) Language [12]
    (c) Pointing [127]
    (d) Manipulative [50]
    Figure 6: Main categories of hand gestures based on type of information conveyed.
  • Camera Point of View: Complementing the previous criteria on gesture content and input modalities, this dimension focuses on the spatial relationship between the camera and the performer [199]. The camera’s perspective influences both the visual characteristics of the gesture and the recognition model design. Three primary viewpoints are considered (see Fig. 7):
    • First-person view (egocentric): The camera is mounted on the performer, capturing the gestures from their own perspective.
    • Second-person view: The camera acts as the direct recipient of the gestures, typically simulating the perspective of a human observer or interacting agent.
    • Third-person view: The camera observes the gesture from an external position, detached from the performer.
    (a) First-person view (egocentric) [199]
    (b) Second-person view [149]
    (c) Third-person view [158]
    Figure 7: Different camera points of view for VHGR.
  • Architecture: Building on the previous criteria, which focused on the nature of input and camera perspective, this criterion concerns the architectural design of the gesture recognition system, regarding the modeling stages performed by the backbone. Specifically, it refers to how spatial and temporal information is processed and integrated to model gesture dynamics. Three main categories are identified:
    • Spatial architectures consider only spatial data, such as isolated RGB frames, and fed them into corresponding DNNs (mainly 2DCNNs) to capture spatial information.
    • Temporal architectures extract only temporal features, utilizing sequential models to process sequences of raw data, such as coordinates of skeleton joints.
    • Spatiotemporal modeling architectures, that simultaneously analyze both the spatial and temporal domain, such as 3DCNNs.
    • Spatial + Temporal pipelines, that firstly extract spatial features (e.g., using 2DCNNs) and secondly pass them through a sequential model (e.g., BiLSTM) to perform temporal modeling.
    • Short-term Spatiotemporal + Long-term Temporal pipelines, which aim, given a sequence of data, to capture local context (that is, they extract spatiotemporal features given small parts of the overall sequence) and then fed them into a temporal model to aggregate the partial information and obtain the global vector representation.
  • Learning Strategy: This criterion explores the different training methodologies adopted by researchers to develop VHGR models, often influenced by data availability, task complexity, and desired generalization. The main strategies identified include:
    • Transfer Learning, where models are pretrained on large-scale datasets (e.g., Jester [117]) and subsequently fine-tuned for specific downstream VHGR tasks [112].
    • Weakly Supervised Learning, often used in Continuous Hand Gesture Recognition (CHGR), where temporal annotations are coarse or unaligned, lacking exact gesture boundaries [188].
    • Knowledge Distillation, in which one part of the network (often a teacher model) transfers knowledge to another (student model) to enhance feature extraction and efficiency [122, 197, 27].
    • Contrastive Learning, used either in a CLIP-based framework to support zero-shot/few-shot recognition by leveraging textual descriptions [22, 146, 23], or to improve the discriminative capacity of learned representations [203, 204].
    • Multi-task Learning, where a single model is optimized to perform multiple related tasks simultaneously, encouraging knowledge sharing and improved generalization [102].
    • Reinforcement Learning, applied to optimize decision-making processes, such as keyframe selection, by rewarding effective representations [46].
    • Few/Zero-shot Learning, which enable the model to generalize to new gestures or languages with limited or no annotated examples, often leveraging semantic or textual guidance [22, 146, 23].
    • Adversarial Learning, used as a data augmentation strategy [13] to improve robustness, as demonstrated in adversarial video synthesis networks [128].
    • Masked Reconstruction Pretraining, inspired by BERT-style architectures, in which models are pretrained to reconstruct masked skeleton sequences, injecting structural prior knowledge (e.g., hand topology) [202, 71, 68].
This comprehensive taxonomy of vision-based hand gesture recognition (VHGR) methods, has been grounded in a diverse set of classification criteria that span from application domains to architectural and training strategies. Among these, the task criterion, distinguishing between static and dynamic gestures based on their temporal properties, emerges as particularly pivotal. It fundamentally shapes not only the selection of appropriate models and input modalities but also the interpretation of gestures in real-world scenarios. Due to its central role in determining the complexity and design of recognition systems, the task criterion serves as a natural point of departure for further analysis. The next three sections delve deeper into this dimension by systematically examining the core types of VHGR tasks—Static Visual Hand Gesture Recognition, Isolated Dynamic VHGR, and Continuous VHGR—each with its own challenges, methodologies, and applications.
4 Methods for Static VHGR
Static VHGR aims to identify gestures relying only on spatial data. As this approach neglects temporal information, is only suitable for recognizing static hand gestures. Proposed methods for static VHGR can be further categorized into single-stage and multi-stage approaches, based on whether the architecture consists of a single model or a multi-model processing pipeline, respectively. Indicative works of each category are briefly analyzed into the two following subsections. The third subsection concludes with the main features of state-of-the-art static VHGR methods, and discusses the open challenges that concern the community.
4.1 Single-stage Approaches
Single-stage approaches rely on a single model, which varies depending on the task. For gesture classification, fine-tuned variants of well-known 2D-CNN architectures are employed, such as VGG and DenseNet, or custom lightweight architectures are trained from scratch. When classification is combined with localization of gestures, variants of YOLO are adopted. As far as it concerns the application, static VHGR methods are used in Sign Language Understanding (SLU) and Human Computer Interaction (HCI). However, a few more general-purpose models have also been proposed:
  • For Sign Language Understanding, the authors in [157] developed G-CNN, a shallow 2DCNN to recognize Indian Sign Language (ISL) signs and slightly outperformed various versions of VGG on a self-collected dataset. The authors in [54] proposed the STN-ND-Net, a combination of a Spatial Transformer Network (STN) that improves robustness to spatial transformations, with a 2D-CNN backbone that uses deconvolution operations to reduce the pixel-wise and channel-wise correlations and eliminates the information redundancy of colored images. The success of vision transformer (ViT) motivated authors of SIGNFORMER[93] to adopt it for ISL recognition, while authors in [161] leveraged both skeleton and RGB data in their dual stream feature extractor and an MLP to perform intermediate fusion. A YOLO variant, namely YOLOv4-CSP[8] has also been developed, to simultaneous detect and identify the hand gesture.
  • For gesture-based Human-Computer Interaction the authors in [198] proposed a lightweight 2D-CNN, named LHGR-Net, which combines a Multiscale Structure (MSS) that aims to make the model scale-invariant via atrous convolution with a Lightweight Attention Structure (LAS) based on CBAM. The authors in [80] proposed 2DPSTPP-Net, a 2D-CNN with four stacked convolutional blocks whose features are pooled using max pooling and a novel method, then concatenated to form the final representation, tailored for HCI scenarios. The authors in [61] modified the YOLOv5 to make it more lightweight and developed a model that can be used for gesture-based user interfaces in smart sports stadiums, improving the experience of spectators and athletes.
  • General-purpose models have also been developed. EDenseNet [171] extends DenseNet by adding convolutional layers between dense blocks. SpAtNet [21] is a lightweight 2DCNN with three streams: two MAFF modules for coarse-grained features and one Interleaved Module for high-level representations; their fused features are fed to an MLP for classification. MMFANet [103] performs gesture detection and recognition using a 2DCNN backbone with task-specific regressor and classifier branches.
4.2 Multi-stage Approaches
Figure 8: Architecture of an indicative multi-stage dual-stream approach for static VHGR named FGDSNet [208].
In contrast to single-stage, multi-stage approaches set up a multi-model processing pipeline. Typically, a hand detector or segmentation model is used to remove the background and normalize the spatial dimensions. These methods are primarily applied in Human Robot Interaction (HRI) and HCI scenarios, as detailed below:
  • Human-Robot Interaction, particularly gesture-based mobile robot navigation, has attracted research interest. URGR [17] is a three-stage approach using YOLOv3 for human cropping, HQ-Net for super-resolution to address long-range scenarios, and a classifier for gesture recognition. Similarly, [153] introduced a two-stage method based on depth data: hand segmentation followed by gesture classification using DRCAM, a 2DCNN with stacked Residual Channel Attention Modules (RCAMs) and dense interconnections.
  • Multi-stage static VHGR methods have also been proposed for Human-Computer Interaction. In [9], a pipeline for egocentric HGR and fingertip localization uses YOLOv2 for hand detection, followed by VGG16 for feature extraction and gesture classification. In [34], a lightweight HCI pipeline segments the hand using DeepLabV3+ or Unet, then classifies the cropped image with a 2D-CNN like MobileNet or EfficientNet. [208] introduced FGDSNet, a two-stage approach with dual-stream segmentation (local detail and semantic branches) and multi-level fusion for feature integration. The segmentation mask and original image are fed into a dual-stream classification module, where mid-level fusion combines features for final classification via an MLP.
4.3 State-of-the-art Methods for Static VHGR
The current SOTA static VHGR methods are summarized in 7, based on input modality, architecture, and performance on public datasets, including HGR1 [88], NUS-I [95], NUS-II [136], ASL-FS [137], Jochen-Triesch II [175], and OUHANDS [118] [^1] For datasets with splits for both subject-independent and subject-dependent modes, only the latter is considered.. Since most methods are evaluated on different or custom datasets, a straightforward performance comparison is not possible.
As shown, most SOTA methods rely on RGB data, with depth-based DRCAM [153] lagging behind RGB methods on ASL-FS. Researchers use attention mechanisms [153, 198, 208], multi-stream approaches [21, 208], and segmentation modules for background removal [208, 34, 153] to improve performance. On OUHANDS [118], ASL-FS [137], and NUS-I [95], single-stage methods outperform multi-stage ones, while on NUS-II [136], FGDSNet [208] with segmentation achieves slightly better accuracy than 2DPSTPP-Net [80]. Thus, background removal does not always improve sign recognition, as single-stage methods often surpass multi-stage models.
Table 3: Static HGR methods and their performance on well-known datasets. The highest accuracies are shown with bold. Methods marked with star (*) perform hand segmentation.
Method Modality Architecture Description Dataset Accuracy
SpAtNet [21] RGB Three-stream 2DCNN, where each stream aims to obtain either high-level or coarse-grained features HGR1 93.33%
ASL-FS 99.67%
NUS-II 96.53%
FGDSNet [208]* RGB Dual stream 2DCNN approach and a custom dual-stream segmentation module to obtain the mask HGR1 94.8%
OUHANDS 97.8%
NUS-II 99.8%
[34]* RGB DL-based Hand Segmenation + Classifier HGR1 98.15%
OUHANDS 89.2%
LHGR-Net [198] RGB BaseNet + MSS (adopting the atrous spatial pyramid pooling) + LAS (based on CBAM) HGR1 93.15%
OUHANDS 98.75%
2DPSTPP-Net [80] RGB Convolutional blocks followed by the proposed pooling method, intermediate features are concatenated to obtain the final representation. NUS-I 100%
NUS-II 99.61%
DRCAM [153]* Depth Attention-based 2CNN with residual connections ASL-FS 93.44%
Gesture-CNN [157] RGB 2DCNN consisting of four convolutional layers Jochen-Triesch II 100%
EDenseNet [171] RGB DenseNet with extra convolution layers ASL-FS 99.99%
NUS-II 99.3%
5 Methods for Isolated Dynamic VHGR
Unlike static VHGR approaches, which often fail to capture temporal dynamics and struggle with motion-based gestures, recent advances in Isolated Dynamic Vision-based Hand Gesture Recognition (IDVHGR) have focused on recognizing individual gesture instances from spatiotemporal inputs such as RGB video sequences or skeleton joint trajectories.
When gestures are part of a sign language, the task is referred to as Isolated Sign Language Recognition (ISLR). A notable distinction is that ISLR models frequently leverage both manual features (e.g., hand movements) and non-manual features (e.g., facial expressions, upper body motion), which can significantly enhance recognition accuracy. As a result, many ISLR approaches go beyond hand-centric analysis to incorporate richer contextual cues.
In the following, we present IDVHGR methods grouped by input modality as (a) unimodal appearance-based, (b) pose-based, and (c) multimodal approaches. Then, we conclude with a summary of key innovations and state-of-the-art (SOTA) performance.
5.1 Unimodal Appearance-based Methods
Unimodal appearance-based methods rely on a single input type—such as RGB, optical flow, infrared (IR), motion history, depth, or depth flow—to classify hand gestures. As shown in Fig. 4, these representations resemble color images, which allows for the application of similar computer vision techniques. These methods can be further categorized based on the architecture of the feature extractor, which typically follows one of three main pipelines: (1) spatiotemporal, (2) spatial + temporal, and (3) short-term spatiotemporal + long-term temporal backbones. The following paragraphs explain each of these architectures in more detail.
5.1.1 Spatiotemporal
Spatiotemporal backbones perform spatial and temporal modeling jointly, typically using 3D CNNs or (2+1)D CNNs—either custom-built or adapted from existing models.
In [182], the authors introduced a custom network, (2+1)D-SLR, composed of five stacked (2+1)D convolutional blocks (i.e., a 2D convolution followed by a 1D convolution). This structure lowers computational cost compared to traditional 3D convolutions, making the model more practical for deployment. To enhance temporal modeling, authors in [20] proposed a trainable key frame selection method based on error correction. VGG16 features are used to identify key frames, which are then preprocessed and fed into a custom 3D CNN for classification.
In [186], a dual-stage multi-stream pipeline for CSL recognition was developed. An adapted EfficientDet model—enhanced with cross-level Bi-FPN connections and a dual-channel spatial attention module (DCSAM)—detects both hands. The full image and cropped hand patches are passed through three parallel 3D-ResNet-18 models (following the P3D scheme), with their features fused and classified via an MLP. Authors in [63] introduced TS3C-Net, a modified 3D ConvNeXt that adds an extra 3×3×3 convolutional branch to each block, enhancing spatiotemporal representation. Finally, [173] addressed limitations in fixed-pipeline designs by combining features from both ResC3D and ConvLSTM using the Selective Spatiotemporal (SeST) module. The fused features are passed through fully connected and softmax layers for classification.
5.1.2 Spatial + Temporal
These backbones first extract spatial features from each frame, typically using 2D CNNs or vision transformers, and then model the temporal dynamics with sequential modules such as GRUs or LSTMs. The resulting sequence embeddings are aggregated, and a classifier (e.g., MLP) outputs the final prediction.
In [65], a method tailored for HCI tasks using only depth data was proposed. The method combined Inception-v3 that extracted spatial features with an LSTM for temporal modeling. For ISLR, authors in [178] introduced a pipeline to recognize Indian Sign Language (ISL) words used by deaf farmers. They combined GoogleNet for spatial encoding with a BiLSTM for sequence modeling.
To reduce inference time, a lightweight ISLR model, called TIM-SLR has been proposed in [183]. The model uses a 2D-CNN for per-frame feature extraction, followed by a Time Interaction Module (TIM)—a zero-parameter unit that fuses frame features into a compact representation. Finally, the Swin Transformer presented in [40] is a two-stage transformer-based model that extracts spatial features per frame, while a causally masked transformer processes the sequence, ensuring that only past and present frame features contribute at each step.
5.1.3 Short-Term Spatiotemporal + Long-Term Temporal
The methods in this category follow a two-stage architecture: the first stage captures short-term spatiotemporal patterns from segments of the input sequence, while the second stage aggregates these features to form a final representation of the entire sequence.
EgoFormer [138] is a model for egocentric dynamic gesture recognition optimized for edge devices, which has been introduced in the context of autonomous driving. EgoFormer combines a custom 3D-CNN with a transformer module and an MLP for final classification. For long-range gesture-based UGV control, authors in [18] proposed a pipeline where a lightweight 2D CNN extracts per-frame features, which are used to select key frames. After cropping the human region, the selected frames are passed through a SlowFast network, followed by a transformer and an MLP.
To address data scarcity in sign language understanding, [22] proposed a zero-shot learning framework. Sign videos are represented using BERT-encoded textual descriptions combined with hand-crafted attributes. This representation guides a dual-stream backbone composed of short-term spatiotemporal modules (e.g., I3D) and long-term models (e.g., BiLSTM).
Figure 9: The zero-shot framework proposed by authors in [22]. Each branch of the dual-stream backbone consists of a spatial-temporal model, such as the I3D, followed by a long-term recurrent model. The BERT is utilized to map textual description into the embedding space. These features are also combined with hand-crafted attributes.
5.1.4 Other
Some appearance-based methods fall outside the previously discussed categories but offer valuable contributions, especially in Sign Language Recogition (SLR). For example, the method proposed in [113] for Indian SLR (ISLR) compresses an entire frame sequence into a single image, which is then processed by a MobileNet. A second MobileNet extracts per-frame features, which are passed through a two-layer LSTM. The outputs of both streams are concatenated and refined using a custom shallow 2D-CNN. A different approach was presented in [94], where the authors introduced a dual-stream ISLR architecture combining Inception V4 with a novel convolutional module called stem.
In the context of pose-based ISLR, [110] proposed a dual-stream model with keyframe selection. A 2D-CNN and attention mechanism filter out less informative frames. The remaining keyframes are processed by two branches: the Hand stream, which crops and analyzes hand regions with attention blocks, and the Full stream, which processes the entire frame. The fused features are used for final classification. To improve signer-independence and reduce overfitting, the SwC GR-MMixer [109], another lightweight model based on MLP-Mixer, incorporates GRU units for temporal modeling and a shifted window concatenation strategy to enhance sequence representation.
5.2 Pose-based Methods
Pose-based methods, also known as skeleton-based approaches, rely on structured motion data—such as skeleton graphs and pose heatmaps—to classify gestures. These methods often incorporate various forms of pose information, including joint coordinates, bones, motion trajectories, and joint angles, sometimes in combination [38]. They also explore joint subsets related to specific body regions [135], often adopting multi-stream architectures to handle these diverse inputs.
Given the natural graph structure of skeletal data, many methods employ Graph Convolutional Networks (GCNs) originally developed for human activity recognition. In the sections that follow, we group the approaches based on the structure of their backbone, following the same organization as in the previous subsection. Notably, pose information alone is often insufficient for accurate sign recognition [125]. Consequently, many models integrate additional cues—typically from RGB inputs—using 3D CNNs and perform score-level fusion to enhance performance.
5.2.1 Spatiotemporal
Pose-based methods that incorporate spatiotemporal representations simultaneously process spatial and temporal data to improve sign language understanding (SLU), sign language recognition (SLR), and human-computer interaction (HCI).
In Sign Language Understanding (SLU), pose-based methods aim to capture dynamic hand and body movements for gesture recognition. SKIM [104] uses a dynamic graph convolutional network (GCN) combined with temporal convolution (TCN) to analyze both global and local body regions, with a feature extraction module focusing on high-motion joints. Similarly, MC-LSTM [133] employs multiple streams (face, hand, and whole body) processed by temporal convolutions and GCN for improved feature ensembling. Another notable method, P3D [98], uses a transformer-based model with part-wise and whole-body encoders to process landmark data for classification. For improved performance, HWGAT [135] utilizes hierarchical spatial windows with attention mechanisms to highlight key body parts. TMS-Net [38] proposes a six-stream GCN-based approach for processing skeleton joints, bones, and motion data, with inter-stream information exchange. Additionally, methods like SLGTformer [168] and VSCT [25] address the challenge of data scarcity by leveraging transformers and few-shot learning, respectively.
In Sign Language Recognition (SLR), pose-based methods often focus on improving the accuracy and robustness of sign classification. Authors in [161] present a dual-stream GCN model for Korean SLR, analyzing joint and motion data. Similarly, MULTISTAGE-CGAR [121] proposes a dual-stream architecture with separable temporal convolution and attention mechanisms for improved feature extraction. For multi-stream fusion, SwC GR-MMixer [109] combines GCN with GRU units for temporal modeling and attention mechanisms to reduce overfitting. Notably, L-TSSI-DenseNet [97] explores skeleton graph representations by converting joint coordinates into image-like matrices processed by DenseNet for classification. Moreover, STST [201] leverages a two-stage transformer-based approach for global and local spatial-temporal feature extraction, offering a dynamic alternative to traditional methods.
In Human-Computer Interaction (HCI), pose-based methods address challenges such as data overfitting and the need for self-supervised learning frameworks. MAE-based frameworks [79] use a masked autoencoder to reconstruct skeleton graphs, focusing on motion and spatial relationships. 3sISTGCN [167] proposes a three-stream model to process joints, bones, and motions, applying modified ST-GCN layers and late fusion for improved classification. Another approach, DG-STA [5], tackles class-incremental learning in HCI by training models on non-overlapping tasks without access to old training data. DSTSA-GCN [30] addresses temporal expression limitations by stacking three stages of graph convolutions and multiscale temporal convolutions. Additionally, DSTA-SLR [77] improves spatiotemporal modeling by introducing dynamic adjacency matrices and multi-stream fusion, enhancing the accuracy of joint, bone, and motion-based recognition tasks.
5.2.2 Spatial + Temporal
Many models use a two-step approach, where the spatial domain is first processed by a spatial perception module to extract spatial features. These features are then passed into a sequential model for temporal aggregation, resulting in a combined spatial-temporal representation of the input sequence. For instance, GCN-BERT [176] utilizes a two-layer GCN for spatial modeling, followed by BERT for temporal modeling.
Other general-purpose methods adopt a multi-stream, two-stage pose-based approach. In [129], three streams, each containing three 1D convolutional layers, are combined with a parallel module of average pooling layers. The extracted features are then concatenated and input into an LSTM, followed by a fully connected layer and softmax for classification. Similarly, STr-GCN [165] employs a two-stage model where a GCN spatial feature extractor is first applied to obtain intra-frame features and then passes the features through a transformer encoder for modeling inter-frame temporal relationships. The final classification scores are generated by an MLP.
5.2.3 Short-Term Spatiotemporal + Long-Term Temporal
In this approach, the backbone first extracts features from smaller segments of the skeleton sequence to capture short-term spatial-temporal relationships. Then, long-term temporal modeling is applied to aggregate global context across the entire sequence.
For example, gesture recognition in autonomous driving has become an important research area. In [59], the authors propose a pose-based approach called MG-GCT for recognizing police traffic gestures. The model uses two graph convolution streams: one to process skeleton joints and another to analyze joint motions. The features extracted from both streams are fused and passed through a transformer encoder for temporal modeling, with the final classification performed by an MLP. In another approach, TD-GCN [107] introduces a pose-based method for IDHGR. Unlike traditional methods that use fixed skeleton topologies, TD-GCN employs temporal-sensitive topology learning to enhance feature quality. First, graph convolution is used to extract high-level spatiotemporal representations, which are then used to calculate temporal and channel-independent adjacency matrices. These matrices help model the skeleton, and the features are fused using the proposed TCF module.
5.2.4 Temporal
Some pose-based methods bypass spatial modeling and focus directly on sequential processing. Typically, keypoints are slightly preprocessed—often by extracting geometric features such as bone angles—and then passed into sequential models like RNNs or TCNs to produce a vector representation of the input sequence.
In [139], the authors proposed a pose-based method for surgical robot teleoperation. Preprocessed skeleton joints are fed into an LSTM, and the resulting features are passed to an MLP for final classification. The model was evaluated on a custom dataset. Similarly, in [151], a cross-domain gesture recognition approach was developed using geometric features (e.g., bone angles), which are processed by a TCN. The resulting features are summarized into a sequence descriptor, and a linear classifier predicts the final class.
To address data scarcity, the authors in [128] introduced AVSN, an adversarial learning-based data augmentation framework. The method jointly optimizes augmentation parameters and the HGR classifier. A BiLSTM serves as the discriminator, while an MLP acts as the generator. The generator produces challenging augmented data that exploits weaknesses in the discriminator, which in turn learns to correctly classify gestures.
5.2.5 Other
Some architectures adopt more complex designs or alternative skeleton representations, placing them outside the previously defined categories.
In [1], the authors combined a Fourier CNN with a BiLSTM to extract spatiotemporal features from 3D skeleton joints and perform ISLR. A trainable binary mask was also used for feature selection to eliminate redundancy. Similarly, authors in [10] performed ISLR using a feature thresholding technique that transforms skeleton features based on their variance and importance. These transformed features undergo exhaustive search for selection and are then processed by a deep CNN for classification.
To address limitations of color-coded feature maps, the authors in [89] proposed the Joint Motion Affinity Map (JMAM) as a new skeleton representation for ISLR. JMAM is input to a custom multi-stream attention-based 2DCNN, which extracts features and performs classification. Key frame selection was explored in [46] through the Frame Selection Network (FS-Net), trained using reinforcement learning. The agent selects frames for classification by DD-Net and receives a reward of +1 or -1 based on the correctness of the classification outcome. Finally, [101] proposed a lightweight dual-stream method to fully capture spatiotemporal features. One stream employs stacked attention-based modules, while the other uses an autoencoder. The outputs of both are classified independently, and their predictions are averaged to yield the final decision.
5.3 Multimodal Methods
Multimodal approaches combine different input types (e.g., RGB, Depth, Skeleton, Text) to improve gesture recognition performance by leveraging complementary information. IDVHGR multimodal methods do not seem to follow any trend, maybe due to the wide spectrum of alternative options (combinations of streams, fusion techniques, modalities, etc.), so we group them by application domain.
5.3.1 General Purpose
In [100], a dual-stream 3DCNN followed by average fusion was proposed for RGB-D gesture recognition. To filter out redundant visual features ($\upsilon$) and retain only the most discriminative components ($z$), a memory bank mechanism was introduced. The training objective maximizes mutual information between $z$ and the gesture class $y$, while minimizing that between $z$ and $\upsilon$. The memory bank is updated via read/write operations during training. Similarly, addressing redundancy, authors in [206] applied Neural Architecture Search (NAS) and attention mechanisms. A modified I3D backbone replaces some inception modules with learnable NAS cells. These progressively transition from small-kernel to dilated convolutions. Dynamic Static Attention (DSA) modules focus on motion and hand/arm regions, aided by a parallel HeatmapNet trained jointly through an auxiliary loss. For RGB-D input, late fusion is used.
To move beyond the limitations of late fusion and exploit modality interactions more effectively, [207] proposed a decoupling-recoupling framework. Independent spatial (with a Focused Region Proposal (FRP) module) and temporal modeling (via attention) are first performed to extract modality-specific features. These are then fused mid-way using a shared MLP-AE module. Multi-level fusion was further studied in [60], where 3DCNN backbones are combined with a Convolutional Transformer Fusion Block (CTFB). Injected at multiple layers, CTFB fuses intermediate features across streams for better spatial-temporal integration.
To address data scarcity, several works explored zero-shot learning (ZSL) by incorporating textual gesture descriptions encoded by pretrained language models (mainly BERT). In [146], a dual-stream model combined a ViT for spatial modeling and an LSTM-MLP for temporal modeling, learning visual features aligned with textual embeddings. Similarly, [147] fused skeleton-derived geometric features with RGB-C3D-LSTM-AE representations into a shared semantic space for classification. To reduce computational cost, [53] replaced the transformer’s self-attention mechanism with convolutional layers and a gated feedforward network, resulting in the lightweight ConvMixFormer architecture.
In [180], knowledge distillation and contrastive learning were employed to enhance multimodal performance. Three 3DCNNs (for RGB, Depth, and Optical Flow) extract features fused via Coordinate Attention (CA). I3D serves as the teacher, and ResNet as the student. CLIP is also integrated for zero-shot gesture representation. The limitations of standard 3DCNNs were explored in [114], where 3D Central Difference Convolution (3D-CDC) replaces standard 3D convolutions. Residual hybrid blocks and a transformer module handle temporal modeling. [195] also adopted 3D-CDC in a multi-stream, multi-rate setup with multi-level fusion. NAS optimizes the configuration of CDC types, inter-stream connections, and frame rates.
5.3.2 Sign Language Understanding
Several recent works have proposed multimodal and pose-centric architectures for Isolated Sign Language Recognition (ISLR), focusing on combining RGB and skeleton data.
A prominent line of research is the SignBERT framework [68], which uses a spatial GCN with temporal and hand chirality encoding to extract gesture features. These are fed into a transformer encoder, while a hand-model-aware decoder reconstructs masked inputs during self-supervised pretraining. Predictions from a parallel RGB network are fused at inference time. An enhanced version, SignBERT+ [71], introduces spatial-temporal position encoding and an improved fusion module for RGB and pose data. The architecture is illustrated in Fig. 10. Similarly, MASA [203] adopts the same GCN+Transformer backbone, trained with contrastive learning on pose reconstruction. BEST [202] extends this by embedding pose triplets (left hand, right hand, body) with BERT and applying late fusion with RGB data.
Figure 10: The pose-based stream of the SignBERT+ [71], an extension of the SignBERT [68], during SSL pretraining. A GCN is first utilized to extract the gesture features, which are after combined with spatial-temporal positing encoding and subsequently pass through a transformer encoder. During pretraining, a MANO-based decoder is placed after the transformer to perform reconstruction of the masked input. The decoder is replaced with a prediction head during downstream fine-tuning.
To improve feature invariance, a contrastive approach was proposed in [204], enforcing consistency among hand, trunk, and motion features. The same GCN+Transformer architecture is used, with late fusion of pose and RGB predictions. The SAM-SLR framework [82] combines six streams (RGB, depth, skeleton) with late fusion of predictions. SAM-SLR-v2 [81] adds a 3D hand graph stream and learnable fusion weights (Fig. 11).
Figure 11: The architecture of SAM-SLR-v2 [81], which is based on the SAM-SLR [82]. The architecture consists of seven streams, regarding various modalities, including RGB frames, optical flow, pose heatmaps and skeleton joints. The predictions of each branch are fused in a learnable late manner.
In VTN-HC/PF [36], hand-cropped images are processed via a 2D-CNN for spatial modeling, followed by a transformer and MLP for temporal and final predictions. The PF variant incorporates geometrical features from pose data. Motion History Images (MHIs) were employed in [119], where a 3DCNN with attention and residual connections analyzes MHIs, while a dual-stream version adds a 2DCNN for RGB data, with predictions fused via weighted summation.
To address data scarcity, ZSSLR [216] adopts a CLIP-based zero-shot framework using RGB and pose data. In [24], a few-shot cross-lingual method is proposed, combining two 3DCNNs and a pose estimator. Features are merged through a Temporal Feature Aggregation Module with self-attention and 1D convolution.
NLA-SLR [214] targets semantically similar signs using a multistream backbone (RGB and pose heatmaps via S3D), followed by a head with language-aware label smoothing and an auxiliary inter-modality mix-up classifier. MSNN [116] uses five I3D streams (RGB and optical flow) plus ST-GCN for pose. hDNN-SLR [142] introduces a hybrid deep network with a 3DCNN for spatial detection, an attention-based BiLSTM for temporal modeling, and a VAE-based module for feature discrimination. Finally, Uni-Sign [102] proposes a unified multimodal architecture with three pose encoders and an RGB encoder fused via an attention-based module. Features are passed through temporal encoders and a pretrained LLM. Training involves pose-only and joint RGB-pose pretraining followed by downstream fine-tuning on ISLR and CSLR tasks.
5.3.3 Human-Computer Interaction
Motivated by the difficulty of recognizing fine-grained gestures in HCI applications, authors in [200] proposed HandSense, a method relying on depth and RGB data. Two 3DCNNs are deployed to extract features for each one modality. These features are fused and fed into a SVM to obtain the final prediction. Authors in [87] proposed an approach relying on RGBD Dynamic Images (RDDIs). These images, are obtained by the combination of RGB and estimated depth data. The resulted RDDI is fed into a modified Xception, resulting from removing the top 4 Xception modules, in order to extract meaningful features. These features are finally used to perform the classification.
5.3.4 Other
Inspired by the success of transformers, GestFormer [51] introduces a lightweight transformer-based framework for gesture-driven HRI. A 2DCNN extracts per-frame features, processed by PoolFormer layers (replacing attention) and two custom blocks for multiscale pooling/WT and token mixing. Each modality is processed independently, and outputs are averaged. In [49], a hybrid 3DCNN/LSTM model fuses RGB, depth, and skeleton data (from an enhanced pose estimator) via weighted sum. The fused input feeds a spatiotemporal model combining 2D/3D convolutions and ConvLSTM layers.
5.4 State-of-the-art Methods for Isolated Dynamic HGR
Recent IDHGR approaches demonstrate high architectural diversity. Pose-based methods often rely on spatial GCNs and temporal transformers [68, 71, 202, 204, 203, 165], or on ST-GCN variants [107, 79, 167]. Due to their limited standalone performance—mainly caused by pose estimation failures in challenging conditions (e.g., motion blur) [203]—pose-based models are frequently fused with simple RGB-based branches (e.g., I3D), using late fusion via score averaging [68, 71, 202, 204, 203, 104].
Appearance-based methods typically follow a multi-stream 3DCNN design to process individual modalities (e.g., RGB, depth, optical flow) [180, 206, 60, 100, 119, 195], employing late or multi-level fusion. Notably, SAM-SLR [82] and SAM-SLRv2 [81] show that adding more modalities—even with simple backbones—can significantly improve accuracy.
Advanced learning strategies further improve performance. Self-supervised learning (SSL) via masked pose reconstruction [202, 68, 71, 79, 203] enhances representation quality by injecting hand topology priors. Neural Architecture Search (NAS) has been used to discover optimal network structures [206, 195], while multi-task learning, as in Uni-Sign [102], boosts performance by leveraging auxiliary tasks and larger training corpora. Performance comparisons across benchmarks are summarized in Tables 4, 5, and 6.
Table 4: Brief description and reported accuracy of state-of-the-art methods on IsoGD [179].
Method Modality Architecture Description Training Setting Accuracy
3SCKI [180] RGBD, OF Three-stream 3DCNN with mid-level fusion Knowledge distillation between two instances of the pipeline, CLIP 65.87%
SlowFast+NAS [195] RGBD, OF Multi-stream and multi-rate 3DCNN with 3D-CDC and multi-level fusion NAS 66.23%
RAAR3DNet [206] RGBD Dual stream I3D with adoptable cells, spatial and temporal attention Multi-task loss and NAS 66.62%
De+Recouple [207] RGBD Dual-stream spatial + temporal with AE for multi-level fusion and mechanisms to reduce redundancy Self-distillation for inter-frame learning, extra supervision to enhance the AE 66.79%
CTFB [60] RGBD Dual-stream I3D with multi-level fusion - 69.02%
[100] RGBD Dual-stream Res3D-18 + average fusion Memory network to enhance the discriminative capability of the extracted features and information-theoretic loss function 74.08%
Table 5: Performance of uni-modal pose-based methods on SHREC [166] and DHG [37].
Method Architecture Split SHREC DHG
[120] Three-stream approach: attention-based spatial-temporal model, attention-based temporal-spatial model and MLP 14 97.01% 92.00%
28 92.78% 88.78%
MS-ISTGCN [167] Three-stream model: ST-GCN with adoptive spatial GC, extended temporal GC and attention 14 96.7% 93.7%
28 94.9% 91.2%
TD-GCN [107] ST-GCN variant with channel-dependent and temporal-dependent adjacency matrices 14 97.02% 93.9%
28 95.36% 91.4%
[101] Two-stream approach consisting of an attention-based dual branch module and MAE 14 96.9% 94.21%
28 94.17% 92.11%
SBI-DHGR [129] Multi-stream 1DCNN + BiLSTM 14 97% 94.64%
28 92.36% 91.79%
STST [201] Two-stage transformer-based approach 14 97.62% 94.82%
28 95.83% 93.18%
DSTSA-GCN [30] Three stacked stages, each one consisting of group channel-wise and temporal-wise GCN, followed by multiscale TCN 14 97.74% 95.04%
28 95.37% 93.57%
DRL-DDNet [46] DD-Net for classification and FS-Net for keyframe extraction 14 96.5% -
28 93.8% -
Table 6: Performance of the current SOTA methods on ISLR. PI, PC, OF and DF stands for per-instance accuracy, per-class accuracy, optical flow and depth flow, respectively. Methods marked with star (*) use additional training data.
Method Modality Architecture Description Training Setting Dataset PI PC
BEST [202] RGB, Pose d-VAE + GCN + Transformer & RGB-based encoder + late fusion SSL pretraining (masked tokens reconstruction) MSASL 71.21% 68.24%
WLASL 54.59% 52.12%
CSL500 97.7% -
SignBERT [68] RGB, Pose Spatial GCN, temporal and chilarity embedding + Transformer & RGB-based stream + late fusion SSL pretraining (MANO-based masked hand reconstruction) MSASL 71.24% 67.96%
WLASL 54.69% 52.08%
CSL500 97.6% -
SignBERT+ [71] RGB, Pose Extension of SignBERT with Spatial-Temporal Position Encoding and enhanced late fusion technique The same as SignBERT MSASL 73.71% 70.77%
WLASL 55.59% 53.33%
CSL500 97.8% -
MASA [203] RGB, Pose Spatial GCN + Transformer & RGB-based stream SSL pretraining (masked input reconstruction), contrastive learning WLASL 55.77% 53.13%
TMS-Net [38] Pose Multi-stream decoupled GCN with multi-level fusion - WLASL 56.4% -
SKIM [104] Pose Multi-stream channel-wise multi-scale GCN + late fusion Hand part mixing for data augmentation WLASL 57.73% 55.40%
[204] RGB, Pose GCN+Transformer and I3D + late fusion Contrastive learning to enhance invariance, consistency constraints between features of hand and trunk parts and between joints and motions MSASL 74.18% 72.18%
WLASL 58.06% 55.66%
CSL500 97.8% -
SAM-SLR-v2 [81] RGBD, Pose, HHA, OF, DF Extension of SAM-SLR-v2 with 3D graph stream and learnable weights for late fusion - WLASL 59.39% 56.63%
CSL500 99.00% -
NLA-SLR [214] RGB, Pose Four-stream network based on S3D with two frame rates and multi-level fusion Language-aware label Smoothing and inter-modality mixup with an auxiliary classifier MSASL 73.80% 70.95%
WLASL 61.26% 58.31%
Uni-Sign [102]* RGB, Pose A pretrained LLM process features extracted by pose, visual and temporal encoders Three-stage training, multi-task learning with additional training data MSASL 78.16% 76.97%
WLASL 63.52% 61.32%
6 Methods for Continuous Dynamic HGR
In contrast to isolated dynamic hand gesture recognition (IDHGR), where each sample contains at most one gesture, continuous hand gesture recognition (CHGR) deals with unsegmented gesture sequences. The goal is to recognize gestures from untrimmed videos and generate the corresponding gesture sequence. CHGR is typically treated as a weakly supervised task, as the precise temporal boundaries of gestures are not provided; the model learns to identify them implicitly, often using Connectionist Temporal Classification (CTC) loss [188, 55].
While CHGR has been explored in domains like autonomous driving [45, 185], most research focuses on Sign Language Understanding (SLU), where it is commonly referred to as Continuous Sign Language Recognition (CSLR) or sign2gloss. CSLR aims to produce a sequence of glosses [^2] Glosses are atomic lexical units in sign language, analogous to words in spoken language. from video input. Due to the importance of non-manual features (e.g., facial expressions and upper-body movements), many approaches also incorporate these regions to improve recognition accuracy. Some CSLR methods are further extended to support gloss-based sign language translation (SLT) by adding a gloss-to-text module [56, 215, 27].
In the following, we review recent advances in CHGR, grouped by input modality. We highlight key architectures, learning strategies, and benchmark results.
6.1 RGB-based
Most state-of-the-art RGB-based Continuous Hand Gesture Recognition (CHGR) models follow a three-stage pipeline architecture: (1) a spatial model (commonly a 2DCNN or a transformer) to extract frame-wise features; (2) a short-term temporal model (mainly 1DCNN or TCN) for local temporal modeling and gloss-level feature extraction; and (3) a contextual module (e.g., BiLSTM) for long-term sequence modeling and final gesture sequence prediction [73, 122, 62, 74, 197, 70, 75, 72, 76]. This standard architecture was introduced in [31]. This decoupled spatial-temporal structure has been favored over unified spatiotemporal models like 3DCNNs due to the latter’s inability to define clear temporal boundaries [2].
Across the various methods that extend the initial three-stage pipeline, six broad enhancement strategies emerge: 1) Auxiliary Supervision & Multi‐Task Objectives, which enrich the standard three‐stage pipeline with extra losses and segmentation tasks; 2) Specialized Temporal Modeling, replacing or augmenting the 1D‐CNN with learned pooling, multi-scale fusion, or hierarchical boundary detectors; 3) Attention & Fusion Modules, inserting correlation, self‐emphasis, or transformer‐fusion blocks into the visual encoder; 4) Backbone Replacements & Hybrids, swapping 2D-CNNs for vision transformers or hybrid CNN–attention architectures; 5) Incorporating Linguistic Priors, aligning visual features with gloss embeddings via cross-modal decoders or contrastive losses; and 6) Data Scarcity & Multilingual Strategies, leveraging self-supervision, prototype memories, and automatic dictionary construction to overcome limited annotations and support multiple sign languages.
Many RGB‐based CHGR methods build on the canonical 2DCNN + 1DCNN + BiLSTM architecture by introducing auxiliary losses and multi‐task objectives. For example, VAC‐CSLR [122] augments the visual and temporal modules with a Visual Enhancement (VE) loss and a Visual Alignment (VA) distillation loss to accelerate convergence and improve representation quality (see Fig. 12). Similarly, SMKD [62] adds two auxiliary classifiers—one after the 1DCNN and one after the BiLSTM—to perform synchronous gloss segmentation alongside recognition, using a three‐stage training schedule to mitigate gradient “spikes.” C2SLR [213] enforces Spatial Attention Consistency (SAC) by focusing the visual module on pose‐informed regions, and Sentence Embedding Consistency (SEC) to align the visual and sequential representations of the same gloss.
Figure 12: An overview of the VAC-CSLR [122]. The feature extractor provides local information, that is processed by the subsequent alignment module to predict the gloss sequence. $F_{a}$ and $F_{b}$ denote auxiliary classifiers, while $L_{VE}$, $L_{VA}$ and $L_{CTC}$ the VE, VA and CTC loss, respectively. This training setting constitutes the baseline for most of the proposed methods.
To better capture local temporal patterns, several works replace or enhance the standard 1D-CNN. Temporal Lift Pooling (TLP) [73] adopts a learned pooling scheme inspired by the lifting transform, eliminating handcrafted kernels. mLTSF‐Net [192] couples content‐aware feature selection with position‐aware convolutions to fuse adjacent‐frame similarities at multiple scales. HS‐I3D [190] extends I3D by extracting intermediate embeddings and deploying a U‐Net–style decoder that upsamples temporally, enabling simultaneous sign detection and boundary localization.
Other methods inject lightweight attention and correlation blocks into the spatial encoder. CorrNet [74] adds parallel correlation and identification modules after each 2DCNN stage to model inter‐frame trajectories and highlight informative regions; CorrNet+ [76] further replaces its correlation operator with a more efficient variant and inserts an attention layer. SEN‐CSLR [75] introduces self‐emphasizing spatial (SSEN) and temporal (TSEN) modules to reweight feature maps and keyframes with minimal overhead. Convolutional Transformer Fusion Blocks (CTFB) [60] fuse intermediate features across 3DCNN branches to enhance cross‐modal integration.
A growing number of approaches swap the 2D-CNN for transformer‐based encoders or hybrid CNN–attention modules. CLIP‐SLA variants [14] leverage a frozen CLIP visual encoder, augmented with LoRA or MLP adapters and an auxiliary classifier for early supervision. Swin‐MSTP [15] pairs a Swin transformer with an MS‐TCN that employs TLP and dilated convolutions, followed by a BiLSTM for global context. TCNet [111] enhances a CNN backbone with trajectory and correlation modules, while SlowFastSign [4] modifies Slow‐Fast‐101 with bidirectional fusion pathways and auxiliary temporal classifiers to refine feature exchange. Most CSLR methods still rely on CNN backbones and thus struggle to model correlations across spatial regions. To address this, SignGraph [^3] Be aware of an IDHGR method with the same name. [47] treats each frame’s patches as nodes in a dynamic graph and alternates between Local Sign Graph (LSG) and Temporal Sign Graph (TSG) modules to capture intra‐frame and inter‐frame relationships; a multiscale variant, MultiSignGraph, further enhances spatial–temporal context (Fig. 13).
Figure 13: An illustration of the SignGraph [47]. The RGB frames are split into patches. Dynamic graphs are utilized to represent the relations between intra-frame and inter-frame patches, that constitute their nodes.
To exploit gloss co-occurrence and language structure, some models fuse linguistic embeddings with visual features. SignBERT‐HKSL [212] combines a (3+2+1)D ResNet with BERT and BiLSTM, pretraining on masked inputs and iteratively fine‐tuning with masked outputs. C2ST [197] replaces the CTC decoder with a Conditional Gloss Decoder (CGD) that models gloss dependencies, and aligns visual and gloss embeddings via cross‐modal match loss. GPGN [58] introduces a Cross‐Modal Match Loss between BERT‐derived gloss features and PDC‐TFE–extracted visual embeddings, followed by a BiGRU for sequence modeling (see Fig. 14). CVT‐SLR [205] pretrains an asymmetric VAE on a pseudo gloss‐to‐gloss task, then bridges the VAE and visual encoder with a video–gloss adapter and contrastive alignment loss (see Fig. 15). Fully transformer‐based PiSLTRc [191] applies content‐aware neighborhood gathering and position‐informed attention for end‐to‐end CSLR. Authors in [72] also introduce a prior-aware cross-modality augmentation that trains a transformer to reconstruct masked gloss sequences and then uses it to generate pseudo video–text pairs by guiding gloss substitutions and insertions.
Figure 14: Authors in GPGN [58] leveraged the linguistic prior from the pretrained BERT. The Cross-Modal Match Loss is introduced, in order to maximize the correlation between the gloss and the visual embeddings.
Figure 15: The approach proposed by authors in CVT-SLR [205]. In the first training stage, a VAE is pretrained on the gloss2gloss pseudo-translation task in order to gain textual knowledge. In the second stage, an adopter is placed between the VAE and the visual encoder, while the consistency of the visual and temporal encoder is maintained via contrastive learning.
Finally, data augmentation and cross‐lingual strategies help alleviate limited annotations. Fangyun Wei et al. [188] build a “dictionary” by using a pretrained CHGR model to automatically segment continuous videos into isolated signs, then mapping German and Chinese sign corpora into a shared embedding space. CSGC [143] maintains a Gloss Prototype memory that stores class prototypes updated during training; a gloss contrastive loss then enforces cross‐sentence consistency.
6.2 Pose-based
Pose-based (or skeleton-based) approaches for CHGR rely exclusively on body pose information, typically in the form of 2D/3D skeleton joints or pose heatmaps. These methods aim to model human motion while discarding appearance-based noise, making them attractive for real-time and privacy-preserving applications.
Several works exploit structured keypoint information by dividing the body into semantic regions such as the face, hands, and upper body. MSKA-SLR [56] processes each region independently via attention-enhanced temporal convolutions across four branches. The resulting features are fused to decode gloss sequences using a CTC loss, complemented by frame-level self-distillation. Similarly, CoSign [83] employs group-specific GCNs for each subset and introduces regularization to preserve inter-stream consistency. Their dual-stream variant integrates joint positions and motion features using multi-level fusion to capture both static and dynamic cues.
Temporal localization of gestures has also been explored in the literature. TY-Net [39] draws inspiration from YOLO to detect gesture instances in time by predicting their boundaries and centers, analogous to spatial bounding boxes.
Beyond joint coordinates, pose heatmaps serve as an alternative input modality. The multi-cue framework in [210] combines spatial and temporal reasoning by localizing facial and hand regions via a pose estimator and subsequently encoding them with a temporal module, before decoding gloss sequences with a CTC head. The FA-STGCN model [45] further improves robustness to motion scale and frequency variability. It normalizes skeleton inputs and applies alternating spatial and temporal GCNs. Spatial modules learn latent joint relationships, while temporal processing is split into global and local paths—extracting coarse-grained and fine-grained motion patterns, respectively.
Finally, in the context of autonomous driving, Wang et al. [185] developed a two-stage pose-based system to classify road intersection gestures. The first stage extracts geometric features and encodes them via LSTM layers. The second stage uses a dual-stream setup, where one branch processes geometric descriptors and the other uses a ConvLSTM to handle co-occurrence matrices. Features are fused with directional orientation indicators for final classification.
6.3 Multimodal
Multimodal approaches for CHGR combine multiple input types—such as RGB frames, pose heatmaps, depth maps, and optical flow—to mitigate the limitations of unimodal methods and capture complementary information from different data sources.
A prominent example is TwoStream-SLR [27], a dual-stream architecture designed for both signed language translation (SLT) and CSLR. It processes RGB and pose heatmaps in parallel using the first four blocks of S3D in each branch, with bidirectional lateral connections enabling multi-level information exchange. Additional supervision is provided through Sign Pyramid Networks and auxiliary prediction heads (video, keypoint, and joint). The model is trained using self-distillation, minimizing KL-divergence between the outputs of the three heads. An illustration of the architecture is shown in Fig. 16. This model has also been applied to ISLR [215], using a sliding window mechanism and leveraging the same framework to extract sign boundaries and construct training dictionaries.
Figure 16: Overview of the TwoStream-SLR [27]. Eight S3D blocks constitute the dual-stream multimodal architecture. Bidirectional Lateral Connection is utilized to exchange the information produced by each branch in a multi-level manner. The Sign Pyramid Networks, Video and Keypoint head are actins as auxiliary classifiers, providing extra supervision. Self-distillation is also employed to share the knowledge between the two streams.
Other CSLR models have adapted multimodal ISLR backbones. Uni-Sign [102] can be used directly for CSLR, while SignBERT+ [71] can be extended by appending temporal pooling and a CTC decoder to generate gloss sequences.
Beyond sign language, general-purpose CHGR systems have explored multimodality across broader visual sensor inputs. For example, [112] proposes a framework that segments RGB, depth, infrared, and optical flow inputs using a shared U-Net and processes them via separate 3D DenseNet-43 backbones. Extracted features are fused through Canonical Correlation Analysis to align shared representations across modalities.
6.4 State-of-the-art Methods for CSLR
This subsection reviews state-of-the-art (SOTA) methods for Continuous Sign Language Recognition (CSLR), focusing on their performance on the three most widely used benchmarks: PHOENIX14 [44], PHOENIX14T [26], and CSL-Daily [209].
As shown in Table 7, the dominant architecture remains the RGB-based Spatial Perception Model+1DCNN+BiLSTM pipeline, where the spatial perception model can be either a 2DCNN or a transformer-based approach. Typically, a ResNet18 backbone extracts spatial features, a 1D CNN or TCN models short-term temporal relations, and a BiLSTM captures long-range context for alignment. VAC-CSLR [122] established a widely adopted training setting to address optimization instability. Many works have since extended this baseline with auxiliary classifiers and additional loss terms to enhance feature learning [83, 4, 14]. Furthermore, the TwoStream-SLR [27] seems to compete worthily the aforementioned pipeline, surpassing it when extra training data are utilized [188] and ranked second in the CSL-Daily [209] benchmark without additional samples.
Recent innovations that improve CSLR include:
  • Data Augmentation Across Datasets: Leveraging visually similar signs from different corpora as training data [188].
  • Modeling Gloss Dependencies: Addressing the independence assumption in vanilla Connectionist Temporal Classification (CTC) loss through conditional modeling [197].
  • Multimodal Inputs: Combining RGB and pose heatmaps to improve robustness [27].
  • Graph-Based Feature Interaction: Using GNNs instead of CNNs to model intra- and inter-frame patch relationships [47].
  • Linguistic Priors: Integrating gloss embeddings or language models to guide learning [58, 57].
  • Pre-trained Vision Backbones: Fine-tuning foundation models like CLIP for richer feature extraction [14].
Table 7: Brief description of current state-of-the-art CSLR methods and their performance on the three most used benchmark (Phoenix 2014 [44], Phoenix 2014 T [26] and CSL-Daily [209]). Methods marked with star (*) have utilized additional training data.
Method Modality Architecture Description Training Setting Dataset WER
SLA-Adapter/LoRA [14] RGB CLIP visual encoder with adopter/LoRA, followed by TCN and BiLSTM VAC-CSLR [122] and early supervision, by adding an extra classifier after the visual encoder. CSL-Daily 25.8%
Phoenix 18.8%/19.3%
Phoenix-T 19.5%/19.4%
Uni-Sign [102]* RGB, Pose A pretrained LLM process features extracted by pose, visual and temporal encoders. Three-stage training, multi-task learning with additional training data CSL-Daily 26%
Swin-MSTP [15] RGB SwinT for spatial feature extraction and a mixture of 1D convolution, convolution with different dilation rates and TLP [73]. VAC-CSLR [122] CSL-Daily 27.1%
Phoenix 18.7%
Phoenix-T 19.7%
SlowFastSign [4] RGB Modified SlowFast with bidirectional feature fusion between the two streams. VAC-CSLR [122] and auxiliary classifiers to provide extra supervision. CSL-Daily 24.9%
Phoenix 18.3%
Phoenix-T 18.7%
MultiSignGraph [47] RGB Multiscale spatial-temporal GNN that considers image patches as the nodes. [47]. VAC-CSLR [122] CSL-Daily 26.4%
Phoenix 19.1%
Phoenix-T 19.1%
CSGC [143] RGB 2DCNN+1DCNN+BiLSTM VAC-CSLR [122], auxiliary losses to maintain inter- and cross-sentence consistency and gloss-prototype. CSL-Daily 26.7%
Phoenix 19%
Phoenix-T 19.5%
CorrNet+ [76] RGB Modification of the CorrNet [74], obtained by adding temporal attention after the correlation and identification modules. VAC-CSLR [122] and TLP [73]. CSL-Daily 28.2%
Phoenix 18.2%
Phoenix-T 19.1%
TCNet [111] RGB Combination of the TLP [73] and the CorrNet [74] VAC-CSLR [122] and an additional regularization term CSL-Daily 29.3%
Phoenix 18.9%
Phoenix-T 19.4%
C2ST [197] RGB SwinT+1DCNN+BiLSTM CGD loss, distillation to maintain the consistency between video- and local-level features and linguistic prior CSL-Daily 25.8%
Phoenix 17.7%
Phoenix-T 18.9%
CoSign-2s [83] Pose Dual stream approach based on ST-GCN, considering both joints and motions. Each stream consists of group-specific branches. Extra supervision via auxiliary classifier and complementary regularization to enhance the robustness to noise. CSL-Daily 27.2%
Phoenix 20.1%
Phoenix-T 20.1%
CorrNet [74] RGB 2DCNN equipped with the proposed identification and correlation modules, followed by 1DCNN+BiLSTM. VAC-CSLR [122] CSL-Daily 30.1%
Phoenix 19.4%
Phoenix-T 20.5%
[188]* RGB, Pose TwoStream-SLR [27] Utilization of extra training data via dictionary construction. Phoenix 16.7%
Phoenix-T 18.5%
TwoStream-SLR [27] RGB, Pose Dual stream (containing S3D blocks) with bidirectional lateral connections. Extra supervision achieved by auxiliary CTC losses and self-distillation applied to additional prediction heads. CSL-Daily 25.3%
Phoenix 18.8%
Phoenix-T 19.3%
7 Datasets
The performance and generalizability of hand gesture and sign language recognition (HGR/SLR) models are heavily influenced by the quality, diversity, and scale of the datasets used during training and evaluation. Over the years, a wide variety of datasets have been developed to support different HGR tasks, ranging from static and isolated gestures to continuous sign language recognition in real-world settings. These datasets vary widely in terms of the number of gestures or glosses, the types of modalities captured (e.g., RGB, depth, pose, IR), the presence of signer diversity, and the complexity of the sequences. In the following subsections, we categorize and review the most prominent datasets according to the specific recognition task they target: (i) Isolated and Continuous Sign Language Recognition, (ii) Semaphoric Gestures, and (iii) Other Gesture Recognition Tasks. This organization aims to provide a clear overview of the resources available for each task type and facilitate their appropriate use in future research.
7.1 Isolated and Continuous SLR
The datasets that have been developed to support video-based Sign Language Recognition (SLR) can be broadly categorized into those designed for Isolated Dynamic SLR (IDSLR) and Continuous SLR (CSLR). The former datasets typically focus on individual signs, recorded in controlled lab environments with a fixed vocabulary, making them well-suited for classification-based models. In contrast, continuous datasets present real-world sequences of signs, often extracted from broadcast material or spontaneous signing, requiring temporal segmentation and sequence modeling approaches. Additionally, some datasets support Sign Language Translation (SLT) by providing sentence-level gloss annotations and parallel spoken language transcripts.
The RWTH-PHOENIX-Weather 2014 [44] and its extended version PHOENIX-2014-T1 [26] are benchmark datasets offering sentence-level annotations in German Sign Language (DGS) captured from television weather forecasts. More recent efforts such as BOBSL [12] and OpenASL [160] have significantly expanded the signer base and vocabulary, with hours of British and American Sign Language footage collected from TV and web sources, respectively. Lab-recorded corpora like KETI [90], CSL-Daily [209], and How2Sign [41] incorporate multimodal inputs (e.g., pose, depth) and structured glosses, aiding both CSLR and SLT tasks. These datasets collectively enable the development and benchmarking of models under varying constraints of signer variability, language, modality, and task complexity.
On the other hand, isolated sign datasets such as MS-ASL [85], WLASL [99], and BSL-1K [11] provide extensive vocabularies (up to 2,000 signs) and diverse signers, facilitating the study of vocabulary scalability and signer generalization. Many of these datasets, such as AUTSL [164] and BosphorusSign22k [217], also include pose or depth modalities, enriching multimodal learning. While isolated datasets are primarily used for gesture classification, continuous datasets address the more complex problem of temporal segmentation, gloss prediction, and SLT. A few hybrid datasets like GSL SD/SI [2] and Corpus FinSL [154] provide data that can be used in both isolated and continuous settings. Together, these datasets serve a broad spectrum of research needs, from foundational gesture recognition to high-level translation across multiple sign languages.
Table 8: Overview of the characteristics of each dataset related to Sign Language Understanding (SLU) and Hand Gesture Recognition (HGR). The sign vocabulary (Sign Voc.) corresponds to the number of distinct classes included in each dataset.
Name Task Year Modality Samples Signers Sign Voc. Language Duration Source
PHOENIX-2014 [44] CSLR/SLT 2014 RGB - 9 1558 DGS 10.73h TV
PHOENIX-2014-T1 [26] CSLR/SLT 2018 RGB 8257 9 1066 DGS - TV
KETI [90] CSLR/SLT 2019 RGB, Pose 14672 14 524 KSL 27.99h Lab
GSL SD/SI [2] CSLR/SLT 2019 RGBD 10295 7 310 GSL 9.59h Lab
Corpus FinSL [154] CSLR/SLT 2020 RGB 343 21 - FinSL 14.37h -
CSL-Daily [209] CSLR/SLT 2021 RGB 20654 10 2000 CSL 23.27h Lab
How2Sign (green studio) [41] CSLR/SLT 2021 RGBD, Pose 2529 11 - ASL 79.12h Lab
BOBSL [12] CSLR/SLT 2021 RGB 1962 39 2281 BSL 1467h TV
OpenASL [160] CSLR/SLT 2022 RGB 98417 220 - ASL 288h Web
Auslan-Daily [158] CSLR/SLT 2023 RGB 25K 67 - AUSLAN - TV, Web
TVB-HKSL-News [130] CSLR/SLT 2024 RGB, Pose 7K 2 6515 HKSL 16.07h TV
QSL [48] CSLR/SLT 2024 RGB 30040 5 654 CSL - Lab
ChicagoFSWild+ [159] CSLR 2019 RGB 55232 260 26 ASL - Web
HKSL [212] CSLR 2021 RGBD 2400 6 55 HKSL - Lab
CSL [78] IDSLR 2018 RGBD, Pose 125K 50 500 CSL - Lab
SMILE [42] IDSLR 2018 RGB - 30 100 DSGS - Lab
ASL Dataset [16] IDSLR 2018 Pose 1200 20 30 ASL - Lab
MSASL [85] IDSLR 2019 RGB 25513 222 1000 ASL 24.65h Web
WLASL [99] IDSLR 2019 RGB 21083 119 2000 ASL 14.11h Web
GSL isol. [2] IDSLR 2019 RGBD 40785 7 310 GSL 6.44h Lab
BosphorusSign22k [217] IDSLR 2020 RGBD, Pose 22542 6 744 TSL 19h Lab
INCLUDE [169] IDSLR 2020 RGB 4287 7 263 ISL 3h Lab
AUTSL [164] IDSLR 2020 RGBD, Pose 38336 43 226 TSL 21h Lab
RKS-PERSIANSIGN [144] IDSLR 2020 RGB 10000 10 100 Persian - Lab
BSL-1K [11] IDSLR 2020 RGB 280K 40 1064 BSL 1060h TV
NMFs-CSL [69] IDSLR 2021 RGB 32010 10 1.067 CSL 24.4h Lab
KArSL [163] IDSLR 2021 RGBD, Pose 75300 3 502 ArSL - Lab
CSLD [184] IDSLR 2022 RGB 120K 30 400 CSL 66.8h Lab
NCSL [182] IDSLR 2022 RGB 90K 30 300 CSL 50.1h Lab
IISL2020 [92] IDSLR 2022 RGB 12100 16 11 ISL 7h Lab
LSA64 [149] IDSLR 2023 RGB 3200 10 64 LSA 1.9h Lab
Signsuisse Lexicon [126] IDSLR 2023 RGB 18221 - - 3 (Mult) - Web
FDMSE-ISL [135] IDSLR 2024 RGBD 40033 20 2002 ISL 36h Lab
7.2 Semaphoric HGR
Semaphoric hand gesture recognition (HGR) research is supported by a diverse range of datasets categorized into three main types based on task formulation: static, isolated dynamic, and continuous dynamic gesture recognition. These datasets vary in their application domains (e.g., Human-Computer Interaction (HCI), Human-Machine Interface (HMI), Human-Robot Interaction (HRI)), modalities (e.g., RGB, depth, pose, infrared), and in the number of gesture classes, subjects, and gesture samples they include. The statistics of datasets for semaphoric HGR are provided in tables 9,10 and 11, regarding the task for which they are used (i.e. static, isolated dynamic and continuous HGR, respectively).
Static HGR datasets focus on recognizing gestures from single-frame input and are generally easier to process due to the absence of temporal complexity. Notable examples include LaRED [67], one of the earliest RGBD datasets with 243,000 samples and 81 classes, and HaGRIDv2 [131], a large-scale RGB dataset designed for HMI with over 1 million samples, covering 34 classes. Smaller yet targeted datasets such as OUHANDS [118] and URGR [17] focus on specific domains like HCI and HRI, while MD-UHGRD [194] offers moderately sized samples and class diversity. Isolated dynamic HGR datasets contain gestures performed in isolation, with start and end points explicitly segmented. Datasets like ChaLearn IsoGD [179] and EgoGesture (isolated) [199] feature rich modality (RGBD, IR, pose) and high gesture diversity, with up to 249 classes. Others like Jester [117] offer large-scale RGB-only data from over 1,300 subjects, making them suitable for data-driven deep learning approaches. More recent contributions such as DATE [35] and ZJUGesture [193] reflect growing interest in real-world HCI scenarios. Continuous dynamic HGR datasets challenge models to identify and segment gestures from continuous video streams. The ChaLearn ConGD [179] dataset remains a benchmark with 22,535 sequences and 249 gesture classes. Similarly, EgoGesture [199] and IPN Hand [19] include both isolated and continuous segments, enabling comprehensive model development and evaluation. Datasets like LD-ConGR [106] and ZJUGesture [193] contribute additional complexity by increasing subject diversity and contextual variability.
Each dataset supports specific sub-tasks within HGR research: static datasets aid in gesture classification, isolated dynamic datasets support temporal gesture modeling and classification, and continuous datasets allow for joint gesture spotting and classification—crucial for real-time and interactive applications. This diversity in dataset structure, scale, and modality reflects the evolving demands and challenges of semaphoric gesture recognition across interactive domains.
Table 9: Datasets for static hand gesture recognition.
Name Year Domain Modality Samples Subjects Classes Resolution
HGR-1 [88] 2014 - RGB 899 12 27 174x131-640x480
LaRED [67] 2014 - RGBD 243000 10 81 640x480
OUHANDS [118] 2016 HCI RGBD 3150 23 10 640x480
LRHG [211] 2021 HRI RGB 4320 8 10 640x480
SHAPE [33] 2022 - RGB 34226 20 32 4128x3096
HaGRID [86] 2024 HCI RGB 552992 34730 19 1920x1080
HaGRIDv2 [131] 2024 HCI RGB 1086158 65977 34 1920x1080
MD-UHGRD [194] 2024 HRI RGB 20000 - 19 -
URGR [17] 2024 HRI RGB 347483 16 6 480x640
Table 10: Datasets for isolated dynamic hand gesture recognition.
Name Year Domain Modality Samples Subjects Classes Frames
SKIG [108] 2013 - RGBD 1080 6 10 -
nvGesture [124] 2016 HMI RGBD, IR, OF 1532 20 25 -
ChaLearn IsoGD [179] 2016 - RGBD 47933 21 249 -
DHG-14/28 [37] 2016 - Depth, Pose 2800 20 14/28 -
SHREC17-14/28 [166] 2017 - Depth, Pose 2800 28 14/28 -
EgoGesture (isol.) [199] 2017 HCI RGBD 24161 50 83 2953224
Briareo [115] 2019 HMI RGBD, Pose, IR 1440 40 12 -
Jester [117] 2019 HCI RGB 148092 1376 27 5331312
IPN Hand (isol.) [19] 2020 HCI RGB 4218 50 13 -
TCG [189] 2020 HMI Pose 250 5 4 839350
LD-ConGR (isol.) [106] 2022 HCI RGBD 44887 30 10 -
ZJUGesture (isol.) [193] 2023 HCI RGB 9892 60 9 -
DATE [35] 2024 HCI RGB 13500 22 27 -
EgoDriving [138] 2024 HMI RGB 400 - 29 70276
Table 11: Datasets for continuous dynamic hand gesture recognition.
Name Year Domain Modality Samples Instances Subjects Classes
ChaLearn ConGD [179] 2016 - RGBD 22535 47933 21 249
EgoGesture (cont.) [199] 2017 HCI RGBD 2081 24161 50 83
IPN Hand (cont.) [19] 2020 HCI RGB 200 4218 50 13
Police Gesture Dataset [66] 2020 HMI RGB 21 3354 - 8
LD-ConGR (cont.) [106] 2022 HCI RGBD 542 44887 30 10
ZJUGesture (cont.) [193] 2023 HCI RGB - 9892 60 9
7.3 Other
In addition to standard semaphoric gesture datasets, several specialized datasets have been introduced to support research on more specific gesture types. The First-Person Hand Action (FPHA) dataset [50], published in 2018, contains 1,175 video sequences captured from a first-person perspective by six participants. It focuses on isolated manipulative hand gestures across 45 action classes, offering RGB, depth, and skeleton data, and is particularly valuable for studies involving egocentric interaction. Another notable example is the DP Dataset [127], released in 2023, which, to the best of our knowledge, is the only publicly available benchmark specifically designed for pointing gesture recognition. Collected from 33 participants in two indoor environments, the dataset comprises approximately 2.8 million frames and 6,335 annotated gesture instances. It includes detailed annotations such as gesture timings and pointing directions, enabling both binary gesture classification (gesture vs. non-gesture) and directional estimation tasks.
8 Evaluation Metrics
In this section, we present the key evaluation metrics adopted in the literature for measuring the performance of hand gesture recognition (HGR) methods. Broadly, metrics fall into two main categories: classification metrics (used in static and isolated dynamic HGR) and sequence-to-sequence metrics (used in continuous dynamic HGR). Additionally, other indicators such as computational resource usage and overfitting-related metrics are also considered.
8.1 Classification Metrics
In multi-class HGR problems—especially when dealing with class imbalance—researchers often report both micro-averaged (per-instance) and macro-averaged (per-class) versions of common classification metrics. Micro-averaged metrics aggregate contributions from all instances across classes, thus being sensitive to class imbalance. Macro-averaged metrics compute the metric independently for each class and then average, treating all classes equally regardless of their frequency.
Micro-averaged metrics:
$\text{Accuracy}_{micro}=\frac{\sum_{i=1}^{N}(TP_{i}+TN_{i})}{\sum_{i=1}^{N}(TP_{i}+TN_{i}+FP_{i}+FN_{i})}$ (1)
$\text{Precision}_{micro}=\frac{\sum_{i=1}^{N}TP_{i}}{\sum_{i=1}^{N}(TP_{i}+FP_{i})}$ (2)
$\text{Recall}_{micro}=\frac{\sum_{i=1}^{N}TP_{i}}{\sum_{i=1}^{N}(TP_{i}+FN_{i})}$ (3)
[…]
Macro-averaged metrics:
$\text{Accuracy}_{macro}=\frac{1}{N}\sum_{i=1}^{N}\frac{TP_{i}+TN_{i}}{TP_{i}+FP_{i}+TN_{i}+FN_{i}}$ (5)
$\text{Precision}_{macro}=\frac{1}{N}\sum_{i=1}^{N}\frac{TP_{i}}{TP_{i}+FP_{i}}$ (6)
$\text{Recall}_{macro}=\frac{1}{N}\sum_{i=1}^{N}\frac{TP_{i}}{TP_{i}+FN_{i}}$ (7)
[…]
F1-Score: For both micro and macro cases, the F1-score is defined as the harmonic mean of Precision and Recall:
$F_{1}=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$ (9)
Here, $TP_{i}$, $FP_{i}$, $TN_{i}$, and $FN_{i}$ denote the number of true positives, false positives, true negatives, and false negatives for the $i$-th class, respectively, and $N$ is the total number of classes.
8.2 Sequence-to-Sequence Metrics
For continuous HGR (CHGR), particularly in Continuous Sign Language Recognition (CSLR), sequence-level metrics are used. Word Error Rate (WER) [122, 2] is the standard metric, measuring the minimum number of substitutions ($S$), deletions ($D$), and insertions ($I$) needed to transform the predicted sequence into the ground truth sequence of length $N$:
$WER=\frac{S+D+I}{N}$ (10)
Edit Distance Accuracy [45, 185] is a complementary metric representing the ratio of correctly predicted elements in the sequence:
[…]
These metrics provide insight not only into classification accuracy but also into temporal and sequential consistency.
8.3 Computational Resource Metrics
In addition to recognition performance, computational efficiency is often evaluated, especially for real-time or embedded HGR applications. Common metrics include:
  • Inference Time: Time (in ms or seconds) taken to process one input sample.
  • Frames Per Second (FPS): Number of frames processed per second; higher values indicate faster processing.
  • Model Size: The number of trainable parameters or total memory footprint of the model.
These metrics are crucial when deploying HGR systems on edge devices or in latency-sensitive environments.
8.4 Overfitting Indicators
To assess generalization and detect overfitting, the following indicators are used:
  • Train-Test Performance Gap: A significant difference between training and testing accuracy (or loss) often signals overfitting.
  • Validation Loss Trend: Increasing validation loss while training loss decreases is a common overfitting pattern.
  • Early Stopping Criteria: Monitored via a plateau or increase in validation loss across epochs.
9 Current Challenges and Future Research Directions
As mentioned in the Methods sections (Sections 4, 5, and 6), researchers have developed sophisticated techniques and large-scale datasets to achieve satisfactory performance. However, as shown in the relevant tables (Tables 3, 4, 5, 6, 7), recognition accuracy remains low in many cases—particularly in more complex tasks such as Continuous Sign Language Recognition (CSLR). In this section, we discuss the current challenges and limitations that hinder efficient Human Gesture Recognition (HGR), posing ongoing concerns for the research community and shaping future research directions. Additionally, we review the techniques proposed to address these issues, offering insights that may guide future efforts to overcome the identified challenges.
9.1 Gesture-Irrelevant Factors
The presence of gesture-irrelevant factors (or redundant information) hinders accurate recognition in diverse real-world scenarios. Some of these factors include the following [171, 9, 80, 17, 141, 206]:
  • Environmental factors: Cluttered background, varying illumination conditions, and self-occlusions make it difficult or even impossible to automatically spot the hand region, which is an important step for efficient HGR;
  • Subject characteristics: Subject independence –—i.e., the ability to correctly identify a gesture regardless of the signer’s characteristics such as skin color, hand shape variations, or clothing—is a key prerequisite for achieving sufficient generalization and enabling real-world deployment.
  • Variations in perception device position: The distance between the performer and the camera angle may reduce the recognition performance;
  • Input data quality: Low resolution and blurriness may corrupt the discriminative characteristics of the hand gesture region.
Aiming to address these challenges, researchers have proposed segmentation models to crop hand regions and remove background interference [208, 34, 153], object detectors for spatial normalization [17, 9], super-resolution techniques to enhance image quality [17], attention mechanisms [206, 207] to enforce the visual encoder to pay attention on the most informative regions, information-theoretic losses [100], and feature selection techniques [173] to reduce information redundancy. Many of these approaches rely on computationally expensive off-the-shelf models, which increase the complexity of the overall HGR methods and set an upper bound to their accuracy. Thus, the presence of gesture-irrelevant factors could concern future research work, while potentially being combined with eXplainable Artificial Intelligence (XAI) approaches [148] to facilitate this task.
9.2 Data Scarcity
Many existing datasets either contain a relatively small number of samples or lack sufficient diversity. For example, among static HGR datasets, only the versions of HaGRID [86, 131] have been collected from more than one hundred participants, despite some others having a plausible number of samples [67, 17]. A similar situation is observed in IDHGR datasets, where only the Jester dataset [117] includes data from over one hundred subjects. Furthermore, the amount of samples for specific modalities, such as IR data, is limited compared with the size of RGB-based datasets, an issue known as data imbalance [181]. The lack of large-scale, diverse datasets leads to overfitting and limited generalization, which is closely related to the challenges discussed earlier. This issue is further exacerbated in tasks that involve costly annotation processes and require data-intensive deep learning models with a large number of parameters, such as CSLR.
Several techniques have been explored to mitigate this challenge:
  • Multilingual Data Aggregation: Combining datasets (e.g., CSL and DGS) to exploit cross-lingual visual similarities increases the overall amount of training data [188].
  • Self-Supervised Learning: In the context of the self-supervised learning setting [91], gloss-to-gloss translation tasks have been used for parameter initialization [205], providing linguistic priors to sign language models. Pretraining with pose reconstruction has also been investigated to incorporate prior knowledge of hand structure [71, 68, 79].
  • Advanced Data Augmentation: Adversarial learning has been employed to optimize augmentation hyperparameters [128].
  • Few-/Zero-shot Learning: Approaches using only a small number of samples to correctly recognize gestures—often based on CLIP—have also been proposed [216, 24, 147, 146, 22].
These approaches have led to improved recognition performance, even when trained on limited data. However, accuracy remains suboptimal, making data scarcity a persistent research challenge and a key direction for future work in the field.
9.3 Optimization Difficulties
Training the visual encoder effectively remains difficult, especially in deep architectures for CSLR. As shown in CTCA [57] (Fig. 17), increasing the depth of the temporal module (e.g., 1D-TCN followed by BiLSTM) improves its generalization but hampers the encoder’s learning due to gradient attenuation in backpropagation. Furthermore, the CTC spike phenomenon—where the model over-focuses on a few keyframes—can cause rapid overfitting and limit encoder training [122, 62].
Figure 17: Experimental results of a study conducted by authors in [57], which aims to measure the impact of the depth of the temporal aggregation module (a 1D-TCN with various number of layers followed by a two-layer BiLSTM) on the effectiveness of training the visual encoder (aka spatial perception module). In both diagrams, the x axis represents the number of layers of the 1D-TCN. In the top chart, the two modules are compared considering the information stored in weights (IIW) [187] (y axis), which can be viewed as an indicator of generalization gap, while in the bottom the corresponding WER is shown. As the number of layers of the 1D-TCN increases, although the generalization ability of the short-term temporal module increases, the overall performance decreases due to the insufficient training of the visual encoder.
To address these issues, several strategies have been proposed:
  • Auxiliary Supervision: Adding early-stage classifiers to the encoder or short-term module has been proposed as a strategy to guide training by providing additional supervision [122, 143, 74, 76, 111, 213].
  • Temporal Module Redesign: Placing the 1D-TCN and BiLSTM in parallel instead of sequentially [57].
  • Linguistic Constraints and Gloss Segmentation: Using gloss boundaries [62] or prior knowledge [58] to regularize learning.
The aforementioned approaches —particularly auxiliary supervision [122]— have served as the foundation for many existing methods. However, the increasing complexity of modern deep architectures often exceeds the effectiveness of these techniques. As a result, optimization remains a significant challenge —one that may be further exacerbated by the continued development of even deeper models.
9.4 Limitations of Architectures
Despite the achievements of DL, the existing architectures suffer from the following limitations:
  • Block-fixed models: The existing architectures are based on hand-crafted design, which may not be optimal[206].
  • GCNs with fixed topology graphs: Standard GCNs use fixed topology graphs based on hand anatomy, which may not capture optimal inter-joint relationships.
  • Ineffective multimodal fusion: Most of the current works utilize late fusion, that does not fully leverage the available information [68, 71, 202, 204, 203, 104, 82, 81].
To address these limitations, researchers have proposed various strategies, including Neural Architecture Search (NAS) to identify optimal model structures [206], learnable or adaptive skeleton adjacency matrices [77, 30, 28, 107], and multi-level fusion for dual-stream approaches [60]. Despite these advances, several aspects of deep learning architectures remain underexplored, leaving a substantial research gap.
10 Conclusions
The undeniable importance of vision-based hand gesture recognition (HGR) stems from its wide range of applications—spanning human-robot interaction to sign language understanding. This broad applicability has made it a vibrant and active research field. Recent advances in deep learning have played a pivotal role in driving progress, enabling the development of increasingly sophisticated architectures and training strategies aimed at tackling more complex and fine-grained tasks, such as continuous sign language recognition.
This survey focused on HGR methods that leverage visual input and deep learning. These methods were highlighted not only for their superior performance compared to traditional hand-crafted approaches, but also for their higher potential for deployment in real-world scenarios, unlike more cumbersome sensor-based techniques. From the surveyed literature, it becomes clear that even relatively simple and shallow deep learning models can achieve high accuracy in static HGR tasks. In such cases, the primary challenges are typical of general computer vision problems and are not inherently tied to HGR itself.
Despite the significant progress achieved in vision-based hand gesture recognition, several open challenges remain that shape future research directions. A major concern is the generalization of models across diverse environments, lighting conditions, camera viewpoints, and user-specific variations, which often lead to performance drops in real-world settings. The lack of large-scale, diverse, and well-annotated datasets, especially for continuous gestures and sign language, also hinders the development of more robust models. Furthermore, model efficiency and scalability are critical for deployment on resource-constrained edge devices, necessitating advances in lightweight architectures and model compression techniques. The integration of multimodal data, such as combining vision with audio or inertial signals, remains underexplored yet promising. Finally, interpretability and trustworthiness of deep learning models are essential, particularly for safety-critical applications like healthcare or assistive technologies. Addressing these challenges will not only enhance the practical applicability of HGR systems but also contribute to more inclusive and accessible human-computer interaction paradigms.
[…]
Acknowledgments The research leading to these results received funding from the European Commission under Grant Agreement No. 101168042 (TRIFFID) and No. 101189557 (TORNADO).
Declarations
Funding The research leading to these results received funding from the European Commission under Grant Agreement No. 101168042 (TRIFFID) and No. 101189557 (TORNADO).
Conflict of interest The authors have no competing interests to declare that are relevant to the content of this article.
Author contribution Konstantinos Foteinos, Jorgen Cani, Manousos Linardakis, Iraklis Varlamis and Georgios Th. Papadopoulos performed the literature review and prepared a draft of the manuscript. Panagiotis Radoglou-Grammatikis, Vasileios Argyriou and Panagiotis Sarigiannidis performed reviewing and editing of the manuscript. Georgios Th. Papadopoulos was responsible for receiving the funding to implement the study.

### v2

@@
-
+Visual Hand Gesture Recognition with Deep Learning: A Comprehensive Review of Methods, Datasets, Challenges and Future Research Directions
 Abstract
-The rapid evolution of deep learning (DL) models and the ever-increasing size of available datasets have raised the interest of the research community in the always important field of vision-based hand gesture recognition (VHGR), and delivered a wide range of applications, such as sign language understanding and human-computer interaction using cameras.
+The rapid evolution of deep learning (DL) models and the ever-increasing size of available datasets have raised the interest of the research community in the always important field of visual hand gesture recognition (VHGR), and delivered a wide range of applications, such as sign language understanding and human-computer interaction using cameras.
 Despite the large volume of research works in the field, a structured and complete survey on VHGR is still missing, leaving researchers to navigate through hundreds of papers in order to find the right combination of data, model, and approach for each task.
-The current survey aims to fill this gap by presenting a comprehensive overview of this aspect of computer vision.
-With a systematic research methodology that identifies the state-of-the-art works and a structured presentation of the various methods, datasets, and evaluation metrics, this review aims to constitute a useful guideline for researchers, helping them to choose the right strategy for delving into a certain VHGR task.
-Starting with the methodology used for study selection, literature retrieval, and the analytical framing, the survey identifies and organizes key VHGR approaches using a taxonomy-based format in various dimensions such as input modality and application domain.
-The core of the survey provides an in-depth analysis of state-of-the-art techniques across three primary VHGR tasks: static gesture recognition, isolated dynamic gestures and continuous gesture recognition.
+The current survey aims to fill this gap by presenting a comprehensive overview of this computer vision field.
+With a systematic research methodology that identifies the state-of-the-art works and a structured presentation of the various methods, datasets, and evaluation metrics, this review aims to constitute a useful guideline for researchers, helping them to choose the right strategy for handling a VHGR task.
+Starting with the methodology used to locate the related literature, the survey identifies and organizes the key VHGR approaches in a taxonomy-based format, and presents the various dimensions that affect the final method choice, such as input modality, task type, and application domain.
+The state-of-the-art techniques are grouped across three primary VHGR tasks: static gesture recognition, isolated dynamic gestures, and continuous gesture recognition.
 For each task, the architectural trends and learning strategies are listed.
-Additionally, the study reviews commonly used datasets —emphasizing on annotation schemes— and evaluates standard performance metrics.
-It concludes by identifying major challenges in VHGR, including both general computer vision issues and domain-specific obstacles, and outlines promising directions for future research.
+To support the experimental evaluation of future methods in the field, the study reviews commonly used datasets and presents the standard performance metrics.
+Our survey concludes by identifying the major challenges in VHGR, including both general computer vision issues and domain-specific obstacles, and outlines promising directions for future research.
 keywords:
-Vision-Based Hand Gesture Recognition; Human-Robot Interaction; Human-Computer Interaction; Sign Language; Deep Learning; Gesture Recognition Taxonomy
+Vision-Based Hand Gesture Recognition , Human-Robot Interaction , Human-Computer Interaction , Sign Language , Deep Learning , Gesture Recognition Taxonomy
+[…]
+[…]
 1 Introduction
-Understanding human behaviors—such as hand gestures [2], actions [156], and facial expressions [150]—is a key area of artificial intelligence research.
-There is growing interest in developing Intelligent User Interfaces (IUIs) that offer more natural and intuitive ways for humans to interact with machines, replacing traditional tools like the keyboard and mouse.
+The recognition and interpretation of hand gestures (ADADOGLOU) play a central role in advancing Human–Computer Interaction (HCI) and assistive communication systems.
+In recent years, deep learning–based visual hand gesture recognition (VHGR) has gained significant momentum as an active subfield of computer vision, building upon progress in related areas such as action (survey-HAR-introduction-shafizadegan2024multimodal) and facial expression recognition (survey-FER-JESWA-2024).
 This has led to increased focus on Hand Gesture Recognition (HGR) across various applications.
-For example, HGR is used in human-robot interaction in industrial settings (e.g., robotic arm control [134, 96]), healthcare (e.g., robot-assisted surgery [139]), and autonomous mobile robot navigation (e.g., UAVs [194] and UGVs [17, 105]).
-It also plays a key role in human-computer interaction, especially in AR/VR environments [32, 199, 5].
-In addition, the large population of deaf and hard-of-hearing individuals has driven research into automatic sign language recognition systems, aiming to improve communication and inclusion for these communities [196, 26].
-Given the importance of the applications mentioned above, Hand Gesture Recognition (HGR) has grown into a broad research field, generally divided into two main approaches: vision-based and sensor-based.
+For example, HGR is widely used in Human-Robot Interaction (HRI) in industrial settings (e.g.
+robotic arm control (Papadopoulos2021Access; ROBOTIC_ARM)), in healthcare (e.g.
+robot-assisted surgery (TELEOPERATED_SURGICAL_ROBOT)), and in the autonomous navigation of mobile robots such as Unmanned Aerial (MD-UHGRD_SA-YOLO) and Ground (ULTRA-RANGE-UGV-Bamani; Linardakis2024Access) Vehicles (UAVs and UGVs).
+It also plays a key role in HCI, especially in AR/VR environments (OO-dMVMT; EgoGesture_DATASET; BOAT-MI_VR).
+In addition, the large population of deaf and hard-of-hearing individuals has driven research into automatic Sign Language Recognition (SLR) systems, aiming to improve communication and inclusion for these communities (survey-deaf-mute-AIR-2024; NEURAL-SIGN-LANGUAGE-TRANSLATION-CVPR-CAMGOZ).
+Given the importance of the applications mentioned above, HGR has grown into a broad research field, generally divided into two main approaches: vision-based and sensor-based.
 Vision-based HGR relies on visual data such as RGB and depth images, as well as features derived from them—like pose heatmaps or optical flow.
-Sensor-based HGR, on the other hand, uses data from devices such as wearable gloves or surface EMG (sEMG) sensors, and even non-wearable sources like acoustic signals.
+Sensor-based HGR, on the other hand, uses data from devices such as wearable gloves or surface electromyography (sEMG) sensors, and even non-wearable sources like acoustic signals.
 Researchers have also explored multi-modal methods that combine inputs from different sources.
 In this study, we focus on vision-based HGR, as sensor-based methods often involve high cost and complex setup, making them less practical for real-world use.
-Vision-based approaches better align with the goal of HGR: to enable more natural, unobtrusive, and accessible human-machine interaction.
-Many of the published surveys consider both sensor-based and vision-based methods together [140, 162, 64, 6].
-As a result, the analysis of the proposed methods in these surveys tends to be superficial, because there are so many methods to cover.
-Other review papers also present the DAQ devices used by researchers [43], further broadening the range of related publications.
-The undeniable importance of sign language recognition has led some researchers to conduct surveys dedicated exclusively to this topic [2, 155, 170, 145, 174].
-In addition, many authors consider both deep learning (DL) and traditional machine learning (ML) methods, or in some cases only traditional ML approaches [174, 140, 43, 132, 64, 6], while others include outdated methods [162, 123, 132, 7, 3, 84].
-Although this broader approach helps lay the foundational understanding of VHGR and facilitates a historical perspective, it often limits the focus on current state-of-the-art methods, which are arguably more critical for driving the field forward.
-Some surveys aim to provide a quantitative overview of the field [132, 64, 3, 84], offering bibliometric analyses on various aspects of VHGR.
-While such surveys present a statistical and objective perspective and help identify major subfields, they fall short in advancing the field due to their limited examination of the existing methodologies.
-In contrast, to the aforementioned surveys, this survey emphasizes very recent publications and provides a detailed analysis of their contributions, aiming to identify current challenges and innovations that could serve as a valuable starting point for future research.
-A comparison with existing reviews is presented in Table 1.
-Table 1: Comparison of this survey with existing ones. ML refers to traditional Machine Learning/Computer Vision techniques.
-Publication Year Range # Papers Input Data Models Scope
-[2] 2021 - - Vision-based DL CSLR
-[123] 2021 2014-2020 98 Vision-based DL & ML General
-[7] 2021 2014-2021 138 Sensor- & Vision-based DL SLR
-[3] 2021 2001-2021 649 Vision-based DL & ML SLR
-[43] 2022 - - Sensor- & Vision-based DL & ML SLR
-[132] 2022 2000-2020 571 Sensor- & Vision-based DL & ML General
-[84] 2022 2010-2021 80 Vision-based DL & ML SLR
-[155] 2023 2015-2023 - Vision-based DL ISLR
-[162] 2024 2014-2024 - Sensor- & Vision-based DL & ML General
-[64] 2024 2018-2023 256 Sensor- & Vision-based ML General
-[6] 2024 2015-2024 58 Sensor- & Vision-based DL & ML SLR
-[174] 2024 - - Sensor- & Vision-based DL & ML SLR
-[140] 2024 - - Sensor- & Vision-based DL & ML General
-This survey 2025 2021-2024 290 Vision-based DL General
-More specifically, the survey focuses on vision-based methods for three key Hand Gesture Recognition (HGR) tasks, which are discussed in detail in the following sections: static HGR [208, 131], isolated dynamic HGR (IDHGR) [82, 68, 214, 60, 206], and continuous dynamic HGR (CHGR) [122, 27, 74].
-Static HGR involves recognizing gestures from single frames or spatial data only.
-Isolated dynamic HGR extends this by analyzing short, pre-segmented sequences to capture both spatial and temporal patterns.
-Continuous dynamic HGR goes a step further by detecting multiple gestures within long, unsegmented video sequences.
+Vision-based approaches better align with the goal of HGR: to enable more natural, unobtrusive, and accessible gesture-based interaction.
+Table 1: Comparison of this survey with indicative others.
+ML refers to traditional Machine Learning/Computer Vision techniques.
+The current survey mainly examines publications from the last four years, focusing on recent advancements regarding architectures, learning strategies, and related challenges, although older datasets that are relevant to this day have also been included.
+Publication Range Papers Input data Models Scope Comments Aspects that this survey aims to complement
+(ADADOGLOU) 2016-2020 - Vision-based DL CSLR Provides an in-depth comparison of existing approaches, loss functions and pretraining schemes for CSLR.
+This survey expands the scope including non-CSLR methods and examines more recent methods.
+(survey-5-9622242) 2014-2020 98 Vision-based DL & ML General Investigates the progress on hand gesture representation and Data Acquisition (DAQ) techniques as well as their performance and future research directions.
+It examines relatively antiquated methods.
+The current work aims to emphasize on recent DL approaches and related challenges and future perspectives.
+(survey-10-al2021deep) 2014-2021 138 Sensor- & Vision-based DL SLR Categorizes Continuous and Isolated SLR methods based on their architectures and discusses pose estimation for ISLR in detail.
+This survey covers more topics including HCI & HRI, and highlights task-specific architectural patterns and challenges for CSLR.
+(survey-11-adeyanju2021machine) 2001-2021 649 Vision-based DL & ML SLR Discusses relationships between research teams and DAQ techniques.
+It includes relatively old methods.
+This survey emphasizes on recently proposed ones in order to facilitate the evolution of the field.
+(survey-6-ojeda2022survey) 2000-2020 571 Sensor- & Vision-based DL & ML General Provides many statistical results and bibliometric charts.
+It does not identify contemporary challenges or future research directions; this work seeks to detect relevant material.
+(survey-3-el2022comprehensive) 1998-2020 - Sensor- & Vision-based DL & ML SLR Discusses in detail existing methods for SLR and provides multiple tables for quantitative comparison.
+This survey analyzes newer approaches for SLR and expands taxonomy and comparison to more application domains, beyond sign language.
+(survey-16-joksimoski2022technological) 2010-2021 80 Vision-based DL & ML SLR A Natural Language Processing (NLP) toolkit was utilized to filter publications.
+Insightful statistics are also provided.
+Only a few methods are discussed and the current trends are not highlighted.
+On the contrary, this survey aims to identify trends and challenges.
+(survey-17-sarhan2023unraveling) 2015-2023 - Vision-based DL ISLR Provides statistical charts and summary tables based on a four-fold taxonomy.
+This review introduces additional taxonomy criteria and covers more tasks.
+(survey-4-shin2024methodological) 2014-2024 - Sensor- & Vision-based DL & ML General Considers a variety of modalities and examines methods for each one.
+The current work aims to discuss further multi-modality fusion strategies and identify more challenges.
+(survey-8-hashi2024systematic) 2018-2023 256 Sensor- & Vision-based ML General Covers various aspects of HGR, including DAQ techniques, hand segmentation and feature extraction.
+Focuses exclusively on traditional ML approaches, which are currently surpassed by DL ones.
+This survey, searching for recent advancements, focuses on DL approaches.
+(survey-9-al2024advancements) 2015-2024 58 Sensor- & Vision-based DL & ML SLR Examines various SLR methods and highlights the superiority of DL methods over older ones.
+Its does not identify challenges and future perspectives; this survey aims to cover this, going deeper into details.
+(survey-13-tao2024sign) 2017-2024 - Sensor- & Vision-based DL & ML SLR Groups methods into RNN-based, CNN-based and transformers whilst discusses well-known benchmarks.
+This survey expands the scope to non-SLU papers and seeks to provide methods comparison and future prospects.
+(survey-1-rahman2024comparative) 2008-2024 - Sensor- & Vision-based DL & ML General Covers a wide spectrum of gesture-based interaction and .
+The considered methods are relatively old.
+This survey aims to identify more recent advancements.
+(linardakis2025_hgr_survey) 2018-2025 125 Vision-based DL & ML General Discuss jointly hand gesture capturing and recognition techniques.
+Many challenges are highlighted, including data bias and privacy issues.
+It covers a limited number of datasets for HGR and relatively old methods; this survey complements it by analyzing more datasets and recent advancements.
+(survey-emporio2025continuous) 2002-2023 - Sensor- & Vision-based DL & ML CDHGR Focuses on online gesture recognition, discussing DAQ technologies and evaluation metrics, among others.
+This survey covers more topics and provides comparison of state-of-the-art methods.
+(survey-khan2025deep) 2010-2024 - Sensor- & Vision-based DL & ML CSLR Discusses in detail methods and datasets for CSLR.
+It does not conduct methods comparison and the challenges are general; this survey highlights optimization difficulties in the context of CSLR.
+The great importance and wide scope of HGR has led researchers to conduct surveys focusing on different aspects of the field, as shown in Table 1.
+Compared to existing review papers, this survey provides an in-depth examination of the recent Deep Learning (DL) methods (published since 2021), covering the current state-of-the-art (SoTA).
+In contrast, other existing reviews either focus on older approaches (published since 1998) (survey-6-ojeda2022survey; survey-3-el2022comprehensive; survey-8-hashi2024systematic) or mainly cover traditional machine learning techniques (survey-1-rahman2024comparative; survey-13-tao2024sign; linardakis2025_hgr_survey).
+While such works are valuable for understanding the fundamental methods and providing statistical insights, they are less suited to reflecting the recent evolution of the field.
+The present survey stands out by its broader and more up-to-date scope, including 238 papers published since 2021—-an average of nearly 47 publications per year.
+By comparison, (survey-11-adeyanju2021machine) and (survey-6-ojeda2022survey) analyze 649 and 571 papers respectively, but span roughly 20 years of research, corresponding to only about 30 publications per year.
+Finally, the current survey focuses exclusively on vision-based HGR.
+Including sensor-based approaches would significantly expand the scope and potentially confuse readers, while such methods generally offer limited expressiveness and reduced applicability in real-world scenarios.
+Nevertheless, the review maintains a broad and inclusive perspective, acknowledging that approaches from different domains may share common principles and can inform each other.
+More specifically, the survey focuses on vision-based methods for three key VHGR tasks, which are discussed in detail in the following sections: static HGR (SHGR) (FGDSNet; HAGRIDv2_ARXIV), isolated dynamic HGR (IDHGR) (SAM_SLR; SignBERT_ICCV; NLA_SLR; CTFB_Hampiholi; RAAR3DNet), and continuous (dynamic) HGR (CDHGR or CHGR) (VAC-CSLR; NEURIPS2022_6cd3ac24_HEATMAPS; CorrNet).
+SHGR involves recognizing gestures from single frames or spatial data only.
+IDHGR extends this by analyzing short, pre-segmented sequences to capture both spatial and temporal patterns.
+CDHGR goes a step further by detecting multiple gestures within long, unsegmented video sequences.
 Each task presents its own unique challenges.
@@
 At the same time, all tasks share common issues such as interference from irrelevant background movements or variations in lighting, viewpoint, and hand appearance.
-Vision-based Hand Gesture Recognition (HGR) has a long history, beginning with traditional computer vision techniques based on hand-crafted features and evolving into complex deep learning pipelines [124].
+VHGR has a long history, beginning with traditional computer vision techniques based on hand-crafted features and evolving into complex deep learning pipelines (NVIDIA_GESTURE_R3DCNN_Molchanov).
 Early approaches typically followed a three-stage process: pre-processing, feature extraction, and classification.
@@
 Modern deep learning architectures have become increasingly deep and sophisticated, especially for dynamic HGR tasks.
-These models first capture short-term motion patterns and then aggregate them to understand long-term context [122, 73, 62, 75, 74].
+These models first capture short-term motion patterns and then aggregate them to understand long-term context (VAC-CSLR; TemporalLiftPooling; SMKD; SEN-CSLR; CorrNet).
 Due to their superior performance, this study focuses primarily on deep learning-based methods.
-Given the broad scope of vision-based Hand Gesture Recognition (HGR) within computer vision, it is crucial to identify the current state-of-the-art methods for the three tasks mentioned above, as well as the key challenges that have shaped the field’s development.
-This survey aims to 1) provide an overview of the field, enabling readers to quickly find the most relevant publications for their needs, 2) offer a detailed analysis of each method, highlighting their key contributions, and 3) present the state-of-the-art solutions, based on the previous analysis, and explain the main challenges that continue to drive research in HGR.
 The main contributions of this study can be summarized as follows:
   1. 1.
-    A structured and comprehensive methodology for reviewing the literature;
+    A structured and comprehensive methodology for reviewing the VHGR literature;
   2. 2.
@@
     A synthesis of current state-of-the-art approaches along with a discussion of the key challenges facing the field.
-The structure of this paper is organized to guide the reader through a comprehensive review of vision-based hand gesture recognition (VHGR).
+The structure of this paper aims to guide the reader through a comprehensive review of VHGR.
 Section 2 presents the literature review methodology, outlining the systematic approach used to collect and evaluate relevant works.
@@
 Special attention is given to sign language recognition and semaphoric gestures in dedicated subsections.
+Section 7 comprises a unified comparison of methods for VHGR regardless the task.
 Section 8 presents the main datasets used for the evaluation of the presented methods and Section 9 summarizes the evaluation metrics commonly used in the field.
@@
 2 Literature Review Methodology
-The review begins with a clear definition of the research methodology, establishing a structured foundation for the comprehensive exploration of the VHGR landscape.
-This systematic approach is intended to serve as a valuable reference for researchers and practitioners seeking to navigate and contribute to the field.
-The initial step involved defining the scope of the study, which informed the formulation of inclusion and exclusion criteria used to guide the selection of relevant publications.
-Subsequently, targeted queries were conducted across major scientific databases, including Scopus and IEEE Xplore, to retrieve works aligned with the established criteria.
-The final stage entailed a thorough examination of the selected literature, with the objective of distilling the key innovations, contributions, and methodological insights of each publication.
-These elements are further elaborated in the following sections.
+The literature review was conducted following a systematic methodology designed to ensure a comprehensive and structured exploration of the VHGR field.
+The study scope was first defined, guiding the development of inclusion and exclusion criteria for selecting relevant publications.
+Subsequently, targeted queries were executed across major scientific databases to retrieve works meeting these criteria.
+Finally, the selected literature was thoroughly examined to extract key innovations, contributions, and methodological insights, which are detailed in the following sections.
 2.1 Survey Scope
@@
 The review is structured around five foundational pillars:
+  • 1.
     Advanced deep learning methods introduced in recent literature;
+  • 2.
     Benchmark datasets employed across various VHGR tasks;
-  • Established evaluation metrics used to measure system effectiveness;
+  • 3.
+    Established evaluation metrics used to measure the effectiveness of the proposed solutions;
+  • 4.
     Key challenges that continue to shape the research agenda;
-  • Innovative solutions and future directions aimed at overcoming existing limitations.
+  • 5.
+    Innovative solutions and future directions designed to address current limitations.
 2.2 Selection Criteria
@@
 The selection process was guided by three key axes:
+  • 1.
     Relevance: Only works that directly address VHGR were considered.
@@
 Studies centered on Human Action Recognition (HAR), sensor-based techniques (e.g., using electromyography or radar), or multimodal approaches combining visual and non-visual data were excluded to maintain a focused scope.
-  • Up-to-date: To reflect the current state of the field, priority was given to publications from 2021 onward, particularly those introducing new methods.
-Exceptions were made for seminal works—defined here as milestone contributions post-2015 with high citation counts—that have significantly shaped VHGR research.
-In the case of datasets, no strict publication date was imposed, but only those actively used by the community or released within the past four years were included.
-All survey papers considered were published after 2021.
-  • Impact: To further focus on seminal and influential works, publications were sourced from highly cited venues (i.e.
-having an h5-index of 90 or above on Google Scholar), with a single exception for the IEEE International Conference on Automatic Face & Gesture Recognition, which is closely aligned with VHGR.
-In addition to the above, and aiming to provide the theoretical context of the study, a small number of milestone papers has also been referenced, though not analyzed in the main methods section.
+  • 2.
+    Up-to-date: To capture the current state of the field, priority was given to publications from 2021 onward, especially those introducing new methods.
+Regarding datasets, no strict publication date was enforced; however, only those actively used by the community or released within the past four years were included.
+All surveyed review papers were published after 2021.
+  • 3.
+    Impact: To further emphasize seminal and influential works, publications were primarily sourced from highly cited venues (i.e.
+those with an h5-index of 90 or higher on Google Scholar), with a single exception for the IEEE International Conference on Automatic Face & Gesture Recognition, due to its strong relevance to VHGR.
+Additionally, to provide the theoretical context of the study, a small number of milestone papers were referenced, although they were not analyzed in detail in the main methods section.
 2.3 Literature Search
-To identify publications that satisfy the previously defined criteria, a comprehensive literature search was conducted using specialized queries across several well-established scientific databases, including Scopus, IEEE Xplore, ACM Digital Library, and arXiv for pre-prints.
-Google Scholar was also used to double-check for publications from major venues that might have been overlooked.
-The search queries were designed based on the previously outlined inclusion/exclusion criteria, utilizing specific commands (e.g., EXACTSRCTITLE, LIMIT-TO) and boolean operators (e.g., AND, OR, NOT) to refine the results.
-A list of the keywords that were applied in the title, abstract and index terms of each one publication, can be found in Table 2.
-This process retrieved over 900 publications in the Scopus database.
-Out of them, only 210 papers were related to vision-based HGR.
-A second query has been performed on IEEE Xplore, excluding the non-IEEE venues, and resulted in 633 articles, most of them already listed in Scopus.
-A few simpler queries were performed on Google Scholar and arXiv, and duplicates were ignored, resulting in 290 papers in total, including review papers, methods, and datasets.
-Table 2: Keywords used in literature search
-Keywords
-gesture recognition, gesture classification, sign language recognition, sign language understanding, pointing recognition, human finger pointing, hand gestures, human gestures, gesture, hand gesture, body gesture
+To identify publications meeting the predefined criteria, a comprehensive literature search was performed using specialized queries across established scientific databases, including Scopus and Web of Science (WoS).
+Google Scholar was additionally consulted to capture potentially overlooked works from major venues, and arXiv was also used to find relevant high-impact pre-prints.
+Queries were constructed according to the inclusion/exclusion criteria, employing specific commands (e.g., EXACTSRCTITLE, LIMIT-TO) and Boolean operators (AND, OR, NOT) to refine results.
+The search targeted keywords in titles, abstracts, and index terms, including gesture recognition, gesture classification, sign language recognition, sign language understanding, pointing recognition, human finger pointing, hand gestures, human gestures, gesture, hand gesture, and body gesture.
+This process retrieved over 900 publications, of which 210 were relevant to vision-based HGR.
+Additional queries on Google Scholar and arXiv, with duplicates removed, produced a final set of 275 publications, encompassing review papers, methods, and datasets.
 2.4 Categorizing the Identified Articles
@@
 Each category was analyzed separately, and the corresponding information was organized into three dedicated tables.
-Publications relevant to more than one category were included in each applicable table as separate entries.
-Key information for each paper was systematically recorded in the appropriate table columns, as detailed below:
-  • Methods: The name of the proposed method, datasets used for evaluation, a brief description of the architecture, the input modality, the learning strategy, and the evaluation metrics;
-  • Datasets: For each dataset, we recorded the following details: name, type of gesture based on temporal characteristics (e.g., static, isolated dynamic, or continuous), data modality, total number of samples, number of subjects, and—when applicable to continuous HGR—the number of gesture instances.
+Publications relevant to more than one category were included in each table as separate entries.
+This procedure resulted in 185 articles presenting methods, 72 presenting datasets, and 21 survey articles.
+Key information for each paper was systematically recorded in three long tables which contain the appropriate table columns, as detailed below:
+  • 1.
+    Methods: For each article, we recorded the name of the method it proposes, the datasets used for evaluation, the evaluation metrics, a brief description of the architecture, the input modality, and the learning strategy.
+  • 2.
+    Datasets: For each dataset, we recorded its name, the type of gesture based on temporal characteristics (e.g., static, isolated dynamic, or continuous), the data modality, the total number of samples, the number of subjects, and—when applicable to CHGR—the number of gesture instances.
 For Sign Language Understanding (SLU) datasets, we also noted the language, vocabulary sizes (both sign and text), and total duration (for video-based datasets).
 Additional attributes include the source of the dataset (e.g., laboratory, television, or web), frame rate, resolution, and the environmental conditions under which the data was recorded.
-  • Surveys: For survey papers, a concise description of the content was collected.
-(a) The distribution of the selected papers on the top-10 venues, in terms of number of publications.
-(b) The number of selected publications grouped by year.
-Figure 1: Bibliometric statistics for the selected papers.
-Additionally, for each publication, metadata such as title, authors’ names, venue, publication year, citation count, keywords, task, application domain, authors’ motivation, challenges, and future directions were also recorded.
-This meticulous process resulted in a curated collection of 290 papers.
-Some bibliometrics are provided in Fig.
+  • 3.
+    Surveys: For survey papers, we provide a concise description of the content.
+Furthermore, the proposed taxonomy, the time range covered, and the scope of the survey were recorded.
+Although we examine reviews considering both sensor-based and vision-based HGR techniques, we focus only on content related to VHGR.
+For each publication, several metadata such as title, authors’ names, venue, publication year, citation count, keywords, task, application domain, authors’ motivation, challenges, and future directions were recorded.
+The tables are not presented directly in this article, due to their size, but are available upon request.
+The most recent surveys have already been presented in Table 1 and discussed in the Introduction section, the methods and datasets are presented in an organised manner in the following Sections, and some bibliometrics are provided in Fig.
 1.
-Notably, the number of review papers is limited, highlighting the research gap that this survey aims to address.
+Papers Range
+Methods 2021-2025
+Datasets 2010-2025
+Reviews 2021-2025
+Figure 1: Bibliometric statistics.
+Figure a) shows the distribution of the selected papers on the top-10 venues, in terms of number of publications.
+Almost 50 of the papers considered in the current survey have been published in IEEE Access, covering the whole spectrum of domains and tasks.
+Figure b) illustrates the number of publications & pre-prints per year.
+As the current survey focuses on recent advancements, most of the selected papers have been published since 2021.
+Methods and reviews published before 2021 were ignored.
+The number of methods, datasets and reviews (which may be overlapped) are shown in figure c).
+The publication range for each category is reported in table d).
 3 Taxonomy of Hand Gesture Recognition Methods
 Figure 2: Taxonomy of hand gesture recognition methods.
-This section performs an overview of the landscape of vision-based hand gesture recognition methods, including various aspects, such as input, representation, applications, etc..
-More specifically, a set of complementary and diverse criteria, each focusing on key aspects of a VHGR system, is considered, resulting into the classification of related works into a corresponding set of main categories (while each class can be further divided into sub-categories, as will be detailed below).
-The different criteria that are used to organize the research works in the field are graphically illustrated in Fig.
-2 and are detailed as follows:
-  • Application Domain: This criterion refers to the various domains in which VHGR (Visual Hand Gesture Recognition) solutions have been applied. The following categories are considered:
-    • Sign Language Understanding (SLU): Involves the automatic interpretation of gestures in either a single sign language (monolingual SLU [102, 12]) or across multiple sign languages (multilingual SLU [126]).
+A diverse set of criteria has been proposed by many works.
+This survey adds a few more, including ”Architecture”, considering it essential for a method-oriented presentation.
+To better understand the diversity and evolution of vision-based hand gesture recognition (VHGR) methods, this section introduces a structured taxonomy of existing approaches.
+It classifies related works according to complementary criteria that capture key aspects of VHGR systems—such as input type, data representation, and application domain—each further divided into subcategories.
+The overall taxonomy is illustrated in Fig.
+2 and described in detail below.
+3.1 Application Domain
+This criterion refers to the various domains in which VHGR solutions have been applied. The following categories are considered:
+  • 1.
+    Sign Language Understanding (SLU): Involves the automatic interpretation of gestures in either a single sign language (monolingual SLU (UNISIGNli2025; BOBSL-BBC)) or across multiple sign languages (multilingual SLU (SIGN-muller-etal-2023-findings)).
+  • 2.
     Human-Computer Interaction (HCI): Encompasses any form of interaction between users and computers.
-Numerous studies focus on HCI [131, 193], with particular emphasis on AR/VR applications [199, 32, 5].
-    • Human-Robot Interaction (HRI): Includes gesture-based control of mobile robots such as UGVs [17, 152] and UAVs [211, 194], industrial robotic arms [96], assistive robots [29], and surgical robots [139].
+Numerous studies focus on HCI (HAGRIDv2_ARXIV; ZJUGesture), with particular emphasis on AR/VR applications (EgoGesture_DATASET; OO-dMVMT; BOAT-MI_VR).
+  • 3.
+    Human-Robot Interaction (HRI): Includes gesture-based control of mobile robots such as UGVs (ULTRA-RANGE-UGV-Bamani; MOBILE_ROBOT_DRCAM) and UAVs (LONG_RANGE_HGR_SSD; MD-UHGRD_SA-YOLO), industrial robotic arms (ROBOTIC_ARM), assistive robots (DIALOG_POINTING), and surgical robots (TELEOPERATED_SURGICAL_ROBOT).
+  • 4.
     Human-Machine Interaction (HMI): Focuses on interfaces between humans and machines, especially in automotive contexts.
-VHGR enables recognition of traffic gestures (e.g., from police officers) [45, 59, 185] and supports the development of more intuitive in-vehicle user interfaces [124, 138, 53, 115, 52].
-  • Task: Complementing the application domains, this criterion focuses on the nature of the gestures themselves, specifically their temporal characteristics.
+VHGR enables recognition of traffic gestures (e.g., from police officers) (FA-STGCN; MG-GCT; CTCX_WANG2022123) and supports the development of more intuitive in-vehicle user interfaces (NVIDIA_GESTURE_R3DCNN_Molchanov; EgoFormer_METHOD_EgoDriving_DATASET; ConvMixFormer; Briareo_dataset; MVTN_ARXIV).
+3.2 Task
+Complementing the application domains, this criterion focuses on the nature of the gestures themselves, specifically their temporal characteristics.
 Gestures are broadly classified as either static or dynamic, depending on whether they involve a fixed hand posture or a sequence of movements over time.
-Dynamic gestures typically consist of three temporal phases: preparation, core (or nucleus), and retraction [124], which reflect the gesture’s evolution from initiation to completion.
-Illustrative examples are provided in Fig.
+Dynamic gestures typically consist of three temporal phases: preparation, core (or nucleus), and retraction (NVIDIA_GESTURE_R3DCNN_Molchanov), which reflect the gesture’s evolution from initiation to completion.
+The recognition of dynamic gestures can be further divided into isolated and continuous, considering only one gesture per data sequence and identifying gestures in a continuous manner, respectively.
+3.3 Input
+In continuation of the previous criteria, which focused on application contexts and gesture types, this criterion examines the nature and variety of visual input data used for VHGR.
+The choice of input modality directly influences the representation architecture and the deep learning models employed.
+While RGB frames remain the most commonly used modality, other forms of visual data have also been explored—such as depth maps (including RGB-like representations, such as HHA (SAM_SLR)), infrared (IR) images, skeleton joints (and derived features like bones and motion (TMS-NET)), pose heatmaps (NEURIPS2022_6cd3ac24_HEATMAPS; NLA_SLR), optical flow and motion history (MOTION_HISTORY_Mercanoglu).
+In terms of the number of input modalities, methods are classified as uni-modal, when they use a single type of input data, or multi-modal, when they integrate two or more modalities.
+In the case of multiple modalities in the input, a multi-modality fusion strategy is chosen (MMTM; CTFB_Hampiholi).
+This could be:
+  • 1.
+    Early (data-level) fusion: Combines modalities at the input level before processing (DHGR_3D_POSE_HRI; 3STREAM_DISTILLATION).
+  • 2.
+    Late (score-level) fusion: Fuses final prediction outputs from independent uni-modal models (MOTION_HISTORY_Mercanoglu; JOINTS_MOTION_VECTORS_Improved_ST-GCN; GestFormer).
   • 3.
-The recognition of dynamic gestures can be further divided into isolated and continuous, considering only one gesture per data sequence and identifying gestures in a continuous manner, respectively.
-    Figure 3: Gesture recognition based on task: a) static [33] and b) dynamic [27] VHGR.
-  • Input Modality: In continuation of the previous criteria, which focused on application contexts and gesture types, this criterion examines the nature and variety of visual input data used for VHGR.
-The choice of input modality directly influences the representation architecture and the deep learning models employed.
-While RGB frames remain the most commonly used modality, other forms of visual data have also been explored—such as depth maps, infrared (IR) images, skeleton joints (and derived features like bones and motion [38]), pose heatmaps [27, 214], optical flow, and motion history [119].
-Representative examples are shown in Fig.
+    Intermediate (feature-level) fusion: Merges features extracted from separate modalities at a single processing level  (tang2021graph).
   • 4.
-    Based on the type of input, methods are broadly categorized as:
-    • Pose-based (or skeleton-based): These approaches exclusively process skeleton data.
-    • Appearance-based: These rely on visual cues from RGB, depth, IR, or similar modalities.
-    (a) From left to right: RGB, Optical Flow, Depth and IR [124]
-    (b) Skeleton joints [102]
-    (c) Pose heatmap [27]
-    (d) Motion history [119]
+    Multi-level fusion: Enables interaction across modalities at multiple stages of processing (NLA_SLR; CTFB_Hampiholi).
+Finally, based on the type of input, the methods are broadly categorized into two groups: pose-based (or skeleton-based) approaches, which exclusively process skeleton data, and appearance-based approaches, which rely on visual cues from RGB, depth, infrared, or similar modalities.
+(a) From left to right: RGB, Optical Flow, Depth and IR (NVIDIA_GESTURE_R3DCNN_Molchanov)
+(b) Skeleton joints (UNISIGNli2025)
+(c) Pose heatmap (NEURIPS2022_6cd3ac24_HEATMAPS)
+(d) Motion history (MOTION_HISTORY_Mercanoglu)
 Figure 3: Modalities that are considered for VHGR.
-    In terms of the number of input modalities, methods can be classified as:
-    • Uni-modal: Utilize a single type of input data.
-    • Multi-modal: Integrate two or more modalities. These methods can be further categorized based on their multi-modality fusion strategy [177, 60] (see Fig. 5):
-        Early (data-level) fusion: Combines modalities at the input level before processing [49, 180].
-        Late (score-level) fusion: Fuses final prediction outputs from independent uni-modal models [119, 167, 51].
-        Intermediate (feature-level) fusion: Merges features extracted from separate modalities at a single processing level  [172].
-        Multi-level fusion: Enables interaction across modalities at multiple stages of processing [214, 60].
-    Figure 5: Multi-modality fusion techniques: a) Early (data-level) fusion, b) Late (score-level) fusion, c) Intermediate (feature-level) fusion, and d) Multi-level fusion. Illustration from [60]
-  • Type of Information Conveyed: Building upon the previous criterion, which focused on the nature of the input, this dimension considers the meaning or intent behind the gestures.
-Specifically, it examines the kind of information gestures are meant to transmit [16], resulting in the following categories (illustrated in Fig.
-6):
-    • Semaphoric gestures convey specific commands or instructions and are commonly used in gesture-based interfaces [124, 131, 17].
-    • Language gestures function similarly to semaphoric ones but are exclusively used in the context of sign languages [12, 158].
-    • Deictic (pointing) gestures are used to indicate objects, directions, or spatial targets [127].
-    • Manipulative gestures aim to interact with or mimic the manipulation of physical objects [50].
-    (a) Semaphoric [86]
-    (b) Language [12]
-    (c) Pointing [127]
-    (d) Manipulative [50]
+3.4 Type of Information Conveyed
+Building upon the previous criterion, which focused on the nature of the input, this dimension considers the meaning or intent behind the gestures.
+Specifically, it examines the kind of information gestures are meant to transmit (TAXONOMY_OF_HAND_GESTURES_2000_ASL_DATASET), resulting in the following categories (illustrated in Fig.
+4):
+  • 1.
+    Semaphoric gestures convey specific commands or instructions and are commonly used in gesture-based interfaces (NVIDIA_GESTURE_R3DCNN_Molchanov; HAGRIDv2_ARXIV; ULTRA-RANGE-UGV-Bamani).
+  • 2.
+    Language gestures function similarly to semaphoric ones but are exclusively used in the context of sign languages (BOBSL-BBC; Auslan-Daily).
+  • 3.
+    Deictic (pointing) gestures are used to indicate objects, directions, or spatial targets (DeePoint).
+  • 4.
+    Manipulative gestures aim to interact with or mimic the manipulation of physical objects (FPHA).
+(a) Semaphoric (WACV_kapitanov2024hagrid)
+(b) Language (BOBSL-BBC)
+(c) Pointing (DeePoint)
+(d) Manipulative (FPHA)
 Figure 4: Main categories of hand gestures based on type of information conveyed.
-  • Camera Point of View: Complementing the previous criteria on gesture content and input modalities, this dimension focuses on the spatial relationship between the camera and the performer [199].
+3.5 Camera Point of View
+Complementing the previous criteria on gesture content and input modalities, this dimension focuses on the spatial relationship between the camera and the performer (EgoGesture_DATASET).
 The camera’s perspective influences both the visual characteristics of the gesture and the recognition model design.
 Three primary viewpoints are considered (see Fig.
-7):
+5):
+  • 1.
     First-person view (egocentric): The camera is mounted on the performer, capturing the gestures from their own perspective.
+  • 2.
     Second-person view: The camera acts as the direct recipient of the gestures, typically simulating the perspective of a human observer or interacting agent.
+  • 3.
     Third-person view: The camera observes the gesture from an external position, detached from the performer.
-    (a) First-person view (egocentric) [199]
-    (b) Second-person view [149]
-    (c) Third-person view [158]
+(a) First-person view (egocentric) (EgoGesture_DATASET)
+(b) Second-person view (Lsa64argentiniansignlanguage)
+(c) Third-person view (Auslan-Daily)
 Figure 5: Different camera points of view for VHGR.
-  • Architecture: Building on the previous criteria, which focused on the nature of input and camera perspective, this criterion concerns the architectural design of the gesture recognition system, regarding the modeling stages performed by the backbone.
+3.6 Architecture
+Building on the previous criteria, which focused on the nature of input and camera perspective, this criterion concerns the architectural design of the gesture recognition system, regarding the modeling stages performed by the backbone.
 Specifically, it refers to how spatial and temporal information is processed and integrated to model gesture dynamics.
 Three main categories are identified:
+  • 1.
     Spatial architectures consider only spatial data, such as isolated RGB frames, and fed them into corresponding DNNs (mainly 2DCNNs) to capture spatial information.
+  • 2.
     Temporal architectures extract only temporal features, utilizing sequential models to process sequences of raw data, such as coordinates of skeleton joints.
+  • 3.
     Spatiotemporal modeling architectures, that simultaneously analyze both the spatial and temporal domain, such as 3DCNNs.
+  • 4.
     Spatial + Temporal pipelines, that firstly extract spatial features (e.g., using 2DCNNs) and secondly pass them through a sequential model (e.g., BiLSTM) to perform temporal modeling.
+  • 5.
     Short-term Spatiotemporal + Long-term Temporal pipelines, which aim, given a sequence of data, to capture local context (that is, they extract spatiotemporal features given small parts of the overall sequence) and then fed them into a temporal model to aggregate the partial information and obtain the global vector representation.
-  • Learning Strategy: This criterion explores the different training methodologies adopted by researchers to develop VHGR models, often influenced by data availability, task complexity, and desired generalization.
+3.7 Learning Strategy
+This criterion explores the different training methodologies adopted by researchers to develop VHGR models, often influenced by data availability, task complexity, and desired generalization.
 The main strategies identified include:
-    • Transfer Learning, where models are pretrained on large-scale datasets (e.g., Jester [117]) and subsequently fine-tuned for specific downstream VHGR tasks [112].
-    • Weakly Supervised Learning, often used in Continuous Hand Gesture Recognition (CHGR), where temporal annotations are coarse or unaligned, lacking exact gesture boundaries [188].
-    • Knowledge Distillation, in which one part of the network (often a teacher model) transfers knowledge to another (student model) to enhance feature extraction and efficiency [122, 197, 27].
-    • Contrastive Learning, used either in a CLIP-based framework to support zero-shot/few-shot recognition by leveraging textual descriptions [22, 146, 23], or to improve the discriminative capacity of learned representations [203, 204].
-    • Multi-task Learning, where a single model is optimized to perform multiple related tasks simultaneously, encouraging knowledge sharing and improved generalization [102].
-    • Reinforcement Learning, applied to optimize decision-making processes, such as keyframe selection, by rewarding effective representations [46].
-    • Few/Zero-shot Learning, which enable the model to generalize to new gestures or languages with limited or no annotated examples, often leveraging semantic or textual guidance [22, 146, 23].
-    • Adversarial Learning, used as a data augmentation strategy [13] to improve robustness, as demonstrated in adversarial video synthesis networks [128].
-    • Masked Reconstruction Pretraining, inspired by BERT-style architectures, in which models are pretrained to reconstruct masked skeleton sequences, injecting structural prior knowledge (e.g., hand topology) [202, 71, 68].
-This comprehensive taxonomy of vision-based hand gesture recognition (VHGR) methods, has been grounded in a diverse set of classification criteria that span from application domains to architectural and training strategies.
-Among these, the task criterion, distinguishing between static and dynamic gestures based on their temporal properties, emerges as particularly pivotal.
-It fundamentally shapes not only the selection of appropriate models and input modalities but also the interpretation of gestures in real-world scenarios.
-Due to its central role in determining the complexity and design of recognition systems, the task criterion serves as a natural point of departure for further analysis.
-The next three sections delve deeper into this dimension by systematically examining the core types of VHGR tasks—Static Visual Hand Gesture Recognition, Isolated Dynamic VHGR, and Continuous VHGR—each with its own challenges, methodologies, and applications.
+  • 1.
+    Transfer Learning, where models are pretrained on large-scale datasets (e.g., Jester (Jester-Dataset)) and subsequently fine-tuned for specific downstream VHGR tasks (pretrained_on_jester_3dDenseNet_LSTM).
+  • 2.
+    Weakly Supervised Learning, often used in Continuous Hand Gesture Recognition (CHGR), where temporal annotations are coarse or unaligned, lacking exact gesture boundaries (Fangyun_Wei_ICCV).
+  • 3.
+    Knowledge Distillation, in which one part of the network (often a teacher model) transfers knowledge to another (student model) to enhance feature extraction and efficiency (VAC-CSLR; C2ST; NEURIPS2022_6cd3ac24_HEATMAPS).
+  • 4.
+    Contrastive Learning, used either in a CLIP-based framework to support zero-shot/few-shot recognition by leveraging textual descriptions (ZERO-SHOT-SLR; ZS-GR; FEW-SHOT-BILGE2024110374), or to improve the discriminative capacity of learned representations (MASA; Weichao-Zhao).
+  • 5.
+    Multi-task Learning, where a single model is optimized to perform multiple related tasks simultaneously, encouraging knowledge sharing and improved generalization (UNISIGNli2025).
+  • 6.
+    Reinforcement Learning, applied to optimize decision-making processes, such as keyframe selection, by rewarding effective representations (KEYFRAMES_DRL_SKELETON).
+  • 7.
+    Few/Zero-shot Learning, which enable the model to generalize to new gestures or languages with limited or no annotated examples, often leveraging semantic or textual guidance (ZERO-SHOT-SLR; ZS-GR; FEW-SHOT-BILGE2024110374).
+  • 8.
+    Adversarial Learning, used as a data augmentation strategy (alimisis2025advances) to improve robustness, as demonstrated in adversarial video synthesis networks (ADVERSARIAL_DATA_AUGM).
+  • 9.
+    Masked Reconstruction Pretraining, inspired by BERT-style architectures, in which models are pretrained to reconstruct masked skeleton sequences, injecting structural prior knowledge (e.g., hand topology) (BEST; SignBERT+; SignBERT_ICCV).
+3.8 Processing
+Existing methods for isolated and continuous dynamic HGR can be further categorized based on whether the models have full access to the entire gesture(s) sequence or not.
+According to (NVIDIA_GESTURE_R3DCNN_Molchanov), two types of processing strategies apply:
+  • 1.
+    Online (real-time) processing: Online HGR models perform gesture recognition in a stream of input frames.
+In other words, these models take as input the gesture sequence recorded until a specific time.
+Thus, online gesture recognition is suitable for real-time applications, as the gesture is recognized before the retraction part, reducing the latency.
+Sliding windows and gesture detector/classifier pipelines are often employed (IPN-HAND; survey-emporio2025continuous).
+  • 2.
+    Offline: Offline processing means that the model treats the entire gesture sequence.
+This approach is limited to applications that do not require real-time recognition.
+Most of the existing methods for VHGR fall into this category, such as those using BiLSTM for temporal modeling (VAC-CSLR).
+The presented taxonomy of VHGR methods highlights the diversity of approaches developed to address different aspects of gesture understanding—from isolated dynamic gestures to sign language recognition and multimodal human–computer interaction.
+It provides a comprehensive foundation for analyzing current trends and identifying research gaps.
+Building on this overview, the following three sections delve deeper into specific VHGR categories, examining in detail the latest methods for Isolated Dynamic, Continuous, and Multimodal gesture recognition, respectively.
 4 Methods for Static VHGR
-Static VHGR aims to identify gestures relying only on spatial data.
+Static VHGR (SHGR) aims to identify gestures relying only on spatial data.
 As this approach neglects temporal information, is only suitable for recognizing static hand gestures.
-Proposed methods for static VHGR can be further categorized into single-stage and multi-stage approaches, based on whether the architecture consists of a single model or a multi-model processing pipeline, respectively.
-Indicative works of each category are briefly analyzed into the two following subsections.
-The third subsection concludes with the main features of state-of-the-art static VHGR methods, and discusses the open challenges that concern the community.
-4.1 Single-stage Approaches
-Single-stage approaches rely on a single model, which varies depending on the task.
-For gesture classification, fine-tuned variants of well-known 2D-CNN architectures are employed, such as VGG and DenseNet, or custom lightweight architectures are trained from scratch.
+For gesture classification, fine-tuned variants of well-known 2DCNN architectures are employed, such as VGG and DenseNet, or custom lightweight architectures are trained from scratch.
 When classification is combined with localization of gestures, variants of YOLO are adopted.
-As far as it concerns the application, static VHGR methods are used in Sign Language Understanding (SLU) and Human Computer Interaction (HCI).
-However, a few more general-purpose models have also been proposed:
-  • For Sign Language Understanding, the authors in [157] developed G-CNN, a shallow 2DCNN to recognize Indian Sign Language (ISL) signs and slightly outperformed various versions of VGG on a self-collected dataset.
-The authors in [54] proposed the STN-ND-Net, a combination of a Spatial Transformer Network (STN) that improves robustness to spatial transformations, with a 2D-CNN backbone that uses deconvolution operations to reduce the pixel-wise and channel-wise correlations and eliminates the information redundancy of colored images.
-The success of vision transformer (ViT) motivated authors of SIGNFORMER[93] to adopt it for ISL recognition, while authors in [161] leveraged both skeleton and RGB data in their dual stream feature extractor and an MLP to perform intermediate fusion.
-A YOLO variant, namely YOLOv4-CSP[8] has also been developed, to simultaneous detect and identify the hand gesture.
-  • For gesture-based Human-Computer Interaction the authors in [198] proposed a lightweight 2D-CNN, named LHGR-Net, which combines a Multiscale Structure (MSS) that aims to make the model scale-invariant via atrous convolution with a Lightweight Attention Structure (LAS) based on CBAM.
-The authors in [80] proposed 2DPSTPP-Net, a 2D-CNN with four stacked convolutional blocks whose features are pooled using max pooling and a novel method, then concatenated to form the final representation, tailored for HCI scenarios.
-The authors in [61] modified the YOLOv5 to make it more lightweight and developed a model that can be used for gesture-based user interfaces in smart sports stadiums, improving the experience of spectators and athletes.
+Hand segmentation is often applied to eliminate the background, reducing the interference of gesture-irrelevant information.
+Hand detection is also utilized in some cases to normalize the spatial dimensions, especially for long-range recognition.
+As far as it concerns the application, static VHGR methods are used in SLU, HCI and HRI.
+However, a few more general-purpose models have also been proposed.
+4.1 Sign Language Understanding (SLU)
+For SLU, the authors in (G-CNN-SHARMA2021115657) developed Gesture-CNN, a shallow 2DCNN to recognize Indian Sign Language (ISL) signs and slightly outperformed various versions of VGG on a self-collected dataset.
+The authors in (STN-ND-Net) proposed the STN-ND-Net, a combination of a Spatial Transformer Network (STN) that improves robustness to spatial transformations, with a 2DCNN backbone that uses deconvolution operations to reduce the pixel-wise and channel-wise correlations and eliminates the information redundancy of colored images.
+The success of vision transformer (ViT) motivated authors of SIGNFORMER (SIGNFORMER) to adopt it for ISL recognition, while authors in (Jungpil-Shin) leveraged both skeleton and RGB data in their dual stream feature extractor and an MLP to perform intermediate fusion.
+A YOLO variant, namely YOLOv4-CSP(YOLOv4-CSP) has also been developed, to simultaneous detect and identify Turkish Sign Language gestures.
+4.2 Human-Computer Interaction (HCI)
+For gesture-based HCI the authors in (LHGR-Net) proposed a lightweight 2DCNN, named LHGR-Net, which combines a Multiscale Structure (MSS) that aims to make the model scale-invariant via atrous convolution with a Lightweight Attention Structure (LAS) based on CBAM.
+The authors in (2DPSTPP) proposed 2DPSTPP-Net, a 2DCNN with four stacked convolutional blocks whose features are pooled using max pooling and a novel method, then concatenated to form the final representation, tailored for HCI scenarios.
+The authors in (WEI_HAN_YOLOV5_COPULA) modified the YOLOv5 to make it more lightweight and developed a model that can be used for gesture-based user interfaces in smart sports stadiums, improving the experience of spectators and athletes.
+In (Mohammad_Mahmudul_static_hgr), a pipeline for egocentric HGR and fingertip localization uses YOLOv2 for hand detection, followed by VGG16 for feature extraction and gesture classification.
+In (tuan_linh_pham_dang2023lightweight), a lightweight HCI pipeline segments the hand using DeepLabV3+ or Unet, then classifies the cropped image with a 2DCNN like MobileNet or EfficientNet.
+(FGDSNet) introduced FGDSNet, a two-stage approach with dual-stream segmentation (local detail and semantic branches) and multi-level fusion for feature integration.
+The segmentation mask and original image are fed into a dual-stream classification module, where mid-level fusion combines features for final classification via an MLP.
+4.3 Human-Robot Interaction (HRI)
+HRI, particularly gesture-based mobile robot navigation, has attracted research interest.
+URGR (ULTRA-RANGE-UGV-Bamani) is a three-stage approach using YOLOv3 for human cropping, High-Quality Network (HQ-Net) for super-resolution to address long-range scenarios, and a classifier, Graph-Vision Transformer (GViT), for gesture recognition.
+The HQ-Net comprises convolutional and deconvolutional pathways.
+Features extracted by Canny Edge algorithm and a convolutional module followed by self-attention are fused with the latent representation.
+GViT utilizes two GCN layers that capture fine-grained details followed by a modified ViT.
+This setting enables effective recognition of gestures in a distance of up to 25m for UGV guidance.
+Similarly, (DRCAM_TIM) introduced a two-stage method based on depth data: hand segmentation followed by gesture classification using DRCAM, a 2DCNN with stacked Residual Channel Attention Modules (RCAMs) and dense interconnections between the cascaded blocks.
+Because of the attention mechanisms, the model focuses effectively on the most informative regions of the image.
+4.4 General Purpose
 General-purpose models have also been developed.
-EDenseNet [171] extends DenseNet by adding convolutional layers between dense blocks.
-SpAtNet [21] is a lightweight 2DCNN with three streams: two MAFF modules for coarse-grained features and one Interleaved Module for high-level representations; their fused features are fed to an MLP for classification.
-MMFANet [103] performs gesture detection and recognition using a 2DCNN backbone with task-specific regressor and classifier branches.
-4.2 Multi-stage Approaches
-Figure 8: Architecture of an indicative multi-stage dual-stream approach for static VHGR named FGDSNet [208].
-In contrast to single-stage, multi-stage approaches set up a multi-model processing pipeline.
-Typically, a hand detector or segmentation model is used to remove the background and normalize the spatial dimensions.
-These methods are primarily applied in Human Robot Interaction (HRI) and HCI scenarios, as detailed below:
-  • Human-Robot Interaction, particularly gesture-based mobile robot navigation, has attracted research interest.
-URGR [17] is a three-stage approach using YOLOv3 for human cropping, HQ-Net for super-resolution to address long-range scenarios, and a classifier for gesture recognition.
-Similarly, [153] introduced a two-stage method based on depth data: hand segmentation followed by gesture classification using DRCAM, a 2DCNN with stacked Residual Channel Attention Modules (RCAMs) and dense interconnections.
-  • Multi-stage static VHGR methods have also been proposed for Human-Computer Interaction.
-In [9], a pipeline for egocentric HGR and fingertip localization uses YOLOv2 for hand detection, followed by VGG16 for feature extraction and gesture classification.
-In [34], a lightweight HCI pipeline segments the hand using DeepLabV3+ or Unet, then classifies the cropped image with a 2D-CNN like MobileNet or EfficientNet.
-[208] introduced FGDSNet, a two-stage approach with dual-stream segmentation (local detail and semantic branches) and multi-level fusion for feature integration.
-The segmentation mask and original image are fed into a dual-stream classification module, where mid-level fusion combines features for final classification via an MLP.
-4.3 State-of-the-art Methods for Static VHGR
-The current SOTA static VHGR methods are summarized in 7, based on input modality, architecture, and performance on public datasets, including HGR1 [88], NUS-I [95], NUS-II [136], ASL-FS [137], Jochen-Triesch II [175], and OUHANDS [118] [^1] For datasets with splits for both subject-independent and subject-dependent modes, only the latter is considered..
-Since most methods are evaluated on different or custom datasets, a straightforward performance comparison is not possible.
-As shown, most SOTA methods rely on RGB data, with depth-based DRCAM [153] lagging behind RGB methods on ASL-FS.
-Researchers use attention mechanisms [153, 198, 208], multi-stream approaches [21, 208], and segmentation modules for background removal [208, 34, 153] to improve performance.
-On OUHANDS [118], ASL-FS [137], and NUS-I [95], single-stage methods outperform multi-stage ones, while on NUS-II [136], FGDSNet [208] with segmentation achieves slightly better accuracy than 2DPSTPP-Net [80].
-Thus, background removal does not always improve sign recognition, as single-stage methods often surpass multi-stage models.
-Table 3: Static HGR methods and their performance on well-known datasets. The highest accuracies are shown with bold. Methods marked with star (*) perform hand segmentation.
-Method Modality Architecture Description Dataset Accuracy
-SpAtNet [21] RGB Three-stream 2DCNN, where each stream aims to obtain either high-level or coarse-grained features HGR1 93.33%
-ASL-FS 99.67%
-NUS-II 96.53%
-FGDSNet [208]* RGB Dual stream 2DCNN approach and a custom dual-stream segmentation module to obtain the mask HGR1 94.8%
-OUHANDS 97.8%
-NUS-II 99.8%
-[34]* RGB DL-based Hand Segmenation + Classifier HGR1 98.15%
-OUHANDS 89.2%
-LHGR-Net [198] RGB BaseNet + MSS (adopting the atrous spatial pyramid pooling) + LAS (based on CBAM) HGR1 93.15%
-OUHANDS 98.75%
-2DPSTPP-Net [80] RGB Convolutional blocks followed by the proposed pooling method, intermediate features are concatenated to obtain the final representation. NUS-I 100%
-NUS-II 99.61%
-DRCAM [153]* Depth Attention-based 2CNN with residual connections ASL-FS 93.44%
-Gesture-CNN [157] RGB 2DCNN consisting of four convolutional layers Jochen-Triesch II 100%
-EDenseNet [171] RGB DenseNet with extra convolution layers ASL-FS 99.99%
-NUS-II 99.3%
+EDenseNet (EDenseNet) extends DenseNet by adding convolutional layers between dense blocks.
+SpAtNet (SpAtNet) is a lightweight 2DCNN with three streams: two Multi-scale Attentive Feature Fusion (MAFF) modules for coarse-grained features and one Interleaved Module for high-level representations; their fused features are fed to an MLP for classification.
+This approach, employing multiple kernel sizes, increases accuracy by capturing various levels of granularity.
+MMFANet (MMFANet) performs gesture detection and recognition using a 2DCNN backbone followed by the proposed feature aggregation pyramid network (FAPN) and task-specific regressor and classifier branches.
+The FAPN module aims to extract multiscale features through upsampling/downsampling operations and fusion of adjacent feature maps, while reducing the computational complexity by utilizing novel expanded convolutional blocks with different dilation rates combined with attention mechanisms that eliminate redundant information.
+4.5 Performance of SHGR Methods
+Most existing methods for Static Hand Gesture Recognition (SHGR) are evaluated on different or custom datasets, which prevents a direct comparison of their performance.
+Table 2 presents the accuracy of current state-of-the-art SHGR approaches.
+Despite their high accuracy and low parameter count—suggesting that SHGR might be a largely solved problem—the datasets used (ASL-FS (ASL-FS-2011-CVPRW), HGR1 (HGR-1-2024-PATTER-RECOGNITION), NUS-I (NUS-I-2010), NUS-II (NUS-II-2013), and OUHANDS (ouhands)) remain relatively small and lack diversity.
+Consequently, model performance would likely degrade in real-world, uncontrolled environments.
+Among the existing methods, LHGR-Net (LHGR-Net) and FGDSNet (FGDSNet) demonstrate strong results on the OUHANDS dataset with lightweight architectures, making them suitable for gesture-based human–computer interaction.
+On the other hand, EDenseNet (EDenseNet) achieves high accuracy on ASL-FS but at a higher computational cost, making it better suited for static sign language recognition.
+Finally, applying background removal through hand segmentation does not appear to significantly improve accuracy compared to end-to-end learning approaches.
+Table 2: Performance of the most accurate SHGR methods.
+Many approaches achieve high accuracy even with small model size, but the limited size and diversity of the utilized datasets may result to performance reduction during real-world deployment.
+Methods marked with star (*) perform hand palm segmentation first.
+Method Year Modality Params ASL-FS HGR1 NUS-I NUS-II OUHANDS
+FGDSNet (FGDSNet) * 2024 RGB 1.20M - 94.80% - 99.80% 97.80%
+SpAtNet (SpAtNet) 2024 RGB 3.8M - 93.33% - - -
+(tuan_linh_pham_dang2023lightweight) * 2023 RGB 5.34M - 98.15% - - 89.20%
+LHGR-Net (LHGR-Net) 2023 RGB 1.81M - 93.15% - - 98.57%
+2DPSTPP-Net (2DPSTPP) 2023 RGB 19.01M - - 100% 99.61% -
+DRCAM (DRCAM_TIM) * 2023 Depth - 93.44% - - - -
+EDenseNet (EDenseNet) 2021 RGB 7.77M 99.99% - - 99.30% -
 5 Methods for Isolated Dynamic VHGR
-Unlike static VHGR approaches, which often fail to capture temporal dynamics and struggle with motion-based gestures, recent advances in Isolated Dynamic Vision-based Hand Gesture Recognition (IDVHGR) have focused on recognizing individual gesture instances from spatiotemporal inputs such as RGB video sequences or skeleton joint trajectories.
+Unlike static VHGR approaches, which often fail to capture temporal dynamics and struggle with motion-based gestures, recent advances in Isolated Dynamic Vision-based Hand Gesture Recognition (IDVHGR or IDHGR) have focused on recognizing individual gesture instances from spatiotemporal inputs such as RGB video sequences or skeleton joint trajectories.
 When gestures are part of a sign language, the task is referred to as Isolated Sign Language Recognition (ISLR).
@@
 In the following, we present IDVHGR methods grouped by input modality as (a) unimodal appearance-based, (b) pose-based, and (c) multimodal approaches.
-Then, we conclude with a summary of key innovations and state-of-the-art (SOTA) performance.
+Then, we conclude with a summary of key innovations and SoTA performance.
 5.1 Unimodal Appearance-based Methods
@@
 As shown in Fig.
-4, these representations resemble color images, which allows for the application of similar computer vision techniques.
+3, these representations resemble color images, which allows for the application of similar computer vision techniques.
 These methods can be further categorized based on the architecture of the feature extractor, which typically follows one of three main pipelines: (1) spatiotemporal, (2) spatial + temporal, and (3) short-term spatiotemporal + long-term temporal backbones.
@@
 Spatiotemporal backbones perform spatial and temporal modeling jointly, typically using 3DCNNs or (2+1)D CNNs—either custom-built or adapted from existing models.
-In [182], the authors introduced a custom network, (2+1)D-SLR, composed of five stacked (2+1)D convolutional blocks (i.e., a 2D convolution followed by a 1D convolution).
+In (2plus1-D-SLR_and_NCSL), the authors introduced a custom network, (2+1)D-SLR, composed of five stacked (2+1)D convolutional blocks (i.e., a 2D convolution followed by a 1D convolution).
 This structure lowers computational cost compared to traditional 3D convolutions, making the model more practical for deployment.
-To enhance temporal modeling, authors in [20] proposed a trainable key frame selection method based on error correction.
+To enhance temporal modeling, authors in (Snehal_Bharti_et_al) proposed a trainable key frame selection method based on error correction.
 VGG16 features are used to identify key frames, which are then preprocessed and fed into a custom 3DCNN for classification.
-In [186], a dual-stage multi-stream pipeline for CSL recognition was developed.
+In (ShiqiWang_et_al), a dual-stage multi-stream pipeline for CSL recognition was developed.
 An adapted EfficientDet model—enhanced with cross-level Bi-FPN connections and a dual-channel spatial attention module (DCSAM)—detects both hands.
 The full image and cropped hand patches are passed through three parallel 3D-ResNet-18 models (following the P3D scheme), with their features fused and classified via an MLP.
-Authors in [63] introduced TS3C-Net, a modified 3D ConvNeXt that adds an extra 3×3×3 convolutional branch to each block, enhancing spatiotemporal representation.
-Finally, [173] addressed limitations in fixed-pipeline designs by combining features from both ResC3D and ConvLSTM using the Selective Spatiotemporal (SeST) module.
+Authors in (TS3C-Net) introduced TS3C-Net, a modified 3D ConvNeXt that adds an extra 3×3×3 convolutional branch to each block, enhancing spatiotemporal representation.
+Finally, (SeST) addressed limitations in fixed-pipeline designs by combining features from both ResC3D and ConvLSTM using the Selective Spatiotemporal (SeST) module.
 The fused features are passed through fully connected and softmax layers for classification.
@@
 The resulting sequence embeddings are aggregated, and a classifier (e.g., MLP) outputs the final prediction.
-In [65], a method tailored for HCI tasks using only depth data was proposed.
+In (DAVID_RICHARD_TOM_HAX), a method tailored for HCI tasks using only depth data was proposed.
 The method combined Inception-v3 that extracted spatial features with an LSTM for temporal modeling.
-For ISLR, authors in [178] introduced a pipeline to recognize Indian Sign Language (ISL) words used by deaf farmers.
+For ISLR, authors in (GoogleNet-BiLSTM) introduced a pipeline to recognize Indian Sign Language (ISL) words used by deaf farmers.
 They combined GoogleNet for spatial encoding with a BiLSTM for sequence modeling.
-To reduce inference time, a lightweight ISLR model, called TIM-SLR has been proposed in [183].
-The model uses a 2D-CNN for per-frame feature extraction, followed by a Time Interaction Module (TIM)—a zero-parameter unit that fuses frame features into a compact representation.
-Finally, the Swin Transformer presented in [40] is a two-stage transformer-based model that extracts spatial features per frame, while a causally masked transformer processes the sequence, ensuring that only past and present frame features contribute at each step.
+To reduce inference time, a lightweight ISLR model, called TIM-SLR has been proposed in (TIM-SLR).
+The model uses a 2DCNN for per-frame feature extraction, followed by a Time Interaction Module (TIM)—a zero-parameter unit that fuses frame features into a compact representation.
+Finally, the Swin Transformer presented in (Full_Transformer_Network) is a two-stage transformer-based model that extracts spatial features per frame, while a causally masked transformer processes the sequence, ensuring that only past and present frame features contribute at each step.
 5.1.3 Short-Term Spatiotemporal + Long-Term Temporal
 The methods in this category follow a two-stage architecture: the first stage captures short-term spatiotemporal patterns from segments of the input sequence, while the second stage aggregates these features to form a final representation of the entire sequence.
-EgoFormer [138] is a model for egocentric dynamic gesture recognition optimized for edge devices, which has been introduced in the context of autonomous driving.
-EgoFormer combines a custom 3D-CNN with a transformer module and an MLP for final classification.
-For long-range gesture-based UGV control, authors in [18] proposed a pipeline where a lightweight 2D CNN extracts per-frame features, which are used to select key frames.
+EgoFormer (EgoFormer_METHOD_EgoDriving_DATASET) is a model for egocentric dynamic gesture recognition optimized for edge devices, which has been introduced in the context of autonomous driving.
+EgoFormer combines a custom 3DCNN with a transformer module and an MLP for final classification.
+For long-range gesture-based UGV control, authors in (SlowFast-Transformer) proposed a pipeline where a lightweight 2DCNN extracts per-frame features, which are used to select key frames.
 After cropping the human region, the selected frames are passed through a SlowFast network, followed by a transformer and an MLP.
-To address data scarcity in sign language understanding, [22] proposed a zero-shot learning framework.
+To address data scarcity in sign language understanding, (ZERO-SHOT-SLR) proposed a zero-shot learning framework.
 Sign videos are represented using BERT-encoded textual descriptions combined with hand-crafted attributes.
 This representation guides a dual-stream backbone composed of short-term spatiotemporal modules (e.g., I3D) and long-term models (e.g., BiLSTM).
-Figure 9: The zero-shot framework proposed by authors in [22].
-Each branch of the dual-stream backbone consists of a spatial-temporal model, such as the I3D, followed by a long-term recurrent model.
-The BERT is utilized to map textual description into the embedding space.
-These features are also combined with hand-crafted attributes.
 5.1.4 Other
 Some appearance-based methods fall outside the previously discussed categories but offer valuable contributions, especially in Sign Language Recogition (SLR).
-For example, the method proposed in [113] for Indian SLR (ISLR) compresses an entire frame sequence into a single image, which is then processed by a MobileNet.
+For example, the method proposed in (HAMZAH_LUQMAN) for Indian SLR (ISLR) compresses an entire frame sequence into a single image, which is then processed by a MobileNet.
 A second MobileNet extracts per-frame features, which are passed through a two-layer LSTM.
-The outputs of both streams are concatenated and refined using a custom shallow 2D-CNN.
-A different approach was presented in [94], where the authors introduced a dual-stream ISLR architecture combining Inception V4 with a novel convolutional module called stem.
-In the context of pose-based ISLR, [110] proposed a dual-stream model with keyframe selection.
-A 2D-CNN and attention mechanism filter out less informative frames.
+The outputs of both streams are concatenated and refined using a custom shallow 2DCNN.
+A different approach was presented in (Ensemble-InceptionV4), where the authors introduced a dual-stream ISLR architecture combining Inception V4 with a novel convolutional module called stem.
+In the context of pose-based ISLR, (Tianyu_Liu) proposed a dual-stream model with keyframe selection.
+A 2DCNN and attention mechanism filter out less informative frames.
 The remaining keyframes are processed by two branches: the Hand stream, which crops and analyzes hand regions with attention blocks, and the Full stream, which processes the entire frame.
 The fused features are used for final classification.
-To improve signer-independence and reduce overfitting, the SwC GR-MMixer [109], another lightweight model based on MLP-Mixer, incorporates GRU units for temporal modeling and a shifted window concatenation strategy to enhance sequence representation.
-5.2 Pose-based Methods
+To improve signer-independence and reduce overfitting, the SwC GR-MMixer (SwC-GR-MMixer), another lightweight model based on MLP-Mixer, incorporates GRU units for temporal modeling and a shifted window concatenation strategy to enhance sequence representation.
+5.2 Unimodal Pose-based Methods
 Pose-based methods, also known as skeleton-based approaches, rely on structured motion data—such as skeleton graphs and pose heatmaps—to classify gestures.
-These methods often incorporate various forms of pose information, including joint coordinates, bones, motion trajectories, and joint angles, sometimes in combination [38].
-They also explore joint subsets related to specific body regions [135], often adopting multi-stream architectures to handle these diverse inputs.
+These methods often incorporate various forms of pose information, including joint coordinates, bones, motion trajectories, and joint angles, sometimes in combination (TMS-NET).
+They also explore joint subsets related to specific body regions (FDMSE-hierarchicalwindowedgraphattention), often adopting multi-stream architectures to handle these diverse inputs.
 Given the natural graph structure of skeletal data, many methods employ Graph Convolutional Networks (GCNs) originally developed for human activity recognition.
 In the sections that follow, we group the approaches based on the structure of their backbone, following the same organization as in the previous subsection.
-Notably, pose information alone is often insufficient for accurate sign recognition [125].
+Notably, pose information alone is often insufficient for accurate sign recognition (moryossef2021evaluating).
 Consequently, many models integrate additional cues—typically from RGB inputs—using 3DCNNs and perform score-level fusion to enhance performance.
 5.2.1 Spatiotemporal
-Pose-based methods that incorporate spatiotemporal representations simultaneously process spatial and temporal data to improve sign language understanding (SLU), sign language recognition (SLR), and human-computer interaction (HCI).
-In Sign Language Understanding (SLU), pose-based methods aim to capture dynamic hand and body movements for gesture recognition.
-SKIM [104] uses a dynamic graph convolutional network (GCN) combined with temporal convolution (TCN) to analyze both global and local body regions, with a feature extraction module focusing on high-motion joints.
-Similarly, MC-LSTM [133] employs multiple streams (face, hand, and whole body) processed by temporal convolutions and GCN for improved feature ensembling.
-Another notable method, P3D [98], uses a transformer-based model with part-wise and whole-body encoders to process landmark data for classification.
-For improved performance, HWGAT [135] utilizes hierarchical spatial windows with attention mechanisms to highlight key body parts.
-TMS-Net [38] proposes a six-stream GCN-based approach for processing skeleton joints, bones, and motion data, with inter-stream information exchange.
-Additionally, methods like SLGTformer [168] and VSCT [25] address the challenge of data scarcity by leveraging transformers and few-shot learning, respectively.
-In Sign Language Recognition (SLR), pose-based methods often focus on improving the accuracy and robustness of sign classification.
-Authors in [161] present a dual-stream GCN model for Korean SLR, analyzing joint and motion data.
-Similarly, MULTISTAGE-CGAR [121] proposes a dual-stream architecture with separable temporal convolution and attention mechanisms for improved feature extraction.
-For multi-stream fusion, SwC GR-MMixer [109] combines GCN with GRU units for temporal modeling and attention mechanisms to reduce overfitting.
-Notably, L-TSSI-DenseNet [97] explores skeleton graph representations by converting joint coordinates into image-like matrices processed by DenseNet for classification.
-Moreover, STST [201] leverages a two-stage transformer-based approach for global and local spatial-temporal feature extraction, offering a dynamic alternative to traditional methods.
-In Human-Computer Interaction (HCI), pose-based methods address challenges such as data overfitting and the need for self-supervised learning frameworks.
-MAE-based frameworks [79] use a masked autoencoder to reconstruct skeleton graphs, focusing on motion and spatial relationships.
-3sISTGCN [167] proposes a three-stream model to process joints, bones, and motions, applying modified ST-GCN layers and late fusion for improved classification.
-Another approach, DG-STA [5], tackles class-incremental learning in HCI by training models on non-overlapping tasks without access to old training data.
-DSTSA-GCN [30] addresses temporal expression limitations by stacking three stages of graph convolutions and multiscale temporal convolutions.
-Additionally, DSTA-SLR [77] improves spatiotemporal modeling by introducing dynamic adjacency matrices and multi-stream fusion, enhancing the accuracy of joint, bone, and motion-based recognition tasks.
+Pose-based methods that incorporate spatiotemporal representations simultaneously process spatial and temporal data to improve sign language understanding (SLU) and human-computer interaction (HCI).
+In SLU, pose-based methods model dynamic hand and body movements for gesture recognition.
+SKIM (SKIM) combines dynamic graph convolutional (GCNs) and temporal convolutional (TCNs) networks to analyze global and local motion, while MC-LSTM (ozdemir2023multi_MC-LSTM) fuses multiple streams (face, hand, body) via temporal convolutions and GCNs.
+P3D (P3D_ICCV) employs transformer-based part-wise and whole-body encoders, and HWGAT (FDMSE-hierarchicalwindowedgraphattention) enhances key body-part focus with hierarchical attention.
+TMS-Net (TMS-NET) introduces a six-stream GCN for skeleton, bone, and motion data with inter-stream fusion.
+To tackle data limitations, SLGTformer (SLGTFORMER) and VSCT (spoter_vsct) use transformers and few-shot learning, respectively.
+Dual-stream approaches such as (Jungpil-Shin) and MULTISTAGE-CGAR (MULTISTAGE-CGAR) integrate joint and motion data through temporal convolutions and attention.
+SwC GR-MMixer (SwC-GR-MMixer) combines GCNs with GRUs for temporal modeling and regularization.
+Other notable advances include L-TSSI-DenseNet (L-TSSI-DenseNet), which converts joint coordinates into image-like matrices for DenseNet processing; STST (STST), a two-stage transformer for local–global spatiotemporal features; and DSTA-SLR (DSTA_SLR_hu2024dynamic), which employs dynamic adjacency matrices and multi-stream fusion for improved recognition accuracy.
+In HCI, pose-based methods address challenges such as data overfitting and the need for self-supervised learning frameworks.
+MAE-based frameworks (omar_ikne) use a masked autoencoder to reconstruct skeleton graphs, focusing on motion and spatial relationships.
+3sISTGCN (JOINTS_MOTION_VECTORS_Improved_ST-GCN) proposes a three-stream model to process joints, bones, and motions, applying modified ST-GCN layers and late fusion for improved classification.
+Another approach, DG-STA (BOAT-MI_VR), tackles class-incremental learning in HCI by training models on non-overlapping tasks without access to old training data.
+DSTSA-GCN (DSTSA-CGN_cui2025dstsa) addresses temporal expression limitations by stacking three stages of graph convolutions and multiscale temporal convolutions.
+Finally, authors in (ST-CGNet-hu2025st) proposed ST-CGNet, a dual-stream architecture comprising a 3DCNN and a convolutional LSTM module to capture different levels of information.
+Attention mechanisms and an advanced fusion strategy are utilized to efficiently incorporate the features of each stream.
 5.2.2 Spatial + Temporal
@@
 These features are then passed into a sequential model for temporal aggregation, resulting in a combined spatial-temporal representation of the input sequence.
-For instance, GCN-BERT [176] utilizes a two-layer GCN for spatial modeling, followed by BERT for temporal modeling.
+For instance, GCN-BERT (GCN-BERT) utilizes a two-layer GCN for spatial modeling, followed by BERT for temporal modeling.
 Other general-purpose methods adopt a multi-stream, two-stage pose-based approach.
-In [129], three streams, each containing three 1D convolutional layers, are combined with a parallel module of average pooling layers.
+In (SBI-DHGR), three streams, each containing three 1D convolutional layers, are combined with a parallel module of average pooling layers.
 The extracted features are then concatenated and input into an LSTM, followed by a fully connected layer and softmax for classification.
-Similarly, STr-GCN [165] employs a two-stage model where a GCN spatial feature extractor is first applied to obtain intra-frame features and then passes the features through a transformer encoder for modeling inter-frame temporal relationships.
+Similarly, STr-GCN (STr-GCN) employs a two-stage model where a GCN spatial feature extractor is first applied to obtain intra-frame features and then passes the features through a transformer encoder for modeling inter-frame temporal relationships.
 The final classification scores are generated by an MLP.
@@
 For example, gesture recognition in autonomous driving has become an important research area.
-In [59], the authors propose a pose-based approach called MG-GCT for recognizing police traffic gestures.
+In (MG-GCT), the authors propose a pose-based approach called MG-GCT for recognizing police traffic gestures.
 The model uses two graph convolution streams: one to process skeleton joints and another to analyze joint motions.
 The features extracted from both streams are fused and passed through a transformer encoder for temporal modeling, with the final classification performed by an MLP.
-In another approach, TD-GCN [107] introduces a pose-based method for IDHGR.
+In another approach, TD-GCN (TD-GCN) introduces a pose-based method for IDHGR.
 Unlike traditional methods that use fixed skeleton topologies, TD-GCN employs temporal-sensitive topology learning to enhance feature quality.
@@
 These matrices help model the skeleton, and the features are fused using the proposed TCF module.
+Authors in (ISLR-STGCN-attention-LSTM-ozdemir2025cross) proposed a dual-stream backbone processing both hand and whole body keypoints using ST-GCN models.
+Cross-attention is then applied to focus on the most important features, which are subsequently pass through a LSTM to model long-term dependencies.
 5.2.4 Temporal
@@
 Typically, keypoints are slightly preprocessed—often by extracting geometric features such as bone angles—and then passed into sequential models like RNNs or TCNs to produce a vector representation of the input sequence.
-In [139], the authors proposed a pose-based method for surgical robot teleoperation.
+In (TELEOPERATED_SURGICAL_ROBOT), the authors proposed a pose-based method for surgical robot teleoperation.
 Preprocessed skeleton joints are fed into an LSTM, and the resulting features are passed to an MLP for final classification.
 The model was evaluated on a custom dataset.
-Similarly, in [151], a cross-domain gesture recognition approach was developed using geometric features (e.g., bone angles), which are processed by a TCN.
+Similarly, in (Sabater_alberto), a cross-domain gesture recognition approach was developed using geometric features (e.g., bone angles), which are processed by a TCN.
 The resulting features are summarized into a sequence descriptor, and a linear classifier predicts the final class.
-To address data scarcity, the authors in [128] introduced AVSN, an adversarial learning-based data augmentation framework.
+To address data scarcity, the authors in (ADVERSARIAL_DATA_AUGM) introduced AVSN, an adversarial learning-based data augmentation framework.
 The method jointly optimizes augmentation parameters and the HGR classifier.
@@
 5.2.5 Other
-Some architectures adopt more complex designs or alternative skeleton representations, placing them outside the previously defined categories.
-In [1], the authors combined a Fourier CNN with a BiLSTM to extract spatiotemporal features from 3D skeleton joints and perform ISLR.
-A trainable binary mask was also used for feature selection to eliminate redundancy.
-Similarly, authors in [10] performed ISLR using a feature thresholding technique that transforms skeleton features based on their variance and importance.
-These transformed features undergo exhaustive search for selection and are then processed by a deep CNN for classification.
-To address limitations of color-coded feature maps, the authors in [89] proposed the Joint Motion Affinity Map (JMAM) as a new skeleton representation for ISLR.
-JMAM is input to a custom multi-stream attention-based 2DCNN, which extracts features and performs classification.
-Key frame selection was explored in [46] through the Frame Selection Network (FS-Net), trained using reinforcement learning.
-The agent selects frames for classification by DD-Net and receives a reward of +1 or -1 based on the correctness of the classification outcome.
-Finally, [101] proposed a lightweight dual-stream method to fully capture spatiotemporal features.
-One stream employs stacked attention-based modules, while the other uses an autoencoder.
-The outputs of both are classified independently, and their predictions are averaged to yield the final decision.
+Several works explore alternative architectures or skeleton representations beyond the main categories.
+In (FS-EFCNN), a Fourier CNN with BiLSTM extracts spatiotemporal features from 3D joints for ISLR, using a trainable binary mask for feature selection.
+Similarly, (SCT-DCNN) applies feature thresholding based on variance and importance, followed by deep CNN classification.
+To overcome color-coded map limitations, (MRCNNSA) introduces the Joint Motion Affinity Map (JMAM) with a multi-stream attention-based 2DCNN for feature extraction and recognition.
+Key frame selection is addressed in (KEYFRAMES_DRL_SKELETON) via FS-Net, a reinforcement learning agent that selects informative frames for DD-Net classification.
+Dual-stream architectures such as STGCN-LSTM (STGCN-LSTM-10965662) combine ST-GCN and self-attention ConvLSTM for capturing global and local features, while (Yangke_patcog) proposes a lightweight variant merging attention-based and autoencoder streams.
+Finally, the Hierarchical Self-Attention Network (HAN) (HAN-liu2025han) employs modular attention across finger joints (J-Att, F-Att, T-Att, Fusion-Att) for efficient spatial–temporal modeling, achieving strong accuracy with just 0.05 GFLOPs.
 5.3 Multimodal Methods
@@
 5.3.1 General Purpose
-In [100], a dual-stream 3DCNN followed by average fusion was proposed for RGB-D gesture recognition.
-To filter out redundant visual features ($\upsilon$) and retain only the most discriminative components ($z$), a memory bank mechanism was introduced.
-The training objective maximizes mutual information between $z$ and the gesture class $y$, while minimizing that between $z$ and $\upsilon$.
-The memory bank is updated via read/write operations during training.
-Similarly, addressing redundancy, authors in [206] applied Neural Architecture Search (NAS) and attention mechanisms.
-A modified I3D backbone replaces some inception modules with learnable NAS cells.
-These progressively transition from small-kernel to dilated convolutions.
-Dynamic Static Attention (DSA) modules focus on motion and hand/arm regions, aided by a parallel HeatmapNet trained jointly through an auxiliary loss.
-For RGB-D input, late fusion is used.
-To move beyond the limitations of late fusion and exploit modality interactions more effectively, [207] proposed a decoupling-recoupling framework.
-Independent spatial (with a Focused Region Proposal (FRP) module) and temporal modeling (via attention) are first performed to extract modality-specific features.
-These are then fused mid-way using a shared MLP-AE module.
-Multi-level fusion was further studied in [60], where 3DCNN backbones are combined with a Convolutional Transformer Fusion Block (CTFB).
-Injected at multiple layers, CTFB fuses intermediate features across streams for better spatial-temporal integration.
+In (Yunan_Li), a dual-stream 3DCNN with average fusion was proposed for RGBD gesture recognition, using a memory bank to retain discriminative features ($z$) while filtering redundant ones ($\upsilon$) by maximizing mutual information with class labels.
+Similarly, (RAAR3DNet) employed Neural Architecture Search (NAS) with Dynamic Static Attention (DSA) modules and a modified I3D backbone to emphasize motion and hand regions, supported by an auxiliary HeatmapNet and late RGBD fusion.
+To better exploit modality interactions, (Benjia_Zhou) introduced a decoupling–recoupling framework combining independent spatial and temporal modeling with mid-level fusion through a shared MLP-AE module.
+Multi-level fusion was further explored in (CTFB_Hampiholi), integrating 3DCNNs with a Convolutional Transformer Fusion Block (CTFB) that fuses intermediate cross-stream features.
 To address data scarcity, several works explored zero-shot learning (ZSL) by incorporating textual gesture descriptions encoded by pretrained language models (mainly BERT).
-In [146], a dual-stream model combined a ViT for spatial modeling and an LSTM-MLP for temporal modeling, learning visual features aligned with textual embeddings.
-Similarly, [147] fused skeleton-derived geometric features with RGB-C3D-LSTM-AE representations into a shared semantic space for classification.
-To reduce computational cost, [53] replaced the transformer’s self-attention mechanism with convolutional layers and a gated feedforward network, resulting in the lightweight ConvMixFormer architecture.
-In [180], knowledge distillation and contrastive learning were employed to enhance multimodal performance.
-Three 3DCNNs (for RGB, Depth, and Optical Flow) extract features fused via Coordinate Attention (CA).
-I3D serves as the teacher, and ResNet as the student.
-CLIP is also integrated for zero-shot gesture representation.
-The limitations of standard 3DCNNs were explored in [114], where 3D Central Difference Convolution (3D-CDC) replaces standard 3D convolutions.
-Residual hybrid blocks and a transformer module handle temporal modeling.
-[195] also adopted 3D-CDC in a multi-stream, multi-rate setup with multi-level fusion.
-NAS optimizes the configuration of CDC types, inter-stream connections, and frame rates.
+In (ZS-GR), a dual-stream model combined a ViT for spatial modeling and an LSTM-MLP for temporal modeling, learning visual features aligned with textual embeddings.
+Similarly, (ZS-DHGR) fused skeleton-derived geometric features with RGB-C3D-LSTM-AE representations into a shared semantic space for classification.
+To reduce computational cost, (ConvMixFormer) replaced the transformer’s self-attention mechanism with convolutional layers and a gated feedforward network, resulting in the lightweight ConvMixFormer architecture.
+Knowledge distillation and contrastive learning were applied in (3STREAM_DISTILLATION), where three 3DCNNs (RGB, Depth, Optical Flow) were fused with Coordinate Attention, using I3D as teacher and ResNet as student, while integrating CLIP for zero-shot recognition.
+Finally, (MFST) and (3D_CDC_NAS_TIP_2021) replaced standard convolutions with 3D Central Difference Convolution (3D-CDC), combining hybrid residual blocks, transformers, and NAS-based optimization for enhanced spatiotemporal modeling.
 5.3.2 Sign Language Understanding
-Several recent works have proposed multimodal and pose-centric architectures for Isolated Sign Language Recognition (ISLR), focusing on combining RGB and skeleton data.
-A prominent line of research is the SignBERT framework [68], which uses a spatial GCN with temporal and hand chirality encoding to extract gesture features.
-These are fed into a transformer encoder, while a hand-model-aware decoder reconstructs masked inputs during self-supervised pretraining.
-Predictions from a parallel RGB network are fused at inference time.
-An enhanced version, SignBERT+ [71], introduces spatial-temporal position encoding and an improved fusion module for RGB and pose data.
-The architecture is illustrated in Fig.
-10.
-Similarly, MASA [203] adopts the same GCN+Transformer backbone, trained with contrastive learning on pose reconstruction.
-BEST [202] extends this by embedding pose triplets (left hand, right hand, body) with BERT and applying late fusion with RGB data.
-Figure 10: The pose-based stream of the SignBERT+ [71], an extension of the SignBERT [68], during SSL pretraining.
-A GCN is first utilized to extract the gesture features, which are after combined with spatial-temporal positing encoding and subsequently pass through a transformer encoder.
-During pretraining, a MANO-based decoder is placed after the transformer to perform reconstruction of the masked input.
-The decoder is replaced with a prediction head during downstream fine-tuning.
-To improve feature invariance, a contrastive approach was proposed in [204], enforcing consistency among hand, trunk, and motion features.
-The same GCN+Transformer architecture is used, with late fusion of pose and RGB predictions.
-The SAM-SLR framework [82] combines six streams (RGB, depth, skeleton) with late fusion of predictions.
-SAM-SLR-v2 [81] adds a 3D hand graph stream and learnable fusion weights (Fig.
-11).
-Figure 11: The architecture of SAM-SLR-v2 [81], which is based on the SAM-SLR [82].
-The architecture consists of seven streams, regarding various modalities, including RGB frames, optical flow, pose heatmaps and skeleton joints.
-The predictions of each branch are fused in a learnable late manner.
-In VTN-HC/PF [36], hand-cropped images are processed via a 2D-CNN for spatial modeling, followed by a transformer and MLP for temporal and final predictions.
-The PF variant incorporates geometrical features from pose data.
-Motion History Images (MHIs) were employed in [119], where a 3DCNN with attention and residual connections analyzes MHIs, while a dual-stream version adds a 2DCNN for RGB data, with predictions fused via weighted summation.
-To address data scarcity, ZSSLR [216] adopts a CLIP-based zero-shot framework using RGB and pose data.
-In [24], a few-shot cross-lingual method is proposed, combining two 3DCNNs and a pose estimator.
-Features are merged through a Temporal Feature Aggregation Module with self-attention and 1D convolution.
-NLA-SLR [214] targets semantically similar signs using a multistream backbone (RGB and pose heatmaps via S3D), followed by a head with language-aware label smoothing and an auxiliary inter-modality mix-up classifier.
-MSNN [116] uses five I3D streams (RGB and optical flow) plus ST-GCN for pose.
-hDNN-SLR [142] introduces a hybrid deep network with a 3DCNN for spatial detection, an attention-based BiLSTM for temporal modeling, and a VAE-based module for feature discrimination.
-Finally, Uni-Sign [102] proposes a unified multimodal architecture with three pose encoders and an RGB encoder fused via an attention-based module.
-Features are passed through temporal encoders and a pretrained LLM.
-Training involves pose-only and joint RGB-pose pretraining followed by downstream fine-tuning on ISLR and CSLR tasks.
+Recent works on Isolated Sign Language Recognition (ISLR) have focused on multimodal, pose-centric architectures combining RGB and skeleton data.
+SignBERT (SignBERT_ICCV) employs a spatial GCN with temporal and hand chirality encoding, feeding features into a transformer encoder and a hand-model-aware decoder for self-supervised pretraining, with parallel RGB predictions fused at inference.
+Its enhanced variant, SignBERT+ (SignBERT+), adds spatial-temporal position encoding and an improved RGB–pose fusion module (Fig.
+6).
+Similarly, MASA (MASA) uses the same GCN+Transformer backbone with contrastive learning on pose reconstruction, while BEST (BEST) embeds pose triplets with BERT and applies late fusion with RGB data.
+Figure 6: The pose-based stream of SignBERT+ (SignBERT+), an extension of SignBERT (SignBERT_ICCV), first uses a GCN to extract gesture features, which are combined with spatial-temporal position encoding and processed by a transformer encoder.
+During self-supervised pretraining, a MANO-based decoder reconstructs masked inputs, and is replaced by a prediction head for downstream fine-tuning.
+To improve feature invariance, (Weichao-Zhao) applies a contrastive approach enforcing consistency among hand, trunk, and motion features, using the same GCN+Transformer backbone with late fusion of pose and RGB predictions.
+SAM-SLR (SAM_SLR) fuses six streams (RGB, depth, skeleton, optical flow, pose heatmaps, HHA), while SAM-SLR-v2 (jiang2021sign-sam-slr-v2) adds a 3D hand graph stream and learnable fusion weights.
+VTN-HC/PF (VTN-HC-PF) processes hand-cropped images via a 2DCNN, transformer, and MLP for final predictions, with PF incorporating pose-based geometrical features.
+Motion History Images are used in (MOTION_HISTORY_Mercanoglu), analyzed by a 3DCNN with attention, and in a dual-stream setup combined with 2DCNN RGB features via weighted summation.
+To tackle data scarcity, ZSSLR (ZSSLR) uses a CLIP-based zero-shot framework, while (FSSLR) proposes a few-shot cross-lingual method combining two 3DCNNs and a pose estimator with a Temporal Feature Aggregation Module.
+NLA-SLR (NLA_SLR) targets semantically similar signs with a multistream S3D backbone and language-aware label smoothing.
+MSNN (MSNN) uses five I3D streams plus ST-GCN, and hDNN-SLR (hDNN-SLR) combines a 3DCNN, attention-based BiLSTM, and a VAE-based module.
+Finally, Uni-Sign (UNISIGNli2025) proposes a unified multimodal architecture with three pose encoders and an RGB encoder fused via attention, followed by temporal encoding and LLM-based pretraining for ISLR and CSLR.
 5.3.3 Human-Computer Interaction
-Motivated by the difficulty of recognizing fine-grained gestures in HCI applications, authors in [200] proposed HandSense, a method relying on depth and RGB data.
+Motivated by the difficulty of recognizing fine-grained gestures in HCI applications, authors in (HandSense) proposed HandSense, a method relying on depth and RGB data.
 Two 3DCNNs are deployed to extract features for each one modality.
 These features are fused and fed into a SVM to obtain the final prediction.
-Authors in [87] proposed an approach relying on RGBD Dynamic Images (RDDIs).
+Authors in (mXception_and_RDDI) proposed an approach relying on RGBD Dynamic Images (RDDIs).
 These images, are obtained by the combination of RGB and estimated depth data.
@@
 5.3.4 Other
-Inspired by the success of transformers, GestFormer [51] introduces a lightweight transformer-based framework for gesture-driven HRI.
+Inspired by the success of transformers, GestFormer (GestFormer) introduces a lightweight transformer-based framework for gesture-driven HRI.
 A 2DCNN extracts per-frame features, processed by PoolFormer layers (replacing attention) and two custom blocks for multiscale pooling/WT and token mixing.
 Each modality is processed independently, and outputs are averaged.
-In [49], a hybrid 3DCNN/LSTM model fuses RGB, depth, and skeleton data (from an enhanced pose estimator) via weighted sum.
+In (DHGR_3D_POSE_HRI), a hybrid 3DCNN/LSTM model fuses RGB, depth, and skeleton data (from an enhanced pose estimator) via weighted sum.
 The fused input feeds a spatiotemporal model combining 2D/3D convolutions and ConvLSTM layers.
-5.4 State-of-the-art Methods for Isolated Dynamic HGR
-Recent IDHGR approaches demonstrate high architectural diversity.
-Pose-based methods often rely on spatial GCNs and temporal transformers [68, 71, 202, 204, 203, 165], or on ST-GCN variants [107, 79, 167].
-Due to their limited standalone performance—mainly caused by pose estimation failures in challenging conditions (e.g., motion blur) [203]—pose-based models are frequently fused with simple RGB-based branches (e.g., I3D), using late fusion via score averaging [68, 71, 202, 204, 203, 104].
-Appearance-based methods typically follow a multi-stream 3DCNN design to process individual modalities (e.g., RGB, depth, optical flow) [180, 206, 60, 100, 119, 195], employing late or multi-level fusion.
-Notably, SAM-SLR [82] and SAM-SLRv2 [81] show that adding more modalities—even with simple backbones—can significantly improve accuracy.
-Advanced learning strategies further improve performance.
-Self-supervised learning (SSL) via masked pose reconstruction [202, 68, 71, 79, 203] enhances representation quality by injecting hand topology priors.
-Neural Architecture Search (NAS) has been used to discover optimal network structures [206, 195], while multi-task learning, as in Uni-Sign [102], boosts performance by leveraging auxiliary tasks and larger training corpora.
-Performance comparisons across benchmarks are summarized in Tables 4, 5, and 6.
-Table 4: Brief description and reported accuracy of state-of-the-art methods on IsoGD [179].
-Method Modality Architecture Description Training Setting Accuracy
-3SCKI [180] RGBD, OF Three-stream 3DCNN with mid-level fusion Knowledge distillation between two instances of the pipeline, CLIP 65.87%
-SlowFast+NAS [195] RGBD, OF Multi-stream and multi-rate 3DCNN with 3D-CDC and multi-level fusion NAS 66.23%
-RAAR3DNet [206] RGBD Dual stream I3D with adoptable cells, spatial and temporal attention Multi-task loss and NAS 66.62%
-De+Recouple [207] RGBD Dual-stream spatial + temporal with AE for multi-level fusion and mechanisms to reduce redundancy Self-distillation for inter-frame learning, extra supervision to enhance the AE 66.79%
-CTFB [60] RGBD Dual-stream I3D with multi-level fusion - 69.02%
-[100] RGBD Dual-stream Res3D-18 + average fusion Memory network to enhance the discriminative capability of the extracted features and information-theoretic loss function 74.08%
-Table 5: Performance of uni-modal pose-based methods on SHREC [166] and DHG [37].
-Method Architecture Split SHREC DHG
-[120] Three-stream approach: attention-based spatial-temporal model, attention-based temporal-spatial model and MLP 14 97.01% 92.00%
-28 92.78% 88.78%
-MS-ISTGCN [167] Three-stream model: ST-GCN with adoptive spatial GC, extended temporal GC and attention 14 96.7% 93.7%
-28 94.9% 91.2%
-TD-GCN [107] ST-GCN variant with channel-dependent and temporal-dependent adjacency matrices 14 97.02% 93.9%
-28 95.36% 91.4%
-[101] Two-stream approach consisting of an attention-based dual branch module and MAE 14 96.9% 94.21%
-28 94.17% 92.11%
-SBI-DHGR [129] Multi-stream 1DCNN + BiLSTM 14 97% 94.64%
-28 92.36% 91.79%
-STST [201] Two-stage transformer-based approach 14 97.62% 94.82%
-28 95.83% 93.18%
-DSTSA-GCN [30] Three stacked stages, each one consisting of group channel-wise and temporal-wise GCN, followed by multiscale TCN 14 97.74% 95.04%
-28 95.37% 93.57%
-DRL-DDNet [46] DD-Net for classification and FS-Net for keyframe extraction 14 96.5% -
-28 93.8% -
-Table 6: Performance of the current SOTA methods on ISLR.
-PI, PC, OF and DF stands for per-instance accuracy, per-class accuracy, optical flow and depth flow, respectively.
-Methods marked with star (*) use additional training data.
-Method Modality Architecture Description Training Setting Dataset PI PC
-BEST [202] RGB, Pose d-VAE + GCN + Transformer & RGB-based encoder + late fusion SSL pretraining (masked tokens reconstruction) MSASL 71.21% 68.24%
-WLASL 54.59% 52.12%
-CSL500 97.7% -
-SignBERT [68] RGB, Pose Spatial GCN, temporal and chilarity embedding + Transformer & RGB-based stream + late fusion SSL pretraining (MANO-based masked hand reconstruction) MSASL 71.24% 67.96%
-WLASL 54.69% 52.08%
-CSL500 97.6% -
-SignBERT+ [71] RGB, Pose Extension of SignBERT with Spatial-Temporal Position Encoding and enhanced late fusion technique The same as SignBERT MSASL 73.71% 70.77%
-WLASL 55.59% 53.33%
-CSL500 97.8% -
-MASA [203] RGB, Pose Spatial GCN + Transformer & RGB-based stream SSL pretraining (masked input reconstruction), contrastive learning WLASL 55.77% 53.13%
-TMS-Net [38] Pose Multi-stream decoupled GCN with multi-level fusion - WLASL 56.4% -
-SKIM [104] Pose Multi-stream channel-wise multi-scale GCN + late fusion Hand part mixing for data augmentation WLASL 57.73% 55.40%
-[204] RGB, Pose GCN+Transformer and I3D + late fusion Contrastive learning to enhance invariance, consistency constraints between features of hand and trunk parts and between joints and motions MSASL 74.18% 72.18%
-WLASL 58.06% 55.66%
-CSL500 97.8% -
-SAM-SLR-v2 [81] RGBD, Pose, HHA, OF, DF Extension of SAM-SLR-v2 with 3D graph stream and learnable weights for late fusion - WLASL 59.39% 56.63%
-CSL500 99.00% -
-NLA-SLR [214] RGB, Pose Four-stream network based on S3D with two frame rates and multi-level fusion Language-aware label Smoothing and inter-modality mixup with an auxiliary classifier MSASL 73.80% 70.95%
-WLASL 61.26% 58.31%
-Uni-Sign [102]* RGB, Pose A pretrained LLM process features extracted by pose, visual and temporal encoders Three-stage training, multi-task learning with additional training data MSASL 78.16% 76.97%
-WLASL 63.52% 61.32%
+Authors in (MDSI-fang2025mdsi) emphasize on two major issues, information redundancy and the difficulty to recognize similar gestures.
+Their approach leverages a CLIP encoder to provide qualitative embeddings and enhance the feature extractor through label smoothing as well as a semantic filter to incorporate multimodal data.
+5.4 Performance of IDHGR Methods
+The performance of IDHGR methods in the most used benchmarks (IsoGD (chalearn-lap), SHREC (SHREC), DHG (DHG-CVPRW), NVIDIA Hand Gesture or nvGesture (NVIDIA_GESTURE_R3DCNN_Molchanov), MSASL (MS-ASL), WLASL (WLASL), CSL500 (CSL500) and AUTSL (AUTSL)) is shown Table 3.
+For ISLR, Uni-Sign (UNISIGNli2025) surpasses relevant methods in terms of recognition accuracy, but requires powerful hardware to be executed fast, due to its large size (592.1M parameters, most of them contained on the language model).
+Thus, smaller but less accurate models like NLA-SLR (NLA_SLR), SignBERT+ (SignBERT_ICCV) and (Weichao-Zhao) are more reasonable solutions for real-world applications.
+Even though the multimodal SAM-SLR (SAM_SLR) and SAM-SLR-v2 (jiang2021sign-sam-slr-v2) are very accurate models for ISLR, they require a large number of modalities that are not fused effectively (using late fusion), which may be suboptimal.
+Pose-based methods for HCI and HRI including TD-GCN (TD-GCN), STST (STST) and DSTSA-GCN (DSTSA-CGN_cui2025dstsa) achieve high recognition rates with relatively low computational cost.
+Although, the computational complexity of pose estimation and the effect of its incorrect predictions in more challenging environments (e.g.
+fast movements) may reduce their overall performance.
+Appearance-based methods are not affected by off-the-shelf pose estimators but they comprise many more parameters (up to 251.4M (3D_CDC_NAS_TIP_2021)).
+Table 3: Performance of IDHGR methods on various datasets.
+Notably the accuracy depends on the datasets’ difficulty; for example, methods with accuracy near to 100% on the AUTSL, do not perform the same way on more challenging datasets with thousands of classes (e.g.
+MSASL and WLASL).
+SHREC DHG MSASL1000 WLASL2000
+Method Year Params IsoGD 14 28 14 28 nvGesture PI PC PI PC CSL500 AUTSL
+STGCN-LSTM (STGCN-LSTM-10965662) 2025 - - - - - - - - - - - 95.2% -
+(ISLR-STGCN-attention-LSTM-ozdemir2025cross) 2025 - - - - - - - - - - - - 86.43%
+HAN-2S (HAN-liu2025han) 2025 0.94M - 95% 92.86% 92.71% 89.15% - - - - - - -
+DSTSA-GCN (DSTSA-CGN_cui2025dstsa) 2025 8.08M - 97.74% 95.37% 95.04% 93.57% - - - - - - -
+Uni-Sign (UNISIGNli2025) 2025 592.1M - - - - - - 78.16% 76.97% 63.52% 61.32% - -
+TMS-Net (TMS-NET) 2024 - - - - - - - - - 56.4% - - -
+MSNN (MSNN) 2024 - - - - - - - 50.59% - 47.26% - - -
+SKIM (SKIM) 2024 - - - - - - - - - 57.73% 55.40% - -
+MASA (MASA) 2024 - - - - - - - - - 55.77% 53.13% - -
+DSTA-SLR (DSTA_SLR_hu2024dynamic) 2024 7.4M - - - - - - 65.74% 62.31% 53.68% 51.17% 98.1%5 -
+(Weichao-Zhao) 2024 - - - - - - - 74.18% 72.18% 58.06% 55.66% 97.8% -
+TD-GCN (TD-GCN) 2024 1.36M - 97.02% 95.36% 93.9% 91.4% - - - - - - -
+STST (STST) 2024 - - 97.62% 95.83% 94.82% 93.18% - - - - - - -
+(Tianyu_Liu) 2024 - - - - - - - - - - - 98.87% 92.53%
+SwC GR-MMixer (SwC-GR-MMixer) 2024 - - - - - - - - - - - 98.54% -
+HWGAT (FDMSE-hierarchicalwindowedgraphattention) 2024 - - - - - - - - - 48.49% - - 95.80%
+(Yangke_patcog) 2024 0.21M - 96.9% 94.17% 94.21% 92.11% - - - - - - -
+(omar_ikne) 2024 - - 94.1% 90.0% - - - - - - - - -
+BEST (BEST) 2023 - - - - - - - 71.21% 68.24% 54.59% 52.12% 97.7% -
+SignBERT+ (SignBERT+) 2023 - - - - - - - 73.71% 70.77% 55.59% 53.33% 97.8% -
+NLA-SLR (NLA_SLR) 2023 - - - - - - - 73.80% 70.95% 61.26% 58.31% - -
+3SCKI (3STREAM_DISTILLATION) 2023 - 65.87% - - - - - - - - - - -
+(Yunan_Li) 2023 - 74.08% - - - - - - - - - - -
+CTFB (CTFB_Hampiholi) 2023 27.2M-37.1M 69.02% - - - - 87.76% - - - - - -
+L-TSSI-DenseNet (L-TSSI-DenseNet) 2023 7.2M - - - - - - - - - - - 93.13%
+TS3C-Net (TS3C-Net) 2023 28.7M - - - - - 87.55% - - - - - -
+(Musa_IEEE_ACCESS) 2023 - - 97.01% 92.78% 92.00% 88.78% - - - - - - -
+SBI-DHGR (SBI-DHGR) 2023 28M - 97% 92.36% 94.64% 91.79% - - - - - - -
+STr-GCN (STr-GCN) 2023 - - 93.39% 89.20% - - - - - - - - -
+DRL-DDNet (KEYFRAMES_DRL_SKELETON) 2023 - - 96.5% 93.8% - - - - - - - - -
+De+Recouple (Benjia_Zhou) 2022 - 66.79% - - - - 91.70% - - - - - -
+MS-ISTGCN (JOINTS_MOTION_VECTORS_Improved_ST-GCN) 2022 7.1M - 96.7% 94.9% 93.7% 91.2% - - - - - - -
+SignBERT (SignBERT_ICCV) 2021 - - - - - - - 71.24% 67.96% 54.69% 52.08% 97.6% -
+SlowFast+NAS+CDC (3D_CDC_NAS_TIP_2021) 2021 251.4M 66.23% - - - - - - - - - - -
+RAAR3DNet (RAAR3DNet) 2021 - 66.62% - - - - 88.59% - - - - - -
+SeST (SeST) 2021 - 60.27% - - - - - - - - - - -
+SAM-SLR-v2 (jiang2021sign-sam-slr-v2) 2021 - - - - - - - - - 59.39% 56.63% 99.00% 98.53%
+SAM-SLR (SAM_SLR) 2021 19.2M - - - - - - - - 58.73% 55.93% - 98.53%
+(Hu_Zhou_Li_2021) 2021 - - - - - - - 69.39% 66.54% 51.39% 48.75% 98.3% -
 6 Methods for Continuous Dynamic HGR
-In contrast to isolated dynamic hand gesture recognition (IDHGR), where each sample contains at most one gesture, continuous hand gesture recognition (CHGR) deals with unsegmented gesture sequences.
+In contrast to IDHGR, where each sample contains at most one gesture, Continuous Hand Gesture Recognition (CHGR) deals with unsegmented gesture sequences.
 The goal is to recognize gestures from untrimmed videos and generate the corresponding gesture sequence.
-CHGR is typically treated as a weakly supervised task, as the precise temporal boundaries of gestures are not provided; the model learns to identify them implicitly, often using Connectionist Temporal Classification (CTC) loss [188, 55].
-While CHGR has been explored in domains like autonomous driving [45, 185], most research focuses on Sign Language Understanding (SLU), where it is commonly referred to as Continuous Sign Language Recognition (CSLR) or sign2gloss.
+CHGR is typically treated as a weakly supervised task, as the precise temporal boundaries of gestures are not provided; the model learns to identify them implicitly, often using Connectionist Temporal Classification (CTC) loss (Fangyun_Wei_ICCV; CTC_graves2006connectionist).
+While CHGR has been explored in domains like autonomous driving (FA-STGCN; CTCX_WANG2022123), most research focuses on Sign Language Understanding (SLU), where it is commonly referred to as Continuous Sign Language Recognition (CSLR) or sign2gloss.
 CSLR aims to produce a sequence of glosses [^1] Glosses are atomic lexical units in sign language, analogous to words in spoken language.
@@
 Due to the importance of non-manual features (e.g., facial expressions and upper-body movements), many approaches also incorporate these regions to improve recognition accuracy.
-Some CSLR methods are further extended to support gloss-based sign language translation (SLT) by adding a gloss-to-text module [56, 215, 27].
+Some CSLR methods are further extended to support gloss-based Sign Language Translation (SLT) by adding a gloss-to-text module (MSKA_ELSEVIER; zuo2024onlinecontinuoussignlanguage; NEURIPS2022_6cd3ac24_HEATMAPS).
 In the following, we review recent advances in CHGR, grouped by input modality. We highlight key architectures, learning strategies, and benchmark results.
 6.1 RGB-based
-Most state-of-the-art RGB-based Continuous Hand Gesture Recognition (CHGR) models follow a three-stage pipeline architecture: (1) a spatial model (commonly a 2DCNN or a transformer) to extract frame-wise features; (2) a short-term temporal model (mainly 1DCNN or TCN) for local temporal modeling and gloss-level feature extraction; and (3) a contextual module (e.g., BiLSTM) for long-term sequence modeling and final gesture sequence prediction [73, 122, 62, 74, 197, 70, 75, 72, 76].
-This standard architecture was introduced in [31].
-This decoupled spatial-temporal structure has been favored over unified spatiotemporal models like 3DCNNs due to the latter’s inability to define clear temporal boundaries [2].
-Across the various methods that extend the initial three-stage pipeline, six broad enhancement strategies emerge: 1) Auxiliary Supervision & Multi‐Task Objectives, which enrich the standard three‐stage pipeline with extra losses and segmentation tasks; 2) Specialized Temporal Modeling, replacing or augmenting the 1D‐CNN with learned pooling, multi-scale fusion, or hierarchical boundary detectors; 3) Attention & Fusion Modules, inserting correlation, self‐emphasis, or transformer‐fusion blocks into the visual encoder; 4) Backbone Replacements & Hybrids, swapping 2D-CNNs for vision transformers or hybrid CNN–attention architectures; 5) Incorporating Linguistic Priors, aligning visual features with gloss embeddings via cross-modal decoders or contrastive losses; and 6) Data Scarcity & Multilingual Strategies, leveraging self-supervision, prototype memories, and automatic dictionary construction to overcome limited annotations and support multiple sign languages.
+Most state-of-the-art RGB-based CHGR models follow a three-stage pipeline architecture: (1) a spatial model (commonly a 2DCNN or a transformer) to extract frame-wise features; (2) a short-term temporal model (mainly 1DCNN or TCN) for local temporal modeling and gloss-level feature extraction; and (3) a contextual module (e.g., BiLSTM) for long-term sequence modeling and final gesture sequence prediction (TemporalLiftPooling; VAC-CSLR; SMKD; CorrNet; C2ST; HezhenHu_TMM; SEN-CSLR; HezhenHu-JunfuPu; corrnetplus_hu2024corrnetsignlanguagerecognition).
+This standard architecture was introduced in (first_2dcnn_tcn_bilstm_TMM_2019).
+This decoupled spatial-temporal structure has been favored over unified spatiotemporal models like 3DCNNs due to the latter’s inability to define clear temporal boundaries (ADADOGLOU).
 Many RGB‐based CHGR methods build on the canonical 2DCNN + 1DCNN + BiLSTM architecture by introducing auxiliary losses and multi‐task objectives.
-For example, VAC‐CSLR [122] augments the visual and temporal modules with a Visual Enhancement (VE) loss and a Visual Alignment (VA) distillation loss to accelerate convergence and improve representation quality (see Fig.
-12).
-Similarly, SMKD [62] adds two auxiliary classifiers—one after the 1DCNN and one after the BiLSTM—to perform synchronous gloss segmentation alongside recognition, using a three‐stage training schedule to mitigate gradient “spikes.” C2SLR [213] enforces Spatial Attention Consistency (SAC) by focusing the visual module on pose‐informed regions, and Sentence Embedding Consistency (SEC) to align the visual and sequential representations of the same gloss.
-Figure 12: An overview of the VAC-CSLR [122].
+For example, VAC‐CSLR (VAC-CSLR) augments the visual and temporal modules with a Visual Enhancement (VE) loss and a Visual Alignment (VA) distillation loss to accelerate convergence and improve representation quality (see Fig.
+7).
+Similarly, SMKD (SMKD) adds two auxiliary classifiers—one after the 1DCNN and one after the BiLSTM—to perform synchronous gloss segmentation alongside recognition, using a three‐stage training schedule to mitigate gradient “spikes.” C2SLR (C2SLR) enforces Spatial Attention Consistency (SAC) by focusing the visual module on pose‐informed regions, and Sentence Embedding Consistency (SEC) to align the visual and sequential representations of the same gloss.
+Figure 7: An overview of the VAC-CSLR (VAC-CSLR).
 The feature extractor provides local information, that is processed by the subsequent alignment module to predict the gloss sequence.
@@
 To better capture local temporal patterns, several works replace or enhance the standard 1D-CNN.
-Temporal Lift Pooling (TLP) [73] adopts a learned pooling scheme inspired by the lifting transform, eliminating handcrafted kernels.
-mLTSF‐Net [192] couples content‐aware feature selection with position‐aware convolutions to fuse adjacent‐frame similarities at multiple scales.
-HS‐I3D [190] extends I3D by extracting intermediate embeddings and deploying a U‐Net–style decoder that upsamples temporally, enabling simultaneous sign detection and boundary localization.
+Temporal Lift Pooling (TLP) (TemporalLiftPooling) adopts a learned pooling scheme inspired by the lifting transform, eliminating handcrafted kernels.
+mLTSF‐Net (mLTSF-Net) couples content‐aware feature selection with position‐aware convolutions to fuse adjacent‐frame similarities at multiple scales.
+HS‐I3D (HS-I3D) extends I3D by extracting intermediate embeddings and deploying a U‐Net–style decoder that upsamples temporally, enabling simultaneous sign detection and boundary localization.
 Other methods inject lightweight attention and correlation blocks into the spatial encoder.
-CorrNet [74] adds parallel correlation and identification modules after each 2DCNN stage to model inter‐frame trajectories and highlight informative regions; CorrNet+ [76] further replaces its correlation operator with a more efficient variant and inserts an attention layer.
-SEN‐CSLR [75] introduces self‐emphasizing spatial (SSEN) and temporal (TSEN) modules to reweight feature maps and keyframes with minimal overhead.
-Convolutional Transformer Fusion Blocks (CTFB) [60] fuse intermediate features across 3DCNN branches to enhance cross‐modal integration.
-A growing number of approaches swap the 2D-CNN for transformer‐based encoders or hybrid CNN–attention modules.
-CLIP‐SLA variants [14] leverage a frozen CLIP visual encoder, augmented with LoRA or MLP adapters and an auxiliary classifier for early supervision.
-Swin‐MSTP [15] pairs a Swin transformer with an MS‐TCN that employs TLP and dilated convolutions, followed by a BiLSTM for global context.
-TCNet [111] enhances a CNN backbone with trajectory and correlation modules, while SlowFastSign [4] modifies Slow‐Fast‐101 with bidirectional fusion pathways and auxiliary temporal classifiers to refine feature exchange.
-Most CSLR methods still rely on CNN backbones and thus struggle to model correlations across spatial regions.
-To address this, SignGraph [^3] Be aware of an IDHGR method with the same name.
-[47] treats each frame’s patches as nodes in a dynamic graph and alternates between Local Sign Graph (LSG) and Temporal Sign Graph (TSG) modules to capture intra‐frame and inter‐frame relationships; a multiscale variant, MultiSignGraph, further enhances spatial–temporal context (Fig.
-13).
-Figure 13: An illustration of the SignGraph [47].
-The RGB frames are split into patches.
-Dynamic graphs are utilized to represent the relations between intra-frame and inter-frame patches, that constitute their nodes.
-To exploit gloss co-occurrence and language structure, some models fuse linguistic embeddings with visual features.
-SignBERT‐HKSL [212] combines a (3+2+1)D ResNet with BERT and BiLSTM, pretraining on masked inputs and iteratively fine‐tuning with masked outputs.
-C2ST [197] replaces the CTC decoder with a Conditional Gloss Decoder (CGD) that models gloss dependencies, and aligns visual and gloss embeddings via cross‐modal match loss.
-GPGN [58] introduces a Cross‐Modal Match Loss between BERT‐derived gloss features and PDC‐TFE–extracted visual embeddings, followed by a BiGRU for sequence modeling (see Fig.
-14).
-CVT‐SLR [205] pretrains an asymmetric VAE on a pseudo gloss‐to‐gloss task, then bridges the VAE and visual encoder with a video–gloss adapter and contrastive alignment loss (see Fig.
-15).
-Fully transformer‐based PiSLTRc [191] applies content‐aware neighborhood gathering and position‐informed attention for end‐to‐end CSLR.
-Authors in [72] also introduce a prior-aware cross-modality augmentation that trains a transformer to reconstruct masked gloss sequences and then uses it to generate pseudo video–text pairs by guiding gloss substitutions and insertions.
-Figure 14: Authors in GPGN [58] leveraged the linguistic prior from the pretrained BERT.
-The Cross-Modal Match Loss is introduced, in order to maximize the correlation between the gloss and the visual embeddings.
-Figure 15: The approach proposed by authors in CVT-SLR [205].
-In the first training stage, a VAE is pretrained on the gloss2gloss pseudo-translation task in order to gain textual knowledge.
-In the second stage, an adopter is placed between the VAE and the visual encoder, while the consistency of the visual and temporal encoder is maintained via contrastive learning.
+CorrNet (CorrNet) adds parallel correlation and identification modules after each 2DCNN stage to model inter‐frame trajectories and highlight informative regions; CorrNet+ (corrnetplus_hu2024corrnetsignlanguagerecognition) further replaces its correlation operator with a more efficient variant and inserts an attention layer.
+SEN‐CSLR (SEN-CSLR) introduces self‐emphasizing spatial (SSEN) and temporal (TSEN) modules to reweight feature maps and keyframes with minimal overhead.
+STNet (STNet-wang2025continuous) applied frame-wise loss to the 2DCNN backbone output to enhance feature extraction and inserted multiple Spatial Resonance Modules to facilitate inter-frame modeling.
+Many recent approaches replace 2DCNNs with transformer-based or hybrid CNN/attention modules.
+CLIP-SLA variants (CLIP_SLA_alyami2025clip_CSLR_2025) use a frozen CLIP visual encoder with LoRA or MLP adapters and an auxiliary classifier for early supervision.
+Swin-MSTP (SWIN_MSTP_Neurocomputing_2025_CSLR) combines a Swin transformer with an MS-TCN using TLP and dilated convolutions, followed by a BiLSTM for global context.
+TCNet (TCNet_Lu_Salah_Poppe_2024) augments a CNN backbone with trajectory and correlation modules, while SlowFastSign (SlowFastSign) modifies Slow-Fast-101 with bidirectional fusion and auxiliary temporal classifiers.
+To better model spatial correlations, SignGraph (SignGraph_CVPR) treats frame patches as nodes in a dynamic graph, alternating between Local Sign Graph (LSG) and Temporal Sign Graph (TSG) modules; its multiscale variant, MultiSignGraph, further improves spatiotemporal context.
+CORE (CORE-CSLR-zhang2025core) proposed a different approach aiming to leverage the relations between regions within an input frame, deploying an off-the-shelf pose estimator to detect fine-grained and contextual cues; the corresponding frame region are fed into a 2DCNN and the extracted features are fused using a GNN to obtain the final embedding.
+These per-frame representations pass through a BiLSTM to perform global modeling.
+SD-Net (SD-Net-CSLR-gao2025structure) enhanced temporal modeling by using three parallel branches.
+Each branch consists of a two-stage temporal context module that captures both frame-wise and gloss-wise features.
+Contrastive loss is applied between two of the branches, based on the proposed Entity and Template scheme.
+Authors in (diffusion-CSLR-geng2026no) formulated CSLR as a cross-modal alignment task using a diffusion model to generate the glosses aiming to mitigate the weak supervision, achieving superior performance.
+DSTEN-CSLR (DSTEN-CSLR-yu2025dsten) replaced the BiLSTM with a custom contextual model, consisting of memory and feed-forward blocks.
+Furthermore, the 2DCNN backbone was enhanced by adding a Spatial-Channel Attention (SCA) mechanism, supervised by ground-truth heatmaps, provided by an off-the-shelf pose estimator.
+Some models integrate linguistic embeddings with visual features to leverage gloss co-occurrence and language structure.
+SignBERT-HKSL (SignBERT-HKSL) combines a (3+2+1)D ResNet with BERT and BiLSTM, pretraining on masked inputs and fine-tuning iteratively on masked outputs.
+C2ST (C2ST) replaces the CTC decoder with a Conditional Gloss Decoder (CGD) to model gloss dependencies and aligns visual and gloss embeddings via cross-modal match loss.
+GPGN (GPGN) similarly applies a cross-modal match loss between BERT-derived gloss features and PDC-TFE visual embeddings, followed by a BiGRU for sequence modeling.
+CVT-SLR (CVT-SLR) pretrains an asymmetric VAE on a pseudo gloss-to-gloss task, then connects it to a visual encoder with a video–gloss adapter and contrastive alignment loss.
+Fully transformer-based PiSLTRc (PiSLTRc) uses content-aware neighborhood gathering and position-informed attention for end-to-end CSLR, while (HezhenHu-JunfuPu) introduces prior-aware cross-modality augmentation, training a transformer to reconstruct masked gloss sequences and generate pseudo video–text pairs via guided gloss substitutions and insertions.
 Finally, data augmentation and cross‐lingual strategies help alleviate limited annotations.
 Fangyun Wei et al.
-[188] build a “dictionary” by using a pretrained CHGR model to automatically segment continuous videos into isolated signs, then mapping German and Chinese sign corpora into a shared embedding space.
-CSGC [143] maintains a Gloss Prototype memory that stores class prototypes updated during training; a gloss contrastive loss then enforces cross‐sentence consistency.
+(Fangyun_Wei_ICCV) build a “dictionary” by using a pretrained CHGR model to automatically segment continuous videos into isolated signs, then mapping German and Chinese sign corpora into a shared embedding space.
+CSGC (CSGC_Rao_Sun_Wang_Wang_Zhang_2024) maintains a Gloss Prototype memory that stores class prototypes updated during training; a gloss contrastive loss then enforces cross‐sentence consistency.
 6.2 Pose-based
@@
 Several works exploit structured keypoint information by dividing the body into semantic regions such as the face, hands, and upper body.
-MSKA-SLR [56] processes each region independently via attention-enhanced temporal convolutions across four branches.
+MSKA-SLR (MSKA_ELSEVIER) processes each region independently via attention-enhanced temporal convolutions across four branches.
 The resulting features are fused to decode gloss sequences using a CTC loss, complemented by frame-level self-distillation.
-Similarly, CoSign [83] employs group-specific GCNs for each subset and introduces regularization to preserve inter-stream consistency.
+Similarly, CoSign (CoSign) employs group-specific GCNs for each subset and introduces regularization to preserve inter-stream consistency.
 Their dual-stream variant integrates joint positions and motion features using multi-level fusion to capture both static and dynamic cues.
 Temporal localization of gestures has also been explored in the literature.
-TY-Net [39] draws inspiration from YOLO to detect gesture instances in time by predicting their boundaries and centers, analogous to spatial bounding boxes.
+TY-Net (TY-Net) draws inspiration from YOLO to detect gesture instances in time by predicting their boundaries and centers, analogous to spatial bounding boxes.
 Beyond joint coordinates, pose heatmaps serve as an alternative input modality.
-The multi-cue framework in [210] combines spatial and temporal reasoning by localizing facial and hand regions via a pose estimator and subsequently encoding them with a temporal module, before decoding gloss sequences with a CTC head.
-The FA-STGCN model [45] further improves robustness to motion scale and frequency variability.
+The multi-cue framework in (MULTI-CUE) combines spatial and temporal reasoning by localizing facial and hand regions via a pose estimator and subsequently encoding them with a temporal module, before decoding gloss sequences with a CTC head.
+The FA-STGCN model (FA-STGCN) further improves robustness to motion scale and frequency variability.
 It normalizes skeleton inputs and applies alternating spatial and temporal GCNs.
@@
 Finally, in the context of autonomous driving, Wang et al.
-[185] developed a two-stage pose-based system to classify road intersection gestures.
+(CTCX_WANG2022123) developed a two-stage pose-based system to classify road intersection gestures.
 The first stage extracts geometric features and encodes them via LSTM layers.
@@
 6.3 Multimodal
-Multimodal approaches for CHGR combine multiple input types—such as RGB frames, pose heatmaps, depth maps, and optical flow—to mitigate the limitations of unimodal methods and capture complementary information from different data sources.
-A prominent example is TwoStream-SLR [27], a dual-stream architecture designed for both signed language translation (SLT) and CSLR.
-It processes RGB and pose heatmaps in parallel using the first four blocks of S3D in each branch, with bidirectional lateral connections enabling multi-level information exchange.
-Additional supervision is provided through Sign Pyramid Networks and auxiliary prediction heads (video, keypoint, and joint).
-The model is trained using self-distillation, minimizing KL-divergence between the outputs of the three heads.
-An illustration of the architecture is shown in Fig.
-16.
-This model has also been applied to ISLR [215], using a sliding window mechanism and leveraging the same framework to extract sign boundaries and construct training dictionaries.
-Figure 16: Overview of the TwoStream-SLR [27].
-Eight S3D blocks constitute the dual-stream multimodal architecture.
-Bidirectional Lateral Connection is utilized to exchange the information produced by each branch in a multi-level manner.
-The Sign Pyramid Networks, Video and Keypoint head are actins as auxiliary classifiers, providing extra supervision.
-Self-distillation is also employed to share the knowledge between the two streams.
+Multimodal CHGR approaches combine inputs such as RGB, pose heatmaps, depth, etc., to overcome unimodal limitations and capture complementary information.
+A key example is TwoStream-SLR (NEURIPS2022_6cd3ac24_HEATMAPS), a dual-stream architecture for CSLR and SLT.
+It processes RGB and pose heatmaps in parallel via the first four S3D blocks in each branch, with bidirectional lateral connections for multi-level information exchange.
+Sign Pyramid Networks and auxiliary heads (video, keypoint, joint) provide additional supervision, while self-distillation minimizes KL-divergence across outputs.
+Fig.
+8 illustrates the architecture.
+The same framework has been applied to ISLR (zuo2024onlinecontinuoussignlanguage) using a sliding window to extract sign boundaries and construct training dictionaries.
+Figure 8: TwoStream-SLR (NEURIPS2022_6cd3ac24_HEATMAPS) is a dual-stream multimodal architecture built from eight S3D blocks.
+Bidirectional Lateral Connections enable multi-level information exchange between branches.
+Auxiliary classifiers, including the Sign Pyramid Network, Video Head, and Keypoint Head, provide additional supervision, while self-distillation transfers knowledge between the two streams.
 Other CSLR models have adapted multimodal ISLR backbones.
-Uni-Sign [102] can be used directly for CSLR, while SignBERT+ [71] can be extended by appending temporal pooling and a CTC decoder to generate gloss sequences.
+Uni-Sign (UNISIGNli2025) can be used directly for CSLR, while SignBERT+ (SignBERT+) can be extended by appending temporal pooling and a CTC decoder to generate gloss sequences.
 Beyond sign language, general-purpose CHGR systems have explored multimodality across broader visual sensor inputs.
-For example, [112] proposes a framework that segments RGB, depth, infrared, and optical flow inputs using a shared U-Net and processes them via separate 3D DenseNet-43 backbones.
+For example, (pretrained_on_jester_3dDenseNet_LSTM) proposes a framework that segments RGB, depth, infrared, and optical flow inputs using a shared U-Net and processes them via separate 3D DenseNet-43 backbones.
 Extracted features are fused through Canonical Correlation Analysis to align shared representations across modalities.
-6.4 State-of-the-art Methods for CSLR
-This subsection reviews state-of-the-art (SOTA) methods for Continuous Sign Language Recognition (CSLR), focusing on their performance on the three most widely used benchmarks: PHOENIX14 [44], PHOENIX14T [26], and CSL-Daily [209].
-As shown in Table 7, the dominant architecture remains the RGB-based Spatial Perception Model+1DCNN+BiLSTM pipeline, where the spatial perception model can be either a 2DCNN or a transformer-based approach.
-Typically, a ResNet18 backbone extracts spatial features, a 1D CNN or TCN models short-term temporal relations, and a BiLSTM captures long-range context for alignment.
-VAC-CSLR [122] established a widely adopted training setting to address optimization instability.
-Many works have since extended this baseline with auxiliary classifiers and additional loss terms to enhance feature learning [83, 4, 14].
-Furthermore, the TwoStream-SLR [27] seems to compete worthily the aforementioned pipeline, surpassing it when extra training data are utilized [188] and ranked second in the CSL-Daily [209] benchmark without additional samples.
-Recent innovations that improve CSLR include:
-  • Data Augmentation Across Datasets: Leveraging visually similar signs from different corpora as training data [188].
-  • Modeling Gloss Dependencies: Addressing the independence assumption in vanilla Connectionist Temporal Classification (CTC) loss through conditional modeling [197].
-  • Multimodal Inputs: Combining RGB and pose heatmaps to improve robustness [27].
-  • Graph-Based Feature Interaction: Using GNNs instead of CNNs to model intra- and inter-frame patch relationships [47].
-  • Linguistic Priors: Integrating gloss embeddings or language models to guide learning [58, 57].
-  • Pre-trained Vision Backbones: Fine-tuning foundation models like CLIP for richer feature extraction [14].
-Table 7: Brief description of current state-of-the-art CSLR methods and their performance on the three most used benchmark (Phoenix 2014 [44], Phoenix 2014 T [26] and CSL-Daily [209]).
-Methods marked with star (*) have utilized additional training data.
-Method Modality Architecture Description Training Setting Dataset WER
-SLA-Adapter/LoRA [14] RGB CLIP visual encoder with adopter/LoRA, followed by TCN and BiLSTM VAC-CSLR [122] and early supervision, by adding an extra classifier after the visual encoder.
-CSL-Daily 25.8%
-Phoenix 18.8%/19.3%
-Phoenix-T 19.5%/19.4%
-Uni-Sign [102]* RGB, Pose A pretrained LLM process features extracted by pose, visual and temporal encoders. Three-stage training, multi-task learning with additional training data CSL-Daily 26%
-Swin-MSTP [15] RGB SwinT for spatial feature extraction and a mixture of 1D convolution, convolution with different dilation rates and TLP [73]. VAC-CSLR [122] CSL-Daily 27.1%
-Phoenix 18.7%
-Phoenix-T 19.7%
-SlowFastSign [4] RGB Modified SlowFast with bidirectional feature fusion between the two streams. VAC-CSLR [122] and auxiliary classifiers to provide extra supervision. CSL-Daily 24.9%
-Phoenix 18.3%
-Phoenix-T 18.7%
-MultiSignGraph [47] RGB Multiscale spatial-temporal GNN that considers image patches as the nodes. [47]. VAC-CSLR [122] CSL-Daily 26.4%
-Phoenix 19.1%
-Phoenix-T 19.1%
-CSGC [143] RGB 2DCNN+1DCNN+BiLSTM VAC-CSLR [122], auxiliary losses to maintain inter- and cross-sentence consistency and gloss-prototype. CSL-Daily 26.7%
-Phoenix 19%
-Phoenix-T 19.5%
-CorrNet+ [76] RGB Modification of the CorrNet [74], obtained by adding temporal attention after the correlation and identification modules. VAC-CSLR [122] and TLP [73]. CSL-Daily 28.2%
-Phoenix 18.2%
-Phoenix-T 19.1%
-TCNet [111] RGB Combination of the TLP [73] and the CorrNet [74] VAC-CSLR [122] and an additional regularization term CSL-Daily 29.3%
-Phoenix 18.9%
-Phoenix-T 19.4%
-C2ST [197] RGB SwinT+1DCNN+BiLSTM CGD loss, distillation to maintain the consistency between video- and local-level features and linguistic prior CSL-Daily 25.8%
-Phoenix 17.7%
-Phoenix-T 18.9%
-CoSign-2s [83] Pose Dual stream approach based on ST-GCN, considering both joints and motions.
-Each stream consists of group-specific branches.
-Extra supervision via auxiliary classifier and complementary regularization to enhance the robustness to noise.
-CSL-Daily 27.2%
-Phoenix 20.1%
-Phoenix-T 20.1%
-CorrNet [74] RGB 2DCNN equipped with the proposed identification and correlation modules, followed by 1DCNN+BiLSTM. VAC-CSLR [122] CSL-Daily 30.1%
-Phoenix 19.4%
-Phoenix-T 20.5%
-[188]* RGB, Pose TwoStream-SLR [27] Utilization of extra training data via dictionary construction. Phoenix 16.7%
-Phoenix-T 18.5%
-TwoStream-SLR [27] RGB, Pose Dual stream (containing S3D blocks) with bidirectional lateral connections.
-Extra supervision achieved by auxiliary CTC losses and self-distillation applied to additional prediction heads.
-CSL-Daily 25.3%
-Phoenix 18.8%
-Phoenix-T 19.3%
+6.4 Performance of CSLR Methods
+Table 4 reports the computational cost and Word Error Rate (WER) of CSLR methods on three benchmark datasets.
+Leveraging large language models, as in (UNISIGNli2025) and (diffusion-CSLR-geng2026no), yields substantial performance gains without dramatically increasing parameters.
+SlowFastSign (SlowFastSign) achieves strong accuracy with a moderate number of parameters, while TwoStream-SLR (NEURIPS2022_6cd3ac24_HEATMAPS) has slightly higher WER, requires twice as many parameters, and involves a more complex training setup, though WER improves with additional training data (Fangyun_Wei_ICCV).
+C2ST (C2ST) is accurate but contains 187.8M parameters, which may limit real-world deployment.
+For resource-constrained scenarios, the pose-based CoSign-2s (CoSign) requires only 30.1 GFLOPs, though pose estimation cost is not included.
+DSTEN-CSLR (DSTEN-CSLR-yu2025dsten) is a lightweight RGB-based model but less accurate.
+Overall, a trade-off between accuracy and computational cost persists.
+Table 4: Comparison of CSLR methods in terms of computational cost and accuracy (Word Error Rate, WER) on three benchmarks: CSL-Daily (Back-Translation-CVPR), Phoenix (PHOENIX-2014), and Phoenix T (NEURAL-SIGN-LANGUAGE-TRANSLATION-CVPR-CAMGOZ).
+”Params” denotes the number of trainable parameters.
+Method Year GFLOPs Params CSL-Daily Phoenix Phoenix T
+(diffusion-CSLR-geng2026no) 2025 - 86.64M 23.7% 16.8% 17.2%
+DSTEN-CSLR (DSTEN-CSLR-yu2025dsten) 2025 366.3 30.19M 28.1% 20.0% 20.3%
+SLA-Adapter (CLIP_SLA_alyami2025clip_CSLR_2025) 2025 - 26.2M 25.8% 18.8% 19.5%
+SLA-LoRA (CLIP_SLA_alyami2025clip_CSLR_2025) 2025 - 30.8M 25.8% 19.3% 19.4%
+Uni-Sign (UNISIGNli2025) 2025 - - 26% - -
+Swin-MSTP (SWIN_MSTP_Neurocomputing_2025_CSLR) 2025 - - 27.1% 18.7% 19.7%
+CORE (CORE-CSLR-zhang2025core) 2025 - - 31.1% 21.7% 21.6%
+SD-Net (SD-Net-CSLR-gao2025structure) 2025 - - 27.9% - 20.3%
+STNet (STNet-wang2025continuous) 2025 - - 27.2% 19.3% 19.8%
+SlowFastSign (SlowFastSign) 2024 - 52.5M 24.9% 18.3% 18.7%
+MultiSignGraph (SignGraph_CVPR) 2024 - - 26.4% 19.1% 19.1%
+CSGC (CSGC_Rao_Sun_Wang_Wang_Zhang_2024) 2024 - - 26.7% 19% 19.5%
+CorrNet+ (corrnetplus_hu2024corrnetsignlanguagerecognition) 2024 - - 28.2% 18.2% 19.1%
+TCNet (TCNet_Lu_Salah_Poppe_2024) 2024 932.1 19.3M 29.3% 18.9% 19.4%
+C2SLR (C2SLR) 2024 1124.6 32.6M - 20.4% 20.4%
+GPGN (GPGN) 2024 - - 30% 20.4% 20.5%
+C2ST (C2ST) 2023 1635 187.8M 25.8% 17.7% 18.9%
+CoSign-2s (CoSign) 2023 30.1 28.2M 27.2% 20.1% 20.1%
+CTCA (CTCA) 2023 - - 29.4% 20.1% 20.3%
+CorrNet (CorrNet) 2023 1035.4 20.5M 30.1% 19.4% 20.5%
+SEN-CSLR (SEN-CSLR) 2023 1144.2 23.1M 30.7% 21% 20.7%
+(Fangyun_Wei_ICCV) 2023 - 105.2M - 16.7% 18.5%
+CVT-SLR (CVT-SLR) 2023 - - - 20.1% 20.3%
+TwoStream-SLR (NEURIPS2022_6cd3ac24_HEATMAPS) 2022 - 105.2M 25.3% 18.8% 19.3%
+STMC (MULTI-CUE) 2022 1742.1 40.71M - 20.7% 21%
+TLP (TemporalLiftPooling) 2022 2950.5 48M - 20.8% 21.2%
+SMKD (SMKD) 2021 1032.3 20.3M - 21% 22.4%
+VAC-CSLR (VAC-CSLR) 2021 1120.4 22.3M - 22.3% -
+7 Summary of SoTA Methods
+After reviewing numerous recent works, this section focuses on state-of-the-art (SoTA) methods across tasks, summarizing the field and comparing key methods in terms of architecture, deployment, strengths, and limitations.
+Table 5 highlights each method’s task, application domain, input modality, architecture, and major strengths and weaknesses.
+One insight from the table is that SLU dominates other fields, likely due to its complexity, which requires advanced visual cues (including facial expressions), extensive temporal modeling for dynamic gestures, and classifiers capable of distinguishing highly similar gestures.
+Consequently, few SoTA methods exist for static SLR, as they struggle to capture hand movements.
+General-purpose IDHGR methods, used as baselines for tasks like SLU, HRI, and HCI, have also been widely studied; the similarity of gestures in HRI and HCI allows the same methods to serve both domains.
+Depth, pose, and other non-RGB modalities are used less frequently and typically in combination with RGB streams.
+This is because they either cannot represent fine-grained gestures (depth, optical flow) or perform poorly in dynamic recognition (pose estimators often fail on blurry images).
+RGB-based models—even older or simpler ones—already achieve strong results on SHGR datasets (HAGRIDv2_ARXIV), so adding pose can incur significant computational overhead without substantial accuracy gains.
+In CSLR, rapid hand movements degrade pose estimator performance, making pose usable only alongside RGB (NEURIPS2022_6cd3ac24_HEATMAPS) and increasing complexity.
+Thus, pose-based approaches are mostly limited to IDHGR, with some exceptions (CoSign), where incorporating multiple modalities can notably improve performance (SAM_SLR; jiang2021sign-sam-slr-v2), even with late fusion.
+Regarding architecture, SHGR’s simplicity allows even small 2DCNNs (EDenseNet; G-CNN-SHARMA2021115657) to achieve very high accuracy ($\geq$95%).
+However, the limited size and diversity of datasets like (NUS-I-2010; NUS-II-2013; ouhands; triesch2001system) can lead to overfitting, reducing real-world performance.
+In contrast, IDHGR and especially ISLR rely on more complex models.
+ST-GCN variants (JOINTS_MOTION_VECTORS_Improved_ST-GCN; TD-GCN) and GCN+Transformer pipelines (SignBERT_ICCV; SignBERT+; BEST; Weichao-Zhao) mainly process skeleton data, often combined with RGB 3DCNNs (e.g., I3D, Res3D) to compensate for pose estimation errors.
+While many studies focus on improving pose-based streams, RGB modalities are less explored.
+SAM-SLR (SAM_SLR) and SAM-SLRv2 (jiang2021sign-sam-slr-v2) propose advanced separate processing for each modality, representing current SoTA for IDHGR.
+RGBD models (RAAR3DNet; CTFB_Hampiholi; Yunan_Li) typically use dual-stream 3DCNNs, increasing computational cost.
+CSLR demands even more complex architectures to capture short-term (gloss-wise) and long-term (global) context.
+The 2DCNN+1DCNN+BiLSTM pipeline remains common, with relatively fewer parameters than SHGR models (e.g., Gesture-CNN (G-CNN-SHARMA2021115657)  67M vs.
+large CSLR models  48M (TemporalLiftPooling)).
+Vision transformers increasingly replace 2DCNN backbones (SWIN_MSTP_Neurocomputing_2025_CSLR; C2ST; CLIP_SLA_alyami2025clip_CSLR_2025) for their attention-based focus on critical regions.
+TwoStream-SLR (NEURIPS2022_6cd3ac24_HEATMAPS) employs a dual-stream 3DCNN with multimodal fusion and pose heatmaps (instead of skeleton joints) alongside RGB frames, achieving high accuracy.
+As mentioned earlier, SHGR is a relatively easy task, so researchers have not proposed any notable learning strategy.
+Pose-based methods for ISLR often (SignBERT_ICCV; SignBERT+; BEST; Weichao-Zhao) insert hand topology prior through SSL.
+NAS is also utilized to determine the optimal structure of 3DCNNs.
+Models for CSLR face serious optimization difficulties compared to other methods as a consequence of their complexity.
+Thus, extra supervision is provided via additional losses attached to various layers of the models (VAC-CSLR; CLIP_SLA_alyami2025clip_CSLR_2025).
+Leveraging additional training data combined with accurate architectures during training (Fangyun_Wei_ICCV) increases the performance, but makes a straightforward comparison with other methods infeasible.
+In summary, LHGR-Net (LHGR-Net) and stand-alone pose-based methods (Yangke_patcog; TD-GCN; CoSign) are suitable for resource-constrained use cases (assuming that the cost of the pose estimation is negligible, which is true especially when the keypoints are extracted by the deployed camera), balancing the computational complexity with the recognition accuracy.
+TwoStream-SLR (NEURIPS2022_6cd3ac24_HEATMAPS), C2ST (C2ST) and (diffusion-CSLR-geng2026no) achieve superior performance for CSLR but come with complex training settings and large computational cost.
+NLA-SLR (NLA_SLR) and SAM-SLRv2 (jiang2021sign-sam-slr-v2) surpass most of the related methods, but they require substantial computational resources.
+Uni-Sign (UNISIGNli2025) introduced a novelty, by combining various SLU tasks into a single one.
+Although it achieves significant performance and unifies a variety of tasks providing a shared framework, the computational complexity makes difficult the real-world deployment and the requirement of training data with detailed annotations prevents the adoption to other sign languages.
+Table 5: Comparison of current SoTA methods.
+They are compared in terms of task, application domain (referred to as ”Domain” for the sake of brevity), input modality (or modalities) and general architecture.
+The strengths of each method or the key innovations introduced as well as some notable limitations of them, are also mentioned.
+Method Task Domain Modality Architecture Strengths or Key Innovation Limitations
+FGDSNet (FGDSNet) SHGR HRI RGB Dual-stream hand segmentation model + dual-stream 2DCNN + mid-level fusion The dual-stream classifier addresses any failures of the task-specific segmentation model.
+The segmentation model and the 2DCNN backbones must be trained separately.
+LHGR-Net (LHGR-Net) SHGR HCI RGB Atrous convolutional layers followed by depth separable convolutions and attention. Suitable for resource-constrained deployment. Very small capacity.
+2DPSTPP-Net (2DPSTPP) SHGR HCI RGB 2DCNN with custom pooling layers and residual connections Spatial pyramid pooling and residual connections obtain robust representations.
+The model is relatively large for SHGR.
+Gesture-CNN (G-CNN-SHARMA2021115657) SHGR SLU RGB 2DCNN consisting of four convolutional layers The simplicity of the architecture enables easy implementation and training.
+Relatively large (67M parameters).
+SAM-SLR (SAM_SLR), SAM-SLR-v2 (jiang2021sign-sam-slr-v2) IDHGR SLU RGBD, Pose, HHA, OF, DF Six/seven streams followed by late fusion Incorporated a large number of modalities, achieving very high accuracy.
+Insufficient fusion and computationally expensive.
+SignBERT (SignBERT_ICCV), SignBERT+ (SignBERT+) (ISLR) IDHGR SLU RGB, Pose Spatial GCN + Transformer & 3DCNN + late fusion The model gains hand topology prior through SSL pretraining (masked hand tokens reconstruction, inspired by BERT) Needs extensive SSL pretraining, the processing and incorporation of the RGB modality may be suboptimal
+SignBERT+ (SignBERT+) (CSLR) CDHGR SLU RGB, Pose Spatial GCN + Transformer & 3DCNN + intermediate fusion + BiLSTM
+BEST (BEST) IDHGR SLU RGB, Pose dVAE + Spatial GCN + Transformer & 3DCNN + late fusion
+NLA-SLR (NLA_SLR) IDHGR SLU RGB, Pose Four-stream network based on S3D with two frame rates and multi-level fusion Novel architecture performing multimodality fusion efficiently.
+Language-aware label smoothing facilitates training.
+Robust to visually similar signs recognition.
+Is affected by the performance of word encoders
+Uni-Sign (UNISIGNli2025) IDHGR, CDHGR SLU RGB, Pose A pretrained LLM process features extracted by pose, visual and temporal encoders Provides a unified framework for multiple SLU tasks, increasing the recognition accuracy.
+Requires additional training data
+TD-GCN (TD-GCN) IDHGR - Pose ST-GCN variant Facilitates temporal modeling through non-fixed time-dependent topology Depends on the pose estimation performance.
+SBI-DHGR (SBI-DHGR) IDHGR - Pose Multi-stream 1DCNN + BiLSTM followed by late fusion Utilizes features of various levels of granularity. Very large compared to other pose-based methods
+DSTSA-GCN (DSTSA-CGN_cui2025dstsa) (4s) IDHGR - Pose Four stream GCN + TCN Group channel-wise and temporal-wise GCNs with dynamic graph topologies improve the performance.
+The static part of the hand topology is initialized randomly, lacking prior.
+SlowFast+NAS (3D_CDC_NAS_TIP_2021) IDHGR - RGBD, OF Multi-stream 3DCNN with 3D-CDC and multi-level fusion Enhanced temporal modeling through central difference convolution and optimal hyperparameter selection through NAS.
+Needs substantials computational resources (251.4M params, 232.1GFLOPs) compared to similar methods
+RAAR3DNet (RAAR3DNet) IDHGR - RGBD Dual stream I3D with adoptable cells, spatial and temporal attention NAS improves the architecture.
+The spatial attention mechanism is supervised by an off-the-shelf pose estimator.
+It relies on the correctness of the pose estimator
+CTFB (CTFB_Hampiholi) IDHGR - RGBD Dual-stream 3DCNN + Multi-level fusion with transformer-based modules Performs efficient multimodality fusion via multiple attention modules placed in different layers.
+The scaling to more than two modalities is not straightforward
+(diffusion-CSLR-geng2026no) CHGR SLU RGB Visual encoder followed by a Diffusion Model (DM). Using a DM for gloss generation makes it very accurate. Increased complexity.
+CLIP-SLA (CLIP_SLA_alyami2025clip_CSLR_2025) CHGR SLU RGB CLIP visual encoder (with MLP adopters or LoRA) + TCN + BiLSTM Early supervision by adding an extra classifier accelerates training, while the utilization of a foundation model enhances the feature extraction.
+The utilization of CLIP implies high computational cost.
+Swin-MSTP (SWIN_MSTP_Neurocomputing_2025_CSLR) CHGR SLU RGB SwinT+1DCNN/TLP Deploying a multiscale TCN, the model is able to recognize continuous signs of varying time durations.
+Computationally expensive.
+C2ST (C2ST) CHGR SLU RGB SwinT+1DCNN+BiLSTM CGD loss (that doesn’t make any independence assumption, compared to CTC) and self-distillation increase the performance.
+The vision transformer as well as the language model utilized during training increase the computational complexity.
+CoSign-2s (CoSign) CHGR SLU Pose ST-GCN variant The computational cost is negligible, thus CoSign-2s perfectly balances efficient execution and accurate recognition.
+The performance depends on the attributes of the pose estimator.
+(Fangyun_Wei_ICCV) CHGR SLU RGB, Pose TwoStream-SLR Utilization of extra training data via dictionary construction. Needs additional training data.
+TwoStream-SLR (NEURIPS2022_6cd3ac24_HEATMAPS) CHGR SLU RGB, Pose Dual-stream 3DCNN with bidirectional lateral connections Extra supervision by auxiliary losses and self-distillation applied to additional prediction heads facilitate training.
+The multi-level fusion incorporates efficiently the two modalities.
+Increased complexity due to the usage of two streams and an off-the-shelf pose estimator.
+Very large.
 8 Datasets
-The performance and generalizability of hand gesture and sign language recognition (HGR/SLR) models are heavily influenced by the quality, diversity, and scale of the datasets used during training and evaluation.
+The performance and generalizability of gesture recognition models are heavily influenced by the quality, diversity, and scale of the datasets used during training and evaluation.
 Over the years, a wide variety of datasets have been developed to support different HGR tasks, ranging from static and isolated gestures to continuous sign language recognition in real-world settings.
 These datasets vary widely in terms of the number of gestures or glosses, the types of modalities captured (e.g., RGB, depth, pose, IR), the presence of signer diversity, and the complexity of the sequences.
-In the following subsections, we categorize and review the most prominent datasets according to the specific recognition task they target: (i) Isolated and Continuous Sign Language Recognition, (ii) Semaphoric Gestures, and (iii) Other Gesture Recognition Tasks.
+In the following subsections, we categorize and review the most prominent datasets according to the specific recognition task they target: (i) Isolated and Continuous Sign Language Recognition (IDSLR and CSLR, respectively), (ii) Semaphoric Gestures (both static and dynamic), and (iii) Other Gesture Recognition tasks.
 This organization aims to provide a clear overview of the resources available for each task type and facilitate their appropriate use in future research.
 8.1 Isolated and Continuous SLR
-The datasets that have been developed to support video-based Sign Language Recognition (SLR) can be broadly categorized into those designed for Isolated Dynamic SLR (IDSLR) and Continuous SLR (CSLR).
+The datasets that have been developed to support video-based Sign Language Recognition (SLR) can be broadly categorized into those designed for IDSLR and CSLR.
 The former datasets typically focus on individual signs, recorded in controlled lab environments with a fixed vocabulary, making them well-suited for classification-based models.
 In contrast, continuous datasets present real-world sequences of signs, often extracted from broadcast material or spontaneous signing, requiring temporal segmentation and sequence modeling approaches.
-Additionally, some datasets support Sign Language Translation (SLT) by providing sentence-level gloss annotations and parallel spoken language transcripts.
-The RWTH-PHOENIX-Weather 2014 [44] and its extended version PHOENIX-2014-T1 [26] are benchmark datasets offering sentence-level annotations in German Sign Language (DGS) captured from television weather forecasts.
-More recent efforts such as BOBSL [12] and OpenASL [160] have significantly expanded the signer base and vocabulary, with hours of British and American Sign Language footage collected from TV and web sources, respectively.
-Lab-recorded corpora like KETI [90], CSL-Daily [209], and How2Sign [41] incorporate multimodal inputs (e.g., pose, depth) and structured glosses, aiding both CSLR and SLT tasks.
-These datasets collectively enable the development and benchmarking of models under varying constraints of signer variability, language, modality, and task complexity.
-On the other hand, isolated sign datasets such as MS-ASL [85], WLASL [99], and BSL-1K [11] provide extensive vocabularies (up to 2,000 signs) and diverse signers, facilitating the study of vocabulary scalability and signer generalization.
-Many of these datasets, such as AUTSL [164] and BosphorusSign22k [217], also include pose or depth modalities, enriching multimodal learning.
-While isolated datasets are primarily used for gesture classification, continuous datasets address the more complex problem of temporal segmentation, gloss prediction, and SLT.
-A few hybrid datasets like GSL SD/SI [2] and Corpus FinSL [154] provide data that can be used in both isolated and continuous settings.
-Together, these datasets serve a broad spectrum of research needs, from foundational gesture recognition to high-level translation across multiple sign languages.
-Table 8: Overview of the characteristics of each dataset related to Sign Language Understanding (SLU) and Hand Gesture Recognition (HGR).
-The sign vocabulary (Sign Voc.) corresponds to the number of distinct classes included in each dataset.
-Name Task Year Modality Samples Signers Sign Voc. Language Duration Source
-PHOENIX-2014 [44] CSLR/SLT 2014 RGB - 9 1558 DGS 10.73h TV
-PHOENIX-2014-T1 [26] CSLR/SLT 2018 RGB 8257 9 1066 DGS - TV
-KETI [90] CSLR/SLT 2019 RGB, Pose 14672 14 524 KSL 27.99h Lab
-GSL SD/SI [2] CSLR/SLT 2019 RGBD 10295 7 310 GSL 9.59h Lab
-Corpus FinSL [154] CSLR/SLT 2020 RGB 343 21 - FinSL 14.37h -
-CSL-Daily [209] CSLR/SLT 2021 RGB 20654 10 2000 CSL 23.27h Lab
-How2Sign (green studio) [41] CSLR/SLT 2021 RGBD, Pose 2529 11 - ASL 79.12h Lab
-BOBSL [12] CSLR/SLT 2021 RGB 1962 39 2281 BSL 1467h TV
-OpenASL [160] CSLR/SLT 2022 RGB 98417 220 - ASL 288h Web
-Auslan-Daily [158] CSLR/SLT 2023 RGB 25K 67 - AUSLAN - TV, Web
-TVB-HKSL-News [130] CSLR/SLT 2024 RGB, Pose 7K 2 6515 HKSL 16.07h TV
-QSL [48] CSLR/SLT 2024 RGB 30040 5 654 CSL - Lab
-ChicagoFSWild+ [159] CSLR 2019 RGB 55232 260 26 ASL - Web
-HKSL [212] CSLR 2021 RGBD 2400 6 55 HKSL - Lab
-CSL [78] IDSLR 2018 RGBD, Pose 125K 50 500 CSL - Lab
-SMILE [42] IDSLR 2018 RGB - 30 100 DSGS - Lab
-ASL Dataset [16] IDSLR 2018 Pose 1200 20 30 ASL - Lab
-MSASL [85] IDSLR 2019 RGB 25513 222 1000 ASL 24.65h Web
-WLASL [99] IDSLR 2019 RGB 21083 119 2000 ASL 14.11h Web
-GSL isol. [2] IDSLR 2019 RGBD 40785 7 310 GSL 6.44h Lab
-BosphorusSign22k [217] IDSLR 2020 RGBD, Pose 22542 6 744 TSL 19h Lab
-INCLUDE [169] IDSLR 2020 RGB 4287 7 263 ISL 3h Lab
-AUTSL [164] IDSLR 2020 RGBD, Pose 38336 43 226 TSL 21h Lab
-RKS-PERSIANSIGN [144] IDSLR 2020 RGB 10000 10 100 Persian - Lab
-BSL-1K [11] IDSLR 2020 RGB 280K 40 1064 BSL 1060h TV
-NMFs-CSL [69] IDSLR 2021 RGB 32010 10 1.067 CSL 24.4h Lab
-KArSL [163] IDSLR 2021 RGBD, Pose 75300 3 502 ArSL - Lab
-CSLD [184] IDSLR 2022 RGB 120K 30 400 CSL 66.8h Lab
-NCSL [182] IDSLR 2022 RGB 90K 30 300 CSL 50.1h Lab
-IISL2020 [92] IDSLR 2022 RGB 12100 16 11 ISL 7h Lab
-LSA64 [149] IDSLR 2023 RGB 3200 10 64 LSA 1.9h Lab
-Signsuisse Lexicon [126] IDSLR 2023 RGB 18221 - - 3 (Mult) - Web
-FDMSE-ISL [135] IDSLR 2024 RGBD 40033 20 2002 ISL 36h Lab
+Additionally, some datasets support Sign Language Translation (SLT) by providing sentence-level gloss annotations and parallel spoken language transcriptions.
+The RWTH-PHOENIX-Weather 2014 (PHOENIX-2014) and its extended version PHOENIX-2014-T1 (NEURAL-SIGN-LANGUAGE-TRANSLATION-CVPR-CAMGOZ) are benchmark datasets offering sentence-level annotations in German Sign Language (DGS) captured from television weather forecasts.
+A semi-automated annotation protocol was proposed, leveraging existing methods for speech recognition and accelerating the process.
+The signers preform gestures in a constrained environment (i.e.
+wearing dark clothes in front of a static gray background), probably limiting the generalization of models trained to it.
+On the contrary, more recent efforts such as BOBSL (BOBSL-BBC) and OpenASL (OpenASL) have significantly expanded the signer base and vocabulary, with hours of British (BSL) and American Sign Language (ASL) footage collected from TV and web sources, respectively.
+The more unconstrained environment and the increased number of signers improve the HGR models real-world performance.
+The lab-recorded KETI (KETI_MDPI) comprises samples captured from two camera angles (i.e.
+front and side), enlarging diversity.
+The aim of the dataset is to interpret Korean sign language gestures that are performed in emergency situations using skeleton and RGB data, reducing the need of professionals.
+Another multimodal dataset collected in a laboratory is the How2Sign (How2Sign-CVPR) for ASL recognition and translation.
+The entire dataset was captured in the termed Green Screen studio with a single RGB camera and depth sensor, while a subset of it in the Panoptic studio using multiple cameras from various viewpoints, enabling accurate 3D pose estimation.
+Pose data and text transcripts are also considered in TVB-HKSL-News (TVB-HKSL-News).
+The samples were collected from TV news with subtitles and Optical Character Recognition was applied to extract the text sequence, while an off-the-shelf pose estimator was utilized to obtain the human keypoints.
+The RGB-based CSL-Daily (Back-Translation-CVPR) provides also glosses and textual transcriptions, aiding both CSLR and SLT tasks.
+A sign dictionary (referred to as SignDict) was also constructed to be used for other tasks, including sign spotting and segmentation.
+The Auslan Daily Dataset (Auslan-Daily) consists of Australian Sign Language (AUSLAN) gestures performed by humans in a television series, TV news and web videos.
+These samples may contain multiple persons interacting with each other, which makes the dataset suitable for in-the-wild scenarios.
+Other datasets for CSLR do not provide text translations.
+ChicagoFSWild+ (ChicagoFSWild-plus-ICCV) was collected from web videos, including ASL signs performed by 260 signers making this dataset the largest in terms of number of subjects.
+HKSL (SignBERT-HKSL) was captured in a laboratory environment, providing RGBD data.
+Datasets for IDLSR are larger compared to datasets for CSLR, as they do not require complex annotation making the data collection easier.
+Many of these datasets, such as CSL (CSL500), GSL iso.
+(ADADOGLOU), BoshporusSign22k (bosphorussign22k), AUTSL (AUTSL), KArSL (KArSL) and FDMSE-ISL (FDMSE-hierarchicalwindowedgraphattention), include pose or depth modalities, enriching multimodal learning.
+All of them, except for the AUTSL, were collected in a laboratory with fixed background.
+Datasets that do not include such modalities, can be extended to them by using off-the-shelf DL pose and depth estimators.
+The ASL Dataset (TAXONOMY_OF_HAND_GESTURES_2000_ASL_DATASET) is the only one consisting of pose sequences exclusively.
+MS-ASL (MS-ASL) and WLASL (WLASL) are two RGB-based datasets isolated ASL recognition collected from publicly available web videos of the deaf community, including educational resources and ASL tutorials and performed by 222 and 119 subjects respectively, increasing the diversity.
+However, they contain relatively few samples compared to large-scale datasets such as BSL-1K (bsl-1K), CSLD (tim-slr_csld_dataseta) and NCSL (2plus1-D-SLR_and_NCSL).
+For isolated Irish Sign Language (ISL) recognition, only laboratory datasets have been proposed (e.g.
+INCLUDE (INCLUDE), IISL2020 (IISL2020_MDPI) and FDMSE-ISL (FDMSE-hierarchicalwindowedgraphattention)), making necessary the collection of an in-the-wild dataset.
+For Vietnamese SLR, authors in (Multi-VSL-dinh2025sign) collecting a large-scale multi-view dataset, considering three different point of views, enhancing diversity.
+Other sign languages have also been considered including Argentinian (e.g.
+LSA64 (Lsa64argentiniansignlanguage)), Persian (e.g.
+RKS-PERSIANSIGN (persian_sl_dataset)) and Iranian (e.g.
+ISLR101 (ISLR101-IRANIAN)).
+The rarity of some sign languages like them mentioned previously poses obstacles to finding relevant annotators, exacerbating the data scarcity challenge, as it will be discussed later.
+Table 6: Overview of the characteristics of each dataset related to SLU. The sign vocabulary (Sign Voc.) corresponds to the number of distinct classes included in each dataset.
+Name Year Task Modality Samples Signers Sign Voc. Language Duration Source
+TVB-HKSL-News (TVB-HKSL-News) 2024 CSLR/SLT RGB, Pose 7000 2 6515 HKSL 16.07h TV
+Auslan-Daily (Auslan-Daily) 2023 CSLR/SLT RGB 25000 67 - AUSLAN - TV, Web
+OpenASL (OpenASL) 2022 CSLR/SLT RGB 98417 220 - ASL 288h Web
+CSL-Daily (Back-Translation-CVPR) 2021 CSLR/SLT RGB 20654 10 2000 CSL 23.27h Lab
+BOBSL (BOBSL-BBC) 2021 CSLR/SLT RGB 1962 39 2281 BSL 1467h TV
+How2Sign (green studio) (How2Sign-CVPR) 2021 CSLR/SLT RGBD, Pose 2529 11 - ASL 79.12h Lab
+Corpus FinSL (Corpus-Finnish) 2020 CSLR/SLT RGB 343 21 - FinSL 14.37h -
+KETI (KETI_MDPI) 2019 CSLR/SLT RGB, Pose 14672 14 524 KSL 27.99h Lab
+GSL SD/SI (ADADOGLOU) 2019 CSLR/SLT RGBD 10295 7 310 GSL 9.59h Lab
+PHOENIX-2014-T1 (NEURAL-SIGN-LANGUAGE-TRANSLATION-CVPR-CAMGOZ) 2018 CSLR/SLT RGB 8257 9 1066 DGS - TV
+PHOENIX-2014 (PHOENIX-2014) 2014 CSLR/SLT RGB - 9 1558 DGS 10.73h TV
+HKSL (SignBERT-HKSL) 2021 CSLR RGBD 2400 6 55 HKSL - Lab
+ChicagoFSWild+ (ChicagoFSWild-plus-ICCV) 2019 CSLR RGB 55232 260 26 ASL - Web
+Multi-VSL (Multi-VSL-dinh2025sign) 2025 IDSLR RGB, Pose 84764 30 1000 Vietnamese - Lab
+BdSLW60 (BdSLW60-rubaiyeat2025bdslw60) 2025 IDSLR RGB 9307 18 60 Bangla - Lab
+ISLR101 (ISLR101-IRANIAN) 2025 IDSLR RGB, Pose 4614 10 101 Iranian - -
+ISW-1000 (ISW-1000-ren2025multi) 2025 IDSLR RGB, Pose, OF 10000 10 1000 CSL - Lab
+FDMSE-ISL (FDMSE-hierarchicalwindowedgraphattention) 2024 IDSLR RGBD 40033 20 2002 ISL 36h Lab
+LSA64 (Lsa64argentiniansignlanguage) 2023 IDSLR RGB 3200 10 64 LSA 1.9h Lab
+Signsuisse Lexicon (SIGN-muller-etal-2023-findings) 2023 IDSLR RGB 18221 - - 3 (Mult) - Web
+CSLD (tim-slr_csld_dataseta) 2022 IDSLR RGB 120000 30 400 CSL 66.8h Lab
+NCSL (2plus1-D-SLR_and_NCSL) 2022 IDSLR RGB 90000 30 300 CSL 50.1h Lab
+KArSL (KArSL) 2021 IDSLR RGBD, Pose 75300 3 502 ArSL - Lab
+NMFs-CSL (NMFs-CSL) 2021 IDSLR RGB 32010 10 1.067 CSL 24.4h Lab
+BosphorusSign22k (bosphorussign22k) 2020 IDSLR RGBD, Pose 22542 6 744 TSL 19h Lab
+AUTSL (AUTSL) 2020 IDSLR RGBD, Pose 38336 43 226 TSL 21h Lab
+INCLUDE (INCLUDE) 2020 IDSLR RGB 4287 7 263 ISL 3h Lab
+IISL2020 (IISL2020_MDPI) 2020 IDSLR RGB 12100 16 11 ISL 7h Lab
+RKS-PERSIANSIGN (persian_sl_dataset) 2020 IDSLR RGB 10000 10 100 Persian - Lab
+BSL-1K (bsl-1K) 2020 IDSLR RGB 280000 40 1064 BSL 1060h TV
+MSASL (MS-ASL) 2019 IDSLR RGB 25513 222 1000 ASL 24.65h Web
+WLASL (WLASL) 2019 IDSLR RGB 21083 119 2000 ASL 14.11h Web
+GSL isol. (ADADOGLOU) 2019 IDSLR RGBD 40785 7 310 GSL 6.44h Lab
+SMILE (SMILE) 2018 IDSLR RGB - 30 100 DSGS - Lab
+ASL Dataset (TAXONOMY_OF_HAND_GESTURES_2000_ASL_DATASET) 2018 IDSLR Pose 1200 20 30 ASL - Lab
+CSL (CSL500) 2018 IDSLR RGBD, Pose 125000 50 500 CSL - Lab
 8.2 Semaphoric HGR
-Semaphoric hand gesture recognition (HGR) research is supported by a diverse range of datasets categorized into three main types based on task formulation: static, isolated dynamic, and continuous dynamic gesture recognition.
-These datasets vary in their application domains (e.g., Human-Computer Interaction (HCI), Human-Machine Interface (HMI), Human-Robot Interaction (HRI)), modalities (e.g., RGB, depth, pose, infrared), and in the number of gesture classes, subjects, and gesture samples they include.
-The statistics of datasets for semaphoric HGR are provided in tables 9,10 and 11, regarding the task for which they are used (i.e.
+Semaphoric HGR research is supported by a diverse range of datasets categorized into three main types based on task formulation: SHGR, IDHGR and CDHGR.
+These datasets vary in their application domains (e.g., HCI, HMI, HRI), modalities (e.g., RGB, depth, pose, infrared), and in the number of gesture classes, subjects, and gesture samples they include.
+The statistics of datasets for semaphoric HGR are provided in tables 7,8 and 9, regarding the task for which they are used (i.e.
 static, isolated dynamic and continuous HGR, respectively).
-Static HGR datasets focus on recognizing gestures from single-frame input and are generally easier to process due to the absence of temporal complexity.
-Notable examples include LaRED [67], one of the earliest RGBD datasets with 243,000 samples and 81 classes, and HaGRIDv2 [131], a large-scale RGB dataset designed for HMI with over 1 million samples, covering 34 classes.
-Smaller yet targeted datasets such as OUHANDS [118] and URGR [17] focus on specific domains like HCI and HRI, while MD-UHGRD [194] offers moderately sized samples and class diversity.
-Isolated dynamic HGR datasets contain gestures performed in isolation, with start and end points explicitly segmented.
-Datasets like ChaLearn IsoGD [179] and EgoGesture (isolated) [199] feature rich modality (RGBD, IR, pose) and high gesture diversity, with up to 249 classes.
-Others like Jester [117] offer large-scale RGB-only data from over 1,300 subjects, making them suitable for data-driven deep learning approaches.
-More recent contributions such as DATE [35] and ZJUGesture [193] reflect growing interest in real-world HCI scenarios.
-Continuous dynamic HGR datasets challenge models to identify and segment gestures from continuous video streams.
-The ChaLearn ConGD [179] dataset remains a benchmark with 22,535 sequences and 249 gesture classes.
-Similarly, EgoGesture [199] and IPN Hand [19] include both isolated and continuous segments, enabling comprehensive model development and evaluation.
-Datasets like LD-ConGR [106] and ZJUGesture [193] contribute additional complexity by increasing subject diversity and contextual variability.
+SHGR datasets target gesture recognition from single frames, making them easier to process due to the lack of temporal complexity.
+Older datasets like HGR-1 (HGR-1-2024-PATTER-RECOGNITION), NUS-I (NUS-I-2010), NUS-II (NUS-II-2013), and OUHANDS (ouhands) contain few samples, limiting deep learning training, yet remain benchmarks for HCI due to their simplicity.
+LaRED (LaRED) offers 243K RGBD samples across 81 gestures, but only from 10 subjects, restricting subject-independent recognition.
+HaGRID (WACV_kapitanov2024hagrid) is much larger and diverse, with nearly twice as many full-resolution RGB frames from  35K subjects in varied scenes, including bounding boxes for hand detection.
+HaGRIDv2 (HAGRIDv2_ARXIV) expands this with 15 additional classes,  66K subjects, and over 1M samples.
+SHAPE (SHAPE_DATASET) has a similar number of gestures (all two-handed) but fewer samples.
+LRHG (LONG_RANGE_HGR_SSD) and MD-UHGRD (MD-UHGRD_SA-YOLO) target UAV-based gesture recognition, with MD-UHGRD offering more samples but fewer classes, captured from varying distances and angles to enhance viewpoint robustness.
+URGR (ULTRA-RANGE-UGV-Bamani) focuses on ultra-range gestures for UGVs, with 6 robot-command gestures performed by 16 signers at distances from 2 to 25 meters.
+IDHGR datasets contain gestures performed in isolation, with start and end points explicitly segmented.
+Most of the existing datasets (e.g.
+SKIG (SKIG), NVIDIA (aka nvGesture) (NVIDIA_GESTURE_R3DCNN_Molchanov), DHG (DHG-CVPRW) and SHREC17 (SHREC)) for IDHGR are relatively small with a few thousands samples and support multimodal settings by providing depth maps, IR frames, optical flow or skeleton joints, accompanying the RGB modality aiming to enhance the accuracy of the models.
+Jester-20bn (Jester-Dataset), proposed for gesture-based HCI, is the largest dataset for (non-SLU) IDHGR containing 1̃50K videos performed by 1376 subjects.
+The remarkable diversity and size of this dataset make it ideal for pretraining DL models.
+Other datasets for HCI include IPN-Hand (IPN-HAND), ZJUGesture (ZJUGesture) and DATE (DATE), comprising 4218, 9892 and 13500 samples respectively.
+LD-ConGR (LD-ConGR), ChaLearn IsoGD (chalearn-lap) and EgoGesture (EgoGesture_DATASET) have sufficient amount of data, constituting adequate baseline for IDHGR.
+EgoDriving (EgoFormer_METHOD_EgoDriving_DATASET), Briareo (Briareo_dataset) and nvGesture (NVIDIA_GESTURE_R3DCNN_Molchanov) have been proposed for human-car interaction, mainly in multimodal setting, including gestures that correspond to car functionalities (e.g.
+CW rotation and audio manipulation [^2] https://youtu.be/NJmk1DUyyB8).
+TCG (TCG) was also proposed for human-car interaction, but for recognizing traffic police gestures in autonomous driving scenarios relying only on pose data.
+EHWGesture (EHWGesture-amprimo2025ehwgesture) is the first dataset for recognizing clinical gestures and providing annotations for many other tasks, including gesture detection & triggering.
+CDHGR datasets challenge models to identify and segment gestures from continuous video streams.
+Many of the datasets mentioned earlier (e.g.
+ChaLearn (chalearn-lap), IPN-Hand (IPN-HAND), LD-ConGR (LD-ConGR), EgoGesture (EgoGesture_DATASET) and ZJUGesture (ZJUGesture)) have been obtained by annotating the temporal boundaries of gestures within continuous sequences, thus these datasets have protocols for both IDHGR and CDHGR.
+Therefore, we will only discuss the Police Gesture Dataset (Police-Gesture-Dataset) as the previous analysis could also be applied here.
+This dataset was introduced for recognizing police traffic gestures, similarly with TCG (TCG), containing a small number of samples (21 RGB videos) and 3354 gesture instances.
 Each dataset supports specific sub-tasks within HGR research: static datasets aid in gesture classification, isolated dynamic datasets support temporal gesture modeling and classification, and continuous datasets allow for joint gesture spotting and classification—crucial for real-time and interactive applications.
 This diversity in dataset structure, scale, and modality reflects the evolving demands and challenges of semaphoric gesture recognition across interactive domains.
-Table 9: Datasets for static hand gesture recognition.
-Name Year Domain Modality Samples Subjects Classes Resolution
-HGR-1 [88] 2014 - RGB 899 12 27 174x131-640x480
-LaRED [67] 2014 - RGBD 243000 10 81 640x480
-OUHANDS [118] 2016 HCI RGBD 3150 23 10 640x480
-LRHG [211] 2021 HRI RGB 4320 8 10 640x480
-SHAPE [33] 2022 - RGB 34226 20 32 4128x3096
-HaGRID [86] 2024 HCI RGB 552992 34730 19 1920x1080
-HaGRIDv2 [131] 2024 HCI RGB 1086158 65977 34 1920x1080
-MD-UHGRD [194] 2024 HRI RGB 20000 - 19 -
-URGR [17] 2024 HRI RGB 347483 16 6 480x640
-Table 10: Datasets for isolated dynamic hand gesture recognition.
+Table 7: Datasets for SHGR.
+Name Year Domain Modality Samples Subjects Classes Resolution Distance
+HaGRID (WACV_kapitanov2024hagrid) 2024 HCI RGB 552992 34730 19 1920x1080 0.5-4m
+HaGRIDv2 (HAGRIDv2_ARXIV) 2024 HCI RGB 1086158 65977 34 1920x1080 0.5-4m
+MD-UHGRD (MD-UHGRD_SA-YOLO) 2024 HRI RGB 20000 - 19 - 3-10m
+URGR (ULTRA-RANGE-UGV-Bamani) 2024 HRI RGB 347483 16 6 640x480 1-25m
+SHAPE (SHAPE_DATASET) 2022 - RGB 34226 20 32 4128x3096 2-5m
+LRHG (LONG_RANGE_HGR_SSD) 2021 HRI RGB 4320 8 10 640x480 1-7m
+OUHANDS (ouhands) 2016 HCI RGBD 3150 23 10 640x480 $\leq$ 1m
+HGR-1 (HGR-1-2024-PATTER-RECOGNITION) 2014 - RGB 899 12 27 174x131-640x480 $\leq$ 1m
+LaRED (LaRED) 2014 - RGBD 243000 10 81 640x480 -
+NUS-II (NUS-II-2013) 2013 - RGB 2000 40 10 160x120 $\leq$ 1m
+ASL-FS (ASL-FS-2011-CVPRW) 2011 - RGBD 48K 4 24 - -
+NUS-I (NUS-I-2010) 2010 - RGB 240 - 10 160x120 $\leq$ 1m
+Table 8: Datasets for IDHGR.
 Name Year Domain Modality Samples Subjects Classes Frames
-SKIG [108] 2013 - RGBD 1080 6 10 -
-nvGesture [124] 2016 HMI RGBD, IR, OF 1532 20 25 -
-ChaLearn IsoGD [179] 2016 - RGBD 47933 21 249 -
-DHG-14/28 [37] 2016 - Depth, Pose 2800 20 14/28 -
-SHREC17-14/28 [166] 2017 - Depth, Pose 2800 28 14/28 -
-EgoGesture (isol.) [199] 2017 HCI RGBD 24161 50 83 2953224
-Briareo [115] 2019 HMI RGBD, Pose, IR 1440 40 12 -
-Jester [117] 2019 HCI RGB 148092 1376 27 5331312
-IPN Hand (isol.) [19] 2020 HCI RGB 4218 50 13 -
-TCG [189] 2020 HMI Pose 250 5 4 839350
-LD-ConGR (isol.) [106] 2022 HCI RGBD 44887 30 10 -
-ZJUGesture (isol.) [193] 2023 HCI RGB 9892 60 9 -
-DATE [35] 2024 HCI RGB 13500 22 27 -
-EgoDriving [138] 2024 HMI RGB 400 - 29 70276
-Table 11: Datasets for continuous dynamic hand gesture recognition.
+EHWGesture (EHWGesture-amprimo2025ehwgesture) 2025 - RGBD 1100 25 11 3.3M
+EgoDriving (EgoFormer_METHOD_EgoDriving_DATASET) 2024 HMI RGB 400 - 29 70276
+DATE (DATE) 2024 HCI RGB 13500 22 27 -
+ZJUGesture (isol.) (ZJUGesture) 2023 HCI RGB 9892 60 9 -
+LD-ConGR (isol.) (LD-ConGR) 2022 HCI RGBD 44887 30 10 -
+IPN Hand (isol.) (IPN-HAND) 2021 HCI RGB 4218 50 13 -
+TCG (TCG) 2020 HMI Pose 250 5 4 839350
+Jester (Jester-Dataset) 2019 HCI RGB 148092 1376 27 5331312
+Briareo (Briareo_dataset) 2019 HMI RGBD, Pose, IR 1440 40 12 -
+EgoGesture (isol.) (EgoGesture_DATASET) 2018 HCI RGBD 24161 50 83 2953224
+SHREC17-14/28 (SHREC) 2017 - Depth, Pose 2800 28 14/28 -
+nvGesture (NVIDIA_GESTURE_R3DCNN_Molchanov) 2016 HMI RGBD, IR, OF 1532 20 25 -
+ChaLearn IsoGD (chalearn-lap) 2016 - RGBD 47933 21 249 -
+DHG-14/28 (DHG-CVPRW) 2016 - Depth, Pose 2800 20 14/28 -
+SKIG (SKIG) 2013 - RGBD 1080 6 10 -
+Table 9: Datasets for CHGR.
 Name Year Domain Modality Samples Instances Subjects Classes
-ChaLearn ConGD [179] 2016 - RGBD 22535 47933 21 249
-EgoGesture (cont.) [199] 2017 HCI RGBD 2081 24161 50 83
-IPN Hand (cont.) [19] 2020 HCI RGB 200 4218 50 13
-Police Gesture Dataset [66] 2020 HMI RGB 21 3354 - 8
-LD-ConGR (cont.) [106] 2022 HCI RGBD 542 44887 30 10
-ZJUGesture (cont.) [193] 2023 HCI RGB - 9892 60 9
+ZJUGesture (cont.) (ZJUGesture) 2023 HCI RGB - 9892 60 9
+LD-ConGR (cont.) (LD-ConGR) 2022 HCI RGBD 542 44887 30 10
+IPN Hand (cont.) (IPN-HAND) 2021 HCI RGB 200 4218 50 13
+Police Gesture Dataset (Police-Gesture-Dataset) 2020 HMI RGB 21 3354 - 8
+EgoGesture (cont.) (EgoGesture_DATASET) 2018 HCI RGBD 2081 24161 50 83
+ChaLearn ConGD (chalearn-lap) 2016 - RGBD 22535 47933 21 249
 8.3 Other
 In addition to standard semaphoric gesture datasets, several specialized datasets have been introduced to support research on more specific gesture types.
-The First-Person Hand Action (FPHA) dataset [50], published in 2018, contains 1,175 video sequences captured from a first-person perspective by six participants.
+The First-Person Hand Action (FPHA) dataset (FPHA), published in 2018, contains 1,175 video sequences captured from a first-person perspective by six participants.
 It focuses on isolated manipulative hand gestures across 45 action classes, offering RGB, depth, and skeleton data, and is particularly valuable for studies involving egocentric interaction.
-Another notable example is the DP Dataset [127], released in 2023, which, to the best of our knowledge, is the only publicly available benchmark specifically designed for pointing gesture recognition.
+Another notable example is the DP Dataset (DeePoint), released in 2023, which, to the best of our knowledge, is the only publicly available benchmark specifically designed for pointing gesture recognition.
 Collected from 33 participants in two indoor environments, the dataset comprises approximately 2.8 million frames and 6,335 annotated gesture instances.
@@
 non-gesture) and directional estimation tasks.
+Finally, the SocialGesture (SocialGesture-CVPR2025-cao2025socialgesture) focuses on the recognition of natural gestures in multi-person environments.
+It comprises 9,889 in-the-wild videos containing 42,533 instances and provides fine-grained annotations including bounding boxes, relations (for example, a subject is pointing to a target) and Visual Question Answering (VQA) details.
+8.4 Datasets Comparison
+After examining the existing datasets for VHGR with significant impact on the current research, in this subsection we aim to determine pros and cons of the most important ones as well as the use cases that they fit.
+For gesture-based HCI, HaGRID (WACV_kapitanov2024hagrid) and its versions (HAGRIDv2_ARXIV) seem to be a reliable solution, containing an enormous number of samples corresponding to 19 distinct classes (that can be mapped to computer commands).
+Regarding that these gestures are also static, even simple architectures - like the ones discussed in sections 4 & 7 - can achieve sufficient performance.
+Furthermore, this dataset could also be used for pretraining, enhancing the backbone’s expression capability.
+However, its size implies that training takes a long time, which constitutes a deterrent in some cases, compared to other datasets for HCI like OUHANDS (ouhands).
+For gesture-based HRI, URGR (ULTRA-RANGE-UGV-Bamani) is the only one for UGV guidance and considers various recognition ranges, which may be crucial for real-world scenarios.
+Even though it contains a large number of samples, the small number of classes may limit the commands given by gestures and, therefore, the functionalities of the entire gesture-based interface.
+Regarding SLU datasets, for ASL recognition, OpenASL (OpenASL) is the largest one, including almost 100K RGB videos (288h in total) collecting from the web.
+MSASL (MS-ASL) comprises almost the same number of subjects, but it is much smaller.
+Furthermore, OpenASL was proposed for CSLR while for MSASL for IDSLR, thus it may be more suitable for real-world use cases, where the temporal boundaries are not known.
+BSL-1K (bsl-1K) is also a very large dataset, proposed for British IDSLR and comprising the largest number of samples (280K, 1060h in total).
+BOBSL (BOBSL-BBC) was proposed for British CSLR and comprises videos of 1467h in total making it the dataset with the longest duration.
+Phoenix (PHOENIX-2014) and Phoenix T (NEURAL-SIGN-LANGUAGE-TRANSLATION-CVPR-CAMGOZ) remain the most adequate dataset for German sign language, covering a large sign vocabulary.
+Nevertheless, the small number of samples and the low diversity - due to the limited number of signers and the constrained environment - make essential the collection of a new, larger and more diverse dataset for DGS.
+For Chinese SLR, NMFs-CSL (NMFs-CSL) is significantly smaller than CSLD (tim-slr_csld_dataseta), CSL (CSL500) and NCSL (2plus1-D-SLR_and_NCSL) but includes much more signs (1064 instead of less than 500).
+CSL-Daily (Back-Translation-CVPR) remains the only one for Chinese CSLR, covering a large sign vocabulary of 2000 signs in total.
 9 Evaluation Metrics
@@
 Macro-averaged metrics compute the metric independently for each class and then average, treating all classes equally regardless of their frequency.
-Micro-averaged metrics:
-$\text{Accuracy}_{micro}=\frac{\sum_{i=1}^{N}(TP_{i}+TN_{i})}{\sum_{i=1}^{N}(TP_{i}+TN_{i}+FP_{i}+FN_{i})}$ (1)
-$\text{Precision}_{micro}=\frac{\sum_{i=1}^{N}TP_{i}}{\sum_{i=1}^{N}(TP_{i}+FP_{i})}$ (2)
-$\text{Recall}_{micro}=\frac{\sum_{i=1}^{N}TP_{i}}{\sum_{i=1}^{N}(TP_{i}+FN_{i})}$ (3)
-[…]
-Macro-averaged metrics:
-$\text{Accuracy}_{macro}=\frac{1}{N}\sum_{i=1}^{N}\frac{TP_{i}+TN_{i}}{TP_{i}+FP_{i}+TN_{i}+FN_{i}}$ (5)
-$\text{Precision}_{macro}=\frac{1}{N}\sum_{i=1}^{N}\frac{TP_{i}}{TP_{i}+FP_{i}}$ (6)
-$\text{Recall}_{macro}=\frac{1}{N}\sum_{i=1}^{N}\frac{TP_{i}}{TP_{i}+FN_{i}}$ (7)
-[…]
-F1-Score: For both micro and macro cases, the F1-score is defined as the harmonic mean of Precision and Recall:
-$F_{1}=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$ (9)
+The formulae for each metric can be found in the Table 10.
+Table 10: Common utilized metrics for evaluation of hand gesture classification.
 Here, $TP_{i}$, $FP_{i}$, $TN_{i}$, and $FN_{i}$ denote the number of true positives, false positives, true negatives, and false negatives for the $i$-th class, respectively, and $N$ is the total number of classes.
+Metrics Micro-averaged formulae Macro-averaged formulae
+Accuracy $\frac{\sum_{i=1}^{N}(TP_{i}+TN_{i})}{\sum_{i=1}^{N}(TP_{i}+TN_{i}+FP_{i}+FN_{i})}$ $\frac{1}{N}\sum_{i=1}^{N}\frac{TP_{i}+TN_{i}}{TP_{i}+FP_{i}+TN_{i}+FN_{i}}$
+Precision $\frac{\sum_{i=1}^{N}TP_{i}}{\sum_{i=1}^{N}(TP_{i}+FP_{i})}$ $\frac{1}{N}\sum_{i=1}^{N}\frac{TP_{i}}{TP_{i}+FP_{i}}$
+Recall $\frac{\sum_{i=1}^{N}TP_{i}}{\sum_{i=1}^{N}(TP_{i}+FN_{i})}$ $\frac{1}{N}\sum_{i=1}^{N}\frac{TP_{i}}{TP_{i}+FN_{i}}$
+Jaccard-Index $\frac{\sum_{i=1}^{N}TP_{i}}{\sum_{i=1}^{N}(TP_{i}+FP_{i}+FN_{i})}$ $\frac{1}{N}\sum_{i=1}^{N}\frac{TP_{i}}{TP_{i}+FP_{i}+FN_{i}}$
+F1-Score $2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$
 9.2 Sequence-to-Sequence Metrics
 For continuous HGR (CHGR), particularly in Continuous Sign Language Recognition (CSLR), sequence-level metrics are used.
-Word Error Rate (WER) [122, 2] is the standard metric, measuring the minimum number of substitutions ($S$), deletions ($D$), and insertions ($I$) needed to transform the predicted sequence into the ground truth sequence of length $N$:
+Word Error Rate (WER) (VAC-CSLR; ADADOGLOU) is the standard metric, measuring the minimum number of substitutions ($S$), deletions ($D$), and insertions ($I$) needed to transform the predicted sequence into the ground truth sequence of length $N$:
 $WER=\frac{S+D+I}{N}$ (1)
-Edit Distance Accuracy [45, 185] is a complementary metric representing the ratio of correctly predicted elements in the sequence:
-[…]
+Edit Distance Accuracy (FA-STGCN; CTCX_WANG2022123) is a complementary metric representing the ratio of correctly predicted elements in the sequence:
+$\text{Edit-Accuracy}=\frac{N-S-D-I}{N}$ (2)
 These metrics provide insight not only into classification accuracy but also into temporal and sequential consistency.
@@
 In addition to recognition performance, computational efficiency is often evaluated, especially for real-time or embedded HGR applications. Common metrics include:
+  • 1.
     Inference Time: Time (in ms or seconds) taken to process one input sample.
+  • 2.
     Frames Per Second (FPS): Number of frames processed per second; higher values indicate faster processing.
-  • Model Size: The number of trainable parameters or total memory footprint of the model.
+  • 3.
+    Model Size in Megabytes (MBs): The total memory footprint of the model.
+  • 4.
+    Number of Parameters: The number of trainable parameters.
+  • 5.
+    Floating Point Operations (FLOPs): Measures the number of numerical operations required to run a model (either during training or inference time).
+Being independent of the hardware utilized, it could be considered - along with the number of parameters - more reliable.
 These metrics are crucial when deploying HGR systems on edge devices or in latency-sensitive environments.
-8.4 Overfitting Indicators
-To assess generalization and detect overfitting, the following indicators are used:
-  • Train-Test Performance Gap: A significant difference between training and testing accuracy (or loss) often signals overfitting.
-  • Validation Loss Trend: Increasing validation loss while training loss decreases is a common overfitting pattern.
-  • Early Stopping Criteria: Monitored via a plateau or increase in validation loss across epochs.
 10 Current Challenges and Future Research Directions
 As mentioned in the Methods sections (Sections 4, 5, and 6), researchers have developed sophisticated techniques and large-scale datasets to achieve satisfactory performance.
-However, as shown in the relevant tables (Tables 3, 4, 5, 6, 7), recognition accuracy remains low in many cases—particularly in more complex tasks such as Continuous Sign Language Recognition (CSLR).
-In this section, we discuss the current challenges and limitations that hinder efficient Human Gesture Recognition (HGR), posing ongoing concerns for the research community and shaping future research directions.
+However, as shown in the relevant tables (Tables 2, 3, 4), recognition accuracy remains low in many cases—particularly in more complex tasks such as CSLR.
+In this section, we discuss the current challenges and limitations that hinder efficient VHGR, posing ongoing concerns for the research community and shaping future research directions.
 Additionally, we review the techniques proposed to address these issues, offering insights that may guide future efforts to overcome the identified challenges.
@@
 The presence of gesture-irrelevant factors (or redundant information) hinders accurate recognition in diverse real-world scenarios.
-Some of these factors include the following [171, 9, 80, 17, 141, 206]:
+Some of these factors include the following (EDenseNet; Mohammad_Mahmudul_static_hgr; 2DPSTPP; ULTRA-RANGE-UGV-Bamani; EnGesto; RAAR3DNet):
+  • 1.
     Environmental factors: Cluttered background, varying illumination conditions, and self-occlusions make it difficult or even impossible to automatically spot the hand region, which is an important step for efficient HGR;
+  • 2.
     Subject characteristics: Subject independence –—i.e., the ability to correctly identify a gesture regardless of the signer’s characteristics such as skin color, hand shape variations, or clothing—is a key prerequisite for achieving sufficient generalization and enabling real-world deployment.
+  • 3.
     Variations in perception device position: The distance between the performer and the camera angle may reduce the recognition performance;
+  • 4.
     Input data quality: Low resolution and blurriness may corrupt the discriminative characteristics of the hand gesture region.
-Aiming to address these challenges, researchers have proposed segmentation models to crop hand regions and remove background interference [208, 34, 153], object detectors for spatial normalization [17, 9], super-resolution techniques to enhance image quality [17], attention mechanisms [206, 207] to enforce the visual encoder to pay attention on the most informative regions, information-theoretic losses [100], and feature selection techniques [173] to reduce information redundancy.
+Aiming to address these challenges, researchers have proposed segmentation models to crop hand regions and remove background interference (FGDSNet; tuan_linh_pham_dang2023lightweight; DRCAM_TIM), object detectors for spatial normalization (ULTRA-RANGE-UGV-Bamani; Mohammad_Mahmudul_static_hgr), super-resolution techniques to enhance image quality (ULTRA-RANGE-UGV-Bamani), attention mechanisms (RAAR3DNet; Benjia_Zhou) to enforce the visual encoder to pay attention on the most informative regions, information-theoretic losses (Yunan_Li), and feature selection techniques (SeST) to reduce information redundancy.
 Many of these approaches rely on computationally expensive off-the-shelf models, which increase the complexity of the overall HGR methods and set an upper bound to their accuracy.
-Thus, the presence of gesture-irrelevant factors could concern future research work, while potentially being combined with eXplainable Artificial Intelligence (XAI) approaches [148] to facilitate this task.
+The approaches mentioned previously may be enhanced by using eXplainable Artificial Intelligence (XAI) (Rodis2024Access) techniques, which could be adopted to understand better the behavior of the networks.
+For instance, Grad-CAM could be further utilized to determine the image region that affects more the predicted gesture class (as has already been done by a few researchers: (CorrNet; SWIN_MSTP_Neurocomputing_2025_CSLR)).
+Such results may indicate the effectiveness of the embedded on the network attention mechanisms or other related modules to emphasize the most important parts of the input image.
+t-SNE visualizations could also be leveraged to evaluate the quality of the extracted features, and in particular, how much they are influenced by the presence of redundant information.
+Thus, guided by the tools provided by the XAI or other similar framework, the interference of gesture-irrelevant factors could be mitigated through the careful design of the future architectures.
+10.2 Computational Cost
+Even lightweight methods for SHGR and IDHGR achieve high accuracy.
+Although, for the more complex task of CHGR, the cost is essential; all RGB-based methods (that are significantly more accurate than pose-based) require more than 366.3 GFLOPs ((DSTEN-CSLR-yu2025dsten)) and reach up to 2950.5 GFLOPs ((TemporalLiftPooling)).
+This issue prevents not only the deployment of such methods but also the efficient training, which takes longer.
+Pose-based methods (CoSign) are significantly lighter, but the pose estimation cost is not taken into account.
+The number of parameters does not change dramatically between IDHGR and CHGR methods, while many SHGR methods require more parameters (e.g.
+Gesture-CNN (G-CNN-SHARMA2021115657) has 67.25M parameters).
+The need of large computational resources does not only prevent the real-time and commercial deployment of VHGR systems, but also put barriers to the evolution of the field, as the conducted experiments take many hours or even days to complete.
+To mitigate this, authors have focused on reducing the complexity of their models by using non-parametric layers (GestFormer) or designing smaller architectures (LHGR-Net).
+Thus, reducing the total complexity of the models, could be further examined in the future.
+Considering also recently proposed approaches for DL models compression (DL_MODEL_COMPRESSION_marino2023deep), such as pruning (i.e., removing less important weights) and quantization (i.e., reducing the arithmetic precision of the parameters), may contribute to this direction.
+Freezing layers of pretrained models could also be investigated in order to reduce the time required for fine-tuning.
 10.3 Data Scarcity
 Many existing datasets either contain a relatively small number of samples or lack sufficient diversity.
-For example, among static HGR datasets, only the versions of HaGRID [86, 131] have been collected from more than one hundred participants, despite some others having a plausible number of samples [67, 17].
-A similar situation is observed in IDHGR datasets, where only the Jester dataset [117] includes data from over one hundred subjects.
-Furthermore, the amount of samples for specific modalities, such as IR data, is limited compared with the size of RGB-based datasets, an issue known as data imbalance [181].
+For example, among static HGR datasets, only the versions of HaGRID (WACV_kapitanov2024hagrid; HAGRIDv2_ARXIV) have been collected from more than one hundred participants, despite some others having a plausible number of samples (LaRED; ULTRA-RANGE-UGV-Bamani).
+A similar situation is observed in IDHGR datasets, where only the Jester dataset (Jester-Dataset) includes data from over one hundred subjects.
+Furthermore, the amount of samples for specific modalities, such as IR data, is limited compared with the size of RGB-based datasets, an issue known as data imbalance (IR_data_imbalance_mobilenet).
 The lack of large-scale, diverse datasets leads to overfitting and limited generalization, which is closely related to the challenges discussed earlier.
@@
 Several techniques have been explored to mitigate this challenge:
-  • Multilingual Data Aggregation: Combining datasets (e.g., CSL and DGS) to exploit cross-lingual visual similarities increases the overall amount of training data [188].
-  • Self-Supervised Learning: In the context of the self-supervised learning setting [91], gloss-to-gloss translation tasks have been used for parameter initialization [205], providing linguistic priors to sign language models.
-Pretraining with pose reconstruction has also been investigated to incorporate prior knowledge of hand structure [71, 68, 79].
-  • Advanced Data Augmentation: Adversarial learning has been employed to optimize augmentation hyperparameters [128].
-  • Few-/Zero-shot Learning: Approaches using only a small number of samples to correctly recognize gestures—often based on CLIP—have also been proposed [216, 24, 147, 146, 22].
-These approaches have led to improved recognition performance, even when trained on limited data.
-However, accuracy remains suboptimal, making data scarcity a persistent research challenge and a key direction for future work in the field.
+  • 1.
+    Multilingual Data Aggregation: Combining datasets (e.g., CSL and DGS) to exploit cross-lingual visual similarities increases the overall amount of training data (Fangyun_Wei_ICCV).
+  • 2.
+    Self-Supervised Learning: In the context of the self-supervised learning setting (konstantakos2025self), gloss-to-gloss translation tasks have been used for parameter initialization (CVT-SLR), providing linguistic priors to sign language models.
+Pretraining with pose reconstruction has also been investigated to incorporate prior knowledge of hand structure (SignBERT+; SignBERT_ICCV; omar_ikne).
+  • 3.
+    Advanced Data Augmentation: Adversarial learning has been employed to optimize augmentation hyperparameters (ADVERSARIAL_DATA_AUGM), making the models robust to spatial deformations.
+  • 4.
+    Few-/Zero-shot Learning: Approaches using only a small number of samples to correctly recognize gestures—often based on CLIP—have also been proposed (ZSSLR; FSSLR; ZS-DHGR; ZS-GR; ZERO-SHOT-SLR).
+Generative AI, particularly diffusion models (DMs) for image augmentation (alimisis2025advances), could help address data scarcity in HGR.
+Existing datasets are often limited in size or focused on common tasks, like well-known sign languages.
+DMs can increase dataset diversity by generating additional samples with varied scenes, subjects, and lighting, or support specialized domains such as rare sign languages and human-robot interaction in firefighting or military scenarios.
+Thus, leveraging DMs represents a promising direction for future work.
 10.4 Optimization Difficulties
 Training the visual encoder effectively remains difficult, especially in deep architectures for CSLR.
-As shown in CTCA [57] (Fig.
-17), increasing the depth of the temporal module (e.g., 1D-TCN followed by BiLSTM) improves its generalization but hampers the encoder’s learning due to gradient attenuation in backpropagation.
-Furthermore, the CTC spike phenomenon—where the model over-focuses on a few keyframes—can cause rapid overfitting and limit encoder training [122, 62].
-Figure 17: Experimental results of a study conducted by authors in [57], which aims to measure the impact of the depth of the temporal aggregation module (a 1D-TCN with various number of layers followed by a two-layer BiLSTM) on the effectiveness of training the visual encoder (aka spatial perception module).
-In both diagrams, the x axis represents the number of layers of the 1D-TCN.
-In the top chart, the two modules are compared considering the information stored in weights (IIW) [187] (y axis), which can be viewed as an indicator of generalization gap, while in the bottom the corresponding WER is shown.
-As the number of layers of the 1D-TCN increases, although the generalization ability of the short-term temporal module increases, the overall performance decreases due to the insufficient training of the visual encoder.
+As shown in CTCA (CTCA), (Fig.
+9), increasing the depth of the temporal module (e.g., 1D-TCN followed by BiLSTM) improves its generalization but hampers the encoder’s learning due to gradient attenuation in backpropagation.
+Furthermore, the CTC spike phenomenon—where the model over-focuses on a few keyframes—can cause rapid overfitting and limit encoder training (VAC-CSLR; SMKD).
+Figure 9: Experimental results in (CTCA) analyze the effect of the temporal aggregation module’s depth—a 1D-TCN with varying layers followed by a two-layer BiLSTM—on training the visual encoder (spatial perception module).
+In both charts, the x-axis shows the number of 1D-TCN layers.
+The top chart plots information stored in weights (IIW) (pac-bayes-wang2022pacbayes), indicating generalization ability, while the bottom shows WER.
+Increasing 1D-TCN layers improves short-term temporal generalization but reduces overall performance due to undertrained visual encoder weights.
 To address these issues, several strategies have been proposed:
-  • Auxiliary Supervision: Adding early-stage classifiers to the encoder or short-term module has been proposed as a strategy to guide training by providing additional supervision [122, 143, 74, 76, 111, 213].
-  • Temporal Module Redesign: Placing the 1D-TCN and BiLSTM in parallel instead of sequentially [57].
-  • Linguistic Constraints and Gloss Segmentation: Using gloss boundaries [62] or prior knowledge [58] to regularize learning.
-The aforementioned approaches —particularly auxiliary supervision [122]— have served as the foundation for many existing methods.
+  • 1.
+    Auxiliary Supervision: Adding early-stage classifiers to the encoder or short-term module has been proposed as a strategy to guide training by providing additional supervision (VAC-CSLR; CSGC_Rao_Sun_Wang_Wang_Zhang_2024; CorrNet; corrnetplus_hu2024corrnetsignlanguagerecognition; TCNet_Lu_Salah_Poppe_2024; C2SLR).
+  • 2.
+    Temporal Module Redesign: Placing the 1D-TCN and BiLSTM in parallel instead of sequentially (CTCA).
+  • 3.
+    Linguistic Constraints and Gloss Segmentation: Using gloss boundaries (SMKD) or prior knowledge (GPGN) to regularize learning.
+The aforementioned approaches —particularly auxiliary supervision (VAC-CSLR)— have served as the foundation for many existing methods.
 However, the increasing complexity of modern deep architectures often exceeds the effectiveness of these techniques.
@@
 10.5 Limitations of Architectures
-Despite the achievements of DL, the existing architectures suffer from the following limitations:
-  • Block-fixed models: The existing architectures are based on hand-crafted design, which may not be optimal[206].
-  • GCNs with fixed topology graphs: Standard GCNs use fixed topology graphs based on hand anatomy, which may not capture optimal inter-joint relationships.
-  • Ineffective multimodal fusion: Most of the current works utilize late fusion, that does not fully leverage the available information [68, 71, 202, 204, 203, 104, 82, 81].
-To address these limitations, researchers have proposed various strategies, including Neural Architecture Search (NAS) to identify optimal model structures [206], learnable or adaptive skeleton adjacency matrices [77, 30, 28, 107], and multi-level fusion for dual-stream approaches [60].
+DL methods for HGR have increased the recognition performance. However, the existing architectures suffer from limitations, including the following:
+  • 1.
+    Block-fixed models: The existing architectures are based on hand-crafted design, consisting of block of layers with hyperparameters (e.g., number of channels, convolution kernel size) that are defined manually by the researchers.
+This approach is suboptimal (RAAR3DNet), as several other possible configurations have not been explored.
+  • 2.
+    Insufficient graph representations of GCNs: The standard GCNs utilized fixed topology graphs, often based on the shape of the hand, to better capture the interactions between the various skeleton joints.
+However, this configuration has been proved to be suboptimal (TD-GCN), unable to efficiently capture the temporal dynamics.
+  • 3.
+    Ineffective multimodal fusion: Out of the existing multimodal methods, most of them perform late fusion to incorporate the information of multiple streams (SignBERT_ICCV; SignBERT+; BEST; Weichao-Zhao; MASA; SKIM; SAM_SLR; jiang2021sign-sam-slr-v2).
+This scheme does not fully leverage the available modalities, as the features extracted by each stream are isolated from the others during the processing on the backbones.
+Thus, the obtained vector representations do not benefit from the utilization of multiple modalities.
+To address these limitations, researchers have proposed various strategies, including Neural Architecture Search (NAS) to identify optimal model structures through searching on a large set of configurations (RAAR3DNet), learnable or time-dependent skeleton adjacency matrices that determine implicitly the optimal way to integrate the features of each joint (DSTA_SLR_hu2024dynamic; DSTSA-CGN_cui2025dstsa; ST-PSM_L-PSM; TD-GCN) and multi-level fusion for dual-stream approaches, which enhances the quality of the extracted features by early incorporating information from various sources (CTFB_Hampiholi).
 Despite these advances, several aspects of deep learning architectures remain underexplored, leaving a substantial research gap.
+For instance, to the best of our knowledge, multi-level and mid-level multimodality fusion have been investigated only in the context of dual-stream approaches.
+Thus, future work could consider the intermediate fusion of three or more streams, e.g.
+using fusion modules (MMTM; CTFB_Hampiholi) and identifying the optimal way to connect the different streams.
 11 Conclusions
-The undeniable importance of vision-based hand gesture recognition (HGR) stems from its wide range of applications—spanning human-robot interaction to sign language understanding.
+The undeniable importance of Vision-based Hand Gesture Recognition (VHGR) stems from its wide range of applications—spanning human-robot interaction to sign language understanding.
 This broad applicability has made it a vibrant and active research field.
-Recent advances in deep learning have played a pivotal role in driving progress, enabling the development of increasingly sophisticated architectures and training strategies aimed at tackling more complex and fine-grained tasks, such as continuous sign language recognition.
-This survey focused on HGR methods that leverage visual input and deep learning.
+Recent advances in deep learning (DL) have played a pivotal role in driving progress, enabling the development of increasingly sophisticated architectures and training strategies aimed at tackling more complex and fine-grained tasks, such as continuous sign language recognition (CSLR).
+This survey focused on VHGR methods that leverage visual input and DL.
 These methods were highlighted not only for their superior performance compared to traditional hand-crafted approaches, but also for their higher potential for deployment in real-world scenarios, unlike more cumbersome sensor-based techniques.
-From the surveyed literature, it becomes clear that even relatively simple and shallow deep learning models can achieve high accuracy in static HGR tasks.
+From the surveyed literature, it becomes clear that even relatively simple and shallow DL models can achieve high accuracy in static HGR tasks.
 In such cases, the primary challenges are typical of general computer vision problems and are not inherently tied to HGR itself.
-Despite the significant progress achieved in vision-based hand gesture recognition, several open challenges remain that shape future research directions.
-A major concern is the generalization of models across diverse environments, lighting conditions, camera viewpoints, and user-specific variations, which often lead to performance drops in real-world settings.
-The lack of large-scale, diverse, and well-annotated datasets, especially for continuous gestures and sign language, also hinders the development of more robust models.
-Furthermore, model efficiency and scalability are critical for deployment on resource-constrained edge devices, necessitating advances in lightweight architectures and model compression techniques.
-The integration of multimodal data, such as combining vision with audio or inertial signals, remains underexplored yet promising.
-Finally, interpretability and trustworthiness of deep learning models are essential, particularly for safety-critical applications like healthcare or assistive technologies.
-Addressing these challenges will not only enhance the practical applicability of HGR systems but also contribute to more inclusive and accessible human-computer interaction paradigms.
-[…]
-Acknowledgments The research leading to these results received funding from the European Commission under Grant Agreement No. 101168042 (TRIFFID) and No. 101189557 (TORNADO).
+For isolated dynamic and continuous VHGR more advanced approaches have been proposed, aiming to capture more complicated context.
+The complexity of the architectures (especially for CSLR) cause serious optimization difficulties, leading researchers to develop sophisticated training setups.
+Nevertheless, the most important challenges regarding VHGR are not yet solved, leaving substantial research gaps; the data scarcity - especially concerning more specialized tasks, such as the recognition of rare sign languages - and the high computational cost remain essentially unresolved.
+Furthermore, the existing methods are limited to a predefined list of gestures, which may not be optimal for real-world interaction with humanoid robots or interpretation of the body language.
+Thus, VHGR could be expected to open-vocabulary settings, enabling deployment to broader applications.
+The potential of leveraging the extensive knowledge of foundation models (kas2025foundation_models) in the context of VHGR enabling zero-shot gesture recognition as well as the utilization of model compression techniques (DL_MODEL_COMPRESSION_marino2023deep) to reduce the computational complexity are also rarely explored, leaving space for future work.
+Acknowledgments
+The research leading to these results received funding from the European Commission under Grant Agreement No. 101168042 (TRIFFID) and No. 101189557 (TORNADO).
 Declarations

### v3

@@
 1 Introduction
-The recognition and interpretation of hand gestures (ADADOGLOU) play a central role in advancing Human–Computer Interaction (HCI) and assistive communication systems.
-In recent years, deep learning–based visual hand gesture recognition (VHGR) has gained significant momentum as an active subfield of computer vision, building upon progress in related areas such as action (survey-HAR-introduction-shafizadegan2024multimodal) and facial expression recognition (survey-FER-JESWA-2024).
+The recognition and interpretation of hand gestures (2) play a central role in advancing Human–Computer Interaction (HCI) and assistive communication systems.
+In recent years, deep learning–based visual hand gesture recognition (VHGR) has gained significant momentum as an active subfield of computer vision, building upon progress in related areas such as action (175) and facial expression recognition (169).
 This has led to increased focus on Hand Gesture Recognition (HGR) across various applications.
 For example, HGR is widely used in Human-Robot Interaction (HRI) in industrial settings (e.g.
-robotic arm control (Papadopoulos2021Access; ROBOTIC_ARM)), in healthcare (e.g.
-robot-assisted surgery (TELEOPERATED_SURGICAL_ROBOT)), and in the autonomous navigation of mobile robots such as Unmanned Aerial (MD-UHGRD_SA-YOLO) and Ground (ULTRA-RANGE-UGV-Bamani; Linardakis2024Access) Vehicles (UAVs and UGVs).
-It also plays a key role in HCI, especially in AR/VR environments (OO-dMVMT; EgoGesture_DATASET; BOAT-MI_VR).
-In addition, the large population of deaf and hard-of-hearing individuals has driven research into automatic Sign Language Recognition (SLR) systems, aiming to improve communication and inclusion for these communities (survey-deaf-mute-AIR-2024; NEURAL-SIGN-LANGUAGE-TRANSLATION-CVPR-CAMGOZ).
+robotic arm control (151; 107)), in healthcare (e.g.
+robot-assisted surgery (156)), and in the autonomous navigation of mobile robots such as Unmanned Aerial (213) and Ground (18; 116) Vehicles (UAVs and UGVs).
+It also plays a key role in HCI, especially in AR/VR environments (34; 220; 5).
+In addition, the large population of deaf and hard-of-hearing individuals has driven research into automatic Sign Language Recognition (SLR) systems, aiming to improve communication and inclusion for these communities (216; 27).
 Given the importance of the applications mentioned above, HGR has grown into a broad research field, generally divided into two main approaches: vision-based and sensor-based.
@@
 Publication Range Papers Input data Models Scope Comments Aspects that this survey aims to complement
-(ADADOGLOU) 2016-2020 - Vision-based DL CSLR Provides an in-depth comparison of existing approaches, loss functions and pretraining schemes for CSLR.
+(2) 2016-2020 - Vision-based DL CSLR Provides an in-depth comparison of existing approaches, loss functions and pretraining schemes for CSLR.
 This survey expands the scope including non-CSLR methods and examines more recent methods.
-(survey-5-9622242) 2014-2020 98 Vision-based DL & ML General Investigates the progress on hand gesture representation and Data Acquisition (DAQ) techniques as well as their performance and future research directions.
+(137) 2014-2020 98 Vision-based DL & ML General Investigates the progress on hand gesture representation and Data Acquisition (DAQ) techniques as well as their performance and future research directions.
 It examines relatively antiquated methods.
 The current work aims to emphasize on recent DL approaches and related challenges and future perspectives.
-(survey-10-al2021deep) 2014-2021 138 Sensor- & Vision-based DL SLR Categorizes Continuous and Isolated SLR methods based on their architectures and discusses pose estimation for ISLR in detail.
+(7) 2014-2021 138 Sensor- & Vision-based DL SLR Categorizes Continuous and Isolated SLR methods based on their architectures and discusses pose estimation for ISLR in detail.
 This survey covers more topics including HCI & HRI, and highlights task-specific architectural patterns and challenges for CSLR.
-(survey-11-adeyanju2021machine) 2001-2021 649 Vision-based DL & ML SLR Discusses relationships between research teams and DAQ techniques.
+(3) 2001-2021 649 Vision-based DL & ML SLR Discusses relationships between research teams and DAQ techniques.
 It includes relatively old methods.
 This survey emphasizes on recently proposed ones in order to facilitate the evolution of the field.
-(survey-6-ojeda2022survey) 2000-2020 571 Sensor- & Vision-based DL & ML General Provides many statistical results and bibliometric charts.
+(146) 2000-2020 571 Sensor- & Vision-based DL & ML General Provides many statistical results and bibliometric charts.
 It does not identify contemporary challenges or future research directions; this work seeks to detect relevant material.
-(survey-3-el2022comprehensive) 1998-2020 - Sensor- & Vision-based DL & ML SLR Discusses in detail existing methods for SLR and provides multiple tables for quantitative comparison.
+(46) 1998-2020 - Sensor- & Vision-based DL & ML SLR Discusses in detail existing methods for SLR and provides multiple tables for quantitative comparison.
 This survey analyzes newer approaches for SLR and expands taxonomy and comparison to more application domains, beyond sign language.
-(survey-16-joksimoski2022technological) 2010-2021 80 Vision-based DL & ML SLR A Natural Language Processing (NLP) toolkit was utilized to filter publications.
+(93) 2010-2021 80 Vision-based DL & ML SLR A Natural Language Processing (NLP) toolkit was utilized to filter publications.
 Insightful statistics are also provided.
@@
 On the contrary, this survey aims to identify trends and challenges.
-(survey-17-sarhan2023unraveling) 2015-2023 - Vision-based DL ISLR Provides statistical charts and summary tables based on a four-fold taxonomy.
-This review introduces additional taxonomy criteria and covers more tasks.
-(survey-4-shin2024methodological) 2014-2024 - Sensor- & Vision-based DL & ML General Considers a variety of modalities and examines methods for each one.
+(174) 2015-2023 - Vision-based DL ISLR Provides statistical charts and summary tables based on a four-fold taxonomy. This review introduces additional taxonomy criteria and covers more tasks.
+(181) 2014-2024 - Sensor- & Vision-based DL & ML General Considers a variety of modalities and examines methods for each one.
 The current work aims to discuss further multi-modality fusion strategies and identify more challenges.
-(survey-8-hashi2024systematic) 2018-2023 256 Sensor- & Vision-based ML General Covers various aspects of HGR, including DAQ techniques, hand segmentation and feature extraction.
+(71) 2018-2023 256 Sensor- & Vision-based ML General Covers various aspects of HGR, including DAQ techniques, hand segmentation and feature extraction.
 Focuses exclusively on traditional ML approaches, which are currently surpassed by DL ones.
 This survey, searching for recent advancements, focuses on DL approaches.
-(survey-9-al2024advancements) 2015-2024 58 Sensor- & Vision-based DL & ML SLR Examines various SLR methods and highlights the superiority of DL methods over older ones.
+(6) 2015-2024 58 Sensor- & Vision-based DL & ML SLR Examines various SLR methods and highlights the superiority of DL methods over older ones.
 Its does not identify challenges and future perspectives; this survey aims to cover this, going deeper into details.
-(survey-13-tao2024sign) 2017-2024 - Sensor- & Vision-based DL & ML SLR Groups methods into RNN-based, CNN-based and transformers whilst discusses well-known benchmarks.
+(192) 2017-2024 - Sensor- & Vision-based DL & ML SLR Groups methods into RNN-based, CNN-based and transformers whilst discusses well-known benchmarks.
 This survey expands the scope to non-SLU papers and seeks to provide methods comparison and future prospects.
-(survey-1-rahman2024comparative) 2008-2024 - Sensor- & Vision-based DL & ML General Covers a wide spectrum of gesture-based interaction and .
+(157) 2008-2024 - Sensor- & Vision-based DL & ML General Covers a wide spectrum of gesture-based interaction and .
 The considered methods are relatively old.
 This survey aims to identify more recent advancements.
-(linardakis2025_hgr_survey) 2018-2025 125 Vision-based DL & ML General Discuss jointly hand gesture capturing and recognition techniques.
+(117) 2018-2025 125 Vision-based DL & ML General Discuss jointly hand gesture capturing and recognition techniques.
 Many challenges are highlighted, including data bias and privacy issues.
 It covers a limited number of datasets for HGR and relatively old methods; this survey complements it by analyzing more datasets and recent advancements.
-(survey-emporio2025continuous) 2002-2023 - Sensor- & Vision-based DL & ML CDHGR Focuses on online gesture recognition, discussing DAQ technologies and evaluation metrics, among others.
+(47) 2002-2023 - Sensor- & Vision-based DL & ML CDHGR Focuses on online gesture recognition, discussing DAQ technologies and evaluation metrics, among others.
 This survey covers more topics and provides comparison of state-of-the-art methods.
-(survey-khan2025deep) 2010-2024 - Sensor- & Vision-based DL & ML CSLR Discusses in detail methods and datasets for CSLR.
+(99) 2010-2024 - Sensor- & Vision-based DL & ML CSLR Discusses in detail methods and datasets for CSLR.
 It does not conduct methods comparison and the challenges are general; this survey highlights optimization difficulties in the context of CSLR.
@@
 Compared to existing review papers, this survey provides an in-depth examination of the recent Deep Learning (DL) methods (published since 2021), covering the current state-of-the-art (SoTA).
-In contrast, other existing reviews either focus on older approaches (published since 1998) (survey-6-ojeda2022survey; survey-3-el2022comprehensive; survey-8-hashi2024systematic) or mainly cover traditional machine learning techniques (survey-1-rahman2024comparative; survey-13-tao2024sign; linardakis2025_hgr_survey).
+In contrast, other existing reviews either focus on older approaches (published since 1998) (146; 46; 71) or mainly cover traditional machine learning techniques (157; 192; 117).
 While such works are valuable for understanding the fundamental methods and providing statistical insights, they are less suited to reflecting the recent evolution of the field.
 The present survey stands out by its broader and more up-to-date scope, including 238 papers published since 2021—-an average of nearly 47 publications per year.
-By comparison, (survey-11-adeyanju2021machine) and (survey-6-ojeda2022survey) analyze 649 and 571 papers respectively, but span roughly 20 years of research, corresponding to only about 30 publications per year.
+By comparison, (3) and (146) analyze 649 and 571 papers respectively, but span roughly 20 years of research, corresponding to only about 30 publications per year.
 Finally, the current survey focuses exclusively on vision-based HGR.
@@
 Nevertheless, the review maintains a broad and inclusive perspective, acknowledging that approaches from different domains may share common principles and can inform each other.
-More specifically, the survey focuses on vision-based methods for three key VHGR tasks, which are discussed in detail in the following sections: static HGR (SHGR) (FGDSNet; HAGRIDv2_ARXIV), isolated dynamic HGR (IDHGR) (SAM_SLR; SignBERT_ICCV; NLA_SLR; CTFB_Hampiholi; RAAR3DNet), and continuous (dynamic) HGR (CDHGR or CHGR) (VAC-CSLR; NEURIPS2022_6cd3ac24_HEATMAPS; CorrNet).
+More specifically, the survey focuses on vision-based methods for three key VHGR tasks, which are discussed in detail in the following sections: static HGR (SHGR) (229; 145), isolated dynamic HGR (IDHGR) (91; 78; 235; 66; 227), and continuous (dynamic) HGR (CDHGR or CHGR) (136; 29; 84).
 SHGR involves recognizing gestures from single frames or spatial data only.
@@
 At the same time, all tasks share common issues such as interference from irrelevant background movements or variations in lighting, viewpoint, and hand appearance.
-VHGR has a long history, beginning with traditional computer vision techniques based on hand-crafted features and evolving into complex deep learning pipelines (NVIDIA_GESTURE_R3DCNN_Molchanov).
+VHGR has a long history, beginning with traditional computer vision techniques based on hand-crafted features and evolving into complex deep learning pipelines (138).
 Early approaches typically followed a three-stage process: pre-processing, feature extraction, and classification.
@@
 Modern deep learning architectures have become increasingly deep and sophisticated, especially for dynamic HGR tasks.
-These models first capture short-term motion patterns and then aggregate them to understand long-term context (VAC-CSLR; TemporalLiftPooling; SMKD; SEN-CSLR; CorrNet).
+These models first capture short-term motion patterns and then aggregate them to understand long-term context (136; 83; 69; 85; 84).
 Due to their superior performance, this study focuses primarily on deep learning-based methods.
@@
   • 1.
-    Sign Language Understanding (SLU): Involves the automatic interpretation of gestures in either a single sign language (monolingual SLU (UNISIGNli2025; BOBSL-BBC)) or across multiple sign languages (multilingual SLU (SIGN-muller-etal-2023-findings)).
+    Sign Language Understanding (SLU): Involves the automatic interpretation of gestures in either a single sign language (monolingual SLU (113; 12)) or across multiple sign languages (multilingual SLU (140)).
   • 2.
     Human-Computer Interaction (HCI): Encompasses any form of interaction between users and computers.
-Numerous studies focus on HCI (HAGRIDv2_ARXIV; ZJUGesture), with particular emphasis on AR/VR applications (EgoGesture_DATASET; OO-dMVMT; BOAT-MI_VR).
+Numerous studies focus on HCI (145; 212), with particular emphasis on AR/VR applications (220; 34; 5).
   • 3.
-    Human-Robot Interaction (HRI): Includes gesture-based control of mobile robots such as UGVs (ULTRA-RANGE-UGV-Bamani; MOBILE_ROBOT_DRCAM) and UAVs (LONG_RANGE_HGR_SSD; MD-UHGRD_SA-YOLO), industrial robotic arms (ROBOTIC_ARM), assistive robots (DIALOG_POINTING), and surgical robots (TELEOPERATED_SURGICAL_ROBOT).
+    Human-Robot Interaction (HRI): Includes gesture-based control of mobile robots such as UGVs (18; 172) and UAVs (232; 213), industrial robotic arms (107), assistive robots (31), and surgical robots (156).
   • 4.
     Human-Machine Interaction (HMI): Focuses on interfaces between humans and machines, especially in automotive contexts.
-VHGR enables recognition of traffic gestures (e.g., from police officers) (FA-STGCN; MG-GCT; CTCX_WANG2022123) and supports the development of more intuitive in-vehicle user interfaces (NVIDIA_GESTURE_R3DCNN_Molchanov; EgoFormer_METHOD_EgoDriving_DATASET; ConvMixFormer; Briareo_dataset; MVTN_ARXIV).
+VHGR enables recognition of traffic gestures (e.g., from police officers) (50; 65; 204) and supports the development of more intuitive in-vehicle user interfaces (138; 155; 58; 128; 57).
 3.2 Task
@@
 Gestures are broadly classified as either static or dynamic, depending on whether they involve a fixed hand posture or a sequence of movements over time.
-Dynamic gestures typically consist of three temporal phases: preparation, core (or nucleus), and retraction (NVIDIA_GESTURE_R3DCNN_Molchanov), which reflect the gesture’s evolution from initiation to completion.
+Dynamic gestures typically consist of three temporal phases: preparation, core (or nucleus), and retraction (138), which reflect the gesture’s evolution from initiation to completion.
 The recognition of dynamic gestures can be further divided into isolated and continuous, considering only one gesture per data sequence and identifying gestures in a continuous manner, respectively.
@@
 The choice of input modality directly influences the representation architecture and the deep learning models employed.
-While RGB frames remain the most commonly used modality, other forms of visual data have also been explored—such as depth maps (including RGB-like representations, such as HHA (SAM_SLR)), infrared (IR) images, skeleton joints (and derived features like bones and motion (TMS-NET)), pose heatmaps (NEURIPS2022_6cd3ac24_HEATMAPS; NLA_SLR), optical flow and motion history (MOTION_HISTORY_Mercanoglu).
+While RGB frames remain the most commonly used modality, other forms of visual data have also been explored—such as depth maps (including RGB-like representations, such as HHA (91)), infrared (IR) images, skeleton joints (and derived features like bones and motion (40)), pose heatmaps (29; 235), optical flow and motion history (133).
 In terms of the number of input modalities, methods are classified as uni-modal, when they use a single type of input data, or multi-modal, when they integrate two or more modalities.
-In the case of multiple modalities in the input, a multi-modality fusion strategy is chosen (MMTM; CTFB_Hampiholi).
+In the case of multiple modalities in the input, a multi-modality fusion strategy is chosen (195; 66).
 This could be:
   • 1.
-    Early (data-level) fusion: Combines modalities at the input level before processing (DHGR_3D_POSE_HRI; 3STREAM_DISTILLATION).
+    Early (data-level) fusion: Combines modalities at the input level before processing (54; 198).
   • 2.
-    Late (score-level) fusion: Fuses final prediction outputs from independent uni-modal models (MOTION_HISTORY_Mercanoglu; JOINTS_MOTION_VECTORS_Improved_ST-GCN; GestFormer).
+    Late (score-level) fusion: Fuses final prediction outputs from independent uni-modal models (133; 186; 56).
   • 3.
-    Intermediate (feature-level) fusion: Merges features extracted from separate modalities at a single processing level  (tang2021graph).
+    Intermediate (feature-level) fusion: Merges features extracted from separate modalities at a single processing level  (190).
   • 4.
-    Multi-level fusion: Enables interaction across modalities at multiple stages of processing (NLA_SLR; CTFB_Hampiholi).
+    Multi-level fusion: Enables interaction across modalities at multiple stages of processing (235; 66).
 Finally, based on the type of input, the methods are broadly categorized into two groups: pose-based (or skeleton-based) approaches, which exclusively process skeleton data, and appearance-based approaches, which rely on visual cues from RGB, depth, infrared, or similar modalities.
-(a) From left to right: RGB, Optical Flow, Depth and IR (NVIDIA_GESTURE_R3DCNN_Molchanov)
-(b) Skeleton joints (UNISIGNli2025)
-(c) Pose heatmap (NEURIPS2022_6cd3ac24_HEATMAPS)
-(d) Motion history (MOTION_HISTORY_Mercanoglu)
+(a) From left to right: RGB, Optical Flow, Depth and IR (138)
+(b) Skeleton joints (113)
+(c) Pose heatmap (29)
+(d) Motion history (133)
 Figure 3: Modalities that are considered for VHGR.
@@
 Building upon the previous criterion, which focused on the nature of the input, this dimension considers the meaning or intent behind the gestures.
-Specifically, it examines the kind of information gestures are meant to transmit (TAXONOMY_OF_HAND_GESTURES_2000_ASL_DATASET), resulting in the following categories (illustrated in Fig.
+Specifically, it examines the kind of information gestures are meant to transmit (17), resulting in the following categories (illustrated in Fig.
 4):
   • 1.
-    Semaphoric gestures convey specific commands or instructions and are commonly used in gesture-based interfaces (NVIDIA_GESTURE_R3DCNN_Molchanov; HAGRIDv2_ARXIV; ULTRA-RANGE-UGV-Bamani).
+    Semaphoric gestures convey specific commands or instructions and are commonly used in gesture-based interfaces (138; 145; 18).
   • 2.
-    Language gestures function similarly to semaphoric ones but are exclusively used in the context of sign languages (BOBSL-BBC; Auslan-Daily).
+    Language gestures function similarly to semaphoric ones but are exclusively used in the context of sign languages (12; 177).
   • 3.
-    Deictic (pointing) gestures are used to indicate objects, directions, or spatial targets (DeePoint).
+    Deictic (pointing) gestures are used to indicate objects, directions, or spatial targets (141).
   • 4.
-    Manipulative gestures aim to interact with or mimic the manipulation of physical objects (FPHA).
-(a) Semaphoric (WACV_kapitanov2024hagrid)
-(b) Language (BOBSL-BBC)
-(c) Pointing (DeePoint)
-(d) Manipulative (FPHA)
+    Manipulative gestures aim to interact with or mimic the manipulation of physical objects (55).
+(a) Semaphoric (95)
+(b) Language (12)
+(c) Pointing (141)
+(d) Manipulative (55)
 Figure 4: Main categories of hand gestures based on type of information conveyed.
 3.5 Camera Point of View
-Complementing the previous criteria on gesture content and input modalities, this dimension focuses on the spatial relationship between the camera and the performer (EgoGesture_DATASET).
+Complementing the previous criteria on gesture content and input modalities, this dimension focuses on the spatial relationship between the camera and the performer (220).
 The camera’s perspective influences both the visual characteristics of the gesture and the recognition model design.
@@
     Third-person view: The camera observes the gesture from an external position, detached from the performer.
-(a) First-person view (egocentric) (EgoGesture_DATASET)
-(b) Second-person view (Lsa64argentiniansignlanguage)
-(c) Third-person view (Auslan-Daily)
+(a) First-person view (egocentric) (220)
+(b) Second-person view (167)
+(c) Third-person view (177)
 Figure 5: Different camera points of view for VHGR.
@@
   • 1.
-    Transfer Learning, where models are pretrained on large-scale datasets (e.g., Jester (Jester-Dataset)) and subsequently fine-tuned for specific downstream VHGR tasks (pretrained_on_jester_3dDenseNet_LSTM).
+    Transfer Learning, where models are pretrained on large-scale datasets (e.g., Jester (131)) and subsequently fine-tuned for specific downstream VHGR tasks (125).
   • 2.
-    Weakly Supervised Learning, often used in Continuous Hand Gesture Recognition (CHGR), where temporal annotations are coarse or unaligned, lacking exact gesture boundaries (Fangyun_Wei_ICCV).
+    Weakly Supervised Learning, often used in Continuous Hand Gesture Recognition (CHGR), where temporal annotations are coarse or unaligned, lacking exact gesture boundaries (207).
   • 3.
-    Knowledge Distillation, in which one part of the network (often a teacher model) transfers knowledge to another (student model) to enhance feature extraction and efficiency (VAC-CSLR; C2ST; NEURIPS2022_6cd3ac24_HEATMAPS).
+    Knowledge Distillation, in which one part of the network (often a teacher model) transfers knowledge to another (student model) to enhance feature extraction and efficiency (136; 217; 29).
   • 4.
-    Contrastive Learning, used either in a CLIP-based framework to support zero-shot/few-shot recognition by leveraging textual descriptions (ZERO-SHOT-SLR; ZS-GR; FEW-SHOT-BILGE2024110374), or to improve the discriminative capacity of learned representations (MASA; Weichao-Zhao).
+    Contrastive Learning, used either in a CLIP-based framework to support zero-shot/few-shot recognition by leveraging textual descriptions (23; 164; 24), or to improve the discriminative capacity of learned representations (223; 225).
   • 5.
-    Multi-task Learning, where a single model is optimized to perform multiple related tasks simultaneously, encouraging knowledge sharing and improved generalization (UNISIGNli2025).
+    Multi-task Learning, where a single model is optimized to perform multiple related tasks simultaneously, encouraging knowledge sharing and improved generalization (113).
   • 6.
-    Reinforcement Learning, applied to optimize decision-making processes, such as keyframe selection, by rewarding effective representations (KEYFRAMES_DRL_SKELETON).
+    Reinforcement Learning, applied to optimize decision-making processes, such as keyframe selection, by rewarding effective representations (51).
   • 7.
-    Few/Zero-shot Learning, which enable the model to generalize to new gestures or languages with limited or no annotated examples, often leveraging semantic or textual guidance (ZERO-SHOT-SLR; ZS-GR; FEW-SHOT-BILGE2024110374).
+    Few/Zero-shot Learning, which enable the model to generalize to new gestures or languages with limited or no annotated examples, often leveraging semantic or textual guidance (23; 164; 24).
   • 8.
-    Adversarial Learning, used as a data augmentation strategy (alimisis2025advances) to improve robustness, as demonstrated in adversarial video synthesis networks (ADVERSARIAL_DATA_AUGM).
+    Adversarial Learning, used as a data augmentation strategy (13) to improve robustness, as demonstrated in adversarial video synthesis networks (142).
   • 9.
-    Masked Reconstruction Pretraining, inspired by BERT-style architectures, in which models are pretrained to reconstruct masked skeleton sequences, injecting structural prior knowledge (e.g., hand topology) (BEST; SignBERT+; SignBERT_ICCV).
+    Masked Reconstruction Pretraining, inspired by BERT-style architectures, in which models are pretrained to reconstruct masked skeleton sequences, injecting structural prior knowledge (e.g., hand topology) (224; 77; 78).
 3.8 Processing
 Existing methods for isolated and continuous dynamic HGR can be further categorized based on whether the models have full access to the entire gesture(s) sequence or not.
-According to (NVIDIA_GESTURE_R3DCNN_Molchanov), two types of processing strategies apply:
+According to (138), two types of processing strategies apply:
   • 1.
@@
 Thus, online gesture recognition is suitable for real-time applications, as the gesture is recognized before the retraction part, reducing the latency.
-Sliding windows and gesture detector/classifier pipelines are often employed (IPN-HAND; survey-emporio2025continuous).
+Sliding windows and gesture detector/classifier pipelines are often employed (20; 47).
   • 2.
@@
 This approach is limited to applications that do not require real-time recognition.
-Most of the existing methods for VHGR fall into this category, such as those using BiLSTM for temporal modeling (VAC-CSLR).
+Most of the existing methods for VHGR fall into this category, such as those using BiLSTM for temporal modeling (136).
 The presented taxonomy of VHGR methods highlights the diversity of approaches developed to address different aspects of gesture understanding—from isolated dynamic gestures to sign language recognition and multimodal human–computer interaction.
@@
 4.1 Sign Language Understanding (SLU)
-For SLU, the authors in (G-CNN-SHARMA2021115657) developed Gesture-CNN, a shallow 2DCNN to recognize Indian Sign Language (ISL) signs and slightly outperformed various versions of VGG on a self-collected dataset.
-The authors in (STN-ND-Net) proposed the STN-ND-Net, a combination of a Spatial Transformer Network (STN) that improves robustness to spatial transformations, with a 2DCNN backbone that uses deconvolution operations to reduce the pixel-wise and channel-wise correlations and eliminates the information redundancy of colored images.
-The success of vision transformer (ViT) motivated authors of SIGNFORMER (SIGNFORMER) to adopt it for ISL recognition, while authors in (Jungpil-Shin) leveraged both skeleton and RGB data in their dual stream feature extractor and an MLP to perform intermediate fusion.
-A YOLO variant, namely YOLOv4-CSP(YOLOv4-CSP) has also been developed, to simultaneous detect and identify Turkish Sign Language gestures.
+For SLU, the authors in (176) developed Gesture-CNN, a shallow 2DCNN to recognize Indian Sign Language (ISL) signs and slightly outperformed various versions of VGG on a self-collected dataset.
+The authors in (60) proposed the STN-ND-Net, a combination of a Spatial Transformer Network (STN) that improves robustness to spatial transformations, with a 2DCNN backbone that uses deconvolution operations to reduce the pixel-wise and channel-wise correlations and eliminates the information redundancy of colored images.
+The success of vision transformer (ViT) motivated authors of SIGNFORMER (105) to adopt it for ISL recognition, while authors in (180) leveraged both skeleton and RGB data in their dual stream feature extractor and an MLP to perform intermediate fusion.
+A YOLO variant, namely YOLOv4-CSP(8) has also been developed, to simultaneous detect and identify Turkish Sign Language gestures.
 4.2 Human-Computer Interaction (HCI)
-For gesture-based HCI the authors in (LHGR-Net) proposed a lightweight 2DCNN, named LHGR-Net, which combines a Multiscale Structure (MSS) that aims to make the model scale-invariant via atrous convolution with a Lightweight Attention Structure (LAS) based on CBAM.
-The authors in (2DPSTPP) proposed 2DPSTPP-Net, a 2DCNN with four stacked convolutional blocks whose features are pooled using max pooling and a novel method, then concatenated to form the final representation, tailored for HCI scenarios.
-The authors in (WEI_HAN_YOLOV5_COPULA) modified the YOLOv5 to make it more lightweight and developed a model that can be used for gesture-based user interfaces in smart sports stadiums, improving the experience of spectators and athletes.
-In (Mohammad_Mahmudul_static_hgr), a pipeline for egocentric HGR and fingertip localization uses YOLOv2 for hand detection, followed by VGG16 for feature extraction and gesture classification.
-In (tuan_linh_pham_dang2023lightweight), a lightweight HCI pipeline segments the hand using DeepLabV3+ or Unet, then classifies the cropped image with a 2DCNN like MobileNet or EfficientNet.
-(FGDSNet) introduced FGDSNet, a two-stage approach with dual-stream segmentation (local detail and semantic branches) and multi-level fusion for feature integration.
+For gesture-based HCI the authors in (218) proposed a lightweight 2DCNN, named LHGR-Net, which combines a Multiscale Structure (MSS) that aims to make the model scale-invariant via atrous convolution with a Lightweight Attention Structure (LAS) based on CBAM.
+The authors in (89) proposed 2DPSTPP-Net, a 2DCNN with four stacked convolutional blocks whose features are pooled using max pooling and a novel method, then concatenated to form the final representation, tailored for HCI scenarios.
+The authors in (67) modified the YOLOv5 to make it more lightweight and developed a model that can be used for gesture-based user interfaces in smart sports stadiums, improving the experience of spectators and athletes.
+In (9), a pipeline for egocentric HGR and fingertip localization uses YOLOv2 for hand detection, followed by VGG16 for feature extraction and gesture classification.
+In (36), a lightweight HCI pipeline segments the hand using DeepLabV3+ or Unet, then classifies the cropped image with a 2DCNN like MobileNet or EfficientNet.
+(229) introduced FGDSNet, a two-stage approach with dual-stream segmentation (local detail and semantic branches) and multi-level fusion for feature integration.
 The segmentation mask and original image are fed into a dual-stream classification module, where mid-level fusion combines features for final classification via an MLP.
@@
 HRI, particularly gesture-based mobile robot navigation, has attracted research interest.
-URGR (ULTRA-RANGE-UGV-Bamani) is a three-stage approach using YOLOv3 for human cropping, High-Quality Network (HQ-Net) for super-resolution to address long-range scenarios, and a classifier, Graph-Vision Transformer (GViT), for gesture recognition.
+URGR (18) is a three-stage approach using YOLOv3 for human cropping, High-Quality Network (HQ-Net) for super-resolution to address long-range scenarios, and a classifier, Graph-Vision Transformer (GViT), for gesture recognition.
 The HQ-Net comprises convolutional and deconvolutional pathways.
@@
 This setting enables effective recognition of gestures in a distance of up to 25m for UGV guidance.
-Similarly, (DRCAM_TIM) introduced a two-stage method based on depth data: hand segmentation followed by gesture classification using DRCAM, a 2DCNN with stacked Residual Channel Attention Modules (RCAMs) and dense interconnections between the cascaded blocks.
+Similarly, (171) introduced a two-stage method based on depth data: hand segmentation followed by gesture classification using DRCAM, a 2DCNN with stacked Residual Channel Attention Modules (RCAMs) and dense interconnections between the cascaded blocks.
 Because of the attention mechanisms, the model focuses effectively on the most informative regions of the image.
@@
 General-purpose models have also been developed.
-EDenseNet (EDenseNet) extends DenseNet by adding convolutional layers between dense blocks.
-SpAtNet (SpAtNet) is a lightweight 2DCNN with three streams: two Multi-scale Attentive Feature Fusion (MAFF) modules for coarse-grained features and one Interleaved Module for high-level representations; their fused features are fed to an MLP for classification.
+EDenseNet (189) extends DenseNet by adding convolutional layers between dense blocks.
+SpAtNet (22) is a lightweight 2DCNN with three streams: two Multi-scale Attentive Feature Fusion (MAFF) modules for coarse-grained features and one Interleaved Module for high-level representations; their fused features are fed to an MLP for classification.
 This approach, employing multiple kernel sizes, increases accuracy by capturing various levels of granularity.
-MMFANet (MMFANet) performs gesture detection and recognition using a 2DCNN backbone followed by the proposed feature aggregation pyramid network (FAPN) and task-specific regressor and classifier branches.
+MMFANet (114) performs gesture detection and recognition using a 2DCNN backbone followed by the proposed feature aggregation pyramid network (FAPN) and task-specific regressor and classifier branches.
 The FAPN module aims to extract multiscale features through upsampling/downsampling operations and fusion of adjacent feature maps, while reducing the computational complexity by utilizing novel expanded convolutional blocks with different dilation rates combined with attention mechanisms that eliminate redundant information.
@@
 Table 2 presents the accuracy of current state-of-the-art SHGR approaches.
-Despite their high accuracy and low parameter count—suggesting that SHGR might be a largely solved problem—the datasets used (ASL-FS (ASL-FS-2011-CVPRW), HGR1 (HGR-1-2024-PATTER-RECOGNITION), NUS-I (NUS-I-2010), NUS-II (NUS-II-2013), and OUHANDS (ouhands)) remain relatively small and lack diversity.
+Despite their high accuracy and low parameter count—suggesting that SHGR might be a largely solved problem—the datasets used (ASL-FS (154), HGR1 (98), NUS-I (106), NUS-II (153), and OUHANDS (132)) remain relatively small and lack diversity.
 Consequently, model performance would likely degrade in real-world, uncontrolled environments.
-Among the existing methods, LHGR-Net (LHGR-Net) and FGDSNet (FGDSNet) demonstrate strong results on the OUHANDS dataset with lightweight architectures, making them suitable for gesture-based human–computer interaction.
-On the other hand, EDenseNet (EDenseNet) achieves high accuracy on ASL-FS but at a higher computational cost, making it better suited for static sign language recognition.
+Among the existing methods, LHGR-Net (218) and FGDSNet (229) demonstrate strong results on the OUHANDS dataset with lightweight architectures, making them suitable for gesture-based human–computer interaction.
+On the other hand, EDenseNet (189) achieves high accuracy on ASL-FS but at a higher computational cost, making it better suited for static sign language recognition.
 Finally, applying background removal through hand segmentation does not appear to significantly improve accuracy compared to end-to-end learning approaches.
@@
 Method Year Modality Params ASL-FS HGR1 NUS-I NUS-II OUHANDS
-FGDSNet (FGDSNet) * 2024 RGB 1.20M - 94.80% - 99.80% 97.80%
-SpAtNet (SpAtNet) 2024 RGB 3.8M - 93.33% - - -
-(tuan_linh_pham_dang2023lightweight) * 2023 RGB 5.34M - 98.15% - - 89.20%
-LHGR-Net (LHGR-Net) 2023 RGB 1.81M - 93.15% - - 98.57%
-2DPSTPP-Net (2DPSTPP) 2023 RGB 19.01M - - 100% 99.61% -
-DRCAM (DRCAM_TIM) * 2023 Depth - 93.44% - - - -
-EDenseNet (EDenseNet) 2021 RGB 7.77M 99.99% - - 99.30% -
+FGDSNet (229) * 2024 RGB 1.20M - 94.80% - 99.80% 97.80%
+SpAtNet (22) 2024 RGB 3.8M - 93.33% - - -
+(36) * 2023 RGB 5.34M - 98.15% - - 89.20%
+LHGR-Net (218) 2023 RGB 1.81M - 93.15% - - 98.57%
+2DPSTPP-Net (89) 2023 RGB 19.01M - - 100% 99.61% -
+DRCAM (171) * 2023 Depth - 93.44% - - - -
+EDenseNet (189) 2021 RGB 7.77M 99.99% - - 99.30% -
 5 Methods for Isolated Dynamic VHGR
@@
 Spatiotemporal backbones perform spatial and temporal modeling jointly, typically using 3DCNNs or (2+1)D CNNs—either custom-built or adapted from existing models.
-In (2plus1-D-SLR_and_NCSL), the authors introduced a custom network, (2+1)D-SLR, composed of five stacked (2+1)D convolutional blocks (i.e., a 2D convolution followed by a 1D convolution).
+In (200), the authors introduced a custom network, (2+1)D-SLR, composed of five stacked (2+1)D convolutional blocks (i.e., a 2D convolution followed by a 1D convolution).
 This structure lowers computational cost compared to traditional 3D convolutions, making the model more practical for deployment.
-To enhance temporal modeling, authors in (Snehal_Bharti_et_al) proposed a trainable key frame selection method based on error correction.
+To enhance temporal modeling, authors in (21) proposed a trainable key frame selection method based on error correction.
 VGG16 features are used to identify key frames, which are then preprocessed and fed into a custom 3DCNN for classification.
-In (ShiqiWang_et_al), a dual-stage multi-stream pipeline for CSL recognition was developed.
+In (203), a dual-stage multi-stream pipeline for CSL recognition was developed.
 An adapted EfficientDet model—enhanced with cross-level Bi-FPN connections and a dual-channel spatial attention module (DCSAM)—detects both hands.
 The full image and cropped hand patches are passed through three parallel 3D-ResNet-18 models (following the P3D scheme), with their features fused and classified via an MLP.
-Authors in (TS3C-Net) introduced TS3C-Net, a modified 3D ConvNeXt that adds an extra 3×3×3 convolutional branch to each block, enhancing spatiotemporal representation.
-Finally, (SeST) addressed limitations in fixed-pipeline designs by combining features from both ResC3D and ConvLSTM using the Selective Spatiotemporal (SeST) module.
+Authors in (70) introduced TS3C-Net, a modified 3D ConvNeXt that adds an extra 3×3×3 convolutional branch to each block, enhancing spatiotemporal representation.
+Finally, (191) addressed limitations in fixed-pipeline designs by combining features from both ResC3D and ConvLSTM using the Selective Spatiotemporal (SeST) module.
 The fused features are passed through fully connected and softmax layers for classification.
@@
 The resulting sequence embeddings are aggregated, and a classifier (e.g., MLP) outputs the final prediction.
-In (DAVID_RICHARD_TOM_HAX), a method tailored for HCI tasks using only depth data was proposed.
+In (72), a method tailored for HCI tasks using only depth data was proposed.
 The method combined Inception-v3 that extracted spatial features with an LSTM for temporal modeling.
-For ISLR, authors in (GoogleNet-BiLSTM) introduced a pipeline to recognize Indian Sign Language (ISL) words used by deaf farmers.
+For ISLR, authors in (196) introduced a pipeline to recognize Indian Sign Language (ISL) words used by deaf farmers.
 They combined GoogleNet for spatial encoding with a BiLSTM for sequence modeling.
-To reduce inference time, a lightweight ISLR model, called TIM-SLR has been proposed in (TIM-SLR).
+To reduce inference time, a lightweight ISLR model, called TIM-SLR has been proposed in (201).
 The model uses a 2DCNN for per-frame feature extraction, followed by a Time Interaction Module (TIM)—a zero-parameter unit that fuses frame features into a compact representation.
-Finally, the Swin Transformer presented in (Full_Transformer_Network) is a two-stage transformer-based model that extracts spatial features per frame, while a causally masked transformer processes the sequence, ensuring that only past and present frame features contribute at each step.
+Finally, the Swin Transformer presented in (43) is a two-stage transformer-based model that extracts spatial features per frame, while a causally masked transformer processes the sequence, ensuring that only past and present frame features contribute at each step.
 5.1.3 Short-Term Spatiotemporal + Long-Term Temporal
 The methods in this category follow a two-stage architecture: the first stage captures short-term spatiotemporal patterns from segments of the input sequence, while the second stage aggregates these features to form a final representation of the entire sequence.
-EgoFormer (EgoFormer_METHOD_EgoDriving_DATASET) is a model for egocentric dynamic gesture recognition optimized for edge devices, which has been introduced in the context of autonomous driving.
+EgoFormer (155) is a model for egocentric dynamic gesture recognition optimized for edge devices, which has been introduced in the context of autonomous driving.
 EgoFormer combines a custom 3DCNN with a transformer module and an MLP for final classification.
-For long-range gesture-based UGV control, authors in (SlowFast-Transformer) proposed a pipeline where a lightweight 2DCNN extracts per-frame features, which are used to select key frames.
+For long-range gesture-based UGV control, authors in (19) proposed a pipeline where a lightweight 2DCNN extracts per-frame features, which are used to select key frames.
 After cropping the human region, the selected frames are passed through a SlowFast network, followed by a transformer and an MLP.
-To address data scarcity in sign language understanding, (ZERO-SHOT-SLR) proposed a zero-shot learning framework.
+To address data scarcity in sign language understanding, (23) proposed a zero-shot learning framework.
 Sign videos are represented using BERT-encoded textual descriptions combined with hand-crafted attributes.
@@
 Some appearance-based methods fall outside the previously discussed categories but offer valuable contributions, especially in Sign Language Recogition (SLR).
-For example, the method proposed in (HAMZAH_LUQMAN) for Indian SLR (ISLR) compresses an entire frame sequence into a single image, which is then processed by a MobileNet.
+For example, the method proposed in (126) for Indian SLR (ISLR) compresses an entire frame sequence into a single image, which is then processed by a MobileNet.
 A second MobileNet extracts per-frame features, which are passed through a two-layer LSTM.
 The outputs of both streams are concatenated and refined using a custom shallow 2DCNN.
-A different approach was presented in (Ensemble-InceptionV4), where the authors introduced a dual-stream ISLR architecture combining Inception V4 with a novel convolutional module called stem.
-In the context of pose-based ISLR, (Tianyu_Liu) proposed a dual-stream model with keyframe selection.
+A different approach was presented in (104), where the authors introduced a dual-stream ISLR architecture combining Inception V4 with a novel convolutional module called stem.
+In the context of pose-based ISLR, (123) proposed a dual-stream model with keyframe selection.
 A 2DCNN and attention mechanism filter out less informative frames.
@@
 The fused features are used for final classification.
-To improve signer-independence and reduce overfitting, the SwC GR-MMixer (SwC-GR-MMixer), another lightweight model based on MLP-Mixer, incorporates GRU units for temporal modeling and a shifted window concatenation strategy to enhance sequence representation.
+To improve signer-independence and reduce overfitting, the SwC GR-MMixer (122), another lightweight model based on MLP-Mixer, incorporates GRU units for temporal modeling and a shifted window concatenation strategy to enhance sequence representation.
 5.2 Unimodal Pose-based Methods
 Pose-based methods, also known as skeleton-based approaches, rely on structured motion data—such as skeleton graphs and pose heatmaps—to classify gestures.
-These methods often incorporate various forms of pose information, including joint coordinates, bones, motion trajectories, and joint angles, sometimes in combination (TMS-NET).
-They also explore joint subsets related to specific body regions (FDMSE-hierarchicalwindowedgraphattention), often adopting multi-stream architectures to handle these diverse inputs.
+These methods often incorporate various forms of pose information, including joint coordinates, bones, motion trajectories, and joint angles, sometimes in combination (40).
+They also explore joint subsets related to specific body regions (152), often adopting multi-stream architectures to handle these diverse inputs.
 Given the natural graph structure of skeletal data, many methods employ Graph Convolutional Networks (GCNs) originally developed for human activity recognition.
 In the sections that follow, we group the approaches based on the structure of their backbone, following the same organization as in the previous subsection.
-Notably, pose information alone is often insufficient for accurate sign recognition (moryossef2021evaluating).
+Notably, pose information alone is often insufficient for accurate sign recognition (139).
 Consequently, many models integrate additional cues—typically from RGB inputs—using 3DCNNs and perform score-level fusion to enhance performance.
@@
 In SLU, pose-based methods model dynamic hand and body movements for gesture recognition.
-SKIM (SKIM) combines dynamic graph convolutional (GCNs) and temporal convolutional (TCNs) networks to analyze global and local motion, while MC-LSTM (ozdemir2023multi_MC-LSTM) fuses multiple streams (face, hand, body) via temporal convolutions and GCNs.
-P3D (P3D_ICCV) employs transformer-based part-wise and whole-body encoders, and HWGAT (FDMSE-hierarchicalwindowedgraphattention) enhances key body-part focus with hierarchical attention.
-TMS-Net (TMS-NET) introduces a six-stream GCN for skeleton, bone, and motion data with inter-stream fusion.
-To tackle data limitations, SLGTformer (SLGTFORMER) and VSCT (spoter_vsct) use transformers and few-shot learning, respectively.
-Dual-stream approaches such as (Jungpil-Shin) and MULTISTAGE-CGAR (MULTISTAGE-CGAR) integrate joint and motion data through temporal convolutions and attention.
-SwC GR-MMixer (SwC-GR-MMixer) combines GCNs with GRUs for temporal modeling and regularization.
-Other notable advances include L-TSSI-DenseNet (L-TSSI-DenseNet), which converts joint coordinates into image-like matrices for DenseNet processing; STST (STST), a two-stage transformer for local–global spatiotemporal features; and DSTA-SLR (DSTA_SLR_hu2024dynamic), which employs dynamic adjacency matrices and multi-stream fusion for improved recognition accuracy.
+SKIM (115) combines dynamic graph convolutional (GCNs) and temporal convolutional (TCNs) networks to analyze global and local motion, while MC-LSTM (148) fuses multiple streams (face, hand, body) via temporal convolutions and GCNs.
+P3D (109) employs transformer-based part-wise and whole-body encoders, and HWGAT (152) enhances key body-part focus with hierarchical attention.
+TMS-Net (40) introduces a six-stream GCN for skeleton, bone, and motion data with inter-stream fusion.
+To tackle data limitations, SLGTformer (187) and VSCT (26) use transformers and few-shot learning, respectively.
+Dual-stream approaches such as (180) and MULTISTAGE-CGAR (134) integrate joint and motion data through temporal convolutions and attention.
+SwC GR-MMixer (122) combines GCNs with GRUs for temporal modeling and regularization.
+Other notable advances include L-TSSI-DenseNet (108), which converts joint coordinates into image-like matrices for DenseNet processing; STST (222), a two-stage transformer for local–global spatiotemporal features; and DSTA-SLR (86), which employs dynamic adjacency matrices and multi-stream fusion for improved recognition accuracy.
 In HCI, pose-based methods address challenges such as data overfitting and the need for self-supervised learning frameworks.
-MAE-based frameworks (omar_ikne) use a masked autoencoder to reconstruct skeleton graphs, focusing on motion and spatial relationships.
-3sISTGCN (JOINTS_MOTION_VECTORS_Improved_ST-GCN) proposes a three-stream model to process joints, bones, and motions, applying modified ST-GCN layers and late fusion for improved classification.
-Another approach, DG-STA (BOAT-MI_VR), tackles class-incremental learning in HCI by training models on non-overlapping tasks without access to old training data.
-DSTSA-GCN (DSTSA-CGN_cui2025dstsa) addresses temporal expression limitations by stacking three stages of graph convolutions and multiscale temporal convolutions.
-Finally, authors in (ST-CGNet-hu2025st) proposed ST-CGNet, a dual-stream architecture comprising a 3DCNN and a convolutional LSTM module to capture different levels of information.
+MAE-based frameworks (88) use a masked autoencoder to reconstruct skeleton graphs, focusing on motion and spatial relationships.
+3sISTGCN (186) proposes a three-stream model to process joints, bones, and motions, applying modified ST-GCN layers and late fusion for improved classification.
+Another approach, DG-STA (5), tackles class-incremental learning in HCI by training models on non-overlapping tasks without access to old training data.
+DSTSA-GCN (32) addresses temporal expression limitations by stacking three stages of graph convolutions and multiscale temporal convolutions.
+Finally, authors in (81) proposed ST-CGNet, a dual-stream architecture comprising a 3DCNN and a convolutional LSTM module to capture different levels of information.
 Attention mechanisms and an advanced fusion strategy are utilized to efficiently incorporate the features of each stream.
@@
 These features are then passed into a sequential model for temporal aggregation, resulting in a combined spatial-temporal representation of the input sequence.
-For instance, GCN-BERT (GCN-BERT) utilizes a two-layer GCN for spatial modeling, followed by BERT for temporal modeling.
+For instance, GCN-BERT (194) utilizes a two-layer GCN for spatial modeling, followed by BERT for temporal modeling.
 Other general-purpose methods adopt a multi-stream, two-stage pose-based approach.
-In (SBI-DHGR), three streams, each containing three 1D convolutional layers, are combined with a parallel module of average pooling layers.
+In (143), three streams, each containing three 1D convolutional layers, are combined with a parallel module of average pooling layers.
 The extracted features are then concatenated and input into an LSTM, followed by a fully connected layer and softmax for classification.
-Similarly, STr-GCN (STr-GCN) employs a two-stage model where a GCN spatial feature extractor is first applied to obtain intra-frame features and then passes the features through a transformer encoder for modeling inter-frame temporal relationships.
+Similarly, STr-GCN (184) employs a two-stage model where a GCN spatial feature extractor is first applied to obtain intra-frame features and then passes the features through a transformer encoder for modeling inter-frame temporal relationships.
 The final classification scores are generated by an MLP.
@@
 For example, gesture recognition in autonomous driving has become an important research area.
-In (MG-GCT), the authors propose a pose-based approach called MG-GCT for recognizing police traffic gestures.
+In (65), the authors propose a pose-based approach called MG-GCT for recognizing police traffic gestures.
 The model uses two graph convolution streams: one to process skeleton joints and another to analyze joint motions.
 The features extracted from both streams are fused and passed through a transformer encoder for temporal modeling, with the final classification performed by an MLP.
-In another approach, TD-GCN (TD-GCN) introduces a pose-based method for IDHGR.
+In another approach, TD-GCN (120) introduces a pose-based method for IDHGR.
 Unlike traditional methods that use fixed skeleton topologies, TD-GCN employs temporal-sensitive topology learning to enhance feature quality.
@@
 These matrices help model the skeleton, and the features are fused using the proposed TCF module.
-Authors in (ISLR-STGCN-attention-LSTM-ozdemir2025cross) proposed a dual-stream backbone processing both hand and whole body keypoints using ST-GCN models.
+Authors in (149) proposed a dual-stream backbone processing both hand and whole body keypoints using ST-GCN models.
 Cross-attention is then applied to focus on the most important features, which are subsequently pass through a LSTM to model long-term dependencies.
@@
 Typically, keypoints are slightly preprocessed—often by extracting geometric features such as bone angles—and then passed into sequential models like RNNs or TCNs to produce a vector representation of the input sequence.
-In (TELEOPERATED_SURGICAL_ROBOT), the authors proposed a pose-based method for surgical robot teleoperation.
+In (156), the authors proposed a pose-based method for surgical robot teleoperation.
 Preprocessed skeleton joints are fed into an LSTM, and the resulting features are passed to an MLP for final classification.
 The model was evaluated on a custom dataset.
-Similarly, in (Sabater_alberto), a cross-domain gesture recognition approach was developed using geometric features (e.g., bone angles), which are processed by a TCN.
+Similarly, in (170), a cross-domain gesture recognition approach was developed using geometric features (e.g., bone angles), which are processed by a TCN.
 The resulting features are summarized into a sequence descriptor, and a linear classifier predicts the final class.
-To address data scarcity, the authors in (ADVERSARIAL_DATA_AUGM) introduced AVSN, an adversarial learning-based data augmentation framework.
+To address data scarcity, the authors in (142) introduced AVSN, an adversarial learning-based data augmentation framework.
 The method jointly optimizes augmentation parameters and the HGR classifier.
@@
 Several works explore alternative architectures or skeleton representations beyond the main categories.
-In (FS-EFCNN), a Fourier CNN with BiLSTM extracts spatiotemporal features from 3D joints for ISLR, using a trainable binary mask for feature selection.
-Similarly, (SCT-DCNN) applies feature thresholding based on variance and importance, followed by deep CNN classification.
-To overcome color-coded map limitations, (MRCNNSA) introduces the Joint Motion Affinity Map (JMAM) with a multi-stream attention-based 2DCNN for feature extraction and recognition.
-Key frame selection is addressed in (KEYFRAMES_DRL_SKELETON) via FS-Net, a reinforcement learning agent that selects informative frames for DD-Net classification.
-Dual-stream architectures such as STGCN-LSTM (STGCN-LSTM-10965662) combine ST-GCN and self-attention ConvLSTM for capturing global and local features, while (Yangke_patcog) proposes a lightweight variant merging attention-based and autoencoder streams.
-Finally, the Hierarchical Self-Attention Network (HAN) (HAN-liu2025han) employs modular attention across finger joints (J-Att, F-Att, T-Att, Fusion-Att) for efficient spatial–temporal modeling, achieving strong accuracy with just 0.05 GFLOPs.
+In (1), a Fourier CNN with BiLSTM extracts spatiotemporal features from 3D joints for ISLR, using a trainable binary mask for feature selection.
+Similarly, (10) applies feature thresholding based on variance and importance, followed by deep CNN classification.
+To overcome color-coded map limitations, (100) introduces the Joint Motion Affinity Map (JMAM) with a multi-stream attention-based 2DCNN for feature extraction and recognition.
+Key frame selection is addressed in (51) via FS-Net, a reinforcement learning agent that selects informative frames for DD-Net classification.
+Dual-stream architectures such as STGCN-LSTM (68) combine ST-GCN and self-attention ConvLSTM for capturing global and local features, while (111) proposes a lightweight variant merging attention-based and autoencoder streams.
+Finally, the Hierarchical Self-Attention Network (HAN) (119) employs modular attention across finger joints (J-Att, F-Att, T-Att, Fusion-Att) for efficient spatial–temporal modeling, achieving strong accuracy with just 0.05 GFLOPs.
 5.3 Multimodal Methods
@@
 5.3.1 General Purpose
-In (Yunan_Li), a dual-stream 3DCNN with average fusion was proposed for RGBD gesture recognition, using a memory bank to retain discriminative features ($z$) while filtering redundant ones ($\upsilon$) by maximizing mutual information with class labels.
-Similarly, (RAAR3DNet) employed Neural Architecture Search (NAS) with Dynamic Static Attention (DSA) modules and a modified I3D backbone to emphasize motion and hand regions, supported by an auxiliary HeatmapNet and late RGBD fusion.
-To better exploit modality interactions, (Benjia_Zhou) introduced a decoupling–recoupling framework combining independent spatial and temporal modeling with mid-level fusion through a shared MLP-AE module.
-Multi-level fusion was further explored in (CTFB_Hampiholi), integrating 3DCNNs with a Convolutional Transformer Fusion Block (CTFB) that fuses intermediate cross-stream features.
+In (112), a dual-stream 3DCNN with average fusion was proposed for RGBD gesture recognition, using a memory bank to retain discriminative features ($z$) while filtering redundant ones ($\upsilon$) by maximizing mutual information with class labels.
+Similarly, (227) employed Neural Architecture Search (NAS) with Dynamic Static Attention (DSA) modules and a modified I3D backbone to emphasize motion and hand regions, supported by an auxiliary HeatmapNet and late RGBD fusion.
+To better exploit modality interactions, (228) introduced a decoupling–recoupling framework combining independent spatial and temporal modeling with mid-level fusion through a shared MLP-AE module.
+Multi-level fusion was further explored in (66), integrating 3DCNNs with a Convolutional Transformer Fusion Block (CTFB) that fuses intermediate cross-stream features.
 To address data scarcity, several works explored zero-shot learning (ZSL) by incorporating textual gesture descriptions encoded by pretrained language models (mainly BERT).
-In (ZS-GR), a dual-stream model combined a ViT for spatial modeling and an LSTM-MLP for temporal modeling, learning visual features aligned with textual embeddings.
-Similarly, (ZS-DHGR) fused skeleton-derived geometric features with RGB-C3D-LSTM-AE representations into a shared semantic space for classification.
-To reduce computational cost, (ConvMixFormer) replaced the transformer’s self-attention mechanism with convolutional layers and a gated feedforward network, resulting in the lightweight ConvMixFormer architecture.
-Knowledge distillation and contrastive learning were applied in (3STREAM_DISTILLATION), where three 3DCNNs (RGB, Depth, Optical Flow) were fused with Coordinate Attention, using I3D as teacher and ResNet as student, while integrating CLIP for zero-shot recognition.
-Finally, (MFST) and (3D_CDC_NAS_TIP_2021) replaced standard convolutions with 3D Central Difference Convolution (3D-CDC), combining hybrid residual blocks, transformers, and NAS-based optimization for enhanced spatiotemporal modeling.
+In (164), a dual-stream model combined a ViT for spatial modeling and an LSTM-MLP for temporal modeling, learning visual features aligned with textual embeddings.
+Similarly, (162) fused skeleton-derived geometric features with RGB-C3D-LSTM-AE representations into a shared semantic space for classification.
+To reduce computational cost, (58) replaced the transformer’s self-attention mechanism with convolutional layers and a gated feedforward network, resulting in the lightweight ConvMixFormer architecture.
+Knowledge distillation and contrastive learning were applied in (198), where three 3DCNNs (RGB, Depth, Optical Flow) were fused with Coordinate Attention, using I3D as teacher and ResNet as student, while integrating CLIP for zero-shot recognition.
+Finally, (127) and (215) replaced standard convolutions with 3D Central Difference Convolution (3D-CDC), combining hybrid residual blocks, transformers, and NAS-based optimization for enhanced spatiotemporal modeling.
 5.3.2 Sign Language Understanding
 Recent works on Isolated Sign Language Recognition (ISLR) have focused on multimodal, pose-centric architectures combining RGB and skeleton data.
-SignBERT (SignBERT_ICCV) employs a spatial GCN with temporal and hand chirality encoding, feeding features into a transformer encoder and a hand-model-aware decoder for self-supervised pretraining, with parallel RGB predictions fused at inference.
-Its enhanced variant, SignBERT+ (SignBERT+), adds spatial-temporal position encoding and an improved RGB–pose fusion module (Fig.
+SignBERT (78) employs a spatial GCN with temporal and hand chirality encoding, feeding features into a transformer encoder and a hand-model-aware decoder for self-supervised pretraining, with parallel RGB predictions fused at inference.
+Its enhanced variant, SignBERT+ (77), adds spatial-temporal position encoding and an improved RGB–pose fusion module (Fig.
 6).
-Similarly, MASA (MASA) uses the same GCN+Transformer backbone with contrastive learning on pose reconstruction, while BEST (BEST) embeds pose triplets with BERT and applies late fusion with RGB data.
-Figure 6: The pose-based stream of SignBERT+ (SignBERT+), an extension of SignBERT (SignBERT_ICCV), first uses a GCN to extract gesture features, which are combined with spatial-temporal position encoding and processed by a transformer encoder.
+Similarly, MASA (223) uses the same GCN+Transformer backbone with contrastive learning on pose reconstruction, while BEST (224) embeds pose triplets with BERT and applies late fusion with RGB data.
+Figure 6: The pose-based stream of SignBERT+ (77), an extension of SignBERT (78), first uses a GCN to extract gesture features, which are combined with spatial-temporal position encoding and processed by a transformer encoder.
 During self-supervised pretraining, a MANO-based decoder reconstructs masked inputs, and is replaced by a prediction head for downstream fine-tuning.
-To improve feature invariance, (Weichao-Zhao) applies a contrastive approach enforcing consistency among hand, trunk, and motion features, using the same GCN+Transformer backbone with late fusion of pose and RGB predictions.
-SAM-SLR (SAM_SLR) fuses six streams (RGB, depth, skeleton, optical flow, pose heatmaps, HHA), while SAM-SLR-v2 (jiang2021sign-sam-slr-v2) adds a 3D hand graph stream and learnable fusion weights.
-VTN-HC/PF (VTN-HC-PF) processes hand-cropped images via a 2DCNN, transformer, and MLP for final predictions, with PF incorporating pose-based geometrical features.
-Motion History Images are used in (MOTION_HISTORY_Mercanoglu), analyzed by a 3DCNN with attention, and in a dual-stream setup combined with 2DCNN RGB features via weighted summation.
-To tackle data scarcity, ZSSLR (ZSSLR) uses a CLIP-based zero-shot framework, while (FSSLR) proposes a few-shot cross-lingual method combining two 3DCNNs and a pose estimator with a Temporal Feature Aggregation Module.
-NLA-SLR (NLA_SLR) targets semantically similar signs with a multistream S3D backbone and language-aware label smoothing.
-MSNN (MSNN) uses five I3D streams plus ST-GCN, and hDNN-SLR (hDNN-SLR) combines a 3DCNN, attention-based BiLSTM, and a VAE-based module.
-Finally, Uni-Sign (UNISIGNli2025) proposes a unified multimodal architecture with three pose encoders and an RGB encoder fused via attention, followed by temporal encoding and LLM-based pretraining for ISLR and CSLR.
+To improve feature invariance, (225) applies a contrastive approach enforcing consistency among hand, trunk, and motion features, using the same GCN+Transformer backbone with late fusion of pose and RGB predictions.
+SAM-SLR (91) fuses six streams (RGB, depth, skeleton, optical flow, pose heatmaps, HHA), while SAM-SLR-v2 (90) adds a 3D hand graph stream and learnable fusion weights.
+VTN-HC/PF (38) processes hand-cropped images via a 2DCNN, transformer, and MLP for final predictions, with PF incorporating pose-based geometrical features.
+Motion History Images are used in (133), analyzed by a 3DCNN with attention, and in a dual-stream setup combined with 2DCNN RGB features via weighted summation.
+To tackle data scarcity, ZSSLR (147) uses a CLIP-based zero-shot framework, while (25) proposes a few-shot cross-lingual method combining two 3DCNNs and a pose estimator with a Temporal Feature Aggregation Module.
+NLA-SLR (235) targets semantically similar signs with a multistream S3D backbone and language-aware label smoothing.
+MSNN (130) uses five I3D streams plus ST-GCN, and hDNN-SLR (159) combines a 3DCNN, attention-based BiLSTM, and a VAE-based module.
+Finally, Uni-Sign (113) proposes a unified multimodal architecture with three pose encoders and an RGB encoder fused via attention, followed by temporal encoding and LLM-based pretraining for ISLR and CSLR.
 5.3.3 Human-Computer Interaction
-Motivated by the difficulty of recognizing fine-grained gestures in HCI applications, authors in (HandSense) proposed HandSense, a method relying on depth and RGB data.
+Motivated by the difficulty of recognizing fine-grained gestures in HCI applications, authors in (221) proposed HandSense, a method relying on depth and RGB data.
 Two 3DCNNs are deployed to extract features for each one modality.
 These features are fused and fed into a SVM to obtain the final prediction.
-Authors in (mXception_and_RDDI) proposed an approach relying on RGBD Dynamic Images (RDDIs).
+Authors in (96) proposed an approach relying on RGBD Dynamic Images (RDDIs).
 These images, are obtained by the combination of RGB and estimated depth data.
@@
 5.3.4 Other
-Inspired by the success of transformers, GestFormer (GestFormer) introduces a lightweight transformer-based framework for gesture-driven HRI.
+Inspired by the success of transformers, GestFormer (56) introduces a lightweight transformer-based framework for gesture-driven HRI.
 A 2DCNN extracts per-frame features, processed by PoolFormer layers (replacing attention) and two custom blocks for multiscale pooling/WT and token mixing.
 Each modality is processed independently, and outputs are averaged.
-In (DHGR_3D_POSE_HRI), a hybrid 3DCNN/LSTM model fuses RGB, depth, and skeleton data (from an enhanced pose estimator) via weighted sum.
+In (54), a hybrid 3DCNN/LSTM model fuses RGB, depth, and skeleton data (from an enhanced pose estimator) via weighted sum.
 The fused input feeds a spatiotemporal model combining 2D/3D convolutions and ConvLSTM layers.
-Authors in (MDSI-fang2025mdsi) emphasize on two major issues, information redundancy and the difficulty to recognize similar gestures.
+Authors in (48) emphasize on two major issues, information redundancy and the difficulty to recognize similar gestures.
 Their approach leverages a CLIP encoder to provide qualitative embeddings and enhance the feature extractor through label smoothing as well as a semantic filter to incorporate multimodal data.
 5.4 Performance of IDHGR Methods
-The performance of IDHGR methods in the most used benchmarks (IsoGD (chalearn-lap), SHREC (SHREC), DHG (DHG-CVPRW), NVIDIA Hand Gesture or nvGesture (NVIDIA_GESTURE_R3DCNN_Molchanov), MSASL (MS-ASL), WLASL (WLASL), CSL500 (CSL500) and AUTSL (AUTSL)) is shown Table 3.
-For ISLR, Uni-Sign (UNISIGNli2025) surpasses relevant methods in terms of recognition accuracy, but requires powerful hardware to be executed fast, due to its large size (592.1M parameters, most of them contained on the language model).
-Thus, smaller but less accurate models like NLA-SLR (NLA_SLR), SignBERT+ (SignBERT_ICCV) and (Weichao-Zhao) are more reasonable solutions for real-world applications.
-Even though the multimodal SAM-SLR (SAM_SLR) and SAM-SLR-v2 (jiang2021sign-sam-slr-v2) are very accurate models for ISLR, they require a large number of modalities that are not fused effectively (using late fusion), which may be suboptimal.
-Pose-based methods for HCI and HRI including TD-GCN (TD-GCN), STST (STST) and DSTSA-GCN (DSTSA-CGN_cui2025dstsa) achieve high recognition rates with relatively low computational cost.
+The performance of IDHGR methods in the most used benchmarks (IsoGD (197), SHREC (185), DHG (39), NVIDIA Hand Gesture or nvGesture (138), MSASL (94), WLASL (110), CSL500 (87) and AUTSL (183)) is shown Table 3.
+For ISLR, Uni-Sign (113) surpasses relevant methods in terms of recognition accuracy, but requires powerful hardware to be executed fast, due to its large size (592.1M parameters, most of them contained on the language model).
+Thus, smaller but less accurate models like NLA-SLR (235), SignBERT+ (78) and (225) are more reasonable solutions for real-world applications.
+Even though the multimodal SAM-SLR (91) and SAM-SLR-v2 (90) are very accurate models for ISLR, they require a large number of modalities that are not fused effectively (using late fusion), which may be suboptimal.
+Pose-based methods for HCI and HRI including TD-GCN (120), STST (222) and DSTSA-GCN (32) achieve high recognition rates with relatively low computational cost.
 Although, the computational complexity of pose estimation and the effect of its incorrect predictions in more challenging environments (e.g.
 fast movements) may reduce their overall performance.
-Appearance-based methods are not affected by off-the-shelf pose estimators but they comprise many more parameters (up to 251.4M (3D_CDC_NAS_TIP_2021)).
+Appearance-based methods are not affected by off-the-shelf pose estimators but they comprise many more parameters (up to 251.4M (215)).
 Table 3: Performance of IDHGR methods on various datasets.
@@
 Method Year Params IsoGD 14 28 14 28 nvGesture PI PC PI PC CSL500 AUTSL
-STGCN-LSTM (STGCN-LSTM-10965662) 2025 - - - - - - - - - - - 95.2% -
-(ISLR-STGCN-attention-LSTM-ozdemir2025cross) 2025 - - - - - - - - - - - - 86.43%
-HAN-2S (HAN-liu2025han) 2025 0.94M - 95% 92.86% 92.71% 89.15% - - - - - - -
-DSTSA-GCN (DSTSA-CGN_cui2025dstsa) 2025 8.08M - 97.74% 95.37% 95.04% 93.57% - - - - - - -
-Uni-Sign (UNISIGNli2025) 2025 592.1M - - - - - - 78.16% 76.97% 63.52% 61.32% - -
-TMS-Net (TMS-NET) 2024 - - - - - - - - - 56.4% - - -
-MSNN (MSNN) 2024 - - - - - - - 50.59% - 47.26% - - -
-SKIM (SKIM) 2024 - - - - - - - - - 57.73% 55.40% - -
-MASA (MASA) 2024 - - - - - - - - - 55.77% 53.13% - -
-DSTA-SLR (DSTA_SLR_hu2024dynamic) 2024 7.4M - - - - - - 65.74% 62.31% 53.68% 51.17% 98.1%5 -
-(Weichao-Zhao) 2024 - - - - - - - 74.18% 72.18% 58.06% 55.66% 97.8% -
-TD-GCN (TD-GCN) 2024 1.36M - 97.02% 95.36% 93.9% 91.4% - - - - - - -
-STST (STST) 2024 - - 97.62% 95.83% 94.82% 93.18% - - - - - - -
-(Tianyu_Liu) 2024 - - - - - - - - - - - 98.87% 92.53%
-SwC GR-MMixer (SwC-GR-MMixer) 2024 - - - - - - - - - - - 98.54% -
-HWGAT (FDMSE-hierarchicalwindowedgraphattention) 2024 - - - - - - - - - 48.49% - - 95.80%
-(Yangke_patcog) 2024 0.21M - 96.9% 94.17% 94.21% 92.11% - - - - - - -
-(omar_ikne) 2024 - - 94.1% 90.0% - - - - - - - - -
-BEST (BEST) 2023 - - - - - - - 71.21% 68.24% 54.59% 52.12% 97.7% -
-SignBERT+ (SignBERT+) 2023 - - - - - - - 73.71% 70.77% 55.59% 53.33% 97.8% -
-NLA-SLR (NLA_SLR) 2023 - - - - - - - 73.80% 70.95% 61.26% 58.31% - -
-3SCKI (3STREAM_DISTILLATION) 2023 - 65.87% - - - - - - - - - - -
-(Yunan_Li) 2023 - 74.08% - - - - - - - - - - -
-CTFB (CTFB_Hampiholi) 2023 27.2M-37.1M 69.02% - - - - 87.76% - - - - - -
-L-TSSI-DenseNet (L-TSSI-DenseNet) 2023 7.2M - - - - - - - - - - - 93.13%
-TS3C-Net (TS3C-Net) 2023 28.7M - - - - - 87.55% - - - - - -
-(Musa_IEEE_ACCESS) 2023 - - 97.01% 92.78% 92.00% 88.78% - - - - - - -
-SBI-DHGR (SBI-DHGR) 2023 28M - 97% 92.36% 94.64% 91.79% - - - - - - -
-STr-GCN (STr-GCN) 2023 - - 93.39% 89.20% - - - - - - - - -
-DRL-DDNet (KEYFRAMES_DRL_SKELETON) 2023 - - 96.5% 93.8% - - - - - - - - -
-De+Recouple (Benjia_Zhou) 2022 - 66.79% - - - - 91.70% - - - - - -
-MS-ISTGCN (JOINTS_MOTION_VECTORS_Improved_ST-GCN) 2022 7.1M - 96.7% 94.9% 93.7% 91.2% - - - - - - -
-SignBERT (SignBERT_ICCV) 2021 - - - - - - - 71.24% 67.96% 54.69% 52.08% 97.6% -
-SlowFast+NAS+CDC (3D_CDC_NAS_TIP_2021) 2021 251.4M 66.23% - - - - - - - - - - -
-RAAR3DNet (RAAR3DNet) 2021 - 66.62% - - - - 88.59% - - - - - -
-SeST (SeST) 2021 - 60.27% - - - - - - - - - - -
-SAM-SLR-v2 (jiang2021sign-sam-slr-v2) 2021 - - - - - - - - - 59.39% 56.63% 99.00% 98.53%
-SAM-SLR (SAM_SLR) 2021 19.2M - - - - - - - - 58.73% 55.93% - 98.53%
-(Hu_Zhou_Li_2021) 2021 - - - - - - - 69.39% 66.54% 51.39% 48.75% 98.3% -
+STGCN-LSTM (68) 2025 - - - - - - - - - - - 95.2% -
+(149) 2025 - - - - - - - - - - - - 86.43%
+HAN-2S (119) 2025 0.94M - 95% 92.86% 92.71% 89.15% - - - - - - -
+DSTSA-GCN (32) 2025 8.08M - 97.74% 95.37% 95.04% 93.57% - - - - - - -
+Uni-Sign (113) 2025 592.1M - - - - - - 78.16% 76.97% 63.52% 61.32% - -
+TMS-Net (40) 2024 - - - - - - - - - 56.4% - - -
+MSNN (130) 2024 - - - - - - - 50.59% - 47.26% - - -
+SKIM (115) 2024 - - - - - - - - - 57.73% 55.40% - -
+MASA (223) 2024 - - - - - - - - - 55.77% 53.13% - -
+DSTA-SLR (86) 2024 7.4M - - - - - - 65.74% 62.31% 53.68% 51.17% 98.1%5 -
+(225) 2024 - - - - - - - 74.18% 72.18% 58.06% 55.66% 97.8% -
+TD-GCN (120) 2024 1.36M - 97.02% 95.36% 93.9% 91.4% - - - - - - -
+STST (222) 2024 - - 97.62% 95.83% 94.82% 93.18% - - - - - - -
+(123) 2024 - - - - - - - - - - - 98.87% 92.53%
+SwC GR-MMixer (122) 2024 - - - - - - - - - - - 98.54% -
+HWGAT (152) 2024 - - - - - - - - - 48.49% - - 95.80%
+(111) 2024 0.21M - 96.9% 94.17% 94.21% 92.11% - - - - - - -
+(88) 2024 - - 94.1% 90.0% - - - - - - - - -
+BEST (224) 2023 - - - - - - - 71.21% 68.24% 54.59% 52.12% 97.7% -
+SignBERT+ (77) 2023 - - - - - - - 73.71% 70.77% 55.59% 53.33% 97.8% -
+NLA-SLR (235) 2023 - - - - - - - 73.80% 70.95% 61.26% 58.31% - -
+3SCKI (198) 2023 - 65.87% - - - - - - - - - - -
+(112) 2023 - 74.08% - - - - - - - - - - -
+CTFB (66) 2023 27.2M-37.1M 69.02% - - - - 87.76% - - - - - -
+L-TSSI-DenseNet (108) 2023 7.2M - - - - - - - - - - - 93.13%
+TS3C-Net (70) 2023 28.7M - - - - - 87.55% - - - - - -
+(135) 2023 - - 97.01% 92.78% 92.00% 88.78% - - - - - - -
+SBI-DHGR (143) 2023 28M - 97% 92.36% 94.64% 91.79% - - - - - - -
+STr-GCN (184) 2023 - - 93.39% 89.20% - - - - - - - - -
+DRL-DDNet (51) 2023 - - 96.5% 93.8% - - - - - - - - -
+De+Recouple (228) 2022 - 66.79% - - - - 91.70% - - - - - -
+MS-ISTGCN (186) 2022 7.1M - 96.7% 94.9% 93.7% 91.2% - - - - - - -
+SignBERT (78) 2021 - - - - - - - 71.24% 67.96% 54.69% 52.08% 97.6% -
+SlowFast+NAS+CDC (215) 2021 251.4M 66.23% - - - - - - - - - - -
+RAAR3DNet (227) 2021 - 66.62% - - - - 88.59% - - - - - -
+SeST (191) 2021 - 60.27% - - - - - - - - - - -
+SAM-SLR-v2 (90) 2021 - - - - - - - - - 59.39% 56.63% 99.00% 98.53%
+SAM-SLR (91) 2021 19.2M - - - - - - - - 58.73% 55.93% - 98.53%
+(79) 2021 - - - - - - - 69.39% 66.54% 51.39% 48.75% 98.3% -
 6 Methods for Continuous Dynamic HGR
@@
 The goal is to recognize gestures from untrimmed videos and generate the corresponding gesture sequence.
-CHGR is typically treated as a weakly supervised task, as the precise temporal boundaries of gestures are not provided; the model learns to identify them implicitly, often using Connectionist Temporal Classification (CTC) loss (Fangyun_Wei_ICCV; CTC_graves2006connectionist).
-While CHGR has been explored in domains like autonomous driving (FA-STGCN; CTCX_WANG2022123), most research focuses on Sign Language Understanding (SLU), where it is commonly referred to as Continuous Sign Language Recognition (CSLR) or sign2gloss.
+CHGR is typically treated as a weakly supervised task, as the precise temporal boundaries of gestures are not provided; the model learns to identify them implicitly, often using Connectionist Temporal Classification (CTC) loss (207; 61).
+While CHGR has been explored in domains like autonomous driving (50; 204), most research focuses on Sign Language Understanding (SLU), where it is commonly referred to as Continuous Sign Language Recognition (CSLR) or sign2gloss.
 CSLR aims to produce a sequence of glosses [^1] Glosses are atomic lexical units in sign language, analogous to words in spoken language.
@@
 Due to the importance of non-manual features (e.g., facial expressions and upper-body movements), many approaches also incorporate these regions to improve recognition accuracy.
-Some CSLR methods are further extended to support gloss-based Sign Language Translation (SLT) by adding a gloss-to-text module (MSKA_ELSEVIER; zuo2024onlinecontinuoussignlanguage; NEURIPS2022_6cd3ac24_HEATMAPS).
+Some CSLR methods are further extended to support gloss-based Sign Language Translation (SLT) by adding a gloss-to-text module (62; 236; 29).
 In the following, we review recent advances in CHGR, grouped by input modality. We highlight key architectures, learning strategies, and benchmark results.
 6.1 RGB-based
-Most state-of-the-art RGB-based CHGR models follow a three-stage pipeline architecture: (1) a spatial model (commonly a 2DCNN or a transformer) to extract frame-wise features; (2) a short-term temporal model (mainly 1DCNN or TCN) for local temporal modeling and gloss-level feature extraction; and (3) a contextual module (e.g., BiLSTM) for long-term sequence modeling and final gesture sequence prediction (TemporalLiftPooling; VAC-CSLR; SMKD; CorrNet; C2ST; HezhenHu_TMM; SEN-CSLR; HezhenHu-JunfuPu; corrnetplus_hu2024corrnetsignlanguagerecognition).
-This standard architecture was introduced in (first_2dcnn_tcn_bilstm_TMM_2019).
-This decoupled spatial-temporal structure has been favored over unified spatiotemporal models like 3DCNNs due to the latter’s inability to define clear temporal boundaries (ADADOGLOU).
+Most state-of-the-art RGB-based CHGR models follow a three-stage pipeline architecture: (1) a spatial model (commonly a 2DCNN or a transformer) to extract frame-wise features; (2) a short-term temporal model (mainly 1DCNN or TCN) for local temporal modeling and gloss-level feature extraction; and (3) a contextual module (e.g., BiLSTM) for long-term sequence modeling and final gesture sequence prediction (83; 136; 69; 84; 217; 76; 85; 75; 82).
+This standard architecture was introduced in (33).
+This decoupled spatial-temporal structure has been favored over unified spatiotemporal models like 3DCNNs due to the latter’s inability to define clear temporal boundaries (2).
 Many RGB‐based CHGR methods build on the canonical 2DCNN + 1DCNN + BiLSTM architecture by introducing auxiliary losses and multi‐task objectives.
-For example, VAC‐CSLR (VAC-CSLR) augments the visual and temporal modules with a Visual Enhancement (VE) loss and a Visual Alignment (VA) distillation loss to accelerate convergence and improve representation quality (see Fig.
+For example, VAC‐CSLR (136) augments the visual and temporal modules with a Visual Enhancement (VE) loss and a Visual Alignment (VA) distillation loss to accelerate convergence and improve representation quality (see Fig.
 7).
-Similarly, SMKD (SMKD) adds two auxiliary classifiers—one after the 1DCNN and one after the BiLSTM—to perform synchronous gloss segmentation alongside recognition, using a three‐stage training schedule to mitigate gradient “spikes.” C2SLR (C2SLR) enforces Spatial Attention Consistency (SAC) by focusing the visual module on pose‐informed regions, and Sentence Embedding Consistency (SEC) to align the visual and sequential representations of the same gloss.
-Figure 7: An overview of the VAC-CSLR (VAC-CSLR).
+Similarly, SMKD (69) adds two auxiliary classifiers—one after the 1DCNN and one after the BiLSTM—to perform synchronous gloss segmentation alongside recognition, using a three‐stage training schedule to mitigate gradient “spikes.” C2SLR (234) enforces Spatial Attention Consistency (SAC) by focusing the visual module on pose‐informed regions, and Sentence Embedding Consistency (SEC) to align the visual and sequential representations of the same gloss.
+Figure 7: An overview of the VAC-CSLR (136).
 The feature extractor provides local information, that is processed by the subsequent alignment module to predict the gloss sequence.
@@
 To better capture local temporal patterns, several works replace or enhance the standard 1D-CNN.
-Temporal Lift Pooling (TLP) (TemporalLiftPooling) adopts a learned pooling scheme inspired by the lifting transform, eliminating handcrafted kernels.
-mLTSF‐Net (mLTSF-Net) couples content‐aware feature selection with position‐aware convolutions to fuse adjacent‐frame similarities at multiple scales.
-HS‐I3D (HS-I3D) extends I3D by extracting intermediate embeddings and deploying a U‐Net–style decoder that upsamples temporally, enabling simultaneous sign detection and boundary localization.
+Temporal Lift Pooling (TLP) (83) adopts a learned pooling scheme inspired by the lifting transform, eliminating handcrafted kernels.
+mLTSF‐Net (210) couples content‐aware feature selection with position‐aware convolutions to fuse adjacent‐frame similarities at multiple scales.
+HS‐I3D (209) extends I3D by extracting intermediate embeddings and deploying a U‐Net–style decoder that upsamples temporally, enabling simultaneous sign detection and boundary localization.
 Other methods inject lightweight attention and correlation blocks into the spatial encoder.
-CorrNet (CorrNet) adds parallel correlation and identification modules after each 2DCNN stage to model inter‐frame trajectories and highlight informative regions; CorrNet+ (corrnetplus_hu2024corrnetsignlanguagerecognition) further replaces its correlation operator with a more efficient variant and inserts an attention layer.
-SEN‐CSLR (SEN-CSLR) introduces self‐emphasizing spatial (SSEN) and temporal (TSEN) modules to reweight feature maps and keyframes with minimal overhead.
-STNet (STNet-wang2025continuous) applied frame-wise loss to the 2DCNN backbone output to enhance feature extraction and inserted multiple Spatial Resonance Modules to facilitate inter-frame modeling.
+CorrNet (84) adds parallel correlation and identification modules after each 2DCNN stage to model inter‐frame trajectories and highlight informative regions; CorrNet+ (82) further replaces its correlation operator with a more efficient variant and inserts an attention layer.
+SEN‐CSLR (85) introduces self‐emphasizing spatial (SSEN) and temporal (TSEN) modules to reweight feature maps and keyframes with minimal overhead.
+STNet (205) applied frame-wise loss to the 2DCNN backbone output to enhance feature extraction and inserted multiple Spatial Resonance Modules to facilitate inter-frame modeling.
 Many recent approaches replace 2DCNNs with transformer-based or hybrid CNN/attention modules.
-CLIP-SLA variants (CLIP_SLA_alyami2025clip_CSLR_2025) use a frozen CLIP visual encoder with LoRA or MLP adapters and an auxiliary classifier for early supervision.
-Swin-MSTP (SWIN_MSTP_Neurocomputing_2025_CSLR) combines a Swin transformer with an MS-TCN using TLP and dilated convolutions, followed by a BiLSTM for global context.
-TCNet (TCNet_Lu_Salah_Poppe_2024) augments a CNN backbone with trajectory and correlation modules, while SlowFastSign (SlowFastSign) modifies Slow-Fast-101 with bidirectional fusion and auxiliary temporal classifiers.
-To better model spatial correlations, SignGraph (SignGraph_CVPR) treats frame patches as nodes in a dynamic graph, alternating between Local Sign Graph (LSG) and Temporal Sign Graph (TSG) modules; its multiscale variant, MultiSignGraph, further improves spatiotemporal context.
-CORE (CORE-CSLR-zhang2025core) proposed a different approach aiming to leverage the relations between regions within an input frame, deploying an off-the-shelf pose estimator to detect fine-grained and contextual cues; the corresponding frame region are fed into a 2DCNN and the extracted features are fused using a GNN to obtain the final embedding.
+CLIP-SLA variants (14) use a frozen CLIP visual encoder with LoRA or MLP adapters and an auxiliary classifier for early supervision.
+Swin-MSTP (15) combines a Swin transformer with an MS-TCN using TLP and dilated convolutions, followed by a BiLSTM for global context.
+TCNet (124) augments a CNN backbone with trajectory and correlation modules, while SlowFastSign (4) modifies Slow-Fast-101 with bidirectional fusion and auxiliary temporal classifiers.
+To better model spatial correlations, SignGraph (52) treats frame patches as nodes in a dynamic graph, alternating between Local Sign Graph (LSG) and Temporal Sign Graph (TSG) modules; its multiscale variant, MultiSignGraph, further improves spatiotemporal context.
+CORE (219) proposed a different approach aiming to leverage the relations between regions within an input frame, deploying an off-the-shelf pose estimator to detect fine-grained and contextual cues; the corresponding frame region are fed into a 2DCNN and the extracted features are fused using a GNN to obtain the final embedding.
 These per-frame representations pass through a BiLSTM to perform global modeling.
-SD-Net (SD-Net-CSLR-gao2025structure) enhanced temporal modeling by using three parallel branches.
+SD-Net (53) enhanced temporal modeling by using three parallel branches.
 Each branch consists of a two-stage temporal context module that captures both frame-wise and gloss-wise features.
 Contrastive loss is applied between two of the branches, based on the proposed Entity and Template scheme.
-Authors in (diffusion-CSLR-geng2026no) formulated CSLR as a cross-modal alignment task using a diffusion model to generate the glosses aiming to mitigate the weak supervision, achieving superior performance.
-DSTEN-CSLR (DSTEN-CSLR-yu2025dsten) replaced the BiLSTM with a custom contextual model, consisting of memory and feed-forward blocks.
+Authors in (59) formulated CSLR as a cross-modal alignment task using a diffusion model to generate the glosses aiming to mitigate the weak supervision, achieving superior performance.
+DSTEN-CSLR (214) replaced the BiLSTM with a custom contextual model, consisting of memory and feed-forward blocks.
 Furthermore, the 2DCNN backbone was enhanced by adding a Spatial-Channel Attention (SCA) mechanism, supervised by ground-truth heatmaps, provided by an off-the-shelf pose estimator.
 Some models integrate linguistic embeddings with visual features to leverage gloss co-occurrence and language structure.
-SignBERT-HKSL (SignBERT-HKSL) combines a (3+2+1)D ResNet with BERT and BiLSTM, pretraining on masked inputs and fine-tuning iteratively on masked outputs.
-C2ST (C2ST) replaces the CTC decoder with a Conditional Gloss Decoder (CGD) to model gloss dependencies and aligns visual and gloss embeddings via cross-modal match loss.
-GPGN (GPGN) similarly applies a cross-modal match loss between BERT-derived gloss features and PDC-TFE visual embeddings, followed by a BiGRU for sequence modeling.
-CVT-SLR (CVT-SLR) pretrains an asymmetric VAE on a pseudo gloss-to-gloss task, then connects it to a visual encoder with a video–gloss adapter and contrastive alignment loss.
-Fully transformer-based PiSLTRc (PiSLTRc) uses content-aware neighborhood gathering and position-informed attention for end-to-end CSLR, while (HezhenHu-JunfuPu) introduces prior-aware cross-modality augmentation, training a transformer to reconstruct masked gloss sequences and generate pseudo video–text pairs via guided gloss substitutions and insertions.
+SignBERT-HKSL (233) combines a (3+2+1)D ResNet with BERT and BiLSTM, pretraining on masked inputs and fine-tuning iteratively on masked outputs.
+C2ST (217) replaces the CTC decoder with a Conditional Gloss Decoder (CGD) to model gloss dependencies and aligns visual and gloss embeddings via cross-modal match loss.
+GPGN (64) similarly applies a cross-modal match loss between BERT-derived gloss features and PDC-TFE visual embeddings, followed by a BiGRU for sequence modeling.
+CVT-SLR (226) pretrains an asymmetric VAE on a pseudo gloss-to-gloss task, then connects it to a visual encoder with a video–gloss adapter and contrastive alignment loss.
+Fully transformer-based PiSLTRc (211) uses content-aware neighborhood gathering and position-informed attention for end-to-end CSLR, while (75) introduces prior-aware cross-modality augmentation, training a transformer to reconstruct masked gloss sequences and generate pseudo video–text pairs via guided gloss substitutions and insertions.
 Finally, data augmentation and cross‐lingual strategies help alleviate limited annotations.
 Fangyun Wei et al.
-(Fangyun_Wei_ICCV) build a “dictionary” by using a pretrained CHGR model to automatically segment continuous videos into isolated signs, then mapping German and Chinese sign corpora into a shared embedding space.
-CSGC (CSGC_Rao_Sun_Wang_Wang_Zhang_2024) maintains a Gloss Prototype memory that stores class prototypes updated during training; a gloss contrastive loss then enforces cross‐sentence consistency.
+(207) build a “dictionary” by using a pretrained CHGR model to automatically segment continuous videos into isolated signs, then mapping German and Chinese sign corpora into a shared embedding space.
+CSGC (161) maintains a Gloss Prototype memory that stores class prototypes updated during training; a gloss contrastive loss then enforces cross‐sentence consistency.
 6.2 Pose-based
@@
 Several works exploit structured keypoint information by dividing the body into semantic regions such as the face, hands, and upper body.
-MSKA-SLR (MSKA_ELSEVIER) processes each region independently via attention-enhanced temporal convolutions across four branches.
+MSKA-SLR (62) processes each region independently via attention-enhanced temporal convolutions across four branches.
 The resulting features are fused to decode gloss sequences using a CTC loss, complemented by frame-level self-distillation.
-Similarly, CoSign (CoSign) employs group-specific GCNs for each subset and introduces regularization to preserve inter-stream consistency.
+Similarly, CoSign (92) employs group-specific GCNs for each subset and introduces regularization to preserve inter-stream consistency.
 Their dual-stream variant integrates joint positions and motion features using multi-level fusion to capture both static and dynamic cues.
 Temporal localization of gestures has also been explored in the literature.
-TY-Net (TY-Net) draws inspiration from YOLO to detect gesture instances in time by predicting their boundaries and centers, analogous to spatial bounding boxes.
+TY-Net (42) draws inspiration from YOLO to detect gesture instances in time by predicting their boundaries and centers, analogous to spatial bounding boxes.
 Beyond joint coordinates, pose heatmaps serve as an alternative input modality.
-The multi-cue framework in (MULTI-CUE) combines spatial and temporal reasoning by localizing facial and hand regions via a pose estimator and subsequently encoding them with a temporal module, before decoding gloss sequences with a CTC head.
-The FA-STGCN model (FA-STGCN) further improves robustness to motion scale and frequency variability.
+The multi-cue framework in (231) combines spatial and temporal reasoning by localizing facial and hand regions via a pose estimator and subsequently encoding them with a temporal module, before decoding gloss sequences with a CTC head.
+The FA-STGCN model (50) further improves robustness to motion scale and frequency variability.
 It normalizes skeleton inputs and applies alternating spatial and temporal GCNs.
@@
 Finally, in the context of autonomous driving, Wang et al.
-(CTCX_WANG2022123) developed a two-stage pose-based system to classify road intersection gestures.
+(204) developed a two-stage pose-based system to classify road intersection gestures.
 The first stage extracts geometric features and encodes them via LSTM layers.
@@
 Multimodal CHGR approaches combine inputs such as RGB, pose heatmaps, depth, etc., to overcome unimodal limitations and capture complementary information.
-A key example is TwoStream-SLR (NEURIPS2022_6cd3ac24_HEATMAPS), a dual-stream architecture for CSLR and SLT.
+A key example is TwoStream-SLR (29), a dual-stream architecture for CSLR and SLT.
 It processes RGB and pose heatmaps in parallel via the first four S3D blocks in each branch, with bidirectional lateral connections for multi-level information exchange.
@@
 8 illustrates the architecture.
-The same framework has been applied to ISLR (zuo2024onlinecontinuoussignlanguage) using a sliding window to extract sign boundaries and construct training dictionaries.
-Figure 8: TwoStream-SLR (NEURIPS2022_6cd3ac24_HEATMAPS) is a dual-stream multimodal architecture built from eight S3D blocks.
+The same framework has been applied to ISLR (236) using a sliding window to extract sign boundaries and construct training dictionaries.
+Figure 8: TwoStream-SLR (29) is a dual-stream multimodal architecture built from eight S3D blocks.
 Bidirectional Lateral Connections enable multi-level information exchange between branches.
@@
 Other CSLR models have adapted multimodal ISLR backbones.
-Uni-Sign (UNISIGNli2025) can be used directly for CSLR, while SignBERT+ (SignBERT+) can be extended by appending temporal pooling and a CTC decoder to generate gloss sequences.
+Uni-Sign (113) can be used directly for CSLR, while SignBERT+ (77) can be extended by appending temporal pooling and a CTC decoder to generate gloss sequences.
 Beyond sign language, general-purpose CHGR systems have explored multimodality across broader visual sensor inputs.
-For example, (pretrained_on_jester_3dDenseNet_LSTM) proposes a framework that segments RGB, depth, infrared, and optical flow inputs using a shared U-Net and processes them via separate 3D DenseNet-43 backbones.
+For example, (125) proposes a framework that segments RGB, depth, infrared, and optical flow inputs using a shared U-Net and processes them via separate 3D DenseNet-43 backbones.
 Extracted features are fused through Canonical Correlation Analysis to align shared representations across modalities.
@@
 Table 4 reports the computational cost and Word Error Rate (WER) of CSLR methods on three benchmark datasets.
-Leveraging large language models, as in (UNISIGNli2025) and (diffusion-CSLR-geng2026no), yields substantial performance gains without dramatically increasing parameters.
-SlowFastSign (SlowFastSign) achieves strong accuracy with a moderate number of parameters, while TwoStream-SLR (NEURIPS2022_6cd3ac24_HEATMAPS) has slightly higher WER, requires twice as many parameters, and involves a more complex training setup, though WER improves with additional training data (Fangyun_Wei_ICCV).
-C2ST (C2ST) is accurate but contains 187.8M parameters, which may limit real-world deployment.
-For resource-constrained scenarios, the pose-based CoSign-2s (CoSign) requires only 30.1 GFLOPs, though pose estimation cost is not included.
-DSTEN-CSLR (DSTEN-CSLR-yu2025dsten) is a lightweight RGB-based model but less accurate.
+Leveraging large language models, as in (113) and (59), yields substantial performance gains without dramatically increasing parameters.
+SlowFastSign (4) achieves strong accuracy with a moderate number of parameters, while TwoStream-SLR (29) has slightly higher WER, requires twice as many parameters, and involves a more complex training setup, though WER improves with additional training data (207).
+C2ST (217) is accurate but contains 187.8M parameters, which may limit real-world deployment.
+For resource-constrained scenarios, the pose-based CoSign-2s (92) requires only 30.1 GFLOPs, though pose estimation cost is not included.
+DSTEN-CSLR (214) is a lightweight RGB-based model but less accurate.
 Overall, a trade-off between accuracy and computational cost persists.
-Table 4: Comparison of CSLR methods in terms of computational cost and accuracy (Word Error Rate, WER) on three benchmarks: CSL-Daily (Back-Translation-CVPR), Phoenix (PHOENIX-2014), and Phoenix T (NEURAL-SIGN-LANGUAGE-TRANSLATION-CVPR-CAMGOZ).
+Table 4: Comparison of CSLR methods in terms of computational cost and accuracy (Word Error Rate, WER) on three benchmarks: CSL-Daily (230), Phoenix (49), and Phoenix T (27).
 ”Params” denotes the number of trainable parameters.
 Method Year GFLOPs Params CSL-Daily Phoenix Phoenix T
-(diffusion-CSLR-geng2026no) 2025 - 86.64M 23.7% 16.8% 17.2%
-DSTEN-CSLR (DSTEN-CSLR-yu2025dsten) 2025 366.3 30.19M 28.1% 20.0% 20.3%
-SLA-Adapter (CLIP_SLA_alyami2025clip_CSLR_2025) 2025 - 26.2M 25.8% 18.8% 19.5%
-SLA-LoRA (CLIP_SLA_alyami2025clip_CSLR_2025) 2025 - 30.8M 25.8% 19.3% 19.4%
-Uni-Sign (UNISIGNli2025) 2025 - - 26% - -
-Swin-MSTP (SWIN_MSTP_Neurocomputing_2025_CSLR) 2025 - - 27.1% 18.7% 19.7%
-CORE (CORE-CSLR-zhang2025core) 2025 - - 31.1% 21.7% 21.6%
-SD-Net (SD-Net-CSLR-gao2025structure) 2025 - - 27.9% - 20.3%
-STNet (STNet-wang2025continuous) 2025 - - 27.2% 19.3% 19.8%
-SlowFastSign (SlowFastSign) 2024 - 52.5M 24.9% 18.3% 18.7%
-MultiSignGraph (SignGraph_CVPR) 2024 - - 26.4% 19.1% 19.1%
-CSGC (CSGC_Rao_Sun_Wang_Wang_Zhang_2024) 2024 - - 26.7% 19% 19.5%
-CorrNet+ (corrnetplus_hu2024corrnetsignlanguagerecognition) 2024 - - 28.2% 18.2% 19.1%
-TCNet (TCNet_Lu_Salah_Poppe_2024) 2024 932.1 19.3M 29.3% 18.9% 19.4%
-C2SLR (C2SLR) 2024 1124.6 32.6M - 20.4% 20.4%
-GPGN (GPGN) 2024 - - 30% 20.4% 20.5%
-C2ST (C2ST) 2023 1635 187.8M 25.8% 17.7% 18.9%
-CoSign-2s (CoSign) 2023 30.1 28.2M 27.2% 20.1% 20.1%
-CTCA (CTCA) 2023 - - 29.4% 20.1% 20.3%
-CorrNet (CorrNet) 2023 1035.4 20.5M 30.1% 19.4% 20.5%
-SEN-CSLR (SEN-CSLR) 2023 1144.2 23.1M 30.7% 21% 20.7%
-(Fangyun_Wei_ICCV) 2023 - 105.2M - 16.7% 18.5%
-CVT-SLR (CVT-SLR) 2023 - - - 20.1% 20.3%
-TwoStream-SLR (NEURIPS2022_6cd3ac24_HEATMAPS) 2022 - 105.2M 25.3% 18.8% 19.3%
-STMC (MULTI-CUE) 2022 1742.1 40.71M - 20.7% 21%
-TLP (TemporalLiftPooling) 2022 2950.5 48M - 20.8% 21.2%
-SMKD (SMKD) 2021 1032.3 20.3M - 21% 22.4%
-VAC-CSLR (VAC-CSLR) 2021 1120.4 22.3M - 22.3% -
+(59) 2025 - 86.64M 23.7% 16.8% 17.2%
+DSTEN-CSLR (214) 2025 366.3 30.19M 28.1% 20.0% 20.3%
+SLA-Adapter (14) 2025 - 26.2M 25.8% 18.8% 19.5%
+SLA-LoRA (14) 2025 - 30.8M 25.8% 19.3% 19.4%
+Uni-Sign (113) 2025 - - 26% - -
+Swin-MSTP (15) 2025 - - 27.1% 18.7% 19.7%
+CORE (219) 2025 - - 31.1% 21.7% 21.6%
+SD-Net (53) 2025 - - 27.9% - 20.3%
+STNet (205) 2025 - - 27.2% 19.3% 19.8%
+SlowFastSign (4) 2024 - 52.5M 24.9% 18.3% 18.7%
+MultiSignGraph (52) 2024 - - 26.4% 19.1% 19.1%
+CSGC (161) 2024 - - 26.7% 19% 19.5%
+CorrNet+ (82) 2024 - - 28.2% 18.2% 19.1%
+TCNet (124) 2024 932.1 19.3M 29.3% 18.9% 19.4%
+C2SLR (234) 2024 1124.6 32.6M - 20.4% 20.4%
+GPGN (64) 2024 - - 30% 20.4% 20.5%
+C2ST (217) 2023 1635 187.8M 25.8% 17.7% 18.9%
+CoSign-2s (92) 2023 30.1 28.2M 27.2% 20.1% 20.1%
+CTCA (63) 2023 - - 29.4% 20.1% 20.3%
+CorrNet (84) 2023 1035.4 20.5M 30.1% 19.4% 20.5%
+SEN-CSLR (85) 2023 1144.2 23.1M 30.7% 21% 20.7%
+(207) 2023 - 105.2M - 16.7% 18.5%
+CVT-SLR (226) 2023 - - - 20.1% 20.3%
+TwoStream-SLR (29) 2022 - 105.2M 25.3% 18.8% 19.3%
+STMC (231) 2022 1742.1 40.71M - 20.7% 21%
+TLP (83) 2022 2950.5 48M - 20.8% 21.2%
+SMKD (69) 2021 1032.3 20.3M - 21% 22.4%
+VAC-CSLR (136) 2021 1120.4 22.3M - 22.3% -
 7 Summary of SoTA Methods
@@
 This is because they either cannot represent fine-grained gestures (depth, optical flow) or perform poorly in dynamic recognition (pose estimators often fail on blurry images).
-RGB-based models—even older or simpler ones—already achieve strong results on SHGR datasets (HAGRIDv2_ARXIV), so adding pose can incur significant computational overhead without substantial accuracy gains.
-In CSLR, rapid hand movements degrade pose estimator performance, making pose usable only alongside RGB (NEURIPS2022_6cd3ac24_HEATMAPS) and increasing complexity.
-Thus, pose-based approaches are mostly limited to IDHGR, with some exceptions (CoSign), where incorporating multiple modalities can notably improve performance (SAM_SLR; jiang2021sign-sam-slr-v2), even with late fusion.
-Regarding architecture, SHGR’s simplicity allows even small 2DCNNs (EDenseNet; G-CNN-SHARMA2021115657) to achieve very high accuracy ($\geq$95%).
-However, the limited size and diversity of datasets like (NUS-I-2010; NUS-II-2013; ouhands; triesch2001system) can lead to overfitting, reducing real-world performance.
+RGB-based models—even older or simpler ones—already achieve strong results on SHGR datasets (145), so adding pose can incur significant computational overhead without substantial accuracy gains.
+In CSLR, rapid hand movements degrade pose estimator performance, making pose usable only alongside RGB (29) and increasing complexity.
+Thus, pose-based approaches are mostly limited to IDHGR, with some exceptions (92), where incorporating multiple modalities can notably improve performance (91; 90), even with late fusion.
+Regarding architecture, SHGR’s simplicity allows even small 2DCNNs (189; 176) to achieve very high accuracy ($\geq$95%).
+However, the limited size and diversity of datasets like (106; 153; 132; 193) can lead to overfitting, reducing real-world performance.
 In contrast, IDHGR and especially ISLR rely on more complex models.
-ST-GCN variants (JOINTS_MOTION_VECTORS_Improved_ST-GCN; TD-GCN) and GCN+Transformer pipelines (SignBERT_ICCV; SignBERT+; BEST; Weichao-Zhao) mainly process skeleton data, often combined with RGB 3DCNNs (e.g., I3D, Res3D) to compensate for pose estimation errors.
+ST-GCN variants (186; 120) and GCN+Transformer pipelines (78; 77; 224; 225) mainly process skeleton data, often combined with RGB 3DCNNs (e.g., I3D, Res3D) to compensate for pose estimation errors.
 While many studies focus on improving pose-based streams, RGB modalities are less explored.
-SAM-SLR (SAM_SLR) and SAM-SLRv2 (jiang2021sign-sam-slr-v2) propose advanced separate processing for each modality, representing current SoTA for IDHGR.
-RGBD models (RAAR3DNet; CTFB_Hampiholi; Yunan_Li) typically use dual-stream 3DCNNs, increasing computational cost.
+SAM-SLR (91) and SAM-SLRv2 (90) propose advanced separate processing for each modality, representing current SoTA for IDHGR.
+RGBD models (227; 66; 112) typically use dual-stream 3DCNNs, increasing computational cost.
 CSLR demands even more complex architectures to capture short-term (gloss-wise) and long-term (global) context.
-The 2DCNN+1DCNN+BiLSTM pipeline remains common, with relatively fewer parameters than SHGR models (e.g., Gesture-CNN (G-CNN-SHARMA2021115657)  67M vs.
-large CSLR models  48M (TemporalLiftPooling)).
-Vision transformers increasingly replace 2DCNN backbones (SWIN_MSTP_Neurocomputing_2025_CSLR; C2ST; CLIP_SLA_alyami2025clip_CSLR_2025) for their attention-based focus on critical regions.
-TwoStream-SLR (NEURIPS2022_6cd3ac24_HEATMAPS) employs a dual-stream 3DCNN with multimodal fusion and pose heatmaps (instead of skeleton joints) alongside RGB frames, achieving high accuracy.
+The 2DCNN+1DCNN+BiLSTM pipeline remains common, with relatively fewer parameters than SHGR models (e.g., Gesture-CNN (176)  67M vs.
+large CSLR models  48M (83)).
+Vision transformers increasingly replace 2DCNN backbones (15; 217; 14) for their attention-based focus on critical regions.
+TwoStream-SLR (29) employs a dual-stream 3DCNN with multimodal fusion and pose heatmaps (instead of skeleton joints) alongside RGB frames, achieving high accuracy.
 As mentioned earlier, SHGR is a relatively easy task, so researchers have not proposed any notable learning strategy.
-Pose-based methods for ISLR often (SignBERT_ICCV; SignBERT+; BEST; Weichao-Zhao) insert hand topology prior through SSL.
+Pose-based methods for ISLR often (78; 77; 224; 225) insert hand topology prior through SSL.
 NAS is also utilized to determine the optimal structure of 3DCNNs.
 Models for CSLR face serious optimization difficulties compared to other methods as a consequence of their complexity.
-Thus, extra supervision is provided via additional losses attached to various layers of the models (VAC-CSLR; CLIP_SLA_alyami2025clip_CSLR_2025).
-Leveraging additional training data combined with accurate architectures during training (Fangyun_Wei_ICCV) increases the performance, but makes a straightforward comparison with other methods infeasible.
-In summary, LHGR-Net (LHGR-Net) and stand-alone pose-based methods (Yangke_patcog; TD-GCN; CoSign) are suitable for resource-constrained use cases (assuming that the cost of the pose estimation is negligible, which is true especially when the keypoints are extracted by the deployed camera), balancing the computational complexity with the recognition accuracy.
-TwoStream-SLR (NEURIPS2022_6cd3ac24_HEATMAPS), C2ST (C2ST) and (diffusion-CSLR-geng2026no) achieve superior performance for CSLR but come with complex training settings and large computational cost.
-NLA-SLR (NLA_SLR) and SAM-SLRv2 (jiang2021sign-sam-slr-v2) surpass most of the related methods, but they require substantial computational resources.
-Uni-Sign (UNISIGNli2025) introduced a novelty, by combining various SLU tasks into a single one.
+Thus, extra supervision is provided via additional losses attached to various layers of the models (136; 14).
+Leveraging additional training data combined with accurate architectures during training (207) increases the performance, but makes a straightforward comparison with other methods infeasible.
+In summary, LHGR-Net (218) and stand-alone pose-based methods (111; 120; 92) are suitable for resource-constrained use cases (assuming that the cost of the pose estimation is negligible, which is true especially when the keypoints are extracted by the deployed camera), balancing the computational complexity with the recognition accuracy.
+TwoStream-SLR (29), C2ST (217) and (59) achieve superior performance for CSLR but come with complex training settings and large computational cost.
+NLA-SLR (235) and SAM-SLRv2 (90) surpass most of the related methods, but they require substantial computational resources.
+Uni-Sign (113) introduced a novelty, by combining various SLU tasks into a single one.
 Although it achieves significant performance and unifies a variety of tasks providing a shared framework, the computational complexity makes difficult the real-world deployment and the requirement of training data with detailed annotations prevents the adoption to other sign languages.
@@
 Method Task Domain Modality Architecture Strengths or Key Innovation Limitations
-FGDSNet (FGDSNet) SHGR HRI RGB Dual-stream hand segmentation model + dual-stream 2DCNN + mid-level fusion The dual-stream classifier addresses any failures of the task-specific segmentation model.
+FGDSNet (229) SHGR HRI RGB Dual-stream hand segmentation model + dual-stream 2DCNN + mid-level fusion The dual-stream classifier addresses any failures of the task-specific segmentation model.
 The segmentation model and the 2DCNN backbones must be trained separately.
-LHGR-Net (LHGR-Net) SHGR HCI RGB Atrous convolutional layers followed by depth separable convolutions and attention. Suitable for resource-constrained deployment. Very small capacity.
-2DPSTPP-Net (2DPSTPP) SHGR HCI RGB 2DCNN with custom pooling layers and residual connections Spatial pyramid pooling and residual connections obtain robust representations.
+LHGR-Net (218) SHGR HCI RGB Atrous convolutional layers followed by depth separable convolutions and attention. Suitable for resource-constrained deployment. Very small capacity.
+2DPSTPP-Net (89) SHGR HCI RGB 2DCNN with custom pooling layers and residual connections Spatial pyramid pooling and residual connections obtain robust representations.
 The model is relatively large for SHGR.
-Gesture-CNN (G-CNN-SHARMA2021115657) SHGR SLU RGB 2DCNN consisting of four convolutional layers The simplicity of the architecture enables easy implementation and training.
-Relatively large (67M parameters).
-SAM-SLR (SAM_SLR), SAM-SLR-v2 (jiang2021sign-sam-slr-v2) IDHGR SLU RGBD, Pose, HHA, OF, DF Six/seven streams followed by late fusion Incorporated a large number of modalities, achieving very high accuracy.
+Gesture-CNN (176) SHGR SLU RGB 2DCNN consisting of four convolutional layers The simplicity of the architecture enables easy implementation and training. Relatively large (67M parameters).
+SAM-SLR (91), SAM-SLR-v2 (90) IDHGR SLU RGBD, Pose, HHA, OF, DF Six/seven streams followed by late fusion Incorporated a large number of modalities, achieving very high accuracy.
 Insufficient fusion and computationally expensive.
-SignBERT (SignBERT_ICCV), SignBERT+ (SignBERT+) (ISLR) IDHGR SLU RGB, Pose Spatial GCN + Transformer & 3DCNN + late fusion The model gains hand topology prior through SSL pretraining (masked hand tokens reconstruction, inspired by BERT) Needs extensive SSL pretraining, the processing and incorporation of the RGB modality may be suboptimal
-SignBERT+ (SignBERT+) (CSLR) CDHGR SLU RGB, Pose Spatial GCN + Transformer & 3DCNN + intermediate fusion + BiLSTM
-BEST (BEST) IDHGR SLU RGB, Pose dVAE + Spatial GCN + Transformer & 3DCNN + late fusion
-NLA-SLR (NLA_SLR) IDHGR SLU RGB, Pose Four-stream network based on S3D with two frame rates and multi-level fusion Novel architecture performing multimodality fusion efficiently.
+SignBERT (78), SignBERT+ (77) (ISLR) IDHGR SLU RGB, Pose Spatial GCN + Transformer & 3DCNN + late fusion The model gains hand topology prior through SSL pretraining (masked hand tokens reconstruction, inspired by BERT) Needs extensive SSL pretraining, the processing and incorporation of the RGB modality may be suboptimal
+SignBERT+ (77) (CSLR) CDHGR SLU RGB, Pose Spatial GCN + Transformer & 3DCNN + intermediate fusion + BiLSTM
+BEST (224) IDHGR SLU RGB, Pose dVAE + Spatial GCN + Transformer & 3DCNN + late fusion
+NLA-SLR (235) IDHGR SLU RGB, Pose Four-stream network based on S3D with two frame rates and multi-level fusion Novel architecture performing multimodality fusion efficiently.
 Language-aware label smoothing facilitates training.
@@
 Is affected by the performance of word encoders
-Uni-Sign (UNISIGNli2025) IDHGR, CDHGR SLU RGB, Pose A pretrained LLM process features extracted by pose, visual and temporal encoders Provides a unified framework for multiple SLU tasks, increasing the recognition accuracy.
+Uni-Sign (113) IDHGR, CDHGR SLU RGB, Pose A pretrained LLM process features extracted by pose, visual and temporal encoders Provides a unified framework for multiple SLU tasks, increasing the recognition accuracy.
 Requires additional training data
-TD-GCN (TD-GCN) IDHGR - Pose ST-GCN variant Facilitates temporal modeling through non-fixed time-dependent topology Depends on the pose estimation performance.
-SBI-DHGR (SBI-DHGR) IDHGR - Pose Multi-stream 1DCNN + BiLSTM followed by late fusion Utilizes features of various levels of granularity. Very large compared to other pose-based methods
-DSTSA-GCN (DSTSA-CGN_cui2025dstsa) (4s) IDHGR - Pose Four stream GCN + TCN Group channel-wise and temporal-wise GCNs with dynamic graph topologies improve the performance.
+TD-GCN (120) IDHGR - Pose ST-GCN variant Facilitates temporal modeling through non-fixed time-dependent topology Depends on the pose estimation performance.
+SBI-DHGR (143) IDHGR - Pose Multi-stream 1DCNN + BiLSTM followed by late fusion Utilizes features of various levels of granularity. Very large compared to other pose-based methods
+DSTSA-GCN (32) (4s) IDHGR - Pose Four stream GCN + TCN Group channel-wise and temporal-wise GCNs with dynamic graph topologies improve the performance.
 The static part of the hand topology is initialized randomly, lacking prior.
-SlowFast+NAS (3D_CDC_NAS_TIP_2021) IDHGR - RGBD, OF Multi-stream 3DCNN with 3D-CDC and multi-level fusion Enhanced temporal modeling through central difference convolution and optimal hyperparameter selection through NAS.
+SlowFast+NAS (215) IDHGR - RGBD, OF Multi-stream 3DCNN with 3D-CDC and multi-level fusion Enhanced temporal modeling through central difference convolution and optimal hyperparameter selection through NAS.
 Needs substantials computational resources (251.4M params, 232.1GFLOPs) compared to similar methods
-RAAR3DNet (RAAR3DNet) IDHGR - RGBD Dual stream I3D with adoptable cells, spatial and temporal attention NAS improves the architecture.
+RAAR3DNet (227) IDHGR - RGBD Dual stream I3D with adoptable cells, spatial and temporal attention NAS improves the architecture.
 The spatial attention mechanism is supervised by an off-the-shelf pose estimator.
 It relies on the correctness of the pose estimator
-CTFB (CTFB_Hampiholi) IDHGR - RGBD Dual-stream 3DCNN + Multi-level fusion with transformer-based modules Performs efficient multimodality fusion via multiple attention modules placed in different layers.
+CTFB (66) IDHGR - RGBD Dual-stream 3DCNN + Multi-level fusion with transformer-based modules Performs efficient multimodality fusion via multiple attention modules placed in different layers.
 The scaling to more than two modalities is not straightforward
-(diffusion-CSLR-geng2026no) CHGR SLU RGB Visual encoder followed by a Diffusion Model (DM). Using a DM for gloss generation makes it very accurate. Increased complexity.
-CLIP-SLA (CLIP_SLA_alyami2025clip_CSLR_2025) CHGR SLU RGB CLIP visual encoder (with MLP adopters or LoRA) + TCN + BiLSTM Early supervision by adding an extra classifier accelerates training, while the utilization of a foundation model enhances the feature extraction.
+(59) CHGR SLU RGB Visual encoder followed by a Diffusion Model (DM). Using a DM for gloss generation makes it very accurate. Increased complexity.
+CLIP-SLA (14) CHGR SLU RGB CLIP visual encoder (with MLP adopters or LoRA) + TCN + BiLSTM Early supervision by adding an extra classifier accelerates training, while the utilization of a foundation model enhances the feature extraction.
 The utilization of CLIP implies high computational cost.
-Swin-MSTP (SWIN_MSTP_Neurocomputing_2025_CSLR) CHGR SLU RGB SwinT+1DCNN/TLP Deploying a multiscale TCN, the model is able to recognize continuous signs of varying time durations.
-Computationally expensive.
-C2ST (C2ST) CHGR SLU RGB SwinT+1DCNN+BiLSTM CGD loss (that doesn’t make any independence assumption, compared to CTC) and self-distillation increase the performance.
+Swin-MSTP (15) CHGR SLU RGB SwinT+1DCNN/TLP Deploying a multiscale TCN, the model is able to recognize continuous signs of varying time durations. Computationally expensive.
+C2ST (217) CHGR SLU RGB SwinT+1DCNN+BiLSTM CGD loss (that doesn’t make any independence assumption, compared to CTC) and self-distillation increase the performance.
 The vision transformer as well as the language model utilized during training increase the computational complexity.
-CoSign-2s (CoSign) CHGR SLU Pose ST-GCN variant The computational cost is negligible, thus CoSign-2s perfectly balances efficient execution and accurate recognition.
+CoSign-2s (92) CHGR SLU Pose ST-GCN variant The computational cost is negligible, thus CoSign-2s perfectly balances efficient execution and accurate recognition.
 The performance depends on the attributes of the pose estimator.
-(Fangyun_Wei_ICCV) CHGR SLU RGB, Pose TwoStream-SLR Utilization of extra training data via dictionary construction. Needs additional training data.
-TwoStream-SLR (NEURIPS2022_6cd3ac24_HEATMAPS) CHGR SLU RGB, Pose Dual-stream 3DCNN with bidirectional lateral connections Extra supervision by auxiliary losses and self-distillation applied to additional prediction heads facilitate training.
+(207) CHGR SLU RGB, Pose TwoStream-SLR Utilization of extra training data via dictionary construction. Needs additional training data.
+TwoStream-SLR (29) CHGR SLU RGB, Pose Dual-stream 3DCNN with bidirectional lateral connections Extra supervision by auxiliary losses and self-distillation applied to additional prediction heads facilitate training.
 The multi-level fusion incorporates efficiently the two modalities.
@@
 Additionally, some datasets support Sign Language Translation (SLT) by providing sentence-level gloss annotations and parallel spoken language transcriptions.
-The RWTH-PHOENIX-Weather 2014 (PHOENIX-2014) and its extended version PHOENIX-2014-T1 (NEURAL-SIGN-LANGUAGE-TRANSLATION-CVPR-CAMGOZ) are benchmark datasets offering sentence-level annotations in German Sign Language (DGS) captured from television weather forecasts.
+The RWTH-PHOENIX-Weather 2014 (49) and its extended version PHOENIX-2014-T1 (27) are benchmark datasets offering sentence-level annotations in German Sign Language (DGS) captured from television weather forecasts.
 A semi-automated annotation protocol was proposed, leveraging existing methods for speech recognition and accelerating the process.
@@
 wearing dark clothes in front of a static gray background), probably limiting the generalization of models trained to it.
-On the contrary, more recent efforts such as BOBSL (BOBSL-BBC) and OpenASL (OpenASL) have significantly expanded the signer base and vocabulary, with hours of British (BSL) and American Sign Language (ASL) footage collected from TV and web sources, respectively.
+On the contrary, more recent efforts such as BOBSL (12) and OpenASL (178) have significantly expanded the signer base and vocabulary, with hours of British (BSL) and American Sign Language (ASL) footage collected from TV and web sources, respectively.
 The more unconstrained environment and the increased number of signers improve the HGR models real-world performance.
-The lab-recorded KETI (KETI_MDPI) comprises samples captured from two camera angles (i.e.
+The lab-recorded KETI (101) comprises samples captured from two camera angles (i.e.
 front and side), enlarging diversity.
 The aim of the dataset is to interpret Korean sign language gestures that are performed in emergency situations using skeleton and RGB data, reducing the need of professionals.
-Another multimodal dataset collected in a laboratory is the How2Sign (How2Sign-CVPR) for ASL recognition and translation.
+Another multimodal dataset collected in a laboratory is the How2Sign (44) for ASL recognition and translation.
 The entire dataset was captured in the termed Green Screen studio with a single RGB camera and depth sensor, while a subset of it in the Panoptic studio using multiple cameras from various viewpoints, enabling accurate 3D pose estimation.
-Pose data and text transcripts are also considered in TVB-HKSL-News (TVB-HKSL-News).
+Pose data and text transcripts are also considered in TVB-HKSL-News (144).
 The samples were collected from TV news with subtitles and Optical Character Recognition was applied to extract the text sequence, while an off-the-shelf pose estimator was utilized to obtain the human keypoints.
-The RGB-based CSL-Daily (Back-Translation-CVPR) provides also glosses and textual transcriptions, aiding both CSLR and SLT tasks.
+The RGB-based CSL-Daily (230) provides also glosses and textual transcriptions, aiding both CSLR and SLT tasks.
 A sign dictionary (referred to as SignDict) was also constructed to be used for other tasks, including sign spotting and segmentation.
-The Auslan Daily Dataset (Auslan-Daily) consists of Australian Sign Language (AUSLAN) gestures performed by humans in a television series, TV news and web videos.
+The Auslan Daily Dataset (177) consists of Australian Sign Language (AUSLAN) gestures performed by humans in a television series, TV news and web videos.
 These samples may contain multiple persons interacting with each other, which makes the dataset suitable for in-the-wild scenarios.
 Other datasets for CSLR do not provide text translations.
-ChicagoFSWild+ (ChicagoFSWild-plus-ICCV) was collected from web videos, including ASL signs performed by 260 signers making this dataset the largest in terms of number of subjects.
-HKSL (SignBERT-HKSL) was captured in a laboratory environment, providing RGBD data.
+ChicagoFSWild+ (179) was collected from web videos, including ASL signs performed by 260 signers making this dataset the largest in terms of number of subjects.
+HKSL (233) was captured in a laboratory environment, providing RGBD data.
 Datasets for IDLSR are larger compared to datasets for CSLR, as they do not require complex annotation making the data collection easier.
-Many of these datasets, such as CSL (CSL500), GSL iso.
-(ADADOGLOU), BoshporusSign22k (bosphorussign22k), AUTSL (AUTSL), KArSL (KArSL) and FDMSE-ISL (FDMSE-hierarchicalwindowedgraphattention), include pose or depth modalities, enriching multimodal learning.
+Many of these datasets, such as CSL (87), GSL iso.
+(2), BoshporusSign22k (150), AUTSL (183), KArSL (182) and FDMSE-ISL (152), include pose or depth modalities, enriching multimodal learning.
 All of them, except for the AUTSL, were collected in a laboratory with fixed background.
 Datasets that do not include such modalities, can be extended to them by using off-the-shelf DL pose and depth estimators.
-The ASL Dataset (TAXONOMY_OF_HAND_GESTURES_2000_ASL_DATASET) is the only one consisting of pose sequences exclusively.
-MS-ASL (MS-ASL) and WLASL (WLASL) are two RGB-based datasets isolated ASL recognition collected from publicly available web videos of the deaf community, including educational resources and ASL tutorials and performed by 222 and 119 subjects respectively, increasing the diversity.
-However, they contain relatively few samples compared to large-scale datasets such as BSL-1K (bsl-1K), CSLD (tim-slr_csld_dataseta) and NCSL (2plus1-D-SLR_and_NCSL).
+The ASL Dataset (17) is the only one consisting of pose sequences exclusively.
+MS-ASL (94) and WLASL (110) are two RGB-based datasets isolated ASL recognition collected from publicly available web videos of the deaf community, including educational resources and ASL tutorials and performed by 222 and 119 subjects respectively, increasing the diversity.
+However, they contain relatively few samples compared to large-scale datasets such as BSL-1K (11), CSLD (202) and NCSL (200).
 For isolated Irish Sign Language (ISL) recognition, only laboratory datasets have been proposed (e.g.
-INCLUDE (INCLUDE), IISL2020 (IISL2020_MDPI) and FDMSE-ISL (FDMSE-hierarchicalwindowedgraphattention)), making necessary the collection of an in-the-wild dataset.
-For Vietnamese SLR, authors in (Multi-VSL-dinh2025sign) collecting a large-scale multi-view dataset, considering three different point of views, enhancing diversity.
+INCLUDE (188), IISL2020 (103) and FDMSE-ISL (152)), making necessary the collection of an in-the-wild dataset.
+For Vietnamese SLR, authors in (41) collecting a large-scale multi-view dataset, considering three different point of views, enhancing diversity.
 Other sign languages have also been considered including Argentinian (e.g.
-LSA64 (Lsa64argentiniansignlanguage)), Persian (e.g.
-RKS-PERSIANSIGN (persian_sl_dataset)) and Iranian (e.g.
-ISLR101 (ISLR101-IRANIAN)).
+LSA64 (167)), Persian (e.g.
+RKS-PERSIANSIGN (163)) and Iranian (e.g.
+ISLR101 (160)).
 The rarity of some sign languages like them mentioned previously poses obstacles to finding relevant annotators, exacerbating the data scarcity challenge, as it will be discussed later.
@@
 Name Year Task Modality Samples Signers Sign Voc. Language Duration Source
-TVB-HKSL-News (TVB-HKSL-News) 2024 CSLR/SLT RGB, Pose 7000 2 6515 HKSL 16.07h TV
-Auslan-Daily (Auslan-Daily) 2023 CSLR/SLT RGB 25000 67 - AUSLAN - TV, Web
-OpenASL (OpenASL) 2022 CSLR/SLT RGB 98417 220 - ASL 288h Web
-CSL-Daily (Back-Translation-CVPR) 2021 CSLR/SLT RGB 20654 10 2000 CSL 23.27h Lab
-BOBSL (BOBSL-BBC) 2021 CSLR/SLT RGB 1962 39 2281 BSL 1467h TV
-How2Sign (green studio) (How2Sign-CVPR) 2021 CSLR/SLT RGBD, Pose 2529 11 - ASL 79.12h Lab
-Corpus FinSL (Corpus-Finnish) 2020 CSLR/SLT RGB 343 21 - FinSL 14.37h -
-KETI (KETI_MDPI) 2019 CSLR/SLT RGB, Pose 14672 14 524 KSL 27.99h Lab
-GSL SD/SI (ADADOGLOU) 2019 CSLR/SLT RGBD 10295 7 310 GSL 9.59h Lab
-PHOENIX-2014-T1 (NEURAL-SIGN-LANGUAGE-TRANSLATION-CVPR-CAMGOZ) 2018 CSLR/SLT RGB 8257 9 1066 DGS - TV
-PHOENIX-2014 (PHOENIX-2014) 2014 CSLR/SLT RGB - 9 1558 DGS 10.73h TV
-HKSL (SignBERT-HKSL) 2021 CSLR RGBD 2400 6 55 HKSL - Lab
-ChicagoFSWild+ (ChicagoFSWild-plus-ICCV) 2019 CSLR RGB 55232 260 26 ASL - Web
-Multi-VSL (Multi-VSL-dinh2025sign) 2025 IDSLR RGB, Pose 84764 30 1000 Vietnamese - Lab
-BdSLW60 (BdSLW60-rubaiyeat2025bdslw60) 2025 IDSLR RGB 9307 18 60 Bangla - Lab
-ISLR101 (ISLR101-IRANIAN) 2025 IDSLR RGB, Pose 4614 10 101 Iranian - -
-ISW-1000 (ISW-1000-ren2025multi) 2025 IDSLR RGB, Pose, OF 10000 10 1000 CSL - Lab
-FDMSE-ISL (FDMSE-hierarchicalwindowedgraphattention) 2024 IDSLR RGBD 40033 20 2002 ISL 36h Lab
-LSA64 (Lsa64argentiniansignlanguage) 2023 IDSLR RGB 3200 10 64 LSA 1.9h Lab
-Signsuisse Lexicon (SIGN-muller-etal-2023-findings) 2023 IDSLR RGB 18221 - - 3 (Mult) - Web
-CSLD (tim-slr_csld_dataseta) 2022 IDSLR RGB 120000 30 400 CSL 66.8h Lab
-NCSL (2plus1-D-SLR_and_NCSL) 2022 IDSLR RGB 90000 30 300 CSL 50.1h Lab
-KArSL (KArSL) 2021 IDSLR RGBD, Pose 75300 3 502 ArSL - Lab
-NMFs-CSL (NMFs-CSL) 2021 IDSLR RGB 32010 10 1.067 CSL 24.4h Lab
-BosphorusSign22k (bosphorussign22k) 2020 IDSLR RGBD, Pose 22542 6 744 TSL 19h Lab
-AUTSL (AUTSL) 2020 IDSLR RGBD, Pose 38336 43 226 TSL 21h Lab
-INCLUDE (INCLUDE) 2020 IDSLR RGB 4287 7 263 ISL 3h Lab
-IISL2020 (IISL2020_MDPI) 2020 IDSLR RGB 12100 16 11 ISL 7h Lab
-RKS-PERSIANSIGN (persian_sl_dataset) 2020 IDSLR RGB 10000 10 100 Persian - Lab
-BSL-1K (bsl-1K) 2020 IDSLR RGB 280000 40 1064 BSL 1060h TV
-MSASL (MS-ASL) 2019 IDSLR RGB 25513 222 1000 ASL 24.65h Web
-WLASL (WLASL) 2019 IDSLR RGB 21083 119 2000 ASL 14.11h Web
-GSL isol. (ADADOGLOU) 2019 IDSLR RGBD 40785 7 310 GSL 6.44h Lab
-SMILE (SMILE) 2018 IDSLR RGB - 30 100 DSGS - Lab
-ASL Dataset (TAXONOMY_OF_HAND_GESTURES_2000_ASL_DATASET) 2018 IDSLR Pose 1200 20 30 ASL - Lab
-CSL (CSL500) 2018 IDSLR RGBD, Pose 125000 50 500 CSL - Lab
+TVB-HKSL-News (144) 2024 CSLR/SLT RGB, Pose 7000 2 6515 HKSL 16.07h TV
+Auslan-Daily (177) 2023 CSLR/SLT RGB 25000 67 - AUSLAN - TV, Web
+OpenASL (178) 2022 CSLR/SLT RGB 98417 220 - ASL 288h Web
+CSL-Daily (230) 2021 CSLR/SLT RGB 20654 10 2000 CSL 23.27h Lab
+BOBSL (12) 2021 CSLR/SLT RGB 1962 39 2281 BSL 1467h TV
+How2Sign (green studio) (44) 2021 CSLR/SLT RGBD, Pose 2529 11 - ASL 79.12h Lab
+Corpus FinSL (173) 2020 CSLR/SLT RGB 343 21 - FinSL 14.37h -
+KETI (101) 2019 CSLR/SLT RGB, Pose 14672 14 524 KSL 27.99h Lab
+GSL SD/SI (2) 2019 CSLR/SLT RGBD 10295 7 310 GSL 9.59h Lab
+PHOENIX-2014-T1 (27) 2018 CSLR/SLT RGB 8257 9 1066 DGS - TV
+PHOENIX-2014 (49) 2014 CSLR/SLT RGB - 9 1558 DGS 10.73h TV
+HKSL (233) 2021 CSLR RGBD 2400 6 55 HKSL - Lab
+ChicagoFSWild+ (179) 2019 CSLR RGB 55232 260 26 ASL - Web
+Multi-VSL (41) 2025 IDSLR RGB, Pose 84764 30 1000 Vietnamese - Lab
+BdSLW60 (168) 2025 IDSLR RGB 9307 18 60 Bangla - Lab
+ISLR101 (160) 2025 IDSLR RGB, Pose 4614 10 101 Iranian - -
+ISW-1000 (165) 2025 IDSLR RGB, Pose, OF 10000 10 1000 CSL - Lab
+FDMSE-ISL (152) 2024 IDSLR RGBD 40033 20 2002 ISL 36h Lab
+LSA64 (167) 2023 IDSLR RGB 3200 10 64 LSA 1.9h Lab
+Signsuisse Lexicon (140) 2023 IDSLR RGB 18221 - - 3 (Mult) - Web
+CSLD (202) 2022 IDSLR RGB 120000 30 400 CSL 66.8h Lab
+NCSL (200) 2022 IDSLR RGB 90000 30 300 CSL 50.1h Lab
+KArSL (182) 2021 IDSLR RGBD, Pose 75300 3 502 ArSL - Lab
+NMFs-CSL (80) 2021 IDSLR RGB 32010 10 1.067 CSL 24.4h Lab
+BosphorusSign22k (150) 2020 IDSLR RGBD, Pose 22542 6 744 TSL 19h Lab
+AUTSL (183) 2020 IDSLR RGBD, Pose 38336 43 226 TSL 21h Lab
+INCLUDE (188) 2020 IDSLR RGB 4287 7 263 ISL 3h Lab
+IISL2020 (103) 2020 IDSLR RGB 12100 16 11 ISL 7h Lab
+RKS-PERSIANSIGN (163) 2020 IDSLR RGB 10000 10 100 Persian - Lab
+BSL-1K (11) 2020 IDSLR RGB 280000 40 1064 BSL 1060h TV
+MSASL (94) 2019 IDSLR RGB 25513 222 1000 ASL 24.65h Web
+WLASL (110) 2019 IDSLR RGB 21083 119 2000 ASL 14.11h Web
+GSL isol. (2) 2019 IDSLR RGBD 40785 7 310 GSL 6.44h Lab
+SMILE (45) 2018 IDSLR RGB - 30 100 DSGS - Lab
+ASL Dataset (17) 2018 IDSLR Pose 1200 20 30 ASL - Lab
+CSL (87) 2018 IDSLR RGBD, Pose 125000 50 500 CSL - Lab
 8.2 Semaphoric HGR
@@
 SHGR datasets target gesture recognition from single frames, making them easier to process due to the lack of temporal complexity.
-Older datasets like HGR-1 (HGR-1-2024-PATTER-RECOGNITION), NUS-I (NUS-I-2010), NUS-II (NUS-II-2013), and OUHANDS (ouhands) contain few samples, limiting deep learning training, yet remain benchmarks for HCI due to their simplicity.
-LaRED (LaRED) offers 243K RGBD samples across 81 gestures, but only from 10 subjects, restricting subject-independent recognition.
-HaGRID (WACV_kapitanov2024hagrid) is much larger and diverse, with nearly twice as many full-resolution RGB frames from  35K subjects in varied scenes, including bounding boxes for hand detection.
-HaGRIDv2 (HAGRIDv2_ARXIV) expands this with 15 additional classes,  66K subjects, and over 1M samples.
-SHAPE (SHAPE_DATASET) has a similar number of gestures (all two-handed) but fewer samples.
-LRHG (LONG_RANGE_HGR_SSD) and MD-UHGRD (MD-UHGRD_SA-YOLO) target UAV-based gesture recognition, with MD-UHGRD offering more samples but fewer classes, captured from varying distances and angles to enhance viewpoint robustness.
-URGR (ULTRA-RANGE-UGV-Bamani) focuses on ultra-range gestures for UGVs, with 6 robot-command gestures performed by 16 signers at distances from 2 to 25 meters.
+Older datasets like HGR-1 (98), NUS-I (106), NUS-II (153), and OUHANDS (132) contain few samples, limiting deep learning training, yet remain benchmarks for HCI due to their simplicity.
+LaRED (74) offers 243K RGBD samples across 81 gestures, but only from 10 subjects, restricting subject-independent recognition.
+HaGRID (95) is much larger and diverse, with nearly twice as many full-resolution RGB frames from  35K subjects in varied scenes, including bounding boxes for hand detection.
+HaGRIDv2 (145) expands this with 15 additional classes,  66K subjects, and over 1M samples.
+SHAPE (35) has a similar number of gestures (all two-handed) but fewer samples.
+LRHG (232) and MD-UHGRD (213) target UAV-based gesture recognition, with MD-UHGRD offering more samples but fewer classes, captured from varying distances and angles to enhance viewpoint robustness.
+URGR (18) focuses on ultra-range gestures for UGVs, with 6 robot-command gestures performed by 16 signers at distances from 2 to 25 meters.
 IDHGR datasets contain gestures performed in isolation, with start and end points explicitly segmented.
 Most of the existing datasets (e.g.
-SKIG (SKIG), NVIDIA (aka nvGesture) (NVIDIA_GESTURE_R3DCNN_Molchanov), DHG (DHG-CVPRW) and SHREC17 (SHREC)) for IDHGR are relatively small with a few thousands samples and support multimodal settings by providing depth maps, IR frames, optical flow or skeleton joints, accompanying the RGB modality aiming to enhance the accuracy of the models.
-Jester-20bn (Jester-Dataset), proposed for gesture-based HCI, is the largest dataset for (non-SLU) IDHGR containing 1̃50K videos performed by 1376 subjects.
+SKIG (121), NVIDIA (aka nvGesture) (138), DHG (39) and SHREC17 (185)) for IDHGR are relatively small with a few thousands samples and support multimodal settings by providing depth maps, IR frames, optical flow or skeleton joints, accompanying the RGB modality aiming to enhance the accuracy of the models.
+Jester-20bn (131), proposed for gesture-based HCI, is the largest dataset for (non-SLU) IDHGR containing 1̃50K videos performed by 1376 subjects.
 The remarkable diversity and size of this dataset make it ideal for pretraining DL models.
-Other datasets for HCI include IPN-Hand (IPN-HAND), ZJUGesture (ZJUGesture) and DATE (DATE), comprising 4218, 9892 and 13500 samples respectively.
-LD-ConGR (LD-ConGR), ChaLearn IsoGD (chalearn-lap) and EgoGesture (EgoGesture_DATASET) have sufficient amount of data, constituting adequate baseline for IDHGR.
-EgoDriving (EgoFormer_METHOD_EgoDriving_DATASET), Briareo (Briareo_dataset) and nvGesture (NVIDIA_GESTURE_R3DCNN_Molchanov) have been proposed for human-car interaction, mainly in multimodal setting, including gestures that correspond to car functionalities (e.g.
+Other datasets for HCI include IPN-Hand (20), ZJUGesture (212) and DATE (37), comprising 4218, 9892 and 13500 samples respectively.
+LD-ConGR (118), ChaLearn IsoGD (197) and EgoGesture (220) have sufficient amount of data, constituting adequate baseline for IDHGR.
+EgoDriving (155), Briareo (128) and nvGesture (138) have been proposed for human-car interaction, mainly in multimodal setting, including gestures that correspond to car functionalities (e.g.
 CW rotation and audio manipulation [^2] https://youtu.be/NJmk1DUyyB8).
-TCG (TCG) was also proposed for human-car interaction, but for recognizing traffic police gestures in autonomous driving scenarios relying only on pose data.
-EHWGesture (EHWGesture-amprimo2025ehwgesture) is the first dataset for recognizing clinical gestures and providing annotations for many other tasks, including gesture detection & triggering.
+TCG (208) was also proposed for human-car interaction, but for recognizing traffic police gestures in autonomous driving scenarios relying only on pose data.
+EHWGesture (16) is the first dataset for recognizing clinical gestures and providing annotations for many other tasks, including gesture detection & triggering.
 CDHGR datasets challenge models to identify and segment gestures from continuous video streams.
 Many of the datasets mentioned earlier (e.g.
-ChaLearn (chalearn-lap), IPN-Hand (IPN-HAND), LD-ConGR (LD-ConGR), EgoGesture (EgoGesture_DATASET) and ZJUGesture (ZJUGesture)) have been obtained by annotating the temporal boundaries of gestures within continuous sequences, thus these datasets have protocols for both IDHGR and CDHGR.
-Therefore, we will only discuss the Police Gesture Dataset (Police-Gesture-Dataset) as the previous analysis could also be applied here.
-This dataset was introduced for recognizing police traffic gestures, similarly with TCG (TCG), containing a small number of samples (21 RGB videos) and 3354 gesture instances.
+ChaLearn (197), IPN-Hand (20), LD-ConGR (118), EgoGesture (220) and ZJUGesture (212)) have been obtained by annotating the temporal boundaries of gestures within continuous sequences, thus these datasets have protocols for both IDHGR and CDHGR.
+Therefore, we will only discuss the Police Gesture Dataset (73) as the previous analysis could also be applied here.
+This dataset was introduced for recognizing police traffic gestures, similarly with TCG (208), containing a small number of samples (21 RGB videos) and 3354 gesture instances.
 Each dataset supports specific sub-tasks within HGR research: static datasets aid in gesture classification, isolated dynamic datasets support temporal gesture modeling and classification, and continuous datasets allow for joint gesture spotting and classification—crucial for real-time and interactive applications.
@@
 Name Year Domain Modality Samples Subjects Classes Resolution Distance
-HaGRID (WACV_kapitanov2024hagrid) 2024 HCI RGB 552992 34730 19 1920x1080 0.5-4m
-HaGRIDv2 (HAGRIDv2_ARXIV) 2024 HCI RGB 1086158 65977 34 1920x1080 0.5-4m
-MD-UHGRD (MD-UHGRD_SA-YOLO) 2024 HRI RGB 20000 - 19 - 3-10m
-URGR (ULTRA-RANGE-UGV-Bamani) 2024 HRI RGB 347483 16 6 640x480 1-25m
-SHAPE (SHAPE_DATASET) 2022 - RGB 34226 20 32 4128x3096 2-5m
-LRHG (LONG_RANGE_HGR_SSD) 2021 HRI RGB 4320 8 10 640x480 1-7m
-OUHANDS (ouhands) 2016 HCI RGBD 3150 23 10 640x480 $\leq$ 1m
-HGR-1 (HGR-1-2024-PATTER-RECOGNITION) 2014 - RGB 899 12 27 174x131-640x480 $\leq$ 1m
-LaRED (LaRED) 2014 - RGBD 243000 10 81 640x480 -
-NUS-II (NUS-II-2013) 2013 - RGB 2000 40 10 160x120 $\leq$ 1m
-ASL-FS (ASL-FS-2011-CVPRW) 2011 - RGBD 48K 4 24 - -
-NUS-I (NUS-I-2010) 2010 - RGB 240 - 10 160x120 $\leq$ 1m
+HaGRID (95) 2024 HCI RGB 552992 34730 19 1920x1080 0.5-4m
+HaGRIDv2 (145) 2024 HCI RGB 1086158 65977 34 1920x1080 0.5-4m
+MD-UHGRD (213) 2024 HRI RGB 20000 - 19 - 3-10m
+URGR (18) 2024 HRI RGB 347483 16 6 640x480 1-25m
+SHAPE (35) 2022 - RGB 34226 20 32 4128x3096 2-5m
+LRHG (232) 2021 HRI RGB 4320 8 10 640x480 1-7m
+OUHANDS (132) 2016 HCI RGBD 3150 23 10 640x480 $\leq$ 1m
+HGR-1 (98) 2014 - RGB 899 12 27 174x131-640x480 $\leq$ 1m
+LaRED (74) 2014 - RGBD 243000 10 81 640x480 -
+NUS-II (153) 2013 - RGB 2000 40 10 160x120 $\leq$ 1m
+ASL-FS (154) 2011 - RGBD 48K 4 24 - -
+NUS-I (106) 2010 - RGB 240 - 10 160x120 $\leq$ 1m
 Table 8: Datasets for IDHGR.
 Name Year Domain Modality Samples Subjects Classes Frames
-EHWGesture (EHWGesture-amprimo2025ehwgesture) 2025 - RGBD 1100 25 11 3.3M
-EgoDriving (EgoFormer_METHOD_EgoDriving_DATASET) 2024 HMI RGB 400 - 29 70276
-DATE (DATE) 2024 HCI RGB 13500 22 27 -
-ZJUGesture (isol.) (ZJUGesture) 2023 HCI RGB 9892 60 9 -
-LD-ConGR (isol.) (LD-ConGR) 2022 HCI RGBD 44887 30 10 -
-IPN Hand (isol.) (IPN-HAND) 2021 HCI RGB 4218 50 13 -
-TCG (TCG) 2020 HMI Pose 250 5 4 839350
-Jester (Jester-Dataset) 2019 HCI RGB 148092 1376 27 5331312
-Briareo (Briareo_dataset) 2019 HMI RGBD, Pose, IR 1440 40 12 -
-EgoGesture (isol.) (EgoGesture_DATASET) 2018 HCI RGBD 24161 50 83 2953224
-SHREC17-14/28 (SHREC) 2017 - Depth, Pose 2800 28 14/28 -
-nvGesture (NVIDIA_GESTURE_R3DCNN_Molchanov) 2016 HMI RGBD, IR, OF 1532 20 25 -
-ChaLearn IsoGD (chalearn-lap) 2016 - RGBD 47933 21 249 -
-DHG-14/28 (DHG-CVPRW) 2016 - Depth, Pose 2800 20 14/28 -
-SKIG (SKIG) 2013 - RGBD 1080 6 10 -
+EHWGesture (16) 2025 - RGBD 1100 25 11 3.3M
+EgoDriving (155) 2024 HMI RGB 400 - 29 70276
+DATE (37) 2024 HCI RGB 13500 22 27 -
+ZJUGesture (isol.) (212) 2023 HCI RGB 9892 60 9 -
+LD-ConGR (isol.) (118) 2022 HCI RGBD 44887 30 10 -
+IPN Hand (isol.) (20) 2021 HCI RGB 4218 50 13 -
+TCG (208) 2020 HMI Pose 250 5 4 839350
+Jester (131) 2019 HCI RGB 148092 1376 27 5331312
+Briareo (128) 2019 HMI RGBD, Pose, IR 1440 40 12 -
+EgoGesture (isol.) (220) 2018 HCI RGBD 24161 50 83 2953224
+SHREC17-14/28 (185) 2017 - Depth, Pose 2800 28 14/28 -
+nvGesture (138) 2016 HMI RGBD, IR, OF 1532 20 25 -
+ChaLearn IsoGD (197) 2016 - RGBD 47933 21 249 -
+DHG-14/28 (39) 2016 - Depth, Pose 2800 20 14/28 -
+SKIG (121) 2013 - RGBD 1080 6 10 -
 Table 9: Datasets for CHGR.
 Name Year Domain Modality Samples Instances Subjects Classes
-ZJUGesture (cont.) (ZJUGesture) 2023 HCI RGB - 9892 60 9
-LD-ConGR (cont.) (LD-ConGR) 2022 HCI RGBD 542 44887 30 10
-IPN Hand (cont.) (IPN-HAND) 2021 HCI RGB 200 4218 50 13
-Police Gesture Dataset (Police-Gesture-Dataset) 2020 HMI RGB 21 3354 - 8
-EgoGesture (cont.) (EgoGesture_DATASET) 2018 HCI RGBD 2081 24161 50 83
-ChaLearn ConGD (chalearn-lap) 2016 - RGBD 22535 47933 21 249
+ZJUGesture (cont.) (212) 2023 HCI RGB - 9892 60 9
+LD-ConGR (cont.) (118) 2022 HCI RGBD 542 44887 30 10
+IPN Hand (cont.) (20) 2021 HCI RGB 200 4218 50 13
+Police Gesture Dataset (73) 2020 HMI RGB 21 3354 - 8
+EgoGesture (cont.) (220) 2018 HCI RGBD 2081 24161 50 83
+ChaLearn ConGD (197) 2016 - RGBD 22535 47933 21 249
 8.3 Other
 In addition to standard semaphoric gesture datasets, several specialized datasets have been introduced to support research on more specific gesture types.
-The First-Person Hand Action (FPHA) dataset (FPHA), published in 2018, contains 1,175 video sequences captured from a first-person perspective by six participants.
+The First-Person Hand Action (FPHA) dataset (55), published in 2018, contains 1,175 video sequences captured from a first-person perspective by six participants.
 It focuses on isolated manipulative hand gestures across 45 action classes, offering RGB, depth, and skeleton data, and is particularly valuable for studies involving egocentric interaction.
-Another notable example is the DP Dataset (DeePoint), released in 2023, which, to the best of our knowledge, is the only publicly available benchmark specifically designed for pointing gesture recognition.
+Another notable example is the DP Dataset (141), released in 2023, which, to the best of our knowledge, is the only publicly available benchmark specifically designed for pointing gesture recognition.
 Collected from 33 participants in two indoor environments, the dataset comprises approximately 2.8 million frames and 6,335 annotated gesture instances.
@@
 non-gesture) and directional estimation tasks.
-Finally, the SocialGesture (SocialGesture-CVPR2025-cao2025socialgesture) focuses on the recognition of natural gestures in multi-person environments.
+Finally, the SocialGesture (28) focuses on the recognition of natural gestures in multi-person environments.
 It comprises 9,889 in-the-wild videos containing 42,533 instances and provides fine-grained annotations including bounding boxes, relations (for example, a subject is pointing to a target) and Visual Question Answering (VQA) details.
@@
 After examining the existing datasets for VHGR with significant impact on the current research, in this subsection we aim to determine pros and cons of the most important ones as well as the use cases that they fit.
-For gesture-based HCI, HaGRID (WACV_kapitanov2024hagrid) and its versions (HAGRIDv2_ARXIV) seem to be a reliable solution, containing an enormous number of samples corresponding to 19 distinct classes (that can be mapped to computer commands).
+For gesture-based HCI, HaGRID (95) and its versions (145) seem to be a reliable solution, containing an enormous number of samples corresponding to 19 distinct classes (that can be mapped to computer commands).
 Regarding that these gestures are also static, even simple architectures - like the ones discussed in sections 4 & 7 - can achieve sufficient performance.
 Furthermore, this dataset could also be used for pretraining, enhancing the backbone’s expression capability.
-However, its size implies that training takes a long time, which constitutes a deterrent in some cases, compared to other datasets for HCI like OUHANDS (ouhands).
-For gesture-based HRI, URGR (ULTRA-RANGE-UGV-Bamani) is the only one for UGV guidance and considers various recognition ranges, which may be crucial for real-world scenarios.
+However, its size implies that training takes a long time, which constitutes a deterrent in some cases, compared to other datasets for HCI like OUHANDS (132).
+For gesture-based HRI, URGR (18) is the only one for UGV guidance and considers various recognition ranges, which may be crucial for real-world scenarios.
 Even though it contains a large number of samples, the small number of classes may limit the commands given by gestures and, therefore, the functionalities of the entire gesture-based interface.
-Regarding SLU datasets, for ASL recognition, OpenASL (OpenASL) is the largest one, including almost 100K RGB videos (288h in total) collecting from the web.
-MSASL (MS-ASL) comprises almost the same number of subjects, but it is much smaller.
+Regarding SLU datasets, for ASL recognition, OpenASL (178) is the largest one, including almost 100K RGB videos (288h in total) collecting from the web.
+MSASL (94) comprises almost the same number of subjects, but it is much smaller.
 Furthermore, OpenASL was proposed for CSLR while for MSASL for IDSLR, thus it may be more suitable for real-world use cases, where the temporal boundaries are not known.
-BSL-1K (bsl-1K) is also a very large dataset, proposed for British IDSLR and comprising the largest number of samples (280K, 1060h in total).
-BOBSL (BOBSL-BBC) was proposed for British CSLR and comprises videos of 1467h in total making it the dataset with the longest duration.
-Phoenix (PHOENIX-2014) and Phoenix T (NEURAL-SIGN-LANGUAGE-TRANSLATION-CVPR-CAMGOZ) remain the most adequate dataset for German sign language, covering a large sign vocabulary.
+BSL-1K (11) is also a very large dataset, proposed for British IDSLR and comprising the largest number of samples (280K, 1060h in total).
+BOBSL (12) was proposed for British CSLR and comprises videos of 1467h in total making it the dataset with the longest duration.
+Phoenix (49) and Phoenix T (27) remain the most adequate dataset for German sign language, covering a large sign vocabulary.
 Nevertheless, the small number of samples and the low diversity - due to the limited number of signers and the constrained environment - make essential the collection of a new, larger and more diverse dataset for DGS.
-For Chinese SLR, NMFs-CSL (NMFs-CSL) is significantly smaller than CSLD (tim-slr_csld_dataseta), CSL (CSL500) and NCSL (2plus1-D-SLR_and_NCSL) but includes much more signs (1064 instead of less than 500).
-CSL-Daily (Back-Translation-CVPR) remains the only one for Chinese CSLR, covering a large sign vocabulary of 2000 signs in total.
+For Chinese SLR, NMFs-CSL (80) is significantly smaller than CSLD (202), CSL (87) and NCSL (200) but includes much more signs (1064 instead of less than 500).
+CSL-Daily (230) remains the only one for Chinese CSLR, covering a large sign vocabulary of 2000 signs in total.
 9 Evaluation Metrics
@@
 For continuous HGR (CHGR), particularly in Continuous Sign Language Recognition (CSLR), sequence-level metrics are used.
-Word Error Rate (WER) (VAC-CSLR; ADADOGLOU) is the standard metric, measuring the minimum number of substitutions ($S$), deletions ($D$), and insertions ($I$) needed to transform the predicted sequence into the ground truth sequence of length $N$:
+Word Error Rate (WER) (136; 2) is the standard metric, measuring the minimum number of substitutions ($S$), deletions ($D$), and insertions ($I$) needed to transform the predicted sequence into the ground truth sequence of length $N$:
 $WER=\frac{S+D+I}{N}$ (1)
-Edit Distance Accuracy (FA-STGCN; CTCX_WANG2022123) is a complementary metric representing the ratio of correctly predicted elements in the sequence:
+Edit Distance Accuracy (50; 204) is a complementary metric representing the ratio of correctly predicted elements in the sequence:
 $\text{Edit-Accuracy}=\frac{N-S-D-I}{N}$ (2)
@@
 The presence of gesture-irrelevant factors (or redundant information) hinders accurate recognition in diverse real-world scenarios.
-Some of these factors include the following (EDenseNet; Mohammad_Mahmudul_static_hgr; 2DPSTPP; ULTRA-RANGE-UGV-Bamani; EnGesto; RAAR3DNet):
+Some of these factors include the following (189; 9; 89; 18; 158; 227):
   • 1.
@@
     Input data quality: Low resolution and blurriness may corrupt the discriminative characteristics of the hand gesture region.
-Aiming to address these challenges, researchers have proposed segmentation models to crop hand regions and remove background interference (FGDSNet; tuan_linh_pham_dang2023lightweight; DRCAM_TIM), object detectors for spatial normalization (ULTRA-RANGE-UGV-Bamani; Mohammad_Mahmudul_static_hgr), super-resolution techniques to enhance image quality (ULTRA-RANGE-UGV-Bamani), attention mechanisms (RAAR3DNet; Benjia_Zhou) to enforce the visual encoder to pay attention on the most informative regions, information-theoretic losses (Yunan_Li), and feature selection techniques (SeST) to reduce information redundancy.
+Aiming to address these challenges, researchers have proposed segmentation models to crop hand regions and remove background interference (229; 36; 171), object detectors for spatial normalization (18; 9), super-resolution techniques to enhance image quality (18), attention mechanisms (227; 228) to enforce the visual encoder to pay attention on the most informative regions, information-theoretic losses (112), and feature selection techniques (191) to reduce information redundancy.
 Many of these approaches rely on computationally expensive off-the-shelf models, which increase the complexity of the overall HGR methods and set an upper bound to their accuracy.
-The approaches mentioned previously may be enhanced by using eXplainable Artificial Intelligence (XAI) (Rodis2024Access) techniques, which could be adopted to understand better the behavior of the networks.
-For instance, Grad-CAM could be further utilized to determine the image region that affects more the predicted gesture class (as has already been done by a few researchers: (CorrNet; SWIN_MSTP_Neurocomputing_2025_CSLR)).
+The approaches mentioned previously may be enhanced by using eXplainable Artificial Intelligence (XAI) (166) techniques, which could be adopted to understand better the behavior of the networks.
+For instance, Grad-CAM could be further utilized to determine the image region that affects more the predicted gesture class (as has already been done by a few researchers: (84; 15)).
 Such results may indicate the effectiveness of the embedded on the network attention mechanisms or other related modules to emphasize the most important parts of the input image.
@@
 Even lightweight methods for SHGR and IDHGR achieve high accuracy.
-Although, for the more complex task of CHGR, the cost is essential; all RGB-based methods (that are significantly more accurate than pose-based) require more than 366.3 GFLOPs ((DSTEN-CSLR-yu2025dsten)) and reach up to 2950.5 GFLOPs ((TemporalLiftPooling)).
+Although, for the more complex task of CHGR, the cost is essential; all RGB-based methods (that are significantly more accurate than pose-based) require more than 366.3 GFLOPs ((214)) and reach up to 2950.5 GFLOPs ((83)).
 This issue prevents not only the deployment of such methods but also the efficient training, which takes longer.
-Pose-based methods (CoSign) are significantly lighter, but the pose estimation cost is not taken into account.
+Pose-based methods (92) are significantly lighter, but the pose estimation cost is not taken into account.
 The number of parameters does not change dramatically between IDHGR and CHGR methods, while many SHGR methods require more parameters (e.g.
-Gesture-CNN (G-CNN-SHARMA2021115657) has 67.25M parameters).
+Gesture-CNN (176) has 67.25M parameters).
 The need of large computational resources does not only prevent the real-time and commercial deployment of VHGR systems, but also put barriers to the evolution of the field, as the conducted experiments take many hours or even days to complete.
-To mitigate this, authors have focused on reducing the complexity of their models by using non-parametric layers (GestFormer) or designing smaller architectures (LHGR-Net).
+To mitigate this, authors have focused on reducing the complexity of their models by using non-parametric layers (56) or designing smaller architectures (218).
 Thus, reducing the total complexity of the models, could be further examined in the future.
-Considering also recently proposed approaches for DL models compression (DL_MODEL_COMPRESSION_marino2023deep), such as pruning (i.e., removing less important weights) and quantization (i.e., reducing the arithmetic precision of the parameters), may contribute to this direction.
+Considering also recently proposed approaches for DL models compression (129), such as pruning (i.e., removing less important weights) and quantization (i.e., reducing the arithmetic precision of the parameters), may contribute to this direction.
 Freezing layers of pretrained models could also be investigated in order to reduce the time required for fine-tuning.
@@
 Many existing datasets either contain a relatively small number of samples or lack sufficient diversity.
-For example, among static HGR datasets, only the versions of HaGRID (WACV_kapitanov2024hagrid; HAGRIDv2_ARXIV) have been collected from more than one hundred participants, despite some others having a plausible number of samples (LaRED; ULTRA-RANGE-UGV-Bamani).
-A similar situation is observed in IDHGR datasets, where only the Jester dataset (Jester-Dataset) includes data from over one hundred subjects.
-Furthermore, the amount of samples for specific modalities, such as IR data, is limited compared with the size of RGB-based datasets, an issue known as data imbalance (IR_data_imbalance_mobilenet).
+For example, among static HGR datasets, only the versions of HaGRID (95; 145) have been collected from more than one hundred participants, despite some others having a plausible number of samples (74; 18).
+A similar situation is observed in IDHGR datasets, where only the Jester dataset (131) includes data from over one hundred subjects.
+Furthermore, the amount of samples for specific modalities, such as IR data, is limited compared with the size of RGB-based datasets, an issue known as data imbalance (199).
 The lack of large-scale, diverse datasets leads to overfitting and limited generalization, which is closely related to the challenges discussed earlier.
@@
   • 1.
-    Multilingual Data Aggregation: Combining datasets (e.g., CSL and DGS) to exploit cross-lingual visual similarities increases the overall amount of training data (Fangyun_Wei_ICCV).
+    Multilingual Data Aggregation: Combining datasets (e.g., CSL and DGS) to exploit cross-lingual visual similarities increases the overall amount of training data (207).
   • 2.
-    Self-Supervised Learning: In the context of the self-supervised learning setting (konstantakos2025self), gloss-to-gloss translation tasks have been used for parameter initialization (CVT-SLR), providing linguistic priors to sign language models.
-Pretraining with pose reconstruction has also been investigated to incorporate prior knowledge of hand structure (SignBERT+; SignBERT_ICCV; omar_ikne).
+    Self-Supervised Learning: In the context of the self-supervised learning setting (102), gloss-to-gloss translation tasks have been used for parameter initialization (226), providing linguistic priors to sign language models.
+Pretraining with pose reconstruction has also been investigated to incorporate prior knowledge of hand structure (77; 78; 88).
   • 3.
-    Advanced Data Augmentation: Adversarial learning has been employed to optimize augmentation hyperparameters (ADVERSARIAL_DATA_AUGM), making the models robust to spatial deformations.
+    Advanced Data Augmentation: Adversarial learning has been employed to optimize augmentation hyperparameters (142), making the models robust to spatial deformations.
   • 4.
-    Few-/Zero-shot Learning: Approaches using only a small number of samples to correctly recognize gestures—often based on CLIP—have also been proposed (ZSSLR; FSSLR; ZS-DHGR; ZS-GR; ZERO-SHOT-SLR).
-Generative AI, particularly diffusion models (DMs) for image augmentation (alimisis2025advances), could help address data scarcity in HGR.
+    Few-/Zero-shot Learning: Approaches using only a small number of samples to correctly recognize gestures—often based on CLIP—have also been proposed (147; 25; 162; 164; 23).
+Generative AI, particularly diffusion models (DMs) for image augmentation (13), could help address data scarcity in HGR.
 Existing datasets are often limited in size or focused on common tasks, like well-known sign languages.
@@
 Training the visual encoder effectively remains difficult, especially in deep architectures for CSLR.
-As shown in CTCA (CTCA), (Fig.
+As shown in CTCA (63), (Fig.
 9), increasing the depth of the temporal module (e.g., 1D-TCN followed by BiLSTM) improves its generalization but hampers the encoder’s learning due to gradient attenuation in backpropagation.
-Furthermore, the CTC spike phenomenon—where the model over-focuses on a few keyframes—can cause rapid overfitting and limit encoder training (VAC-CSLR; SMKD).
-Figure 9: Experimental results in (CTCA) analyze the effect of the temporal aggregation module’s depth—a 1D-TCN with varying layers followed by a two-layer BiLSTM—on training the visual encoder (spatial perception module).
+Furthermore, the CTC spike phenomenon—where the model over-focuses on a few keyframes—can cause rapid overfitting and limit encoder training (136; 69).
+Figure 9: Experimental results in (63) analyze the effect of the temporal aggregation module’s depth—a 1D-TCN with varying layers followed by a two-layer BiLSTM—on training the visual encoder (spatial perception module).
 In both charts, the x-axis shows the number of 1D-TCN layers.
-The top chart plots information stored in weights (IIW) (pac-bayes-wang2022pacbayes), indicating generalization ability, while the bottom shows WER.
+The top chart plots information stored in weights (IIW) (206), indicating generalization ability, while the bottom shows WER.
 Increasing 1D-TCN layers improves short-term temporal generalization but reduces overall performance due to undertrained visual encoder weights.
@@
   • 1.
-    Auxiliary Supervision: Adding early-stage classifiers to the encoder or short-term module has been proposed as a strategy to guide training by providing additional supervision (VAC-CSLR; CSGC_Rao_Sun_Wang_Wang_Zhang_2024; CorrNet; corrnetplus_hu2024corrnetsignlanguagerecognition; TCNet_Lu_Salah_Poppe_2024; C2SLR).
+    Auxiliary Supervision: Adding early-stage classifiers to the encoder or short-term module has been proposed as a strategy to guide training by providing additional supervision (136; 161; 84; 82; 124; 234).
   • 2.
-    Temporal Module Redesign: Placing the 1D-TCN and BiLSTM in parallel instead of sequentially (CTCA).
+    Temporal Module Redesign: Placing the 1D-TCN and BiLSTM in parallel instead of sequentially (63).
   • 3.
-    Linguistic Constraints and Gloss Segmentation: Using gloss boundaries (SMKD) or prior knowledge (GPGN) to regularize learning.
-The aforementioned approaches —particularly auxiliary supervision (VAC-CSLR)— have served as the foundation for many existing methods.
+    Linguistic Constraints and Gloss Segmentation: Using gloss boundaries (69) or prior knowledge (64) to regularize learning.
+The aforementioned approaches —particularly auxiliary supervision (136)— have served as the foundation for many existing methods.
 However, the increasing complexity of modern deep architectures often exceeds the effectiveness of these techniques.
@@
     Block-fixed models: The existing architectures are based on hand-crafted design, consisting of block of layers with hyperparameters (e.g., number of channels, convolution kernel size) that are defined manually by the researchers.
-This approach is suboptimal (RAAR3DNet), as several other possible configurations have not been explored.
+This approach is suboptimal (227), as several other possible configurations have not been explored.
   • 2.
     Insufficient graph representations of GCNs: The standard GCNs utilized fixed topology graphs, often based on the shape of the hand, to better capture the interactions between the various skeleton joints.
-However, this configuration has been proved to be suboptimal (TD-GCN), unable to efficiently capture the temporal dynamics.
+However, this configuration has been proved to be suboptimal (120), unable to efficiently capture the temporal dynamics.
   • 3.
-    Ineffective multimodal fusion: Out of the existing multimodal methods, most of them perform late fusion to incorporate the information of multiple streams (SignBERT_ICCV; SignBERT+; BEST; Weichao-Zhao; MASA; SKIM; SAM_SLR; jiang2021sign-sam-slr-v2).
+    Ineffective multimodal fusion: Out of the existing multimodal methods, most of them perform late fusion to incorporate the information of multiple streams (78; 77; 224; 225; 223; 115; 91; 90).
 This scheme does not fully leverage the available modalities, as the features extracted by each stream are isolated from the others during the processing on the backbones.
 Thus, the obtained vector representations do not benefit from the utilization of multiple modalities.
-To address these limitations, researchers have proposed various strategies, including Neural Architecture Search (NAS) to identify optimal model structures through searching on a large set of configurations (RAAR3DNet), learnable or time-dependent skeleton adjacency matrices that determine implicitly the optimal way to integrate the features of each joint (DSTA_SLR_hu2024dynamic; DSTSA-CGN_cui2025dstsa; ST-PSM_L-PSM; TD-GCN) and multi-level fusion for dual-stream approaches, which enhances the quality of the extracted features by early incorporating information from various sources (CTFB_Hampiholi).
+To address these limitations, researchers have proposed various strategies, including Neural Architecture Search (NAS) to identify optimal model structures through searching on a large set of configurations (227), learnable or time-dependent skeleton adjacency matrices that determine implicitly the optimal way to integrate the features of each joint (86; 32; 30; 120) and multi-level fusion for dual-stream approaches, which enhances the quality of the extracted features by early incorporating information from various sources (66).
 Despite these advances, several aspects of deep learning architectures remain underexplored, leaving a substantial research gap.
@@
 Thus, future work could consider the intermediate fusion of three or more streams, e.g.
-using fusion modules (MMTM; CTFB_Hampiholi) and identifying the optimal way to connect the different streams.
+using fusion modules (195; 66) and identifying the optimal way to connect the different streams.
 11 Conclusions
@@
 Thus, VHGR could be expected to open-vocabulary settings, enabling deployment to broader applications.
-The potential of leveraging the extensive knowledge of foundation models (kas2025foundation_models) in the context of VHGR enabling zero-shot gesture recognition as well as the utilization of model compression techniques (DL_MODEL_COMPRESSION_marino2023deep) to reduce the computational complexity are also rarely explored, leaving space for future work.
+The potential of leveraging the extensive knowledge of foundation models (97) in the context of VHGR enabling zero-shot gesture recognition as well as the utilization of model compression techniques (129) to reduce the computational complexity are also rarely explored, leaving space for future work.
 Acknowledgments
@@
 Conflict of interest The authors have no competing interests to declare that are relevant to the content of this article.
-Author contribution Konstantinos Foteinos, Jorgen Cani, Manousos Linardakis, Iraklis Varlamis and Georgios Th.
+Author contribution Konstantinos Foteinos, Manousos Linardakis, Iraklis Varlamis and Georgios Th.
 Papadopoulos performed the literature review and prepared a draft of the manuscript.